A Complete Bibliography of Publications in Studies in History and Philosophy of Science Part C: Studies in History and Philosophy of

Biological and Biomedical Sciences”

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
     WWW URL: https://www.math.utah.edu/~beebe/

08 August 2024 Version 1.36

Title word cross-reference

+ [Cla07a]. / = [Nic13]. $19.95 [Tur18]. $25 [Oka03]. $27.95 [Har13a]. 3 [Bra11b]. $39.95 [DiM18]. $40.00 [Ayl18]. $45.00 [Ren15]. $58.99 [Tor13]. 1 [Fis15].

-dimensions [Bra11b]. -I [Ste08]. -no [Ste08].

0 [Dea10, Oka03, Ren15, Tor13, Whi04]. 0-14-017309-9 [Whi04]. 0-19-508820-4 [Sar98]. 0-19-508821-2 [Sar98]. 0-226-20791-9 [Dea10]. 0-226-90134-3 [Oka03]. 0-262-01167-0 [Aga01]. 0-521-49544-X [Gur98]. 0-7619-5425-2 [Mur01]. 0-7619-5426-0 [Mur01]. 0-7914-1260-1 [Wil98].

1 [Alt07, Cob15]. 17.99 [Mur01]. 1840s [BL07]. 1859-1914 [Coh17]. 1875-1944 [FM07]. 1876 [FS15]. 19.95 [Sar98]. 1920s [Erl09]. 1930s [OG14]. 1940s [Tur20]. 1950s [Löw12, Mor14, Qui05, Iid20]. 1960s [Hag01, Kel11b, Qui05, SD14]. 1970s [Fay12]. 1980s [Ser11, WL08, Wyl10]. 1990s [WL08, Wyl10]. 19th [Gis20]. 1The [Hag01].

2 [Alt08b, DiM18, Hal18, Qui04, TB16]. 2-7071-3607-7 [Qui04]. 2005 [Sha14]. 2006 [Ano07a]. 2009 [Ano09a]. 2019 [Ano19a, Ano19b, Ano19i]. 2020 [Ano20a, Ano20b, Ano20c, Ano20j, Ano20k]. 20th [BP15, Gis20, Sko11]. 20th-century [BP15]. 21st [Ode14].

3 [Gut18, Oka03, Tay15]. 31.95 [Aga01]. 33C [Wil03]. 33C/2 [Wil03]. 34-2 [Ano03a].

5 [Ano13a, Ayl18].

7 [Qui04, Tur18].

8 [Har13a].

9 [Dea10, Whi04]. 978 [Ayl18, Cob15, DiM18, Gut18, Hal18, Har13a, Ren15, Tay15, Tor13, Tur18]. 978-0-00-822631-2 [Hal18]. 978-0-19-870459-1 [Cob15]. 978-0-226-10849-0 [Ren15]. 978-0-262-02737-3 [Tay15]. 978-0-262-53328-7 [Tur18]. 978-0-674-97696-2 [DiM18]. 978-0-87140-413-8 [Har13a]. 978-1-4214-0479-0 [Tor13]. 978-1-4214-2074-5 [Ayl18]. 978-1-84792-381-3 [Gut18].

A. [Smi13, vW16]. abc [CBV11]. Aberdeen [Ols11]. Abhorreas [Sto12]. abhorrent [Lew20]. abnormal [All01]. abortion [Dav99, Ell14]. abounds [Vai15]. Abrams [Wal13]. absences [Meb14]. abstraction [Tee18]. abstraction-by-aggregation [Tee18]. Academic [GS11a, RPLBMA20]. acceleration [Sch14b]. acceptance [Tha98a]. according [vL37]. account [Bon13, Rei12, Sch13a, Ver14]. accounts [dC09b]. Achieving [Ams04]. acids [Hal11]. acquire [Leo13]. across [Bor16, Bra07]. Act [FS15]. action [Gro10]. acts [Dac12]. Adamite [Sno01]. adaptability [BP15]. Adaptation [Ber20, Cre07, Dep11, Bar15, Oka03, Wal00, Sha08]. adapters [Lea19]. Adaptive [Bak05, Kap13, Pei15, Pol15, RB13, SC12]. adds [Mor16b]. adept [Edd08]. ADHD [Haw10]. administration [Fio19]. Admiralty [vW13]. ado [Har18]. Adolf [Gau05a]. Adrian [Whi04]. Adriano [CC02]. advantageousness [Bro09b]. advent [Lan02]. adventure [Hun06]. advice [Juk19, Lou03]. advocate [Fel09]. aesthetic [Fio11]. aesthetics [Pan12]. affairs [Dix02]. affecting [LaB14]. affinities [Tur20]. Africa [Pri12]. African [Bri16]. After [Sol20, Bae12, Ban14, BdC14, Ber14b, Cre14, Hod11, Lin16, Lip14, Tab16, vH14]. Afterword [Rhe16, Ove08]. AG [Gau05a]. Against [Amu00, Gam20, Gro11, Men17, Nor07, Pol15]. Age [MT17, BD19, Man15]. Age-Table [MT17]. Agency [Mil19, Bar12]. agent [Pat20, vdB02]. aggregation [Tee18]. Aging [MP11]. Agricultural [Ber14b, Cas03, Cur17, GSM19, Rou19]. agriculture [Cla07a, Dut18, Fra07, Wil07a]. agro [vdHP14]. agro-industrialism [vdHP14]. agrobiodiversity [Dut18]. AIC [MSM16]. aid [Fin08, Lew02]. Air [Mil20]. Albany [Wil98]. Albert [She12a]. Alberto [San00]. Albumasare [Smo10]. alchemy [Pri13, Roo13, Roo13]. Aldolphe [Ari07]. Alex [Mor19]. Alexander [Det99]. Alfred [FK07, Mit14]. alienation [Hor08]. alienists [Fin08]. Alimentary [Pen12]. alive [For16, KS16, vR16]. Allergy [Jac03a, Jac03b, Jam10, Mit03, Par03, Tan03]. allosterism [ADF17]. allow [Güt13]. almost [Ric12]. alone [Gre20b, Sta13]. alternative [BA05, Cle07, Kid13]. alternatives [KR13]. Although [Whi04]. Altruism [Sch16b, CC16, GS16b, Ram16b, Sti16]. Altruistic [Bir19a]. altruists [Wal08]. Am [McL07, Ste08]. ambiguities [Bow14]. ambiguity [Pod17]. ambiguous [Slo06]. America [Gur98, Smi13, Abr12, Cow09, Dav14b, Fab10, Gau02a, Gol96, Her11, Hey15, Kro03, Mit03, Qui04, Rus17, VR20]. American [Ams04, Bor05, Bow17, Cai02, Cre02, End09, Fay12, Rou19]. Amérique [Qui04, Gau02a]. among [Dar05, Har13a, Ken20, Tay09]. Amphioxus [Low16]. Analogies [Alt08a, She99]. Analogy [Gol13, Hol19, Gil04, Pat20, Ste02b]. Analysing [Hut12]. analysis [KHDW03, Lem17, Mau16b, Mer14, Nic14, Ste11b, Tay09]. analyzed [Soc20]. analyzing [RP16]. anaphylactic [Kro99]. anaphylaxis [Löw03, Tan03]. anarchism [Cle08]. Anarchy [Cla06]. anatomical [Kno15, Mae12, Van00]. Anatomy [Kel11b, Cun02, Cun03, Hun06, Kno15]. Ancestor [Rie13]. ancestry [Gan14]. anchor [Ken20]. Ancient [Jon18]. Anderson [Aga01, Kle18]. Andrew [Ayl18]. angels [Sno01]. Animal [Lem17, Wil02a, Wil03, Bak18, Cla98, Fel09, GSM19, GCT17, GS11b, Hal20, Kir08, LH10, Mat17, Pen20, She10, The14, Tho16, Wil07b, vdB18, Pol07]. Animals [FS15, Bak18, LH10, Lil04a, She10]. animats [Man04]. Annie [Ger04b]. Annotated [Tur18]. Announcement [Ano04b, Ano14a, Ano15a, Ano03a]. answer [Bru13, Hea20]. answers [Ree06]. antebellum [Rou19]. Anthony [Pei15]. anthrax [Cas05]. anthropological [McM20]. anthropology [Bac15, Coh06, KL15, SD14, VR20, Zim01, dC14]. anti [Far06, FK07, FS15, Han14, Kra14]. anti-evolution [Kra14]. anti-mechanism [Han14]. anti-rational [Far06]. anti-vaccination [FK07]. anti-vivisectionism [FS15]. Antibodies [BC01]. Anticipatory [Mor19]. Anticoagulant [LaB14]. antivivisection [Fel09]. ants [Cla98, Har13a]. Anya [Gre20a, Plu20]. ape [Far06, Tho06]. ape-language [Far06]. apes [Etk08]. apparent [Sme13]. appearance [Sch03a]. appetite [Olb00, Wil12a]. Applying [Sha16, Edm13]. appointment [vW13]. apportionment [ER15]. appraisal [Bou17]. Appreciation [Ing99]. approach [BE18, Gra13, Hou05, Pâs14, Roo13, Wim13, YT10]. approaches [All01, FD19, Gil05, Gri13, YH17]. appropriating [Swa05]. Appropriation [Cas05]. appropriations [Don20]. Approximations [Row11]. April [Ano20a]. arbitrariness [Lea19]. archaeological [Lym19]. archetypal [Rie13]. Archipelago [Fic14]. architecture [Bau08, Qui16, Top07]. archival [Sha16]. archive [Mox18]. Archives [Lin16, AGS16, GS16a, Per16, dC16]. Arctic [SS15a]. Arend [The14]. Argentina [Pod17]. Argument [Hym12, Bar13b, Gli00, Gro11]. arguments [Bus20, De 08a, FK07, Pol15]. Aristotelian [Mau00]. Aristotle [Dep08, Gre01]. ark [Wei00]. Army [Juz09]. ars [Roc11]. art [Bru11, De 16]. artefact [Lew00]. Articulating [Wim13]. artifact [Hol13, HP13]. artifacts [BS13]. Artificial [Ric08, Ben07, Bra07, DPS13, Ste02b, Wil07b, Wil06]. artist [Ger04d]. Aryan [Hoc15]. Aschheim [OG14]. Asexual [Bab20]. Asia [Kit20]. Asian [Bre02, Chi09, Jia17]. ask [Gra19]. assault [Rus14]. assembling [KCN16]. Assessing [Hun13, SM06]. assessment [Pol15]. association [Sap94, Sar98, She12a]. Astbury [Cob15, Hal11, Hal14b]. Aster [Gro11]. asthma [Kei03]. astrological [Hay10, Sch10a]. Astrology [Bro10, Azz10, Bou10, Bur10, Car10a, Kas10, Rut10, Smo10]. astronomical [Lou16]. asylum [Hac10]. atherosclerosis [Par16]. Atlantic [Whi14]. Atomic [Iid20]. atomism [Nic10]. atoms [Iid20]. ATP [Web02]. attachment [Mon16]. attempt [Fio11]. attraction [Rut10]. attributions [SLR12]. August [Ano19a, Ano20b]. Augustin [Dro01]. Augustin-Pyramus [Dro01]. australe [R81, R77, LoT03]. Austrian [Cha07]. Austro [Top07]. authentic [Ree06]. Author [Ano02a, Ano05a, Ger20a, Nel20, Ano04a, Bau20, Plu20, Set20, Ste20]. Authority [Ell08, Mae12, Wex08]. Authors [Ano01a, WSP20]. authorship [mKC10]. Autism [Fel17, Pan19, Ver14]. autoimmunity [Par03]. autonomous [San14b]. Autonomy [Ash03, Top07]. autoprint [GvA19]. Auzoux [Ols11]. Avery [Cre08, Hal16]. Aviation [Mil20]. avocet [Dav11]. Axes [Bra11a]. axiology [De 19].

B [Fal13, Pri14, Smi13, HS17]. Babel [Wim13]. bacilli [Tul07]. back [Don15]. backward [Esp18]. Bacteria [Sha07, Swe15, Tha98a, Tha98b]. bacterial [Cre07, Lyo07, Ned03]. Bacteriological [Wor07]. bacteriology [GF09, Sha07, Sta13]. bacteriophage [San10]. bacteriophages [San14a]. bad [McM04]. Bahia [Mon18]. balance [Hod11, Ish17]. Balancing [Dac12, De 16]. Balfour [Bel05, Bla04]. Baltimore [Ayl18, Tor13, MP11]. Bank [Tab16]. banking [CWJ06, Ers18]. banks [Per16]. Barbara [Nas99]. Barr [Mil06]. barrier [MH14]. barriers [CA16]. based [AR20, Mat17, Sch16b]. Basic [KL13]. Basis [Joh01]. battle [Whi04]. Baum [Edi20, Kit20, Pol20, Bau20]. Baumgarten [Pan12]. Bayle [Eas11]. BCG [Bon05]. be [Bec15, Bud16, Can17, Dea10, FD19, For16, Gre15a, Kro15, McG16, Mor12, Oka03, Ram16b, Sti16]. Beagle [vW13]. bear [Dom06]. Beating [Bil17]. beautiful [Col14]. beauty [Der20, Efs12]. became [Bon19]. bed [Löw04a]. Beecher [Low16]. been [Sag17, Ste08]. beer [Bud14]. bees [Cla98, Dav04]. before [Bac15, Cun02, Cun03, Jon18]. beginning [GS20]. beginnings [BM16]. behavior [Bec13, Dri09, Pen20, Sch16b, Sme13, Gir20, Nel20, Sul20]. behavioral [Dri09]. behaviors [Sul20]. behaviour [Hal20, MM20, OB12]. Behavioural [Bol12]. Behind [Nor07]. being [Erl09, McG16, Wel21, Wel26, Wel31, Wel56, WPW71, W^{+}76]. beings [Mil20]. Belgium [PM11]. belief [Dix02, Rat02]. Belt [Lin02]. Bence [Ste10a]. benefits [Sch13a]. Berlin [vH00]. Best [MM20, Cre08, Dri09, Gli00]. Better [Gau05a, Kei03, Lan08]. Between [GSM19, Rob13, Tho16, Wil07a, dC14, Ata12, Bud14, De 19, Gel13, GvA19, Löw04a, Mat16, Mor11b, MZ00, NM17, OMW10, Pie20, PM11, RPLBMA20, Soc20, Ste02b, Too03, VR20, Wol14]. Beyond [BV16, Cra05, Etk08, Ger20b, Lem16, Zim01, Tay15]. bias [Juk17]. big [Aic16, AGS16, Cal12a, Ped15, dC09b, Mun15]. Bildungstrieb [Ric00]. bimodal [MN13]. binary [MM20]. Binmore [Sme13]. bio [Smi09]. bio-political [Smi09]. biobanks [Han06, Shi06]. BioBricks^{T M} [Lew13]. biochemist [San00]. biochemistry [ADF17]. Biodiversity [Fra14, Mor09]. bioethical [CWJ06]. bioethicists [Lil04b]. Bioethics [Ash04, Joh04, Ash03, Bon13, Ell04, Hol16, Lew04b, McM04]. biofuels [Mac13]. Biogeographical [Gan14]. biogeography [O’M08, Pâs14]. Biography [Dav14b, Cla98]. Biological [Ano01d, Ano02b, Bar13a, Nic10, Tem13, Vle14, Wil03, Won20, dMM03, Ano98c, ABM20, Bon19, Bru11, Coh06, Dep08, Dev18, Don20, DO07, Gau02b, GL01, Gre15a, Hal11, Hul01, KR13, Leo12, LB08, Lou16, Löw03, Lym19, MN19, Mam06, MN12, NM17, Ped12, Qua06, RP16, Sac14, Sam13, Sch16b, Ste06a, Teu11, TS20, WL08, Wou03]. biologist [Nil15]. biologists [Ano15h, Cre02, SGK04]. Biology [GB17, Key99a, Key99b, RH17, Spe13, AGS16, All05, Bae20, BS13, Ber20, BP13a, BM15, Bri13a, Bri13b, Bro15b, Bue11, Cal12a, CF11, Cal12b, Car14, Cre09, CDA^{+}15, Dei02, Del00, Dil13, Don15, Dup12, Esp18, Fal13, FM07, Fra14, Fre11, Fri13, Gam14, GS16b, Gau09, Gel13, Ger20b, GW13, GKG13, Gra13, Gre13, Gre16, Gro11, Güt13, HP13, Hol16, Hon15, Kas13, KL13, KR13, KA03, Lan02, LA12, Lew13, Lew19, Liu17, MN13, MN15, Mac13, Mau00, McL11, McS16, Meu12, Mor11a, Mor11b, Le 18, MW07b, Mun15, MM20, Mye16, Nic12, OMW10, OS12a, Ode14, OP13, Pea12, Poo11, Pre13, Qui16, Rhe09, Rie13, Row11, Rus16, San02, Sch03b, Sch13b, Ser11, Ser16, She99, Slo12, SL14, SG15, Str02, SSH14]. biology [SC12, Tee19, Toe12, UH14, VR20, Vin13, Web02, WL08, Zam06, ZDSF11, dCS02, dC02b, Bar13a, Fal13, Kap13]. biology-inspired [KL13]. biology/culture [Del00]. Biomachine [Rad13]. biomédecine [Qui04]. Biomedical [Ano01d, Ano02b, Wil03, Ano98c, Gau04, KW08, Leo12, LB15, Löw04b, Pie98, Sch14b, Sok02, Tho16]. biomedicine [Bon13, Fra07, Gau02a, HK13, Qui04]. bioméedecine [Gau02a]. Biophysics [CC02]. biopolitical [Spe11]. biopolitics [Bar20, Kar18]. bioscience [Aic16]. biospheres [Cle07]. biostatistical [Tre20a]. biotech [Gau01]. biotechnological [Gau07]. Biotechnology [GSM19, HS17, Cas03, CBV11, Gau09, Yi18]. Biotherapies [Löw05]. birds [Soc20]. birth [AG14, Dav14a, Kri02]. black [Hal16]. black-box [Hal16]. blackmail [Cra18]. blending [Por14]. blindness [BV16]. Block [Ram06, Nic09, TB16]. Blockbusters [Her11]. Blood [Ban14, Cre99, BM16, Rad14, Wid14]. blooded [Whi04]. bloodletting [Car10b, Car12b, Gil13]. blooming [vdB18]. blot [Hey15]. Blumenbach [Ric00, Zam12]. bo [Ste04]. Board [Ano18a, Sol00, Ano03c, Ano05d, Ano05e, Lin16, Ano03d, Ano98d, Ano02c, Ano02d, Ano03b, Ano03e, Ano04d, Ano04e, Ano04f, Ano04g, Ano18b, Ano18c, Ano18d, Ano18e, Ano19c, Ano19d, Ano19e, Ano19f, Ano19g, Ano19h, Ano20d, Ano20e, Ano20f, Ano20g, Ano20h, Ano20i]. Boccone [GvA19]. Bodley [Gut18]. Body [Ros14, Fio19, Mil06, Ral10, Sle09, Ste08]. Boerhaave [Kle03, Orl12]. Bohr [Dom06]. Bomb [Iid20, Lin16, Löw12]. Bones [Ger99]. Book [Aga01, Ano98a, Ano98b, Ano99a, Ano99b, Ano99c, Ano99d, Ano99e, Ano99f, Ano99g, Ano00a, Ano00b, Ano00c, Ano00d, Ano01b, Ano01c, Ano02b, Ayl18, Bau20, Bra20, DiM18, Edi20, Góm20, Gur98, Gut18, Hal18, Iid20, Kit20, Mur01, Oka03, Pol20, Sar98, Set20, Sie20, Tur18, Wil98]. Books [Ano98c, Ano01d, Ano07a, Ano09a]. borderlands [Brä12]. Borelli [Des05]. borrowings [Don20]. botanical [Dea10, End01, HA09, Nic06, Ols11]. botanists [Kle18, Nic06]. Botany [Coo12, AH06a, Bel05, GvA19, MW07a, Pug09, Ber14b]. both [KHDW03, PD09, San12]. bottle [Bor04a, Gol96, Gur98]. bottles [Gau09]. bound [Ble10, CC10, Coo10]. boundaries [LB08, Wol14]. Box [Mil16, Hal16, Mil20]. Boyle [Ans02]. Bradshaw [Pei15]. Brain [Kai04, YH17, Abr12, BM16, Bas12, Bea04, Cav04, Chi19, Ers18, Ger04a, Ger04c, GG04, Hak06, Par04, Rad00, Sma04, Tre04, WCB17]. Brain-mind [Kai04]. brainbank [Bea04]. Brains [Ger04c, Man04, Har99b, SM04]. Brains-in-vats [Ger04c]. brambles [All01]. Brazilian [dSV14]. Breaking [San13, Tul07, CA16]. breast [Gol96, Gur98, KCN16]. breed [Ank20]. Breeding [Lig20, Rus17, Bar15, Ber14b, Der20, Gra07, The14, Wil07b]. Brenner [dC09a]. Bretonne [LoT03]. bricolage [Lew13]. bridge [NM17, SP16]. Bridging [Hol16]. Bringing [Dom06]. Britain [Cla07b, Bru03, Bur06, Dav11, Dav14a, Del14, Ell08, Ell14, FS15, GCT17, Gis20, Hom08, Jac03b, Kir08, Lee05, Man17, OG14, Qui05, Wal01, Wil07b, Wil02a, Wil03, dC02b]. British [Bra20, De 16, Ams04, Bac15, Esp11, Góm20, Joh12, Kra14, Qui05, SS15b, Tei13, vdHP14]. broader [LaB14]. Brown [Smi13]. Brünn [Ber14a]. Bruno [Ber14a]. Buffon [vdB18]. Bühler [Gun12, Stu12]. building [Don14, Ink16, Nic09, Vin13]. Buller [Kla14]. Burgeoning [Wil00]. Burgess [Bar09, Bry08]. Burkitt [Cla14b]. Burnet [San10]. Business [Ona14, Gau01, Gol13, Wel14]. Butenandt [Gau05a]. Buzzati [CC02]. Buzzati-Traverso [CC02]. buzzing [vdB18].

C [GS12, Gil13, Rad99, The00, Dav14a, Sto12, Tho01, Ano12e]. C. [Rad99, Sta13]. c.1900s [Kel11b]. ca [Cla07a, Kir08]. cabinets [Kno15, Str12]. cadavers [Van00]. Caenorhabditis [dC98]. cage [Mil16]. Cain [Win01]. Cajal [Fio11, GL15]. Calculating [CF11]. Caldasia [Soc20]. Calender [Ano02a]. calibrating [Ban14]. California [Ste17]. Californian [Aic14]. Callebaut [Cal12b]. called [RH98]. calls [Rad06]. Calsbeek [Kap13]. Cambrian [Pri17]. Cambridge [Aga01, Bra20, DiM18, Gur98, Tay15, Bla04, Pol07]. came [Löw14a]. Campbridge [Tur18]. Can [Bec15, Gre15a, Gre20b, Kro15, McG16, Ram16b, San13, SP16, Bon13, Can17, KPB15, Pol20, Ste10b, Ste05a, Tei19]. Canada [LB18]. Cancer [Fan17, Gre20a, Plu13, vH14, Bid20, Cre14, Gar14a, Ger20b, KC12, KCN16, Sch14a, vH00, Gre20a, Plu20]. Candolle [Dro01]. Canguilhem [Chi03]. can’t [BMG13]. Capital [BC01]. capture [Bri16]. Capturing [Mit14, Drö11]. care [KHDW03, MZ00]. career [Bur06, Erl09]. careful [Noa14]. Carefully [Edm13]. carers [Mil19]. Carl [MWR07]. Carpenter [Del14]. carrier [Gra10]. Carter [Gil13]. Cartesian [Bas12, Eas11]. Case [Tji01, ADF17, Bae14, BP19, Blo16, Bor05, BM15, Bro14, Bue11, DL12, De 08b, Far06, Gil05, Gla12, Gli18, Hal16, HA09, Haw10, Hod11, Hun19, Hut12, HS17, Ken20, Lem17, MSM16, MWC12, Pie98, Poo11, Por14, Pri17, Rac16, Rey06, Rhe09, Sec12, Sin04, SC14, Suá01, Tam17, Tul11, Ule14, Ver04, Ver05, WEHT20, Wal12, WCB17, vdHP14]. case-study [Far06, Suá01, Ver04, Ver05]. cast [San12]. castration [vdM08]. Casualty [Iid20]. Catalonia [Cle08]. Catalytic [BC01]. categories [Bar13a, Mag14, Smi05]. cathedral [Oka03, Wil02b]. Cathy [Ger20a, Gue20, Lew20]. Cattrell [Ger04b]. causae [Cla16]. Causal [Cla12, De 08b, Dro12, Hoo12, Sch16a, Bog05, Gre15b, Lea19, Mag14, Mar19a, Mau16a, Meb14, Men12, Per09, Pla15, Pri17, SC13, SLR12, Tso12]. causality [Bog05]. Causally [NG11]. Causation [Bro09a, Bro15a, Rei12, Dro12, FD15, Hun13, OP13, RBF11]. causative [vdB02]. cause [Teu11, Vec20, Luq16]. causes [Abr13, Kin16, Sch15]. CCTV [Edm13]. Celebrating [Del11a]. celibate [Ric08]. Cell [MW10, Bec10, CK00, Drö11, Fag07, Ker16, MS14, Nic10, OMW10, O’M10, Rey10, Rey14, Ser11, Ser16, Too03]. Cells [Coe12, Ser11, Bra11b, Gro10, Ned03, Pow13b]. Cellular [Drö11, Lan02]. Central [Key99a, Key99b, Bri16]. Centre [Spa12, BdC14]. centuries [Gis20]. Century [Car99, Ste02a, Abr12, Ach11, All05, BP15, Bog01, Bra20, Bro10, Bud14, Buk07, Bur10, Car12a, mKC10, Cla98, Cow09, Dac12, Des05, Gan03, GSM19, Gau01, Hak06, Krá12, LH10, Lee15, Mae12, Mét16, Mil16, MW07b, Mye16, Nic06, Ode14, Orl12, Par03, Pea12, PS11, Ral10, SS15b, Sko11, Sle09, Spa12, Sta12, Tam17, Tho16, Wil07a, Wil07b, Wil99, Wor07, Yu17, Amb15, Rus17, DiM18]. cerebral [Gam20]. certainty [Spe11]. CFS [Kar18]. CFS/ME [Kar18]. chair [Bel05]. Challenge [Key99a, Key99b, Cal12a, For16, Kid13, Mae12, MSM16]. challenges [Car14, Efs16, Fro16, HP13, DE16, Efs16, Kle16]. Chambers [Meg16]. chance [Tan03]. chances [Pen15]. Chang [Roc12]. Change [Car10b, Coe12, Col15, JA08, Mét16, OB05, Pow13b]. changing [All01, Chu03, Cre02, Hen11, Kei03, Røs17]. character [Gue12]. characteres [vL37]. characters [vL37, Krá12]. Charité [BH11]. Charles [Bou10, Car04, Del11a, Gra98, Kro99, Pen15, vW13]. Chasing [Wal00]. Chemical [Cle12, Lea19]. Chemically [Lan16]. chemiosmotic [Web02]. chemistry [Kle03]. Chicago [Dea10, Oka03, Ren15, Sul20]. chicken [San12, San14a]. child [Wri11]. childbirth [Dav14a]. Children [San14b, Hac10]. Chimpocentrism [Vae14]. China [Pol20, Edi20]. Chinese [Slo14b, Yu17]. chlorpromazine [PM11]. choice [Fra14, Ric08]. cholera [Tul11]. choose [DAC^{+}20]. Chrétien [Bou10]. Christian [Wex08]. Chromosome [dC14]. chronic [Löw05]. chronicity [Hen11]. chronobiology [Bec13]. cinchona [vdHP14]. cipher [WH17]. circle [Pow13b]. circulation [Orl12]. circumstance [Pea10]. civilization [Bea03]. Civilizations [Pow13b, Coe12]. civilizing [Cha07]. clade [Lym19]. cladistics [End01]. claim [Hut12]. Claiming [She14]. claims [Can20]. clarification [De 19, Kle16]. Clarifying [Sar13]. Classes [Rap04]. classical [Hol17, Web98]. classification [Coo15, Dea10, Dup01, Ful18, McM20, Qui17, Var18]. classifications [Con19]. classificatory [Con19]. classifying [All01, End09]. clear [Bro15b]. clear-eyed [Bro15b]. Clement [Gig14]. Clements [Eli07]. clever [Fis15, DM14]. Climate [Kel17]. clinic [Bec13, Gau07, Mit03, Sch14a, Wil07a, Woo07, BH11]. clinical [Ash02, GMST15, Gra04, Jac03b, KC12, SC14, Tei13, Tre20b, Tre20a]. clinics [Lin16]. close [Zaw15]. closed [Mil20]. closing [Mat16]. closure [Vle14]. cloth [Ren15, Sar98]. Club [Aic14, RH17]. Clustering [LB08]. co [OB12, Sme13]. co-operation [OB12, Sme13]. Coat [Cob15, Hal14b]. code [WH17]. code-script [WH17]. Coen [Pow13b]. coevolution [Alt07, Alt08b]. cognition [Den04, Pan12, Sha07, vdB18]. Cognitive [Has97, Wri11, Aic14, Cha13, Kid13, MN19, Sch16b, Vle14, Mur01]. cognitive-efficiency-based [Sch16b]. coherence [Too03]. Cohn [Cre99]. coin [Kas13]. Cold [Gli18, Hey15, Lin14, Mun15]. collaboration [End01]. Collaborative [mKC10]. collapse [Car12b, Gil13]. collation [MW07a]. Collecting [Ban14]. Collection [MW07a, Bud18]. collective [Ona14]. collectors [Fab10, Smi13]. college [Bou10, Bou10]. Collins [Hal18]. Colombia [Soc20]. Colonial [BD19, Góm20, Gra17, Pri12, Pug09, vdHP14]. Combining [Gre13, Han06]. Comfort [Mor19]. comment [Juk17, Rey12]. Commentary [Col15, Smo15, Cal12b, Sul20]. commercial [Lew04a]. commercialization [Cas05, Der20]. Commission [Iid20]. Committee [Ram14, Cai02]. commodification [Edd08]. commodities [Ram16a]. Common [Cai02, RBOH20]. commonplaces [Yal11]. Communicating [She17]. Communication [RBOH20, GvA19]. communities [TS20]. community [Cai02, Iid20, Aga01]. comparative [Hun06, LB18, Man15, Pen20, Str02]. comparing [IW10]. comparison [GvA19]. competing [Sno01]. complaints [Edd08]. complete [Cla98]. Complex [Gir20, GS11b, Can17, Don14, MN19, Web11]. Complexity [RB13, BM15, Gre15a, MSM16]. Composite [Hut12]. compositional [AG19]. compromise [Ker16]. compulsory [FK07]. computation [GS12]. computer [GS12]. computers [Hag01]. concept [Ank20, Bro14, Gan13, Haw10, Hol17, Mac07, NSY07, Nic12, Pig13, Pla15, Ste06b, Str14, Val12, Wil02a, Wil03]. concept-the [Haw10]. conception [Nic14]. conceptional [AG14]. conceptions [Dro12]. Concepts [Kap13, BP15, Kov17, Mac07, Mat16, SD09, SSH14, Vel08]. Conceptual [Kle16, HP13, Mét16, Pre10, Wol14]. conceptualization [MM20]. conceptualize [SGK04]. concern [Wil02a, Wil03]. concerns [Han06]. conclusions [Tul11]. concussion [Bel05]. conditions [Tre20b]. Confirmation [For10]. conflict [Vel08]. Conflicts [Pie20, Ash04]. Confused [Sil01, Bir09]. confusion [Lym19, vdB18]. Confusions [Ses13]. Conjecture [For12]. conjoined [Dre98]. Connection [Wil00]. connections [OMW10, Tho16]. connective [GS11a]. conquest [Har13a, Wil12b]. Conrad [Bae13]. Consent [O’N01, Shi06]. Consequence [Dep08, McL11]. consequences [Lew20]. consequentialism [Lew20]. conservation [Bri16, Fra14, Ode14, Whi14]. consilience [Qui17]. consistency} [Gra19]. constitution [Bec15, Fag09, Har15]. constitutive [Ful18, Har15, Toe12]. constraints [Sha08, Vle14]. Constructing [Kra14, JA08, Nic06]. construction [Man15, Mon18, Nic09, Rad03, SDAM08]. constructions [Ser16]. consumer [Her11]. consumers [Mil19]. contained [Gis17]. containment [Mil16]. contemporary [Ers18, Gra07, McM07, Rie13, Wal06, Zam06, dC16]. content [SB19]. Contents [Ano01a, Ano04c, Ano04a, Ano05a, Ano10e, Mam06]. contestation [O’M10]. contested [Mül12]. contests [She14]. Context [Hed06, All05, Bar15, Bir09, Güt13, Hog18, Hym12]. context-sensitivity [Güt13]. Contexts [BP15, Fee07]. contextual [Sha01]. continents [SP16]. Contingency [Car16, Hes16, Lou16, Blo16, Kin16, Won20, Har16]. contingent [Bin18a]. continuing [Gis14]. continuity [San13]. contours [Fay12]. contraries [Chi19]. contrasts [Sch16a]. contributions [Hoc15]. Control [Gut18, Lee15, Par02]. controlled [Her11, Man04]. Controlling [All01]. controversies [Wil99, Zam06]. controversy [Fel09, Hic15, Pan19, Tho15]. conundrums [Ger08]. Convalescent [Alb12]. Convenience [Kro12]. convergence [Hes16, LB18, Lou16]. converging [Chi09]. cookery [Alb12]. cooperation [Lyo07, Mar19b, UH14]. coordinating [Wei20]. cordon [Pal00]. corporate [McM04]. Corporation [Har13a]. corporeality [Sle09]. Correctives [Eng20]. Correlation [WEHT20]. correspondence [Coo12, Yal11]. Corticosteroids [Qui05]. Cortisone [Qui05]. cosmetic [Efs12]. cost [GG06, Wil02a, Wil03]. costs [Erl09]. Could [Oka03, Wal08]. counselling [Ger08]. counterfactual [Abr06, Dag19, Tam20]. Counterfactuals [Hes16, Kin16, Dag17]. Coupling [MN13]. course [Bae13]. cows [Gra07]. Crack [Gut18]. craniological [Juz09]. Creating [Güt13, ZDSF11, Hor11, Fal13]. Creation [Gut18, Rad99, DPS13]. creationists [Kra14]. Creative [RB13]. Creators [Tur18]. Credibility [Hat00, vH00]. creed [Sar13]. creole [Cow09]. Crick [Aic14, Aic16]. cries [Car12a]. crime [Jer13]. criminal [Ell14, Rob13]. criminalistics [BP13b]. Crisis [Gun12, SM12, Bor04a, Car12a, Fay12, Fee12, Hat12, Mül12, Stu12, All12, Fee12, Hym12]. criterion [Bir19b]. critical [Bou17, Nic14, Slo06]. criticisms [Don17]. critics [Sar14b]. critique [Dev20, Fee07, Gue10, Kla14, Pil07, Rad01, SC13, Swa05, Zim01, Bre06]. Crook [Fra07]. Crookes [Kid14]. crops [Hic15]. cross [Aic16, OCG16]. cross-disciplinary [OCG16]. cross-worlds [Aic16]. crossings [Gau02b]. Cruelty [FS15]. CSI [BP13b]. cult [Tre04]. cultivation [vdHP14]. Cultural [Ste17, AH17, Ang02, SP16, Wim13]. Culture [BKP13, Ble10, CC10, Col13, Coo10, Del00, Ger04c, Mar20]. culture-bound [Ble10, CC10, Coo10]. cultured [Ned03]. Cultures [Bra11b, Ham13, Kas13, Lan02, Lyn13, MW07b, Rob13]. cum [Smo10]. curious [Por14]. Curiously [Alt08a]. current [Pâs14]. curriculum [Hay10]. Cutting [Bra07, Löw04a]. cyberpunk [Cav04]. Cycle [Gre01, Str14]. cytogenetics [LB18, San14b]. cytological [Dar05]. Czech [Sle09]. Czechoslovakia [Shm20].

D [Can17, Swe15]. dairy [Gra07, Woo07]. Dale [Coo12, Tan03]. dance [Mon18]. dancers [Mon18]. Darbishire [Woo15]. D’Arcy [Hol19]. dark [Gol13]. Darkly [Sol00]. Darwin [Smi13, Alt07, Alt08a, Alt08b, Ari07, Bed15, Bil17, Bro11b, Bru13, Cla16, Dav04, Daw14, Del11a, Dep08, Fel09, Gil04, Hau12, Mel17, Mil12, Pen15, Rad08b, Ren15, Ric12, Rus14, ST17, Sha11, Sha14, She14, Ste02b, Tab16, Tam17, Tho17, Wil06, Wil02b, vW13, vWK15, Oka03]. Darwinian [Bro11a, Cla06, De 08a, Gre15b, Hes15, Hun19, Rad08b, RR06b, Web11]. Darwinism [Ber20, Del09, Bec11, Bil17, Bou11, Bow17, Bro11a, Coh17, Dep11, Gra19, Hod11, Hul11, Man17, Rus05, Rus11, Yu17]. Darwins [Amu05]. Data [Ben16, Bur16, Str12, Ban14, Bau08, Cal12a, Cal12b, GS12, KPB15, Leo12, Leo13, Ser16]. Data-driven [Str12, Cal12b, Leo12]. database [Bea04]. databases [AH06b, LA12, Str12]. date [Rus11]. David [Oka03, Pei15, Tur18, Gri13]. Dawkins [Gli00, Ist13, Sha01]. Dawkins/Gould [Sha01]. dead [Fab10, Smi13, Ste08]. Deal [Ram08]. Death [Gig14, Azz10, Bla04, Rey14, Sle98]. deaths [Rey14]. debate [Bar09, Del14, KPB15, Lem16, Mat16, Nic15, Rou19, Sha04, Too03, YH17]. debates [Ben07, Cre07, FK07, Rad00]. decades [Bra14]. December [Ano19b, Ano20c]. decentred [Daw14]. Deception [Mar19b, Bir19a]. decision [Mar19a, PT13]. Decisions [Bat01, Sme13, KHDW03, OB12]. decisive [KS16]. declaration [Mül12]. decline [Car10b, Car12b, Gil13]. decomposition [Don14]. Deconstructing [Ona14]. decontextualised [Dev20]. Découverte [Qui04, LoT03, R81, R77]. dédale [R81, R77]. Deep [Col14, SS15b]. defects [AG14]. defence [Dup01, Mit14]. Defending [AG19]. defense [Bro15b, LaP07]. defined [Lan16]. Defining [Ere09, Hul11, Wyl10]. definitely [Ule14]. definitions [For16]. definitive [Zaw15]. Deflate [Hoc16]. Defusing [Löw12, Don17]. degenerating [Hal20]. degeneration [Bar06]. degrees [Wol14]. Delbrück [Dom06]. della [Rut10]. delusion [Ger04d, Mit14]. demand [Her11, OG14]. democracy [Kel17, Mot08]. demography [Ram08]. demons [Ral10]. demonstrations [Hay10]. Dennett [Zaw15]. departing [Hed07a]. dependence [Coo15]. depth [Hic15]. Descartes [KR13]. Descent [Ren15, Hal14a, Alt08a]. Describing [Pen11]. Descriptions [Wal13]. descriptive [She17]. Design [Bat01, Ada05, Gli00, Güt13, Hun19, Ink16, Lew13]. designers [Rus14]. desire [Rut10]. desires [Sch13a]. Desmond [Whi04, Whi04]. Desmond/Huxley [Whi04]. detect [Fis15]. determination [Ger99, Soo12]. determinism [Esp17, O’M08, RR06b]. Development [Bat01, Lew02, Man12, Ada05, BP19, Gau04, Hak06, KHDW03, KA03, Low16, Meu12, Nic14, Qui05, WH17, dC98]. Developmental [Joh01, CDA^{+}15, FM07, Meu12, Rie13, SG15, Sun11, UH14, Vec20]. devil [Des98, Whi04]. devo [BM18]. diabetology [Sin04]. diagnoses [Mau16b]. Diagnosis [Mau16a, Hog16, Hog18, Löw14b]. Diagnostic [Ano13a, Mau16b, OG14, Coo15]. diagrams [Lym19, Sch16a, Tee18]. dialogues [Bed15]. diaspora [Hel13]. dichotomy [Sar13, Wit15, Wit16]. Dickinson [Ger04d]. Dictyostelium [Sun11]. Did [Bru13, Gil13, Mon16, MZ00]. Diet [Smi12, Gue12]. difference [Sch16a, Tay15, Bra20, Góm20, Set20, Sie20]. Differences [Ata12, LB18, PD08, Pie98]. Different [Kov17, VR20, KA03]. Differentiating [Don17]. difficult [CA16, Kri02]. difficulties [Man15]. difficulty [Drö11]. digestion [Cle12]. Digital [Fri13, Van00, Ger04c]. dilemma [Gli00]. Dilthey [Fee07]. dimensional [Mau16b]. Dimensions [LL13, Bra11b, Drö11, Mor09]. directed [Ule14]. directional [McG16]. disagreement [Del11a]. disappearance [Lan08]. disbelief [Ams04]. disciple [Des98, Whi04]. Disciplinary [Bri16, Cun02, Cun03, OCG16, Sep14]. Discipline [Vin13, Esp18]. Discipline-building [Vin13]. disciplines [Abr12, KHDW03, Mét16, SD09]. Disciplining [Smo14]. discourses [CF11]. discoveries [For16, LaB14, Mor14]. Discovery [BC01, Mét16, DC02a, Pre10, San12, Tha98a]. discrimination [Dol01, Hoc13]. Discussion [Ano00e, Ano00f, Ano00g, She12b, KHDW03]. discussions [SM12]. Disease [Ayl18, Bro14, Coo02, GMST15, Bro09a, Bru03, Ere09, Ful18, GF09, Hut12, SM06, Sta13, Tre20a, Wyl10, Bra20, Góm20, Set20, Sie20]. Disease-mongering [GMST15]. diseases [Ful18, KCN16, Löw05]. Disentangling [De 19, Mel17]. disorder [PS11, Ver14]. disorders [Ano13a, Coo10, De 08a, Tso12, Coo15]. Displaying [BN00]. displays [Mat17]. dispute [Sha01, vH00]. disputed [Bre02]. disputes [Bri16]. Dissecting [SG15]. dissections [Kel11b]. dissemination [CBV11]. dissertation [mKC10]. distinction [De 19, Chu03]. Distinctions [Tay15, Lem16]. Distinctive [Ric01, Vec20]. Distribution [Lip14, O’M08]. divergence [Ric12]. divergent [Dom06]. diversification [CDA^{+}15]. diversity [Bon19, ER15, Lym19]. divide [Bak18, Rac16]. divina [Rau04]. DNA [Fis15, Cob15, CWJ06, Fis15, Jon18, Lyn13, Ona14, Shi06, Suá01, Vec20]. do [BMG13, Hoc13, Mau16b, O’M14, Rad08b, Vel08, Vle14]. doable [BP19]. Dobzhansky [Dep11, Gan13, Wit15, Wit16]. doctor [Eas11]. Documenting [Sha16, GS16a]. does [Mor16b, Ste02b]. doesn’t [MH14]. Dog [Fel09, Lig20, WSP20, Ank20, Der20]. Dogma [Key99a, Key99b]. dogs [FB15, Löw03]. Dolphin [Gla08]. Domesticating [Whi14]. Donders [The00]. Donhauser [Don17, Sag17]. don’t [Har99b, Sha08]. dopamine [Col14]. dose [Ful20]. double [Hal14b, Cob15]. double-helix [Hal14b, Cob15]. Doudna [Gut18]. down [Smi20, San14b]. dowsing [Ver05]. Dr [Edd08, Cre99]. Dr. [Ols11]. drama [Jer13, Lou03]. Draughtsmen [Nic06]. drawing [Cla12]. Dreading [Gar14a]. dream [Mic08, PT13]. dreams [Rad13, Ram16a]. Driesch [All12]. drift [Luq16]. drive [Sch13b]. driven [Cal12b, Leo12, Str12]. drug [Ada05, Cre07, Gau05b, PS11]. drugs [BH11, Gau05a, Her11, Hor11, PS11, Tso12]. DSM [Ano13a, CB18]. DSM-5 [Ano13a]. DSM-III [CB18]. Dual [Car14]. dualism [Kai04]. Duelling [CF11]. Dugald [Edd06]. d’un [R81]. Duncan [Sno01]. Dupré [Bar13a]. during [Del11b, Hag01, Mil20, Smi09, WL08]. Dutch [KHDW03, vdHP14]. dynamic [Rie13]. dynamics [Bec15, Bri13a, Drö11, KHDW03, MN15, Web11]. dysfunctions [De 08a].

E. [Bar06, Kap13, Meg16, Pow13b]. Early [Bak18, Mox18, MW05, Alb12, Ata12, Buk14, Can17, Car12a, Cle12, Cra04, Dav14b, Edd06, Ell14, Fal13, Gis20, Kre09, LH10, LB06, Par16, San14b, Sko11, Sle09, Slo12, Sto12, Tei13, Wil07b, Yal11, Yu17, ZDSF11]. Earth [Wil12b, Har13a]. East [vdHP14, BH11, Chi09, Jia17]. East-Side [BH11]. eats [Lan16]. ECHO [Too03]. Ecological [TS20, Col15, Eli07, Hou05, Ink16, O’M08, O’M16, Pea11, Swi17]. Ecology [Don17, Ben16, Boc04, Bol12, Bre02, Don16, Dri09, Dus18, Mar98, Mit03, MSM16, Sag17, Str14, Tau06]. economy [Wel20]. Ed [Tur18]. Edgar [Kle18]. edge [Rud11]. Edinburgh [Ash03, Bel05, GSM19, Tho01]. Editorial [Ano99h, Ano03c, Ano05d, Ano05e, Ano05b, Ano05c, Ano06a, Ano06b, Ano06c, Ano06d, Ano07b, Ano07c, Ano07d, Ano07e, Ano08a, Ano08b, Ano08c, Ano08d, Ano09b, Ano09c, Ano09d, Ano09e, Ano10a, Ano10b, Ano10c, Ano10d, Ano11a, Ano11b, Ano11c, Ano11d, Ano12e, Ano12a, Ano12b, Ano12c, Ano12d, Ano13b, Ano13c, Ano14b, Ano14c, Ano14d, Ano14e, Ano14f, Ano14g, Ano15b, Ano15c, Ano15d, Ano15e, Ano15f, Ano15g, Ano16a, Ano16b, Ano16c, Ano16d, Ano16e, Ano16f, Ano17b, Ano17c, Ano17d, Ano17e, Ano17f, Ano17a, Ano18a, Ano03d, Ano98d, Ano02c, Ano02d, Ano03b, Ano03e, Ano04d, Ano04e, Ano04f, Ano04g, Ano18b, Ano18c, Ano18d, Ano18e, Ano19c, Ano19d, Ano19e, Ano19f, Ano19g, Ano19h, Ano20d, Ano20e, Ano20f]. Editorial [Ano20g, Ano20h, Ano20i]. editors [Ano09f, GG04]. Edmund [Low16]. Eds [Fal13, Kap13, Tur18, Aga01, Sme13]. Education [Bar06, Ols11, RPLBMA20, Rud11]. Edward [Aga01, Har13a]. Edwardian [Pet02]. Effectiveness [Ste15a, Ash02, Gab20, Ste15b]. effects [DM14]. efficacy [Bon05, SC13]. efficiency [Han06, Sch16b]. efficient [Bru08]. egg [San12]. egoism [Sch16b]. Eighteenth [Ste02a, Ach11, Bra20, mKC10, Cow09, Dac12, Krá12, Mae12, Mil16, Nic06, Spa12, Sta12]. Eighteenth-Century [Ste02a, Ach11, Bra20, Cow09, Krá12, Nic06, Spa12]. El-Haj [Hel13]. electrochemistry [Gro10]. electron [Drö11]. electronic [Str12]. elegans [GS12, dC98]. eliminated [Slo12]. elimination [Lud14]. eliminativism [Bae20]. elite [RO00]. Ellis [Cro08]. Elsie [Buk14]. Embedding [Haw10]. EMBL [Kri02]. EMBO [Kri02, San02]. embodiment [Smi12]. embrace [Pet02]. Embracing [Tur20]. embryogenesis [Bra11a]. embryology [Low16]. embryonic [Fra07, Ker16]. Embryos [Mai16, Bra11b, Mac15]. emergence [Ada05, Smi09, Sta12, Wil02a, Wil03, dSV14]. emergentist [Sar13]. emerging [Gis20]. Emigration [Dei02]. emigrations [Fra07]. Emil [FM07]. Emily [Edi20, Kit20, Pol20]. emotion [Eng20]. emotions [Mam06]. Empire [Bra20, Spa12, BL07, Cha07, Góm20]. Empires [Boc04]. Empirical [Mét14, SGK04]. empiricists [Bri17]. ENCODE [Kai18]. encounters [Pri14, Soc20]. End [Lig16, KHDW03, Mox18]. end-of-life [KHDW03]. Endersby [Dea10]. engineered [Gre15a]. engineering [BP13a, KL13, KR13, Sha07]. Engineers [Tur18]. England [Cla98, FK07, Hal14a, KA03, Ral10, Ren15]. English [Gue12]. Enlightenment [Slo14a, Brä12, Wil12a]. enough [Gre13, Mil06, Mos12, Pil07]. Enrico [Bra14]. entail [Vle14]. entangled [LB18]. Entanglements [Ste16]. entities [BS13]. entity [Win01]. Entry [Tho01]. environment [Ari03, BV16, Lil04a, Mel17, O’M08, Pea10, Tay09]. Environmental [Ode14, Sar14b, Sar14a, Bre05, Nan10, Ste10a]. environmentalism [Mar98, Vai15]. enzymatic [San00]. eons [BM18]. eorumque [vL37]. epic [Hes15, Meg16]. epidemic [Mon18]. Epidemics [Löw02]. epidemiological [Bau08, FD15, Rad14]. epidemiology [Ams04, Bro09a, Bro15a, Can20, FBF15, Gra10, Kos01, LB15, Plu14, Wid14, dC14]. epigenesis [Coh06]. epigenetic [Bae13]. Epigenetics [NM17, RH17, BD19, Dei16]. Epistemic [Bar09, Bid20, Cai02, Mag14, Col13, DiF19, Fag07, Gre13, Kas13, Sam13, Tul07]. epistemically [Kid14]. Epistemological [Cle07, Hic15, Bri16, De 19, CA16, Del09, MW05]. epistemologically [PM02]. epistemology [Bor04a, Chi03, EH12, Fio11, Löw04b, Hel13]. epithelial [Ned03]. Equal [KA03]. equilibria [Dre19]. equivalence [DiF17]. equivocation [FJEZM20]. era [Alb12, Drö11]. Ernst [Chu03, Pow13a]. erotica [Roc11]. Erratum [Ano08e]. error [Col14]. Escaping [Cra18]. eschews [Ari07]. Essay [Cob15, Har13a, Ren15, Tay15, Tor13]. Essays [Bar13a, Dup12, Sch03b]. essence [Lew12, Slo14a]. Essentialism [Ped12, Dev18]. ethical [Gla08, Wil02a, Wil03]. Ethics [Bro11b, Ode14, Bid20, Fos19, Gir20, Hed07a, Hed06, KHDW03, Lew02, Lew04a, Lil04a, McM07, Rus02]. ethnic [Fos19]. ethnobiological [VR20, Wei20, FJEZM20]. Ethnobiology [Pie20]. ethnography [Gra07]. Ethnology [SS11, McM20]. ethological [Bol12]. ethology [Wil02a, Wil03]. etiological [Don16, Sag17]. Etiology [Lin01, Dep08, Ful18, Lin02]. Étude [Cla07b]. Eugenic [vdM08, Rad01, Tho06]. eugenicist [Cro08]. Eugenics [Cle08, Hal08, JA08, KL15, Lil01, Mot08, Ove08, Ram08, Etk08, Gaw08, Ger08, Moo07, Ric08, Sin08, The14, Wal01]. eukaryote [O’M10]. Europe [Car10a, Smo10, Str02, dCS02]. European [Cre02, GCT17, PT13]. Eutelegenesis [McM07, Ric08]. evaluation [Gil17, Hed07b]. even [Can17]. event [Soc20]. ever [Cre14]. every [Cla98]. everyone [Amu05]. Everything [O’M08]. everywhere [Fee07, O’M08]. Evidence [FD15, Gil17, AR20, Bog01, Cla12, Ell08, End01, Hay10, Lyn13, Par16, Ste11b]. evidence-based [AR20]. evidential [Can20]. evidentiary [Juk19]. eviscerating [Mar98]. evitability [Won20]. evo [BM18]. evo-devo [BM18]. Evolution [Ano12e, Hal18, HH16, OB12, Pie20, Rat02, Ren15, Rus17, Sap94, Sme13, Smo14, Ågr18, AH17, Ang02, Bec16, CC16, Del00, Des98, Dix02, Ewe19, GS11b, Hal14a, Hod11, Kle18, Kra14, Liv04, Lou16, McG16, Mor12, Nic14, Nic15, Oka03, Pat20, Pie12, Rad00, Rad03, Sch13a, SP16, Smo15, Ste17, SD09, Suá01, Ule14, Vae14, Whi04, Wil02b, Wim13, Sme13, Gut18, Tor13, Sar98]. Evolutionary [Rus16, Spe13, Won20, Abr06, Abr13, Ber20, Blo16, Bor04b, Bus20, Cai02, Col15, Del09, Del11b, Den04, Esp11, FB15, Gan03, Gri13, Hes15, Hol17, Kap13, Kla14, LL13, Lyo07, Meg16, Mét14, Mor11b, OB05, Pot12, Rie13, Rus02, Sap07, Sch16b, Sha01, Sha08, She14, She99, SM05, Ste05a, Sti16, SC12, The14, Tho06, UH14, Vel08, Wal06]. evolutionists [Bru08]. evolves [Lev06]. evolving [Dre19, Kro15, Sha14]. evoText [RP16]. examination [FS15]. Examining [Fay12]. example [Bur10]. except [Dil13]. exceptional [Lyn13]. exceptions [Hog16]. exchange [FM07, HA09, Yal11]. exemplar [Mag05]. exemplars [Sko11]. Exemplary [O’M14]. exhibition [BN00]. existence [Bil17]. exorcism [Tos02]. expanding [Web11]. expands [Woo15]. Expectation [Esp17]. expectations [Ben07]. experience [Bud14, Kel11b, Pen11]. experiment [Cur17, Hol19, MN13, Ned03, Row11, RBOH20, Wil06]. Experimental [HK13, Kle03, Bon05, mKC10, DL12, DM14, Gam20, Gra04, Hol17, Ink16, Kir08, Low16, Mor19, Noa14, Par16, Pil07, Roo13, Sch03a, Suá01, Tee19, Tho06, Tso12, WEHT20, Web02]. experimentalist [Noa14]. experimentation [GCT17, Kro12]. experiments [Mil20, Par14, Sch03c, Tha98b, Tho16]. expert [Lou03]. expertise [RO00, Woo07]. explain [Sha04, Ste10b]. explained [Gre20a, Kro15]. Explaining [Mar20, For10, Gre20a, Plu20]. Explanation [BA05, ARR17, Bec15, Bre06, Bri13b, Bur16, Car10b, Cre08, DiF17, Eas11, For12, Ger20b, Gli00, HH16, Har16, Hun19, Lew12, MN15, Man12, Mau16a, Meb14, Mil12, MN12, Sti16, Tur00]. explanations [AG19, Bar13a, Bog05, De 08b, Eli07, GW13, LL13, McL13, Mor11a, Rey12, dMM03]. explanatorily [Pea07]. Explanatory [BMG13, Gre16, Bir09, Fee07, NG11, RBF11, Too03, Tul11]. exploitation [Lew04a]. exploration [Rhe09]. explosion [Pri17]. export [Cre02]. exposome [Can20]. exposure [Mit14]. expression [NSY07]. extended [Bus20]. Extending [Web11]. Extension [Cur17]. Extimatrix [Bot04]. extinction [Del07]. extract [Löw05]. Extrapolation [Ste11a, Lem17]. eyed [Bro15b].

F [Mur01, Pen11, Qui04]. F. [Bla04, The00]. Fabian [Smi13]. Face [Det99, Ing99, Fri13]. faces [Laf07, Pea11]. factor [LaB14]. Factors [LaB14, Sha01, Ste11a]. facts [Sin04]. fails [O’M13]. failure [Cla14a]. fairy [McL01]. faith [Moo07]. fall [Hoc15, Nor07, Win13]. fallacy [Mic06, Smi10]. falsification [NG11]. Family [Kuk12]. far [Tei19]. farm [Gau07, Wil07a, Woo07]. farming [Gra07, Woo07]. fascinating [vH14]. Fat [Sto12]. fauna [Bry08]. favoured [Dar59, Dar90]. fear [Nil15]. features [PM11]. February [Ano20j]. federal [SC14]. Feminist.Epistemology [Hat00]. Fertilité [Cla07b]. Fertility [Cla07b]. feticide [Ell14]. fetu [Löw14b]. Feud [Kuk12]. fever [Jac03b, Kei03]. Feyerabend [Kid13]. fiction [Kre09]. Fictional [Tee19]. fictions [Kir13]. fidelity [Gar14b]. field [Ree06, WGEN15]. fields [Dar05, NM17]. Fieldwork [Wid14]. Fifteenth [DiM18, Ral10]. fifteenth-century [Ral10]. Fifth [Ona14]. fight [Fel09]. figuram [vL37]. figure [vL37]. files [dC01]. film [Wex08]. fin [BP13b]. fin-de-siècle [BP13b]. Finding [Pow13a]. Finn [Tur18]. First [Gre17, Mül12, O’M10, Par03, Pea12]. Fisher [Ber20, IR20, Moo07, Pil07, Plu06]. fishing [Bre05]. fit [Moo07]. Fitness [Ram06, Bir19b, Bro09b, Hun19, Pea11, Sob13, Swe15]. Fittest [Bou11]. flatness [Bra11b]. Fleck [Bor04a, Fag09, Hed07a, Hed07b, Lin01, Lin02, Löw04b, vdB02]. Fleckian [Ver14]. flesh [Cav04]. flexible [Bau08, Rad14]. Flies [Par14]. Florentine [Mae12]. fluid [Orl12]. flying [FB15]. Flynn [MT17]. Flyway [Whi14]. Foam [Löw12]. Fodor [Ric12]. Fogel [Fal13]. foil [Sha14]. Folk [Mur01, Has97]. folklores [vdM08]. Following [Sch14a, Sch13a]. Food [Alb12, Buk14]. foods [BPS20]. Forber [Rey12]. Force [Mil20]. forces [Gam14]. Foreign [RO00]. Forensic [BKP13, Col13, Ham13, Kir13, Edm13, Jer13, Lyn13, Rob13, Win13]. Foreword [Del11a]. forgotten [Cob15, Etk08, Hal14b]. Form [Hol19]. Formal [Ber20]. formaldehyde [NG11]. formation [Laf07, dSV14]. forms [Pod17]. Forster [Sta13]. fossils [Tur00]. Foucault [Roc11, Swa05]. found [Car12a]. foundation [CC02, Del09, Bal00, LZ00, MZ00, Pal00, RO00, Sol00, Wei00, Wil00]. Four [Wou03, Rey05]. Fractionation [Cre99]. frame [Has97, Mur01]. framework [BPS20, LaB14, OCG16]. frameworks [BMG13]. Framing [Pri17]. francais [R81, R77]. Francaise [Cla07b]. France [Qui04, Coh17, Fio19, Gau02a, Gau02b, Gis20, LH10, Tho16]. Francesco [Par14]. Francis [Aic14, Aic16, Bow14]. Francois [Eas11]. Frankenstein [Tur18]. Freaks [Nas99]. Frederik [Kno15]. Free [WEHT20, Bec15]. Free-viewing [WEHT20]. Freedom [McS16]. Freeman [Jac03b]. French [Car12a, Gau02a, Hen11, Kra14, MZ00, Qui04, Tre04]. frequencies [Abr06]. frequency [Lym19]. Freud [Bra06, Mic08, Tji01]. Friedrich [Gam14]. friends [Alt07]. frightening [Brä12]. frozen [Rad14]. fructification [vL37]. fructificationis [vL37]. full [Gol13]. Function [IW10, Lew00, Amu00, Bin18a, Bin18b, BMG13, BS20, Dus18, Rad00, Rat00, Wou03, WCB17, Zam06]. Functional [DiF17, Dus18, Clo02, CBV11, Mer14, Mor11b, Tur00]. functionalist [Bra06]. functionality [Kro15]. Functions [Sch03b, CC16, Gre20b, Hol13, Mam06, Mau00, Tur00]. Functions-New [Sch03b]. Fundamental [Ber20, Plu06]. funded [Ell04]. fuss [Olb15]. Fustigating [Ros14]. future [AGS16, Dep11, Fro16, Gis17, Per16, Rad14, Sch14b, dC16]. futurity [Esp17].

G [Bar06, Mur01]. Gaia [Tho15]. Galton [ARR17, Bow14]. Galtonian [Rad11]. galvanism [Sle98]. game [Sch10b]. games [Pot12]. gap [Mat16, NM17]. gardening [Mot08]. gathering [Bau08]. Gaudillière [Qui04]. gender [FS00, Fio19, Mil06]. Gene [NSY07, Pla15, Ari03, Bab20, BMG13, Cre08, Fal10, Fal13, McL01, Olb15, Per16]. Gene-concept [Pla15]. genealogical [EH12, Hel13]. Genealogy [Nic15]. Genera [Kle18, MWR07, vL37]. general [FB15]. Generalizations [Rey06, Bog05]. generation [BV16, Par14]. generative [WEHT20]. Generic [Pod15]. Genes [Bae12, Lil01, Lip14, BRAB16, Bon19, Joh12, Sil01, SGK04, Tay09, Web98]. Genesis [Gau04]. Genetic [AH06b, Ayl18, Hel13, Lip14, O’N01, Ric01, Wil01, Ban14, Bon19, Dix02, Dol01, ER15, Esp17, Gan13, GS12, Lee05, OEB10, OP13, RR06b, Sha07, WH17, Wid14, McM20]. genetical [Web98]. Genetics [Cai02, CC02, Rad99, Bau14, Ber14a, Hal20, Hol17, IR20, Lee05, Lee10, Lin14, Lin16, Löw14a, MW07b, Nic15, Ram08, Sko11, Spe15, The14, Woo15, dSV14, Ayl18]. genie [McL01]. Genome [Sha16, Bae12, Gan03, Kai18]. genomes [Bea03, BC08]. genomic [Bon13, KCN16, OEB10]. Genomics [WMN15, BC08, GS16a, Sha16]. genuine [Rey12]. geographic [Bak05]. Geographical [Lip14]. geography [Coo19]. geometry [Bra11a]. Georges [Chi03]. Gere [Ger20a, Gue20, Lew20]. German [Ben07, Tam17, Zim01]. Germany [BN00, DM14, Dei02, Ste02a]. germinal [Ric08]. germs [Bru03, Sta13]. gerontology [Mor19]. Gervais [Bou10, Bou10]. gestation [Esp18]. Gesture [Ing99]. get [Gre17, Zaw15]. Getting [Reu12]. Giant [CA16, Ger04c]. Gifford [Ash03]. Gillies [Hut12]. Giovanni [Rut10]. give [Efs16]. giving [Dav14a]. Glasgow [Ols11]. Glass [Sol00, Den04]. Glaxo [Qui05]. Gleason [Eli07]. Global [Bed15, Dut18, Gra17, Pop20, ST17, Wil00, Men17]. glocal [dCS02]. glycolysis [Pre10]. GM [Hic15]. go [Ste02b, Tei19, Wol14]. go-between [Wol14]. goals [Har99a, Plu13]. Godfrey [Hal18]. Godfrey-Smith [Hal18]. Godlewski [FM07]. gods [Dix02]. goes [Bre05, Law12, Plu07]. Goethe [Rie13]. Going [vWK15, Zaw15]. Golden [Gur98]. Goltz [Gam20]. Gonadex [Nor07]. good [BS13, Bea16, Bog01, Hes15, McL01, Mil06, Ger20a, Lew20, Gue20]. Gordon [SS11]. Gould [Bou17, Dre19, Ist13, KPB15, Sep16, Sha01, She14]. Gourhan [Ing99]. governmentality [JA08]. GP [FF15]. gradient [BPS20]. Gradman [Sta13]. grain [Men17]. Grains [Men17]. grand [Efs16, Fro16, DE16, Efs16, Kle16]. graphics [She17]. graphs [Bur16, Lym19]. grasp [Sch03a]. gray [Pan12]. Great [Ben07, Pen12, Ram08]. Greater [Ger20a, Lew20, Gue20]. Greece [Bar20, KL15]. Green [Bar15]. Gregory [Alt08a]. Grene [Hon15]. grey [Ken20]. Gross [BP13b, Mor14]. ground [Ker16]. grounds [Lin16]. group [All01, Dol01, Oka03, Sha04, Cre09]. group-level [Oka03]. groups [Ban14, Fos19, Sac14]. growing [Hor08]. Growth [Cre99, Ari07, Buk14, Web98]. Guest [GG04]. guide [Bro15a, Ere01, Maz16]. guinea [Kir08, Löw03, Rad08a]. Gulik [Roc11]. guns [Sha08]. Guston [Tur18]. GUTs [Mar98]. gynaecological [Tot07].

H [Tur18, Wil98]. H. [Bar06]. Habsburg [BL07, Sec12, Spa12]. Hagedoorn [The14]. Haj [Hel13]. Hal [Bae13]. Hale [Ren15]. half [Par03]. Hall [Cob15, Car04]. Hamilton [Can17, Mar19a, Swe15, Tho06]. Hamiltonian [Ågr18]. Hand [Ing99]. handicap [Bar13b]. handling [Kno15, MN19]. Hans [All12, BP13b, DM14, LH10]. Happily [Cre14]. hard [Coo15, End01]. Hardback [Aga01, Mur01, Ayl18, Cob15, Dea10, Oka03, Tay15, Gur98]. Hardcover [Har13a, Tor13, DiM18, Gut18]. hardening [Ish17]. Harmony [Too03, Gra04]. Harvard [DiM18]. Harvey [Gol13, Gre01]. Haselager [Mur01]. Haunted [Noa14]. Havelock [Cro08]. hay [Jac03b, Kei03]. Hayek [Ang02]. Head [Gut18, BM16]. heads [Kre09]. healing [Alb12]. Health [De 16, Gue12, KHDW03, Sol00, Bau14, Ere09, Fin08, Gra12, Gra17, Gue10, MZ00, Pal00, Pen12, Pop20, Wil00, YT10, Bir00, Sol00, SC14]. heart [Bin18a, Bin18b]. Hebrew [Bur10]. hedonism [Gar16]. Hegel [Mic12]. Helicobacter [Hut12]. helix [Ari03, Hal14b, Cob15]. Helmholtz [Aic14]. helpful [Rey12]. helping [Sch16b]. hematopoietic [Fag07]. Hempel [CB18]. Hempelian [Gil05]. Henk [Lin02]. Henry [Tan03]. hepatitis [HS17]. herbarium [GvA19]. Herbert [Pea10]. Hercules [Ink16]. hereditary [LB06]. heredity [BdC14, Lee05, Rad11, Wal01, Woo15]. heritability [Pea07]. heritage [Bud16, Slo06]. Herman [Kle03, Orl12]. herring [Pre13]. Herschel [Gil04]. heterogeneity [Tay09]. heteronomy [Pra16]. heuristic [BE18]. heuristics [Rey12]. hierarchical [Eng20]. hierarchies [BM18, Hoo12]. high [Des98, Sep14, Whi04]. higher [Sha11].

Hippocratic [Góm20, Tot07]. histamine [Tan03]. historian [Cha16, GS16a, Whi04, Win01, dC16]. historians [Chi09, Pol20]. Historical [Ber20, Dev18, Gab20, Pie20, Ric00, Smo15, All05, Chi03, Cre09, Fio19, Gam17, Gar14b, Ham13, HH16, Har16, Jef08, Kai04, KHDW03, SM12]. historically [Bin18a]. historicisation [SS15a]. historicity [GS12, Sep16]. historicization [Qui16]. historicize [Aic16]. historicizing [Tho17]. Histories [Ayl18, MH14, Dag19, Hes16, LB18, Røs17, San14a, Tam20, Vai15]. historiographic [Del11b]. Historiographical [Low18, Cra18]. historiographies [AGS16]. historiography [Dag17, GL01, GL15, Roo13]. History [Aga01, Ano01d, Ano02b, Ber10, Fel17, Fos19, Góm20, Gue20, Lig16, Rud11, SDAM08, Wil03, AR98, Ang02, Ano98c, Ano09f, Bor16, Bra07, Bru11, Car04, Chi03, Cla07b, Cra18, Dav11, Dav14b, De 08b, Edm13, Fin08, Gam17, Gau01, Gau04, Gau09, Gol96, Gur98, Har16, Hes16, Hol17, Jac03a, Jam10, Jon18, Kas10, Kin16, Kle03, Lan16, Man17, Mar98, Mét16, Mil19, Mit03, MWC12, OMW10, O’M10, Ona14, Ped12, Pie12, Rad08a, Sap94, Sar98, ST17, SS15b, Sim99, Sko11, Slo06, Slo13, Slo14b, Smi05, Som14, SC14, Str14, SSH14, Tan03, Tor13, Vel08, Wel21, Wel26, Wel31, Wel56, WPW71, W^{+}76, Wol14, Zim01, dC01, vWK15, De 16, Bar06]. HIT [WCB17]. HMS [vW13]. Hofstätter [Gun12]. Hogan [Ayl18]. Holcomb [Wil98]. holistic [Rac16]. hollow [Bil17]. Holobionts [TS20]. Holocaust [Hed07a]. homme [R81, R77]. homme-volant [R77]. Homo [Oka03]. homology [Eng20]. homunculus [Smi10]. Hooker [Dea10, Bel05, End08]. Hookworm [Wil00]. Hopkins [Ayl18, Tor13]. Hormone [Mag05, Nor07]. Hormones [Nor07, Lav05]. hot [Whi04]. hot-blooded [Whi04]. How-possibly [Rey12]. Howard [She12a]. HPV [Bro14]. HPV-itis [Bro14]. Hughlings [Chi19]. Hull [Hon15]. Human [BdC14, Cre99, Ker16, Lin14, Lin16, SS15b, Van00, Bae12, Bak18, Ban14, Coh06, Del00, Dix02, Don15, Dri09, ER15, Gla08, Han06, Kai18, Kla14, Lip14, Lud14, Mat17, Mil20, Rad00, San14b, Smi13, Ste05a, Vae14, Val12, dC14, dSV14, Sha16]. human/animal [Bak18]. humanism [Zim01]. humanity [Cla98, Don15]. humans [Ben07, Etk08, LB08, She10, Tem13]. Humboldt [Det99]. Hume [Cla16]. hundred [Bog01]. Hungarian [Laf07, Pal00, Top07]. Hungary [Krá12, KL15]. Husserl [Fee12]. Hutton [Bel05]. Huxley [Des98, Erl09, Lig16, Sto04, Whi04, Whi04]. Hybridity [Wol14]. hybridization [Fis15]. hybridizing [Etk08]. Hybrids [MW07b, TS20]. Hydra [Ink16]. Hygiene [Bal00, Tho01, Wil00]. hyper [Lou16]. hyper-astronomical [Lou16]. Hypnosis [Gra14, Win13]. Hypotheses [Fee07]. hypothesis [Col14, WEHT20]. Hysteria [Lud14].

ice [Par04]. Icones [GvA19]. icons [Ser11]. idea [Efs16, Hoc15, KL13, KR13, Pea10]. ideal [Bar15, She17]. idealizations [Row11]. ideals [Juk17]. Ideas [Lee05, Wal01]. ideational [Spe11]. identified [LB15]. identities [Kai04]. Identity [DiM18, Joh01, Bab20, Bon13, Buk07, Cun02, Cun03, MHH10, Sar13, WCB17].

Ideologies [Dol01]. II [vH00, Cre99, Cun03, Key99b, Mil20, Tha98b, Ver05]. III [CB18]. illness [Azz10, FD19, FD15, Kid17, YT10]. illnesses [YT10]. illusions [Gra04]. illustrations [Nic06]. image [Wex08]. Images [LoT03, Sam13, Edm13, Nas99, Ser16]. imaginaries [Sin08]. imagination [Ver04, Ver05]. Imaging [Bor16, BM16]. Imagining [Jam10, End09]. Imitating [Hol19]. immortality [Par04]. immune [Gre20b]. Immunity [Kro99, Swi17]. immunohistochemical [Ned10]. immunology [Bae14, Fag07, Jam10, San10, San13]. impact [Bae13, LA12, vW16]. impartiality [Tei13]. Imperial [End08, Ram16a, BN00, Dea10]. Implications [Bid20, ER15, Bae20, LR06, Low18, Mat17, Spe11]. importance [All05, Pen12, PD09]. Imposture [Mit14]. improve [Ned03]. improvement [Bir19b, Sec12]. Inbred [Rad99]. Inclusive [Bir19b, Bou11, Hun19, Swe15]. incommensurability [Kit20, Mau20]. increasing [Mor11a]. indentations [Ere01]. independence [Bro09b]. independent [vH00]. Index [Ano98e, Ano99i, Ano00h, Ano01a, Ano03f, Ano05a, Ano02a, Ano04a]. India [Löw12]. Indian [Bar15]. Indies [vdHP14]. Indigenous [Ach11, SD14, Ken20, Swa05]. individual [Pat20]. individual-as-maximising-agent [Pat20]. individualism [Wri11]. individuality [Dre98, Kov17, Löw03]. Individuals [Rap04, Sac14, TS20]. Inducing [Fio11]. industrialism [vdHP14]. industry [Ada05, Ber14b, Gau01, Sok02]. inertia [Sha11]. infections [Gra10]. Infectious [Spe11]. inference [Cla12, Gli00, Har15, Tur00]. Inferential [GW13]. inferentialist [Rei12]. Infinite [Abr06]. influence [Ari07, Bir00, Rey14]. influencer [Aic16]. influences [Coh17]. Information [O’N01, Ric01, Ano05b, Ano05c, Ano06a, Ano06b, Ano06c, Ano06d, Ano07b, Ano07c, Ano07d, Ano07e, Ano08a, Ano08b, Ano08c, Ano08d, Ano09b, Ano09c, Ano09d, Ano09e, Ano10a, Ano10b, Ano10c, Ano10d, Ano11a, Ano11b, Ano11c, Ano11d, Ano12a, Ano12b, Ano12c, Ano12d, Ano13b, Ano13c, Ano14b, Ano14c, Ano14d, Ano14e, Ano14f, Ano14g, Ano15b, Ano15c, Ano15d, Ano15e, Ano15f, Ano15g, Ano16a, Ano16b, Ano16c, Ano16d, Ano16e, Ano16f, Ano17b, Ano17c, Ano17d, Ano17e, Ano17f, Ano17a, Art20, Bau08, GS20, MWC12, Pla15]. informational [Bea04]. Informed [O’N01, Kel11a]. inheritance [Por14]. inherited [Woo15]. Inhibition [Smi20]. ink [Hey15]. innate [PD08]. Innateness [Bir09]. Innovation [Dut18, Rus17, HS17]. inquiry [Bec10, Gau07, She17]. Inscribing [Ste02a]. insemination [Ben07, Bra07, Ric08, Wil07b]. Inseminator [McM07]. inspired [KL13]. installment [Ses13]. Institute [Bal00, Ber14b, Mil20, vH00, PM02]. Institutes [SC14, Tho16]. institutional [Ger13, KA03]. Institutionalizing [Str02, WL08]. institutions [De 16, MW05, Røs17, Top07]. instrument [LZ00]. Instruments [Hay10, Rhe16, Ste02a, Tha98b, Ste16, Tso12]. integrated [The14]. Integrating [Ada05, Leo13]. Integration [Bec13, Bri13a, Ger13, Gri13, Bri13b, Cra05, De 16, Del00, LaB14, Leo13, LL13, OS12a, O’M13, OCG16, Plu13, VR20]. Integrative [Par02, Kid13, MN13, MN15]. integrity [Han06, Swa05]. intellect [Erl09]. intellectual [Hod11, Yi18]. intelligence [LH10, MT17, Ram14, Val12]. Intelligent [Hal18, Rus14]. intensive [Mic06]. intentional [Bru08]. Intentionality [Had99]. inter [Löw05]. inter-war [Löw05]. interaction [Fag07, Mor11b, Pea10, Tay15]. interactions [Tha98b]. interactive [Ned10]. Intercultural [RPLBMA20]. Interdisciplinarity [DiF19, Fro16, Bri13a]. Interdisciplinary [Poo11, Sch15, Tji01, Bri16, CF11, LL13]. interest [Ash04, RBOH20]. interface [Row11]. interferons [Pie98]. interfield [DC02a]. Intermediate [Pod17]. International [Bar15, San02, CC02, Sol00]. interpretation [FJEZM20, Hed07b, Qua06, SC13]. interpretations [Ber20, Far06, Mic08]. intertwining [Mar20]. Intervention [Tso12, Per09]. interventionist [RW16]. interventions [Bor04a, Reu12, Ste15a]. Interview [dC09a, Ger04d]. interwar [Gaw08]. intestine [San00]. Introducing [BKP13, De 19]. Introduction [AGD14, Ano15h, BP15, CC10, CD05, DE16, GS16b, Gau05b, HH16, IRR12, Leo12, Löw04b, OS12b, RR06a, ST17, Spa12, Spe13, Ste06a, WMN15, dCR09, AH06b, Don20, Hag01, PM11, SS15b, Som14, MWR07]. introductory [Hal08]. introspection [Kro03]. Inuit [SS15a]. invasion [Ned03]. Inventer [Gau02a, Qui04]. Inventing [Gau02a]. Invention [Bau20, Edi20, Kit20, Lig20, Pol20, WSP20, Ank20, Der20, Fis15, Gra10]. investigating [Ste16]. Investigation [Ram14]. investigations [GS12, PKD16]. investment [CWJ06]. Invisible [BC08, BN00]. Irish [Kel11b]. irrationality [Sme13]. irresistible [Löw14b]. Irretrievably [Bir09]. Isabelle [Sno01]. ISBN [Aga01, Ayl18, Cob15, Dea10, DiM18, Gur98, Gut18, Hal18, Har13a, Mur01, Oka03, Qui04, Ren15, Sar98, Tay15, Tor13, Tur18, Whi04, Wil98]. ISBN-13 [Tor13]. Island [Pâs14]. isolated [Kre09]. isolating [Wid14]. isolation [Dei02]. issue [Don20, Ist13]. Issues [Bar20, Cle07, Cre07, Spe11]. Italian [Azz10]. Italy [Gra07, Bra14, Fio19]. Itapagipe [Mon18]. iterativity [OEB10]. itis [Bro14].

J. [Bar13a, Pri14]. Jackson [Chi19]. Jacob [Ful20, Gab20, Hea20, Sol20, Ste20]. Jacques [FM07]. James [Aga01, Tay15]. Jamieson [Lil04a]. Jan [Sar98]. Janet [Gur98]. Japan [Iid20, Cra18, Hom08]. Japanese [Iid20]. Jason [Tur18]. Jay [Dre19, Sep16, She14]. Jean [Qui04, vdB02]. Jean-Paul [Qui04]. Jennifer [Gut18]. Jewish [Hel13, EH12]. Jim [Dea10]. Johannes [Sch03a]. John [Bel05, HA09, Jac03b, PD08, Tul11]. Johns [Ayl18, Tor13]. joint [Ano09f]. jointly [Haw10]. Jolan [Roc12]. Joseph [Dea10, Bel05, End08, Roc12]. journal [Soc20]. Journals [Ano09f]. judgement [Ste06b]. Judgment [Ste02a, Ham13, Bre06]. judgments [Juk17]. Judicial [Car10a]. Julian [Erl09]. Julie [Ank20, Der20, Lig20, WSP20]. Julie-Marie [Ank20, Der20, Lig20]. June [Ano20k]. Jurassic [Jon18]. Just [Edm13, De 08a, Fis15]. justice [Ham13, Rob13]. Justifying [Ser16].

Juveniles [Har13b].

K. [Gil13]. Kamikazes [AH17]. Kant [Bre05, Bre06, Coh06, Coo19, Han14, KR13, Lel20, Qua06, Ric00, Slo06, Slo14a, Ste06b, Wel20, Zam06, Zam12, vdB13, vdB18]. Kantian [Gam17, Han14, Ste06a, Stu12]. Karl [DM14, Gam14, Pen11]. karyotype [San14b]. keeping [Ers18]. Kersten [Cob15]. key [De 19, The14]. Kielmeyer [Gam14]. Kind [Gla12, Eng20]. Kind-Making [Gla12]. Kindred [Pet02]. kinds [Mag14, Rey06, Var18, VR20, WGEN15, Tur18]. king [Dea10, SS15a]. Kingdom [Bra07]. kings [Azz10]. Kitcher [Rad01, Tji01]. knife [Hak06]. know [Har99b]. Knowledge [Car99, Amb15, Bae14, Bre02, Cra18, De 16, Fro16, Gel13, Krá12, Leo13, Mar20, Mon18, Mox18, Ned10, Pen12, Pop20, Pri12, Pug09, RPLBMA20, Sec12, Swa05, TB16, Web98, Wei20]. knowledge-making [Mox18]. knowledge/power [Swa05]. Knox [SS11]. Koch [Sta13, Gra04, GF09, Gra10, RW16]. Koffka [Hat12]. Köhler [Hat12]. Korea [DiM18]. Krall [DM14]. Krohs [Cal12b]. Kuhnian [Gil05].

L [The14]. L. [Roo13]. lab [WGEN15]. labor [mKC10]. laboratorie [MP11]. laboratories [KA03]. Laboratory [CC02, GF09, Rad99, Löw04a, MP11, OG14, RH98, Sch14a, Sta13]. Labs [Jer13]. lac} [Rac16]. Laing [Edd08]. Lamarckism [Dei16]. Lamarckisms [Gis20]. Lancisi [Gig14]. landscape [Bae13, Kap13, SC12]. landscapes [Kap13]. Language [DiM18, Ing99, Rad00, Alt07, Alt08a, Alt08b, Bru08, Edd06, Far06, Rad08b, Rey14]. languages [Bed15, Sec12]. Lankester [Bar06]. largest [Des18]. largest-scale [Des18]. late [All05, Buk07, GSM19, Hak06, Hes15, KA03, Mae12, Pug09, Tam17, Tur20, Wor07, Slo14a, Ste02a]. late-eighteenth-century [Mae12]. late-twentieth [GSM19]. late-Victorian [KA03]. later [Car10a]. Latham [SS11]. Latin [Bur10, VR20, vL37]. law [Ari07, Har13b, KHDW03, Löw04a, Rob13, Sha11, Luq16, vW16]. laws [Gre15b, Hau12]. leader [Bla04]. learn [Pol20]. Learning [Ank20]. Lectures [Ash03]. left [Cla06]. legacy [Bra06, Dep11]. legal [Fio19, Lyn13]. legitimacy [Con19]. Leiden [GvA19]. lemons [BRAB16]. Lenoir [Zam12]. lens [The00]. Leroi [Ing99]. Leroi-Gourhan [Ing99]. lesson [Mat16]. Lessons [Ber14b, Bri16, Tho15, Chi09, De 16, Lyo07]. letters [Pri14, Coo12]. Letting [Bak18]. Lettre [R81]. Leukemia [Sch14b, Mor14]. level [MN15, Oka03, Qua06]. levels [BE18, De 19]. Leviathan [Hey15]. Liberals [Rus14]. lichen [Ken20]. Life [Bru11, GR04, RH17, Sep16, Sle98, Bae13, Bar12, Bar13a, Ben16, CF11, Cha13, Cle07, Coe12, Dar59, Dar90, Des05, Dom06, DPS13, Dup12, Efs16, Esp18, For16, Gau02a, Gig14, Gro10, Hes16, KHDW03, Kid17, Kos16, Lan02, Man17, Mel17, Nor07, O’M13, Pie12, Pow13b, Pra16, Rou19, San12, Sil01, Ste16, Ste05b, Str14, Tor13, Wel21, Wel26, Wel31, Wel56, WPW71, W^{+}76, dC02b, Hal18, Pow13b, Ayl18]. light [Dag17, Dil13, KS16, Tay09]. lightly [Den04]. Like [Ink16, BRAB16]. limitations [Nan10]. limited [Ste10a]. Limits [Wil01, Bre06, Dev20, Dre98, KPB15, Rad06, Slo14b]. Lindenmann [vdB02]. line [Ber14b]. lineages [Bry08]. lines [Bru11, MW07b]. linguists [Alt07, Alt08b]. link [Del00]. Linking [Man17]. links [Pie20]. Linnaean [Ere01, MW07a]. Linnaeus [MWR07, MWC12, Wel20, Win01]. Linnean [GvA19]. Lionel [Val12]. Lisker [SD14]. List [Ano10e, Mox18]. lists [HA09]. literature [Sin08]. little [Ped15, Rad99]. Liveright [Har13a]. lives [Dre98, GL01]. Living [Coo19, DG16, Har16, Pet02, vR16]. Local [DiM18, Hun13, Pop20, Pug09, SS11, YH17]. localization [Gam20]. location [Bur10]. lock [Coo15]. lock-in [Coo15]. Locke [AH06a, HA09]. locus [Bec10]. Loeb [FM07]. logic [FK07]. Logical [Bri17]. logit [Cra04]. Lombardy [Gra07]. London [Aga01, Gut18, Hal18, Mur01, Whi04, Wil00, Sie20]. long [Dac12, Esp18, Sta12]. Longitudinal [MP11]. look [Mac02]. Looking [Zim01, PD09]. loss [Ell14]. lost [Pen20, Whi04]. love [MH14, Nil15, Whi04]. low [GG06]. luck [FB15]. Ludwik [Bor04a, Hed07a, Hed07b, Löw04b, Mor14, vdB02]. lymphocytes [CK00]. lymphoma [Cla14b]. Lysenko [Bau14]. lysogeny [PM02].

M [Bla04, Roo13]. MA [Aga01, DiM18, Tay15, Tur18]. Maître [Bou10]. Macfarlane [San10]. Machine [McS13, BP13a, HP13, Nic14, Sha08]. Machines [Nic13, Gau00, KR13]. MacIntyre [Gla08]. Macroevolution [Dre19, Bou17]. macroevolutionary [Dre19]. Made [Rus17, Fis15, Rad14, Wal08]. Madness [Edi20, Bau20, Kit20, Pol20]. Madrid [San14b]. Magic [SM04]. magnitudes [Mic06]. Magyar [Laf07]. make [Pig13, Sch13b, Tay15]. makers [LR06]. makes [Dil13, Mac02]. Making [AG14, BP13b, Can20, Cla14a, Gla12, Hog16, Leo12, Qui05, Sto04, Ule14, Wid14, WL08, Bur06, Gau07, Gra98, Hes15, Jia17, Mox18, Ree06, San02, MP11]. malaria [Com09]. Malay [Fic14]. Malinowskian [Bac15]. Malthus [Ren15, Ari07, Hal14a, Ren15]. mammal [Pod17]. Man [Lan08, Sno01, Swe15, Cob15, Dea10, Fal13, Hal14b, Ste04, Sto04, Man15, ZDSF11, Alt08a]. Managing [Gra07, Pie98, Sch14b, WL08]. Manchester [GS11a]. Manifest [Pod17]. mankind [Wel21, Wel26, Wel31, Wel56, WPW71, W^{+}76]. Manual [Coo15, Ano13a]. manuscripts [Bou10]. many [KC12, Mor09, Pen15]. Mapping [Cla14b, Det99, VR20, Web98]. March [Ano09a]. Marginalia [Yal11]. Marie [Ank20, Der20, Lig20, WSP20]. market [Bec16, Kir08]. marriage [Ger08]. Marshall [Hal16]. Mary [Tur18]. mass [Bud16]. Mastodon [Bro11a]. materia [Ach11]. Material [Bas12]. materialised [HK13]. materialism [Kai04]. materiality [Mar20]. mathematical [Bri13b, Don14]. mathematics [Mor16b]. mathematization [SL14]. mating [Kla14]. matrices [GS11a]. Matter [Car99, MS14, Ger99, Kel11a, Pan12, Pen12]. Matter/Mattering [Car99]. Mattering [Car99]. matters [Dei16]. maximising [Pat20]. may [Bor16]. Mayr [Chu03, Pow13a, Wit15, Wit16]. McCance’s [Buk14]. McClintock [Nas99]. MD [Ayl18]. ME [Kar18]. meaning [ABM20, Art20]. meanings [Hen11, Rey05, Sil01]. means [Dar59, Dar90, Dar08, Mil12]. measure [Mor19]. Measurement [Det99, Mor19]. Measuring [Ste15b]. meat [Par14]. Mechanical [Bre06, Cha13, Cle12, Smi12]. mechanics [Orl12]. Mechanism [All05, Tee18, Wal12, DC02a, Gil17, Han14, Mat16, Meb14, MN12, Mos12, Nic12, Rus05, Web02]. Mechanisms [Des05, Soo12, ADF17, Bak05, BMG13, Cra05, Dar05, Gle05, IW10, Men12, MS14, Par16, Sch15, SM05, Tei19]. mechanist [BA05]. mechanistic [Bec15, Bri13b, Bur16, Des16, Har15, MN15, Man12, Meb14, Meu12, Pâs14, Rac16]. mechanistic-holistic [Rac16]. media [Ber10, Krá12, Lan16, Rhe16, Ste16]. medica [Ach11]. Medical [Ayl18, Dev20, FS15, Hed07a, Juz09, Sec12, SD14, Tho01, Ata12, Bau14, Bru03, Cha07, Dre98, GF09, Gra14, Hom08, Juk17, Krá12, Lee05, Lee10, Mae12, Mil20, Mon18, Ols11, Sin04, Sol20, Spa12, Sta13, Ste15a, Wei00, Ful20, Gab20, Hea20, Sol20, Ste20]. medicalisation [Gig14]. Medicine [Bra20, DiM18, Efs12, Fio19, Góm20, LZ00, Rad99, Ano09f, Bae20, Ber10, Bou10, BL07, Car12b, Car14, Cha07, Cle12, De 08b, Dev20, Edd08, FBF15, FF15, Gau00, Gil05, Gra10, Hut12, Jia17, KCN16, Kid13, LB18, Löw04b, Mét14, Par02, Slo14b, Ste08, Sto12, Wor07, Mil20, Wil00, Gil13]. medieval [Car10a, Smo10]. medium [Edd06]. mediumship [Gra14]. meet [Ano15h, Bue11]. meets [GB17, Mar19a, Mur01, Spe15]. melancholy [Bro11b]. Melanesian [Wid14]. membranes [Gro10]. memory [Ona14, Woo15]. men [GG06, Löw03, Sno01]. Mendel [Pil07]. Mendelian [Ber14a, Dar05, Woo15]. Mendelism [Gra19, Hod11, MW05]. meningococcal [BP19]. menstruation [Clo02]. Mental [Coo15, Pop20, Ano13a, De 08a, Fin08, Tso12]. merit [Bel05, Tam20]. meritocracy [Ram14]. merits [Dag19]. mesmerism [Sch10a]. Mesoscopic [MN19]. Message [Bor04a]. meta [Ste11b]. meta-analysis [Ste11b]. metabolic [Don14]. metabolism [Sta12]. Metagenomics [DO07]. metaphor [Ger04c, Mil16, Rey14, Rus05, Smi20, vR16]. Metaphors [Mye16, BP13a, Kap13, MS14, Sep16]. metaphysical [Del09, Sar13, Sma04, TS20]. metaphysically [Fry12]. metaphysics [Bou17, Eli07, Lem16, Soo12]. metatheory [MT17]. Method [Eli07, MT17, Man15, Wil99]. Methodological [Bro15a, GS16a, Sha01, Bae13]. methodologies [Cla14b]. Methodology [Gam20, Tji01, Har15, Per09]. methods [Bak18, Car12a]. metropole [SS11]. Mexican [Lav05]. Mexico [Ach11, Bir00, LB18, SD14]. Mice [Rad99, GSM19, GG06, Har18, HK13]. Michael [Ank20, Der20, Lig20, WSP20]. Microarrays [KC12]. microbes [Ped15]. microbial [Ams04, Cle07, Gre20b, Lee15, O’M08, Sap07]. microbiology [OB05, OD07, Str14]. microRNA [OEB10]. microscopes [Mai16]. microscopical [Sch03a]. microscopy [Drö11, Sch03a]. Microstudies [dC09b]. mid [Abr12]. mid-twentieth [Abr12]. Middle [Ker16].

Middle-ground [Ker16]. might [Mor12, Sti16]. migratory [Whi14]. Militarised [Dav11]. milk [Spe11]. Mill [PD08]. Miltown [Her11]. Mind [Car99, Alt07, Alt08b, Dyd16, Edd06, Has97, Hey15, Kai04, Sta12, Tho06, Tho16, WGEN15]. Minding [Car99]. Minds [Hal18, Har13b, Rou19]. minimizing [Gar14a]. Mining [Bau08]. Minister [Law12]. Mirandola [Rut10]. mirror [Qui04]. mis [De 08b]. Miscarriage [Ell14]. misguided [KS16]. mismeasure [BP13a]. missing [MHH10, Rat02]. mission [Cha07]. mistaken [WCB17]. Misunderstanding [Lym19, Ric00]. misunderstandings [FJEZM20]. mobility [Ram14]. Mode [TB16]. Model [DL12, Hal20, Aic16, Com09, Cra04, FF15, Ful18, Gre13, Gri13, Har18, Ink16, IR20, LA12, Lew00, Lud14, Meu12, Nic09, Pei15, Rie13, Sul20, FF15, Gir20, Nel20, Sul20]. model-building [Ink16]. model-taxon [Gri13]. Modeling [Gle05, MN15, Pot12, GB17, Hou05, KL13, MN19, Tee19]. Modelling [Dim17]. Models [ADF17, Art20, Bae14, Bro09a, mKC10, Cha13, Don14, Dri09, GSM19, GS11b, Gro11, Hoo12, HK13, Mae12, Meu12, NG11, Ols11, She10, Wal13]. Modern [Kle18, Lig20, WSP20, Alb12, Ank20, Ata12, Bak18, Cle12, Cur17, Der20, Ish17, Jia17, Kir13, Mox18, Per09, Røs17, Smo14, Sto12, Yal11]. Modernity [Edi20, Gra12, Cra18, Gau07]. modes [Leo13]. Molecular [Key99a, Key99b, SD09, dCS02, dC02b, BMG13, Cha13, Cre09, Dar05, Dei02, Fal13, GS11a, Gau00, Gau09, Gro11, Lea19, OS12a, Rhe09, San02, Ser16, Slo12, Str02, SDAM08, WL08, ZDSF11]. Molecularizing [Ser11]. molecules [Bec13, PD09]. monarchy [Top07, Sec12]. money [McM04]. mongering [GMST15]. monism [BS20]. Monkey [Wel14]. Monkeynut [Cob15, Hal14b]. monogenesis [Coh06]. monogenism [SS15a]. moral [Cas03, Gue20, Ker16, Ver05]. morality [Lev06]. morally [DPS13]. Morgan [Cla16]. Morphogenesis [Sun11, GB17]. morphology [Gri13, Tur00]. Morselli [Bra14]. Morton [KPB15]. mosaic [Bae14]. most [Hog16]. motherhood [Ric08]. Motives [Dag19]. motor [Chi19]. mountaineering [Bla04]. Mouse [Rad99]. moustache [Laf07]. movement [Ben16, Gaw08]. movements [Kra14]. Moving [BdC14]. mtEve [DiM20]. Much [Har18]. Müller [Sch03a]. multi [GB17]. multi-scale [GB17]. multicultural [YT10]. multifactorial [Ful18]. multifield [Cra05]. Multiple [She99, CA16]. multispecies [TS20]. murine [Mor14]. Museum [Juz09]. museums [KA03, Mat17]. music [Pan12]. Mutant [HK13]. Mutations [Bau14]. Mutualism [Ren15, Hal14a]. Mutualistic [Pra16]. My [vW13, Dyd16, Pen12]. mycology [Hom08]. mycorrhizal [She12a]. Mysteries [Rut10, Rad03]. Mystery [Rad03]. myth [Bec11, CB18, Sno01].

N. [Hel13]. name [Hor11, Lee10, Rad06, Shm20]. Names [Ere01]. Naming [Hor11, Sim99, Ken20, DiM18]. Nanay [Ste10a]. nanoscale [Bue11]. Naples [CC02]. Narrative [Dim17, Aic16, LB06, MHH10]. Narratives [Dav14a, Bea16, Kra14, SS15a, dSV14]. nation [Shm20, Slo14b, dSV14]. National [San02, SC14, Buk07, Dut18, Gue12, Tau06, Bal00, Ber14b]. nationalism [BL07]. nationality [Coh17]. Natural [BC01, Chr12, Des16, Des18, Mit03, MWC12, Nan10, Bak05, Ber20, Bir12, Car04, Coo19, Dar59, Dar90, Dar08, Dav11, Dav14b, Dim17, Eng20, Ewe19, Fin08, Gam17, Har16, Hun13, IW10, Lig16, LoT03, Mar98, Mat16, Mic12, Plu06, RBF11, RB13, Rey06, Ric12, RH98, Sha07, SM05, Slo06, Ste10b, Ste06b, Ste02b, SC13, Var18, Wal00, Wal06, Wal12, dC01, vWK15, vL37, Ste10a, Sha14]. naturales [vL37]. naturalisation [Ver05]. Naturalising [Hun06]. Naturalism [Lig16, FD19, Gla08, Hoc14, Spe14]. naturalist [Gra98, Gue10, McL08, vW13]. Nature [Det99, Pet02, Bon19, Bra07, Bre06, Bre02, Car16, Coh06, Cra18, Cre08, Dea10, Del07, Dut18, End08, End09, Fal13, Gam14, Gre15b, Hol19, Kei03, MN12, Nas99, Nic06, OCG16, Oka03, Rei04, Rey05, RH98, Slo06, Ste10a, Ste05a, Tau06, Tay15, Toe12, Tor13, Wel20, Whi14, Wil02b, Røs17]. natures [FJEZM20, Gla08]. Nazi [Efs12, Mil20]. Nazis [Hoc15]. Nazism [Sok02]. necessary [Bon13]. Needham [Roc11, Roc12]. needs [Lil04b]. negative [Bir12, KCN16]. Neil [Ank20, Der20, Lig20]. Neither [Sar13, Sch16b]. Nelson [Gir20, Hal20, Nel20, Sul20]. nematode [GS12]. Nemesis [Rau04]. neo [Bro11a, Del09, Lew19]. neo-Darwinism [Del09, Bro11a]. Neo-Paleyan [Lew19]. nerve [Lan02, Too03]. Nervennahrung [Sta12]. Nerves [Sta12]. nervous [Edd08]. Net [Aga01]. Netherlands [PM11, vdM08]. Nets [Aga01, AR98]. network [Lin14]. Networks [Aga01, AR98, Kro15, YH17]. Neural [Aga01, AR98]. neurally [Man04]. neurally-controlled [Man04]. neuroanatomy [Hak06]. neurobiology [Had99, Tso12]. neurodevelopmental [Ver14]. neuroleptics [Hen11]. neuron [Too03]. neuroscience [Aic14, Cra05, DL12, Har13b, Maz16, Soo12]. neurosciences [Bor16]. neuroscientific [Bog01]. neutral [Fry12]. neutrality [Gla12]. never [Sag17, Whi04]. nexus [OMW10]. Nicole [Gir20, Hal20, Nel20, Sul20]. Nihilism [Hea20, Dev20, Sol20, Ful20, Gab20, Hea20, Sol20, Ste20]. Nineteenth [Bud14, Car99, All05, Bog01, Buk07, Cla98, Hak06, MW07b, SS15b, Tam17, Wil99, Wor07]. Nineteenth-Century [Car99, Bog01, Buk07, Cla98, Hak06, MW07b, SS15b, Tam17, Wil99, Wor07]. Nirenberg [Hal16]. No [Joh12, Hor11, Ste08, vR16]. noise [Sch03c]. Nomadic [SSH14]. nomenclatures [Ken20]. nominalism [Edd06]. Non [Pug09, Bak05, BS13, Dus18, Hun13, MM20, She17, Won20]. non-binary [MM20]. Non-colonial [Pug09]. non-geographic [Bak05]. non-ideal [She17]. non-local [Hun13]. non-selectionist [Dus18]. non-sentient [BS13]. non-trivial [Won20]. nor [Sar13]. Normal [Low16, All12, Amu00, Gan03, Val12]. normative [HP13, SL14]. Normativity [Bar12, Chr12, MN12, Pop20, She17]. Norms [Mam06, Chi03, Con19]. North [VR20]. Northern [KA03]. Norwegian [Hoc15]. Note [Ano19j]. notebook [Mox18]. Nothing [Dil13, Tay09, De 08a]. notion [Mac07, Mic12]. notions [Wou03]. nouvelle [R81, R77]. novel [LaB14]. novelty [End01, LL13]. now-now-I [Ste08]. nucleic [Hal11]. number [vL37]. numbers [Ere01, KC12, Lou16, Mor16b].

Numerical [Ver01]. numerum [vL37]. nursing [Gol96, Gur98]. nurture [Tay15]. nutrition [BD19, Juk19, Orl12]. NY [Wil98].

O [Har13a, Meg16]. obesity [Sto12]. object [Bec10, Gau04]. objective [Won20]. objectivity [Car12a, Ell08, Fag09, Gla12, Juk17, SDAM08]. objects [Smi09]. Observation [Bac15]. observational [SS11]. observations [Ree06]. observer [Mac02]. observers [SS15a]. obstacles [Bri16, MW05]. obvious [Bin18b]. occurring [Coo10]. October [Ano07a, Ano19i, Ano20l]. Octopus [Hal18]. Off [Kre09, Pre10]. off-shoot [Pre10]. offenders [vdM08]. offs [Ink16]. often [Vel08]. Okanagan [Ken20]. Okasha [Sme13]. Old [Cun02, Cun03, Gau09, Jia17, KS16, PM11, Cre07, Man15]. O’Malley [Cal12b]. omnium [vL37]. oncology [BMG13]. One [Ros14, Gre13]. One-Sex-Body [Ros14]. only [Fee07]. ontogeny [Lew02]. ontological [Lud14, Wol14]. ontologically [Vec20]. Ontology [Ken20, Bar13a, DO07, Per09]. Onward [Wex08]. open [Bor04a, PM02]. operation [OB12, Sme13]. Operationalism [Hea20]. operationalization [Mac07]. operon [Rac16]. Opinion [Coh17, Kle18]. opportunities [GS16a, Mor11a]. optics [Zim01]. optimality [Ber20, Dri09]. options [Gar14a]. Oral [Aga01, AR98]. orang [vWK15]. orangutans [vWK15]. Orchis [Tab16]. Order [Rus17, Bec16, Car16, PS11, Swe15, Tor13]. ordinary [SLR12]. Organic [RH17, Ste02a, Teu11, De 19, Ste16]. organicism [All05]. Organism [Hol13, HP13, Ari03, BV16, Ben16, Com09, DAC^{+}20, Har18, LA12, Meu12, Nic14, Pea10, Sul20, Wol14]. organism-environment [Pea10]. organismal [McL20]. Organisms [BS13, Nic13, Wal06, Bab20, GL01, Hol13, HK13, Lel20, LA12]. organization [BE18, Gam14, IW10, Smi20]. organizational [Ger13]. organized [Toe12]. organs [Kre09]. Origin [Dar98, Rus11, Chu03, Dar59, Dar90, Dar08, Gil04, Kos16, LL13, San12, SD09, Tab16]. origin-of-life [San12]. origins [Cra04, DM14, EH12, Hel13, IR20, Jac03b, McL20, Pea10, Pei15, Rad00, San14b, Smi13, Wit15, Wit16]. ornithological [Law14]. ornithology [Mac02]. orthodoxy [She14, Whi04]. orthogenesis [Ule14]. Osijek [Ata12]. osteoporosis [Wyl10]. Oswald [Hal16]. Other [Kro99, Bed15, FJEZM20, Mat17, Tau06, Hal18]. Otto [Nic09]. Ottoman [Cha07]. our [Dri09, Gau07, Pen11]. outcome [Fis15]. Outline [Bar06, Wel21, Wel26, Wel31, Wel56, WPW71, W^{+}76]. overload [MWC12]. overloaded [Wei00]. oversimplification [Wit15, Wit16]. own [Kov17, Löw04a]. Oxford [Cob15, Sar98]. oxidative [Pre10].

P. [Fal13]. p53 [BMG13]. Pain [Bae20, Bur10, Ger20a, Gue20, Lew20]. Palaeolithic [Man15]. Palaeontology [Bow17, Coh17, Man17]. paleodemography [Ger99]. paleontological [ST17]. Paleontology [Cai02, Sep14, Pod17, ST17, Tam17, Yu17]. Paley [Bru13, Gli00, Hol13, McL08, Sha14]. Paleyan [Lew19]. pandemics [Lee15].

Pandora [Mil20]. Paolo [GvA19]. paper [Fal13, Hed07a, Ned10, vW16, ZDSF11]. Paperback [Mur01, Hal18, Qui04, Sar98, Tur18, Whi04, Wil98]. Paradigm [OB05, Hut12, KS16]. paradigms [Hut12]. paradise [Men17]. Paradox [Rap04, Law12, PS11]. parallel [Alt08a]. Parallels [San14a, Rad08b]. parapsychology [DM14]. parents [Sil01]. Paris [Qui04, Gau02b, GvA19, Ste17]. Park [Jon18]. parks [Tau06]. Parsing [Die16]. part [Wit15, Wit16, Alt07, Alt08b, Key99a, Key99b, Ver04, Ver05]. partium [vL37]. partly [Sha04]. partnership [Par02]. parts [Güt13, Kai18, Teu11, vL37]. passions [Law14]. Past [Fos19, Gis17, Ist13, McM07, Pri17, Rus11]. Pasteur [Cas05, PM02, Tho16]. Pasteurization [Ber14a]. patchwork [Ger20b]. patenting [BC08]. patents [Cha16]. path [Car12a, Car14, Coo15]. Path-dependence [Coo15]. pathogenic [Spe11]. Pathological [Hac10]. pathology [Par03, Tul11, vdB02]. patient [SM06]. Patients [Mil19]. patriotism [Cow09]. patron [Aic14]. pattern [Ned10]. Patterns [Ayl18, Lip14]. Paul [Qui04]. Peaceful [Iid20]. Pearson [Pen11]. peasants [Lav05]. pedagogy [JA08]. Pemberton [Lig20, Ank20, Der20, WSP20]. pen [Cun02, Cun03]. Penguin [Whi04]. penguins [Wex08]. Penicillium [Löw05]. Penitentiary [Com09]. Penrose [Val12]. peoples [Swa05]. peptone [Löw05]. perception [Pan12, She14]. Performing [PT13]. period [Löw05]. Peripheral [Cow09]. periphery [Spa12]. permanent [Sle09]. perplex [FS00]. Perrault [Des05]. persistence [Cav04]. personae [Bor05]. personhood [Dol01]. perspective [Den04, FD15, Fio19, Ham13, Kai04, Pop20, RW16, Sch16b, TS20, Wal06]. perspectives [Ber20, Bri13a, GKG13, McL20, Smo15, SM12, Tay09]. perspectivism [Cal12a]. perspiration [Dac12]. persuasion [End01]. Peru [Bro10]. perverse [Can17]. pesticides [Vai15]. Peter [Hal18, Gun12]. phage [San10, Cre09]. pharma [Ell04]. pharma-funded [Ell04]. pharmaceutical [Ada05, Gau01]. pharmaceuticals [Mil19]. Pharmacogenetics [CWJ06, LR06, AH06b, Hed06, SM06]. Pharmacological [Tso12]. phenomenological [FD19, Pan19]. Phenomenology [Kid17]. phenomenon [Dom06]. phenotype [Mun15, NSY07]. phenotypic [ER15, Nic15, Pei15]. philosopher [Cal12a, Pig13, Plu07]. philosopher-scientist [Pig13]. Philosophers [Ano15h, Chi09]. Philosophical [Bri13a, GKG13, PKD16, Ber20, Low18, Mat16, OCG16, Ped15, Sch03a, SM12]. philosophies [Pen11]. philosophique [R81, R77]. Philosophy [Ano01d, Ano02b, Gam14, Pea12, Plu14, Spe15, Wil03, Ano98c, Bar13a, Bro15b, Chi03, Dup12, Edd06, Edm13, Gil05, Kid17, Mos12, OD07, OMW10, O’M14, Ode14, Qui16, Sar14b, Sar14a, Sch03b, Slo06, Smi12, Som14, WMN15]. phlogiston [Lud14]. Phosphorus [Cre09]. Phosphorus-32 [Cre09]. phosphorylation [Pre10]. Photography [Lef17, Ell08, Mit14]. Phylogenetic [Sha11]. phylogenies [SDAM08]. phylogeny [O’M13]. physical [Coo19, Fal13, KL15, Mor11a, Ver04, Ver05, ZDSF11]. Physics [McL11, Rad11, Bue11, Dom06, Fre11, GB17, Noa14, Poo11, Row11]. Physiology [Del14, Tho01, Cun02, Cun03, Mun15, Orl12]. phytotronist [Mun15]. Pico [Rut10]. picture [dCS02, dC09b]. Picturing [Amb15, Dac12, Rei04]. Piers [Ren15]. Pietsch [Tor13]. pig [Rad08a]. pigeons [Abr13]. pigs [Bra07, Kir08, Löw03]. pinguedinem [Sto12]. pipettes [Fra07]. pixels [CK00]. place [Ree06, Smi13, Tre04]. Placebo [Tei19]. placental [MH14]. places [Krá12]. plain [Wel21, Wel26, Wel31, Wel56, WPW71, W^{+}76]. planes [Bra11a]. Planning [De 16]. Plant [Mun15, Rus17, BP15, Bar15, Ber14b, Leo13, MZ00]. plantarum [vL37, MWR07]. plants [Kle03, Smo15, vL37]. Plasma [Cre99]. Plasmids [Cha16]. Plastic [Sle09]. Plasticity [Pei15, BP15, Des18, Nic15]. platform [Gri13, Meu12]. platinum [Ste11b]. Plato [Pow13a]. platypus [Law12]. please [Joh12]. Pleasure [Ger20a, Lew20, Gue20, Had99]. Plumed [Law14]. Pluralism [Mau20, Abr13, AG19, Con19, Del09, DiF19, Gre16, Mor16a, Pla15]. pluralisms [Mar20]. pluralist [Kid13]. pluralistic [Hou05]. Plutynski [Gre20a, Plu20]. pneumococcal [Cre08]. Po-bo [Ste04]. point [Per09]. Poland [Bal00, Gaw08]. policies [Kra14]. Policing [Fio19]. policy [AR20, Bid20, LR06]. Political [Hal14a, Azz10, Gla12, RO00, Smi09, Ren15]. Politics [Ren15, AG14, Cre02, Cur17, EH12, Ger04d, GCT17, Hal14a, Hel13, Hey15, Lan08, Mot08, San02, Tho06, Top07]. pollen [Mil16]. polyoma [Mor14]. pool [Per16]. poorly [Bog01]. pop [Hoc16]. Pope [Gig14]. popes [Azz10]. Popper [Kid13, Stu12]. popular [Erl09]. popularization [Ule14, Sep14]. popularizing [Smo14]. Population [McL20, MP11, Ari07, Chu03, IR20, Löw12, Ram08, Spe15, Wit15, Wit16, dSV14, Ram14]. Populations [Abr13, Abr06, BdC14, Bou11, CK00, Gro11, Löw02, Mil15, SD14, WGEN15, dC14]. positions [Dom06, Ker16]. Possession [Tos02]. possibility [Tay09]. possible [Efs16, For10, Pie20]. possibly [Rey12]. Post [Bau14, Bar20, Bon13, Ere01, Gau02b, KCN16, Str02, Swe15]. post-genomic [Bon13, KCN16]. post-Linnaean [Ere01]. Post-Lysenko [Bau14]. post-war [Bar20, Gau02b, Str02, Swe15]. postgenomic [Mel17]. Postgenomics [BS20, Die16]. posthumous [MHH10]. postmodern [Mar98]. postnatal [Buk14]. postscript [Edw16]. postulates [RW16]. Postwar [Ayl18, Cre02, Cre07, Dav11, Hed07a, Her11, dCS02, dC02b]. pour [Cla07b]. Power [Gut18, Røs17, Ste06b, Swa05, Tul11]. powers [Del14, Ver04, Ver05]. pp [Aga01, Ayl18, Cob15, Dea10, DiM18, Gur98, Hal18, Har13a, Mur01, Oka03, Qui04, Ren15, Sar98, Tay15, Tor13, Whi04, Wil98]. Practical [De 16, The14]. practice [Ata12, BC08, Bru03, Car10a, Hol17, Mat16, Mer14, MW07a, Pen12, Sar14a, WEHT20, WGEN15]. practices [Cre02, Dea10, End08, Gam20, Gra07, Ken20, PS11, Sam13, Soc20]. Practicing [Lef17]. pragmatism [RO00]. praxis [BM15, Bud14]. Pre [AG14, Sno01, GvA19, McM20, vH00]. pre-1945 [McM20]. Pre-Adamite [Sno01]. Pre-conceptional [AG14]. pre-Linnean [GvA19]. pre-World [vH00]. precautionary [Res03, Ste11a]. Precise [Det99]. preclinical [Lem17]. precognition [Pri14]. predator [Rad06]. predicates [Clo02]. predicting [Azz10]. Prediction [FBF15, Wil01, Bro15a, Col14, DL12, FF15]. Preface [Ano15i]. pregnancy [AGD14, Dav14a, Ell14, OG14]. pregnant [San14b]. Prelude [Key99a]. premenstrual [KW08]. premises [Coh06]. Prenatal [Hog18, Löw14b, Buk14, Hog16, San14b]. preparations [Kno15]. prepared [Gau05a]. prescription [The00]. Present [Fel17, Smo14]. presented [GS16a]. preservation [Dar59, Dar90]. Preserving [Tau06, Sha16]. Press [Aga01, Ayl18, Bra20, Cob15, Dea10, DiM18, Gur98, Oka03, Ren15, Sar98, Tay15, Tor13, Tur18, Wil98, Rou19]. presuppositions [Meg16]. preventable [AG14]. Prevention [Ayl18]. preventive [Gar14a]. Price [Ayl18, Cob15, Dea10, DiM18, Gut18, Hal18, Har13a, Oka03, Qui04, Ren15, Tay15, Tor13, Tur18, Whi04]. priest [Des98, Whi04]. Priestley [Pri14]. primary [Ser16]. primatological [Bur06, Ree06]. Prime [Law12]. princes [Azz10]. Principe [Roo13]. Principle [Gam17, Bar13b, Res03, Ste11a, Luq16]. Principles [Dro01, Sko11, Ans02, Coe12, Pow13b, Ric12]. Prion [Key99a, Key99b]. Prions [Key99a, Key99b]. prisoner [Com09]. prisoners [SC14]. proactive [GS16a]. Probabilistic [RBF11]. probabilities [Dro12]. probability [Won20]. Problem [Key99b, BP19, Brä12, CWJ06, Del00, DiF17, Gau09, Güt13, Hor11, Mur01, Rey05, Shi06, Slo13]. problematic [Man15, Alt07]. problématique [Gis14]. Problems [Cai02, All12, Ell04, Ped15, Ram14]. process [Dep11]. Processes [Dup12, Ste02a, BM18, DG16, Sun11, Bar13a]. Producing [DiM20]. Production [BC01, CK00, mKC10, Gau02a, Kir13, Qui04, TB16]. products [Coo19]. Profession [Tho01]. professional [RO00, Wil02a, Wil03]. professionalism [Krá12]. progeny [CBV11]. program [GS12, Hal20, Wri11, Sch14b]. programmes [dC98]. Progress [Man17, Ble10, De 19, Kro03, Sha01]. project [Bae12, Ber20, Pol15, Sha16, Van00, Bra06]. projectibility [Clo02]. projects [Kov17]. Prokaryote [O’M13]. promiscuous [Fre11]. promise [SM06]. proof [Bon05]. propagandist [Fan17]. propensity [Sob13]. Proper [Hol13, Mau00, Mil06]. property [Bre05, Yi18]. prophecy [Ver04]. prophet [Ver04, Ver05]. prophylactic [Fre11]. proportion [vL37]. proportionem [vL37]. proposal [Del11b, MM20]. Propositional [SB19]. Prosaic [Abr13]. prospects [Ble10, SM06]. protein [DC02a, IW10]. Protocol [Ned10]. protocols [DL12]. provincial [FS15]. proving [Lin16]. Prüm [PT13]. Psychiatric [BH11, Top07, Mau16b, Var18]. Psychiatry [Bra14, Hen11, Kit20, Mau16a, Pol20, Pri14, Tho06]. psychic [Del14]. Psychical [Noa14, Pen20, Som14, Bra14]. Psychoanalysis [Bra06, Tos02]. psychological [CC16, GS16b, Gar16, Smi05, Sti16, Tho16, Wal08]. psychologies [Gis20]. Psychology [Bra14, Car99, All12, Bae20, Bor05, Bru08, Car12a, Cla16, Fay12, Fee07, Fee12, Has97, Hat12, Kla14, Mac07, Mül12, Mur01, Pen20, Sch03b, Stu12, SM12, Wil99, Wil02a, Wil03, All12, Gun12, Mur01, Bra06]. psychometrician [Mic06]. Psychophysics [Mic06]. psychosomatic [Eas11]. psychotropic [BH11, PS11]. Public [Bir00, De 16, Liv04, Sol00, Bau14, Gra12, Lou03, Pal00, She14, Wil00, Wil07b]. publication [Ano05b, Ano05c, Ano06a, Ano06b, Ano06c, Ano06d, Ano07b, Ano07c, Ano07d, Ano07e, Ano08a, Ano08b, Ano08c, Ano08d, Ano09b, Ano09c, Ano09d, Ano09e, Ano10a, Ano10b, Ano10c, Ano10d, Ano11a, Ano11b, Ano11c, Ano11d, Ano12a, Ano12b, Ano12c, Ano12d, Ano13b, Ano13c, Ano14b, Ano14c, Ano14d, Ano14e, Ano14f, Ano14g, Ano15b, Ano15c, Ano15d, Ano15e, Ano15f, Ano15g, Ano16a, Ano16b, Ano16c, Ano16d, Ano16e, Ano16f, Ano17b, Ano17c, Ano17d, Ano17e, Ano17f, Ano17a]. Publications [Mur01, Slo14b]. publics [Smo14]. published [GvA19]. Publisher [Ano19j]. Publishing [Har13a, Soc20, Sok02]. Punctuated [Dre19]. Punnett [Edw12, Edw16]. pure [Ber14b, MW07b, Sar13]. purpose [Hun06, McS16, Mic12, Mor16b, Ste06b, Wal12]. purposes [Wal06]. pursuit [NG11]. Putnam [SM04]. Putting [Don15, Fel17, Smi13]. puzzle [Sac14]. pylori [Hut12]. Pyramus [Dro01]. pyrrhic [Tam17].

Quaalude [Her11]. Quackery [Krá12]. quality [Ram08]. Quantifying [Ewe19]. quarter [Tho16]. Quasi [Bro09b]. Quasi-independence [Bro09b]. Queen [Dag17]. quest [Jer13]. question [Bel05, Bru13, Dav99, Gab20, KS16, LH10, Rat02, Rus16, Sma04, Zam06]. questions [Bre02, Hed07b, Pri17, Ree06]. Quetelet [Ari07]. quiet [Too03]. quincunx [ARR17]. quorum [Lyo07].

R [Fal13, Gun12, Kap13, Pen11, Smi13, Wil98, vW16]. R. [Ber20, Moo07]. Race [BD19, Góm20, Hoc16, Lef17, Rad08b, Shm20, Smi13, Spe13, Alt08a, Bra20, Bri17, Fab10, Gan13, Gan14, Gis08, Hoc15, Lud14, Mar16, McM20, Pig13, Ses13, Slo14a, Spe15, WMN15, Smi13]. races [Dar59, Dar90, LB15, Mil12, Mil15, Tem13]. Racial [Hoc13, Hoc14, Lem16, Smi13, Spe14]. raciology [McM20]. Radical [Con19, Mor16a]. Radick [Alt08a]. Radioactivity [Mye16]. Radioisotopes [Iid20, Cre02, Cre09]. Rafinesque [End09]. Ramón [Fio11, GL15]. rare [FB15]. Rational [Far06, Ger08, Ada05, Fra14, Güt13, Lew13]. rationale [Des18]. rationality [OB12, Pol15, Sme13]. Ray [Bar06, Ona14]. Raymond [Tre04]. Re [LA12, DiM20, FS15, KCN16]. re-examination [FS15]. Re-thinking [LA12]. reaction [OG14, Sch03c]. reactivity [Jam10]. read [Ren15]. Reade [Hes15]. reading [Hed07b, Liv04, Men17]. real [Can17, Coo10]. realism [Bar13a, Fre11, Jer13, Lem16, RH98, Sma04, YH17]. reality [Bec11, Bor04a, Teu11, Wol14]. really [Bru13, Del11b, Fis15, vW13]. realm [End01]. reason [Hun06]. reasoning [Dro12, Rad01, Tso12]. reassessed [vW16]. Received [Ano01d, Ano07a, Ano09a, vW13]. reception [Bow17, ST17, Tam17, Yu17]. reclassification [Bry08, Hut12]. recognition [Fio19]. recognize [Gre20b]. recombining [Bau08]. reconciled [Bec15]. reconsideration [LB15]. Reconstructing [dC02b]. reconstruction [Gau02b]. reconstructions [Vae14]. records [Sha16]. Recovering [Kit20, Cun02, Cun03]. red [Pre13, Dag17]. Red-Queen [Dag17]. Redefinition [Tho17]. Redi [Par14]. redoubtable [Rey10]. reduction [AG14, Cra05]. Reductionism [GB17, Gro11, RR06b]. reductively [Kro15]. Redux [KPB15]. Reevaluating [Par14]. refer [Mau16b]. reference [Hed07b]. reflecting [Ste06b]. Reflections [Cla07a, dC16, Poo11, Tan03, RH98, Slo14a]. reform [Cha07, Gaw08, The00, WL08]. Reforming [Rou19]. refracting [The00]. refugee [Hac10, Wei00]. Refugees [Sok02]. refutes [RR06b]. regimes [Spe11]. regional [Mit03, Top07]. Régis [Des05]. Regularities [Bog05, Jef08]. regularity [Des16, Har15]. regulation [Ata12, OEB10, PD08, San00]. Reimarus [vdB18]. rejoinder [Don17, vdB02]. Relations [Dar05, Bud14]. relationship [Alt08b]. relatives [Tay09]. Relaxed [Var18]. relevance [Hym12, Pâs14, Zam06]. reliability [Ned03]. reliable [Tay09]. religion [Oka03, Wil02b]. remarkable [Esp17]. remarks [Hal08]. remedy [Nor07]. reminder [Vae14]. Renaissance [Azz10]. Renegotiating [Rob13]. reopened [Gli18]. Replaying [Sep16]. replicated [Bog01]. replicator [KS16]. Reply [Bau20, Ger20a, Nel20, Plu20, Set20, Ste20, Ste10a, WSP20, Alt08a, Don17, For12, Sag17]. representation [Chi19, Lel20, Tee19]. representational [Mam06]. representations [Sle09]. representatives [FM07]. Representing [Bud16, Web98, Wei20]. Reproduction [Cla07b, Bar20, MW10, PD08, Teu11, Wal01]. Reproductive [Löw10, Cla07a, Fra07, Gra07, McM07, Wil07a, Wil07b]. reproductively [All01]. Republic [Coo12, JA08, Soc20]. Republican [Edi20]. require [Kov17]. Rereading [All12]. Rescue [Pal00]. Research [Cre99, Pol07, Swa05, Ada05, Bog01, Bra14, Bri16, BE18, Bru11, Buk14, Cal12b, Clo02, Com09, Cre02, DAC^{+}20, Efs16, Fos19, GS11a, Gri13, Hag01, Hal20, Har18, Juk17, KC12, KCN16, Ker16, Kov17, Leo12, LB15, Mag05, Mau20, Noa14, OEB10, Par16, Pen20, Pie98, San10, Sch14b, Sch03a, Som14, SC14, Sul20, Wil07b, vH00]. Reshaping [Hen11]. resistance [Cre07]. Resister [FK07]. resolution [Pan19, Too03]. resolutions [Pie20]. resource [Bon19, Rad14]. resources [Bre05, Iid20, Nan10, Ste10a]. Response [Sar14b, Ano09f, Cal12a, Kla14, Lin02, SC13]. responses [Car04, Wal13]. responsibility [Fos19, dMM03]. responsible [JA08]. restless [Erl09]. restoration [Gar14b]. restored [Str14]. result [Mor11b]. results [Pil07]. Resurecting [McM20]. resurgence [Kra14]. Rethinking [Chi09, Eng20, MS14, SLR12]. reticular [Too03]. Rétif [LoT03]. return [Dav11, Hel13, McM07, Mel17]. reveal [KPB15]. Reverend [Edd08, McL08]. reverse [Gre15a]. reversion [ARR17]. Review [Aga01, Ayl18, Bau20, Bra20, Cob15, DiM18, Edi20, Góm20, Gur98, Gut18, Hal18, Har13a, Iid20, Kit20, Mur01, Oka03, Pol20, Ren15, Sar98, Set20, Sie20, Tay15, Tor13, Tur18, Wil98, Ano98a, Ano98b, Ano99a, Ano99b, Ano99c, Ano99d, Ano99e, Ano99f, Ano99g, Ano00a, Ano00b, Ano00c, Ano00d, Ano01b, Ano01c, Som14]. Reviewing [Bar13a, Fal13, Kap13, Gil13, Hel13, Pow13b, Roo13, Sme13, Smi13, Sta13]. revise [Coo15]. Revisited [Rap04, Gun12, Rac16, Vaz16, Zam12, Fal10]. Revisiting [Hun19, Pil07]. revival [McL08]. Revolution [Bac15, Bro15a, Wor07, BC08, Too03, Ver01, Bar15]. revolutionary [Dav99]. revolutions [Löw10]. reward [Col14, WCB17]. Reydon [For12]. rhetoric [Gra98]. ribozymes [San12]. Richard [SS15a]. Richards [SC13]. Richet [Kro99]. rid [Reu12]. riddle [Mar16]. right [Bre05, Ful20, Gre17, Has97, Ric12]. rights [Swa05]. rise [ARR17, Hoc15, Löw14b, Nor07, Pug09, Ver14, Win13]. risk [AG14, Gar14a, FF15]. risks [Bid20]. Ritual [Dre98]. river [Löw04a]. RNA [Fis15, Fis15, San12]. Road [Cob15, Hal14b, Kri02]. Robert [Sta13, Tur18, Buk14, Gra04, GF09, Gra10, Meg16, SS11]. Robertson [Liv04]. Robustness [Par16]. Rockefeller [Pal00, RO00, Wei00, Wil00, Bal00, Bir00, Gau00, LZ00, MZ00, Sol00]. Rod [Ger04d]. rodent [McL07]. role [Bak05, BV16, FK07, Gil04, Hol17, Hul01, Lea19, LB15, Meb14, Mor11a, RBF11, San10, SLR12, Tan03, Toe12, Web11]. roles [Kos16, OS12a, San12]. Romania [Tur20]. Romantic [Gre15b]. Rome [Gig14]. Ronald [Pil07]. Room [Hun13, Bro11a]. rooms [Lin16]. Roots [RH17, Bar15, Bol12, Dre19, Sch10a, vdB13]. Rosenfeld [Aga01]. roundtrip [Gau02b]. Roussel [Tre04]. Royal [Røs17, Mil20]. Ruben [SD14]. rule [Pri12, Sch13a]. rush [BM16]. Russel [FK07, Mit14]. Russia [Dav99]. Russo [Cla12, FD15]. rustics [Ach11]. Ruysch [Kno15].

S. [Sme13]. sacrifice [Dre98]. safety [Bon05]. Sage [Mur01]. Sagoff [Don17]. saltationism [Bow14]. Salvador [Mon18]. same [Kas13]. Samí [Ken20]. Samuel [Gut18]. sanction [vW13]. sands [Lee10]. sanitaire [Pal00]. sank [Whi04]. Santiago [Fio11, GL15]. sapiens [Oka03]. Sapp [Sar98]. Sarah [Mor14]. Sarawak [vW16]. sarcoma [San14a]. Satellite [Suá01]. Satellite-DNA [Suá01]. savants [Ach11]. savers [Gau02a]. Saving [Per16]. savoirs [Gau02a, Qui04]. scale [Bec15, Bud16, Des18, GB17]. scale-free [Bec15]. scales [BM18]. Scaling [Lin14]. scandal [Etk08]. scenarios [Kos16]. Schaudinn [Lin01, Lin02]. Schering [Gau05a]. scholarship [Gab20]. School [Wil00]. Schrödinger [Dom06, WH17]. Science [Bac15, BL07, Kir13, Lou03, Lyn13, Mur01, Pri12, Rud11, vdHP14, Ano98c, Ano09f, BP15, Bau14, Bra06, Bre02, Bud14, Bud16, Cal12a, Chi09, Col13, Cow09, Cur17, Dag17, Dag19, Del00, Edm13, Efs16, EH12, End08, Fab10, Far06, FS15, Fry12, GSM19, Gau01, Gre15b, GL15, Har16, Has97, Haw10, Hel13, Hes16, Hol93, Jon18, Juz09, Kel11a, Kel17, Kin16, KL13, Kos01, Kre09, Leo13, Men17, Mil06, OMW10, O’M14, Pen20, Rey06, Rey14, Rhe09, Rob13, RPLBMA20, RH98, Rou19, Rus05, Shm20, Smi13, Som14, SC14, Sto04, Str14, Tam20, The00, TB16, Toe12, Tur20, Wil07b, WL08, Yal11, vdB02, Mun15, Dea10, Wil98, Hol19]. Sciences [Ano01d, Ano02b, Wil12a, Wil03, Bae13, Cla07a, Cla12, Gau02b, GL01, Gro10, Hey15, Jef08, Leo12, Löw04b, Lud14, Rad11, RP16, Rei12, Sam13, SS15b, Ste06a, Str12, dC16]. Scientia [Roc11]. Scientific [Abr12, BC01, Ble10, Bor05, Cal12a, FM07, Lig16, Ram16a, Soc20, Bae14, Bak18, Bed15, Bou10, Bra06, BE18, Cla07b, De 08b, Efs16, Ell08, Erl09, Esp18, Fag09, Gau00, Gis20, Har99a, Hed07b, Iid20, LaB14, Liv04, Mat16, Nic09, Røs17, San10, Sch16a, SP16, She17, The00, Wex08, WGEN15, YH17, vH00]. scientist [Mac02, Pig13, Win01, Wri11]. Scientists [Tur18, Wei00]. Scotland [Fin08, Liv04]. Scots [Ken20]. Scott [Tur18]. screening [Bid20]. Scribal [Yal11]. script [WH17]. search [Cre08, De 08b, EH12, Fag07, Hel13, Lav05, Lee15, MHH10, Sun11]. Searle [Le 18]. Second [JA08, Bry08, Dre19, Tho16]. seconds [BM18]. Secrets [Roo13, Pri13]. sectioning [Hak06]. secundum [vL37]. Seed [Per16, HA09]. seedbed [Aic14]. Seeds [MZ00]. Seeing [Bon19, Bra11b, Le 18]. seeking [Hac10]. segregation [Shm20]. Selection [McL13, Tab16, Bak05, Ber20, Bir12, Bow14, Cre07, Dar59, Dar90, Dar08, Des16, Des18, Dim17, Ewe19, Hun13, IW10, Ist13, Mat16, Mil12, Nan10, Plu06, RBF11, RB13, Ric12, Sha08, Sha04, SM05, Ste10a, Ste10b, Ste02b, SC13, Wal08, Wal00, Wil06]. selectionism [Mel17]. selectionist [Dus18]. selects [O’M08]. Self [Vaz16, Gra98, LB15, San13, Teu11]. self-identified [LB15]. self-making [Gra98]. self-reproduction [Teu11]. Self-tolerance [Vaz16]. sell [Rus11]. sell-by [Rus11]. selves [Kro99]. semantic [Mau16b]. semantics [SM04]. seminal [Ans02]. Semmelweis [Gil05, Per09]. senescence [Mor19]. sense [Cla14a, Dil13, Leo12, MP11]. Sensibility [Det99]. sensitivity [Güt13]. sensory [Chi19]. sensory-motor [Chi19]. sentient [BS13]. sequence [Ben16]. Sequencing [Low18, Ona14]. serial [Hak06]. seriation [Lym19]. serology [Löw04a]. service [Efs12, Wil07b, vdHP14]. services [OG14]. Serving [Shm20]. Seth [Bra20, Sie20, Góm20, Set20]. Setting [Wyl10, Har18]. seventeenth [Bro10, Des05, Orl12]. seventeenth-century [Bro10, Orl12]. Sewall [Hod11, Ish17]. Sex [Ger99, Ros14, Tot07, FS00, Gau04, Gau05a, Gaw08, Hal08, Hol93, Mil06, Ove08, Roc12, Wil98, vdM08]. sex/gender [FS00]. sexual [JA08, Mil12, Sch10b, Wal08, vdM08]. sexualis [Roc11]. sexuality [Chi09]. shadows [BD19, Gol13, Wal00]. Shakespeare [Har13a]. Shale [Bar09, Bry08]. shape [Coe12, Pow13b]. shaped [San10]. shapes [Sch13b]. shared [Iid20, Sun11]. shed [Dag17]. sheds [KS16]. sheep [GSM19]. Shelley [Tur18]. Sherpa [Ken20]. shift [Hut12]. Shifting [Fre11, Sha04, Hod11, Ish17, Lee10]. shoot [Pre10]. shorebirds [Whi14]. Should [FD19, Gra19, Bud16, Vel08]. Sibylla [Smo10]. Sibyls [Smo10]. sick [Mon18]. Side [BH11]. sides [Kas13]. siècle [BP13b]. Siegel [Lin01, Lin02]. signaling [RBOH20]. Signals [BHZ20, ABM20, SB19]. signé [R81]. significance [BE18, Cas03, Gam17, Hal11, Tul07]. significant [DPS13]. signs [Edd06, Kas10]. silent [Wyl10, Vai15]. simian [SC14]. similarities [Ata12]. similarity [Tay09]. Simpson [Wit15, Wit16]. simulation [Gra13, MN13]. Simulations [Sep16]. since [AGD14, DiM18]. Sir [Kid14]. site [vL37]. sites [Ree06, SS11]. situm [vL37]. Six [Ell04]. Sixteenth [Amb15]. skull [Fab10, Smi13]. Skulls [Juz09]. slabs [Jer13]. sleeping [Ste08]. sleeping-and [Ste08]. Sloan [Fal13, Oka03]. slow [Dei02]. small [DE16, Ped15, Sha07]. Smith [Hal18, Liv04]. Snow [Tul11]. so-called [RH98]. Social [Bec11, Don20, Pat20, Sin08, Ågr18, Bil17, Bra06, Can17, Cla12, Fag07, Fag09, Fay12, Fio19, Gol96, Gur98, Har13a, JA08, LoT03, LZ00, Mot08, Pot12, Rad03, Swe15, Tha98b, Wil12b, dMM03]. socialism [Cla06]. sociality [Lyo07]. Société [Cla07b]. societies [Cla07b]. Society [Cla07b, Haw10, Löw04a, Mai16, Sch14a, Wil02b, Cla07b, Ram08, Røs17, Oka03]. socio [Sha07]. socio-bacteriology [Sha07]. Sociobiology [Hol93, Wil98]. sociology [Col13, Edm13, vdB02]. sociopolitical [Iid20]. soil [Str14]. Solastalgia [Gla12]. solitary [Tre04]. Solly [Bur06]. Sols [San00]. solutions [Efs16]. Solving [Mar16]. Some [Bae20, Wal13, Wil99, Bog01, De 08a, Fee07, Hal08]. Something [MH14, Lew04b]. soul [Pen20]. Sound [Pan12]. sources [Chr12]. South [Aic14, Men17]. South-Californian [Aic14]. Soviet [Sol00, Bau14, Kre09, Tur20]. Soyer [Cal12b]. Soyoung [DiM18]. spa [Edd08]. Space [Spe13, BP13b, Drö11]. Spain [Gra14, RO00, San02]. Spanish [Cow09, GL15, JA08, San00]. speak [Bak18]. Special [Don20, Bud16, Hed07b, Sch14b]. Specialization [Hom08]. specialties [Ger13]. speciation [Bak05]. specie [Rey05]. Species [Dar98, Lew12, Vel08, Bou17, Chu03, Con19, Dar59, Dar90, Dar08, LaP07, Mor16a, Rad08b, Rey05, Rey06, Rou19, Slo13, Rus11]. specificity [MW10, Pla15, Sin04]. Spectacle [Lig20, Gra14, Liv04]. spectrum [Ver14]. Speech [Ing99, Bak18]. Spencer [Lig16, Pea10]. Spinoza [Eng20]. spirit [Mit14]. Spiritism [Bra14]. spirits [Kas10]. spiritual [Bla04]. spiritualism [Del14]. spoils [Juz09]. spontaneous [Bec16, Par14]. Spot [Sch16a]. Spreading [Bru03]. Spring [Vai15]. square [Edw12, Edw16]. stability [Pei15]. stabilization [Ned10]. Stabilizing [Ver14]. stable [Bon05, Sch03c]. stage [BdC14, Lou03]. Stages [Meu12]. stakes [Cre07]. stand [Per09]. stand-point [Per09]. Standard [Har18, FF15, Kir08, Ste11b, SC13, Wyl10]. Standard-setting [Har18]. standardisation [BH11, Kir08]. standardization [KW08]. standardized [Hor11]. Standardizing [PS11, Pie98]. standards [Bon05, Juk19]. Stanovich [Pol15]. Stars [Kas10, Ral10]. start [Dei02, Don16, Mox18]. Starting [Ped15]. Stasis [Luq16]. State [Wil98, Cle08, Gra10, Gra12, Hal08, Jia17, Laf07, Mot08, Nor07, Ove08, Slo14b]. States [Juz09, Hag01, Kei03, Soc20]. Stateville [Com09]. Station [Pol07]. Statistical [Coo15, Ano13a, Ari07, ARR17, Hun13, Pen11, SLR12]. Statistics [Pil07, LB06]. status [Lyn13, Sep14]. Stegenga [Ful20, Gab20, Hea20, Juk17, Sol20, Ste20]. stem [Bry08, Fag07, Ker16]. stemness [LMR09]. Stephen [Dre19, Sep16, She14]. Stepping [TB16]. steps [DE16]. sterility [Nor07]. Sternberg [Gut18]. steroid [Lav05]. steroids [Gau04, Gau05a]. Stewart [Edd06, Mor14]. stick [Mon16]. Still [vH14, Meb14]. stochastic [IR20]. Stone [Man15, TB16]. stones [Sno01]. Stop [GCT17]. store [Bon19]. stored [Bea03]. stories [GR04, Men17, Rad14]. story [BH11]. Storytelling [LB06, Kir13].

Stowaways [SC14]. Strange [Ank20, Lig20, Der20]. strategic [OB12, Sme13]. Strategies [DC02a, Ned03, Can20, Gau00, Ink16, Sch10b]. strategy [Don14, MN13, MN19, Nic09]. stratification [SM06]. streptomycin [Gil17]. structural [YH17]. structure [ADF17, Bor04b, Hoo12, Men12, Sap07, SL14, Too03]. structures [Cha07, Fre11, Gre20b]. struggle [Bil17, Dar59, Dar90, Tay15, vH00]. strychnine [Sim99]. Stuart [PD08]. student [Kel11b]. Studie [vH14]. Studies [Wil03, Abr12, Bau08, Bor05, BM15, Cai02, Chi09, Juz09, Lem17, Lip14, Per09, Ram14, RH98, Ano12e]. Study [Cla07b, MP11, Bea04, Blo16, Can20, Cle07, DL12, De 08b, Far06, GS11a, HA09, Ken20, LB18, Löw03, MSM16, PM02, Pre10, Pri17, Sec12, SGK04, Str02, Suá01, Tho06, Tho16, Tul11, Ver04, Ver05, dC98]. Studying [Gro11, RH98]. stumbling [TB16]. stupid [Sha07]. style [Chi09, San10, San14a]. styles [Abr12, Ams04, Gau04, VR20]. Subject [Car99, JA08, MHH10]. subjects [Ger08]. substances [Her11]. subversion [Bar13b, MW05]. Subversive [Tur20]. succeed [San13]. success [Esp17, Fag07]. Suh [DiM18]. suicide [Mau20]. suivie [R81]. Suman [Bra20, Góm20, Sie20]. superior [Hoc15]. supplements [AG14]. support [LB06]. Surfaces [Gro10]. surgeon [Mai14]. Surgery [Buk07]. surroundings [Sch03c]. surrounds [Lan16]. Surveillance [Whi14]. Survey [Kle18]. Surveying [Ram14]. surveys [dC14]. Survival [Bou11]. sustainability [TB16]. Svensson [Kap13]. Sweden [Hac10]. Swedish [Nor07]. Swiss [Mot08]. sword [Cun02, Cun03]. Sydney [dC09a]. symbiosis [Sap94, Sar98]. symbolically [Bea03]. symbols [Ser11]. syndrome [KW08, San14b]. syndromes [Ble10, CC10, Coo10]. Synthesis [Chi19, LB18, Bus20, DC02a, Del09, Del11b, Dre19, Esp11, Hod11, IW10, Ish17, Orl12, Rad06, Smo14, Web02, Web11]. Synthesising [Bus20]. synthesized [Del11b, Gau05a]. Synthetic [BP13a, Gel13, Güt13, KR13, Lew13, Mac13, OP13, Pre13, BS13, Cal12b, Fri13, GKG13, Gra13, HP13, Kas13, Liu17, Sch13b, Vin13, KL13]. Syphilis [Lin01, Lin02, Löw04a, vdB02]. Syros [Qui04]. system [Gre20b, PM02, WEHT20]. systematic [Hag01, SL14]. Systematics [Rap04, Hul01, Cai02]. Systems [Bri13b, Cal12b, CF11, Car14, Col15, Don14, GS11b, Gre13, Kas13, MN13, MN15, MN19, OS12a, PD09, Toe12, TS20, Web11]. systems-level [MN15].

T [Cob15]. Tabery [Tay15]. Table [MT17, Sep14]. tablets [Löw12]. tackle [Ped15]. takes [Sch10b]. Tales [Dav11]. talk [Lew00]. Talking [Aga01, AR98, Ist13]. Taming [Yi18]. tango [PM11]. Tape [Sep16]. targets [Hor11]. taste [Pan12]. taxa [Ped12]. Taxidermy [Gra98]. taxon [Gri13]. taxonomic [Ver01]. taxonomy [Dro01, Ere01, Mor16a, MW05, Ver01]. teach [Bon13, Ste05a]. teaching [Don15, Ols11]. technical [Iid20]. technicity [Mac13]. techniques [Ari07, Suá01, Web98, Wid14]. techno [CK00, Kas13]. techno-visual [CK00]. Technological [Rus17, De 19].

Technologies [Ham13, Par04, Bor16, Ger04d]. technology [Ano09f, Gau01, Gra07, Wil07a]. technoscience [Gel13]. Tegetmeier [Dav04]. TEK [VR20]. teleological [McG16]. Teleology [Rus00, Toe12, Zam06, BHZ20, Dep08, Hun19, MN12, Qua06, Rus16, Slo12, Ste06a, Teu11, Tho17, Wal12, vdB13]. Television [Jer13, Kir13]. Telling [Men17]. Tempo [Col15]. Temporal [Don14, WEHT20]. temporary [Wit15, Wit16]. temptations [Rus02]. tension [Gre15b]. terms [Mau16b]. terran [Cle07]. test [She10, WEHT20]. Teste [Smo10]. Testimonies [Pri14]. testimony [Ham13]. Testing [Jef08, DL12, Gra04, OG14, San14b, Web02]. textbooks [Ser16]. texts [Hed07b]. TGN1412 [Lem17]. thalidomide [Gil17]. theater [Van00]. their [FK07, Kov17, Mat16, Slo14b, Spe11, Teu11, Top07, vL37]. themes [Tau06]. Theodore [Tor13]. Theodosius [Dep11, Gan13]. Theological [Meg16]. Theology [Sha14, Dil13, Gol13]. theorem [Ber20, Plu06]. Theoretical [Don16, Sag17, Tre20a, Bro14, Don17, Mac07, RH17]. Theories [Hou05, Bru03, Cle12, Fic14, Hul01, MS14, San12]. Theorizing [Kit20]. Theory [Pâs14, Web02, Abr06, ADF17, Ang02, Ber14b, Bor04b, Car10a, Dag17, FB15, Gue10, Har15, Hod11, Hol17, Ish17, IR20, Liv04, Lyo07, Mar19a, MW07a, MW10, Ned03, Nic10, Sap07, Sar14a, Sch10b, Sha08, Sha16, Ste05a, Swe15, The14, Too03, Tre20a, WEHT20, WCB17]. theory-generative [WEHT20]. therapeutic [Car10b, Car12b, Gil13]. therapeutics [Fin08]. therapies [Stu12]. therapy [Hen11]. There [Spe13, GS20, Har99a, MH14, Sti16, Sul20, Wor07]. these [vH14]. Theses [Cla12]. thesis [Ros14, Zam12, FD15]. thick [Low18]. thin [Low18]. thing [Gel13]. things [Har16, Mac02]. think [Bor16]. Thinking [Hal16, Mil15, SM05, Ger04b, Hon15, LA12, Sch15, VR20]. third [Rus14]. Thomas [Coo12, Sto04]. Thompson [Hol19]. though [KA03]. Thought [Ger04b, Alt08b, Ams04, Edd06, Gan03, Gau04, LB06, ST17]. thoughts [Noa14, Pig13]. threat [Ano09f]. Three [Can20, Leo13, WGEN15, Bor05, Cla07b, Fal13, Pea11, ZDSF11]. three-man [Fal13, ZDSF11]. Tibetan [YT10]. ties [Tul07]. Time [BM18, Sch03c, Bae13, Buk14, Drö11, Ers18, Gis17, Mil15, SS15b]. timely [Vae14]. times [Gun12, Jef08, Lan02]. tissue [GS11a]. today [Rus00]. token [McL13]. Tokugawa [Cra18]. Told [Cre99]. tolerance [Vaz16]. tomorrow [Rus00]. Too [KC12]. tool [Bru08, RP16]. Tools [Ing99, Gre13, Hou05, Kap13, Sam13]. top [Smi20]. top-down [Smi20]. Touching [Kno15]. tourism [Tau06]. Town [De 16]. Toxicity [Vai15, Sim99]. tracers [Cre09, Mye16]. Tracing [Cre02, Lee10]. Trackable [Ben16]. Trade [Ink16]. Trade-offs [Ink16]. trading [RPLBMA20]. tradition [Rad08b]. traditional [Car12b, For16, Gil13, RPLBMA20]. traditions [Bre02, Gra17]. tragic [Lem17]. training [SS11]. Trait [Sob13, McL20]. traits [McL13]. trajectories [Gau05b, Lav05, Won20]. trans [Sta13]. transatlantic [Coo12, FM07, Gau02b]. Transcending [Abr12]. Transfer [Gis20, Bab20]. transformation [Cre08, Gau05a, HS17, Woo07]. transformations [GS11a]. Transforming [AGD14]. transition [Bea03, Cai02]. Transitions [Gra17, FB15]. translation [Chi09, MWR07]. Translational [KCN16]. translations [Bas12]. transmission [Tul11]. transmutation [Rou19]. transplantation [Mai14]. transport [San00]. Transylvania [Sec12]. trap [Mil16]. travel [SS15a]. travels [Fic14]. Traverso [CC02]. Treasuring [Hog16]. treat [She10]. treatises [Tot07]. Treatment [Pan19, Dre98, Gra14, Kei03, PM11]. tree [O’M13]. Trees [Pie12, Par14, Tor13]. trends [Des18, San02]. très [R81]. trial [BV16]. trials [GMST15, Tei13, Tei19]. tribunal [Chi03]. Triple [KCN16, Ari03]. trivial [Won20]. Tropical [Wil00, Gra10, Rei04]. Trouble [Par03, BP13a]. troubles [Noa14]. true [Car12a, Mil06]. truly [Ver01]. trust [Ash03]. truth [Edm13, Kel17, Lyn13]. tuberculin [Gra04]. tubes [Bra11a]. Turkish [Bil17]. Turks [Bil17]. turn [Bea04]. Turning [The00]. twelfth [Bur10]. twentieth [Abr12, All05, Car12a, Gan03, GSM19, LH10, Lee15, Mét16, MW07b, Mye16, Par03, PS11, Sle09, Tho16, Wil07a, Wil07b, Yu17, Rus17]. twentieth-century [All05, Car12a, Gan03, LH10, Lee15, MW07b, Mye16, Tho16, Wil07b, Rus17]. twenty [Pea12]. twenty-first [Pea12]. twins [Dre98]. Two [Bog01, BM15, Gar16, Kas13, SLR12, De 19, mKC10, Cla12, Mau16b, MZ00, Qua06, Sch10b, Ban14]. two-dimensional [Mau16b]. two-level [Qua06]. type [Car10b, Hol17]. types [Cla12, Gar16, Hon15, SLR12]. typicality [SLR12]. typological [Chu03, Hon15]. typological/population [Chu03]. typology [Rie13, Tre20b, Wit15, Wit16]. typology/population [Wit15, Wit16].

U.S. [Mat17]. ubiquitous [O’M08]. UK [Cla07a, WL08]. Ulcers [Tha98a, Tha98b]. ultimately [Whi04]. unanalysed [Win01]. unappreciated [Tam20]. unborn [Löw14a]. unburied [Fab10, Smi13]. uncertain [Del09]. uncertainty [Hog16, LR06, Ste11a]. Uncommon [Lav05]. uncultivated [Rou19]. undeath [Ste08]. Underdetermination [AR20, Mag05, Tul07]. underlying [Tay09]. understand [Tay15]. Understanding [MN15, PKD16, Bou11, De 08b, Dus18]. Understandings [Ell14]. unexpected [Tho16]. unexplained [Gre20a]. unfinished [O’M10]. Unfolding [Rad14]. Unformed [Har13b]. unificationism [DiF19]. unified [Ram16b]. Union [PT13]. unit [GG06, Ist13]. United [Bra07, Hag01, Juz09, Kei03, Soc20]. unity [Cra05]. Universal [Ful18, Bon19]. Universality [BM15]. universally [Coo10]. universally-occurring [Coo10]. universe [LoT03]. universities [Kel11b, Ols11]. University [Ash03, Ayl18, Bra20, Cob15, Dea10, DiM18, Gur98, Oka03, Ren15, Sar98, Tor13, Wil98, WL08, GS11a, Hay10]. unnatural [Spe14]. Unnaturalised [Hoc14]. Unraveling [Lee15]. unreason [Tre04]. unscientific [Res03]. unstable [Bud18]. Until [Smi13]. US$35.00 [Dea10]. US$40.00 [Tay15]. USA [Cla07b]. usage [Mat16]. use [Edm13]. useful [Pea07]. uses [Azz10, Dro01, Gra07, Sha14, Sko11]. Using [Ped15, Han06]. utility [CC16]. Utopianism [Esp11].

v [Oka03, Bou10, LaB14]. vaccination [FK07]. Vaccine [BP19, Bon05, Cas05, HS17]. vaccines [BP19, Gau00, HS17]. vagaries [End09]. valence [Haw10]. Validation [Car14]. validity [Mic08, She10]. value [Ers18, Gar14b, Kid13]. values [Bar09, De 16, Haw10, Ken20, Sha04]. vampire [Brä12]. vapours [Ver04]. variability [Car14]. variables [Bau08]. variation [Ams04, GL01, Lip14, Pod17, Smo15, Sob13, Val12]. varied [GL01]. variety [Smi13]. Various [Lip14]. vat [Cav04, Ger04a, GG04, Ger04b, Sma04, SM04]. vats [Ger04c, Man04]. vegetables [Tot07]. verae [Cla16]. vernacular [Sec12]. versatility [OG14]. versus [Far06, Krá12, Roc11, Rus14, Tay15, dC09b]. vertical [Bab20]. vervet [Rad06]. ves [Bir00]. Vesalian [Maz16]. veterinary [Gra10, Woo07]. Victorian [Dea10, Ren15, Del14, Ell08, End08, Fin08, GCT17, Hal14a, Hes15, KA03, Lig16, Liv04, Pet02, Rad00]. victory [Tam17]. Vienna [Ata12, Buk07, Hay10, Mae12]. View [Sie20, Ågr18, Bir12, Ger13, Mac15, Mai14, Mor12, Sep16, Ste05b, Swi17, The14, Whi04, RB13]. viewing [WEHT20]. views [Chu03, Hun13]. Vikings [Hoc15]. vindication [Dei16]. viral [Mor16a]. virology [Mét16]. virtue [Smi12]. virtues [Gla08]. virtuous [Kid14]. Virus [Sch14b, Mor14, Mor16a, O’M16, SC14, vH14]. Viruses [DG16, CA16, Cre14, For16, Gau00, KS16, Kos16, Pra16, PKD16, San14a, Sch14a, vR16]. visibilities [Fio11]. visible [Löw14b, Van00]. vision [Bro11b, Cow09, Woo15]. visions [Wil00]. visiting [Car04]. visual [Aic14, CK00, GvA19, Pie12, Sam13, Tor13]. Visualizing [Tor13]. Vital [Gam14]. vitalism [All05, Gis14]. vitality [Gis14]. vitamin [AG14]. vitamins [Smi09]. vitro [Lan02]. vivant [Gau02a, Qui04]. Vivisection [GCT17]. vivisectionism [FS15]. vivo [Tee19]. volant [R81, R77]. Volkskrankheiten [BN00]. Vygotsky [Hym12].

W [Tor13]. W. [Can17, Mur01, Pen11, Swe15]. Wa [Bir00]. Waddington [Bae13]. wake [Hed07a, Kra14, Gri13]. wall [MH14]. Wallace [FK07, Fic14, Mit14, vWK15, vW16]. Walton [Car04]. wandering [Gra98]. Wanted [Kir08]. wanting [McS13]. War [Ban14, Cre99, Hey15, Mil20, Smi09, vH00, Bar20, Gau02b, Gun12, Juz09, Law12, Löw05, Str02, Swe15, Lin14, Mun15]. Warburg [Nic09]. Wars [MZ00]. Wartime [Dav14a]. wary [FD19]. Was [Kid14, Wor07, Del11b, GS20, Har16, McL11, Plu06, Ric12, Slo12, vW13]. Washburnian [Rad06]. wasps [Can17, Par14]. Waste [Ers18]. water [Edd08]. Waterton [Car04, Gra98]. way [Fis15, Mac02, NM17, Roc12, WL08]. ways [PD09]. weakened [Meb14]. Weather [Gre01]. weird [Bry08]. Weismann [Smi10]. Weldon [Pen11]. welfare [Gra12, Nor07]. well [vR16]. Wells [Bar06]. we’re [Joh12, Zaw15, San14a]. wet [Gol96, Gur98]. wheat [Bar15]. Where [Dyd16, Mox18]. whether [Fis15]. Whetting [Olb00]. Whewell [Qui16, Qui17]. Who [Lil04b, Dea10, Rad14]. whole [Mun15, Pen11, vWK15]. Wholes [Teu11]. whom [Pan19]. Whose [Mac15, Rad08a]. wicked [McL01]. Widdowson [Buk14]. Wide [Bar15]. wild [BPS20, Hol17, Lav05, Plu07]. wildlife [Wex08]. will [Mit14, Mor11b]. William [Cob15, Edd08, Hal18, Kid14, Del14, Gol13, Hal11, Hal14b, Liv04, Qui16, Sha14]. Williamson [Cla12, FD15]. Wilson [Har13a, Oka03, Low16, Meg16]. wind [YT10]. wine [Gau09]. Winwood [Hes15]. wish [Mox18]. wish-list [Mox18]. witch [McL01]. withdrawal [Hac10]. within [Bar13b, Jam10, Pan19, Par03, Shi06, Spe11]. without [BHZ20, Bou11, Cle08, Gro11, Hut12, MN15, McG16, RH98, RBOH20, Tei19]. witness [Ham13]. Witte [Löw05]. Wittgenstein [Smi10]. Wolffian [vdB13]. Women [Tho01, Dav14a, Pet02, San14b]. wonder [LH10, Str12]. wonders [Bry08, Law14]. Worboys [Ank20, Der20, Lig20]. words [Dim17]. Work [Det99, Can17, Gau04, Gue20, Pol07, SD14, Aga01]. works [Dro01, San00]. world [Can17, Ger04c, Mar98, Pet02, Whi04, BD19, Ban14, Cre99, Mil20, MZ00, Smi09, vH00]. worlds [Aic16]. worm [GS12]. worms [dC98]. worth [Ers18]. would [Dea10]. Wright [Hod11, Ish17, IR20]. Writing [Mét16, Bor16]. writings [Hed07b]. written [Gre17]. wrong [Lew04b]. wronging [MHH10]. Wu [Ona14]. Wundt [Mül12].

X [Gur98, Mur01, Wil98]. xi [Aga01, Gig14]. xiii [Tay15]. xvii [Sar98]. xxii [Whi04].

yam [Lav05]. years [Mac07, vH14]. Yerkes [Tho06]. Yes [Ste08]. yesterday [Rus00]. yield [Pei15]. York [Har13a, Sar98, Wil98, Gau02b].

Zebrafish [Meu12]. zebras [Sha08]. Zero [Luq16]. Zero-Cause [Luq16]. Zondek [OG14]. zone [RPLBMA20]. zoology [Bla04]. Zou [LH10]. Zuckerman [Bur06].

References

Artiga:2020:MBS

[ABM20] Marc Artiga, Jonathan Birch, and Manolo Martínez. The meaning of biological signals. Studies in History and Philosophy of Biological and Biomedical Sciences, 84(??):Article 101348, December 2020. CODEN ???? ISSN 1369-8486 (print), 1879-2499

            (electronic). URL http://www.sciencedirect.com/science/
            article/pii/S136984862030162X.

Abrams:2006:IPC

[Abr06] Marshall Abrams. Infinite populations and counterfactual frequencies in evolutionary theory. Studies in History and Philosophy of Biological and Biomedical Sciences, 37(2):256–268, June 2006. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848606000161.

Abraham:2012:TDS

[Abr12] Tara H. Abraham. Transcending disciplines: Scientific styles in studies of the brain in mid-twentieth century America. Studies in History and Philosophy of Biological and Biomedical Sciences, 43 (2):552–568, June 2012. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848612000258.

Abrams:2013:PPP

[Abr13] Marshall Abrams. Populations and pigeons: Prosaic pluralism about evolutionary causes. Studies in History and Philosophy of Biological and Biomedical Sciences, 44(3):294–301, September 2013. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848613000939.

Achim:2011:RSI

[Ach11] Miruna Achim. From rustics to savants: Indigenous materia medica in eighteenth-century Mexico. Studies in History and Philosophy of Biological and Biomedical Sciences, 42(3):275– 284, September 2011. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848611000215.

Adam:2005:IRD

[Ada05]  Matthias Adam.  Integrating research and development:

the emergence of rational drug design in the pharmaceutical industry. Studies in History and Philosophy of Biological and Biomedical Sciences, 36(3):513–537, September 2005. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848605000488.

Alleva:2017:MTS

[ADF17] Karina Alleva, José Díez, and Lucia Federico. Models, theory structure and mechanisms in biochemistry: the case of allosterism. Studies in History and Philosophy of Biological and Biomedical Sciences, 63(??):1–14, June 2017. CODEN ???? ISSN 1369-8486 (print), 1879-2499 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S1369848617300328.

Al-Gailani:2014:MBD

[AG14] Salim Al-Gailani. Making birth defects ‘preventable’: Preconceptional vitamin supplements and the politics of risk reduction. Studies in History and Philosophy of Biological and Biomedical Sciences, 47 (part B)(??):278–289, September 2014. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848613001349.

Aizawa:2019:DPA

[AG19] Kenneth Aizawa and Carl Gillett. Defending pluralism about compositional explanations. Studies in History and Philosophy of Biological and Biomedical Sciences, 78(??):Article 101202, December 2019. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S136984861830116X.

Agar:2001:BRC

[Aga01] Jon Agar. Book review: Community (Net) Work: James A. Anderson and Edward Rosenfeld (eds), Talking Nets: An Oral History of Neural Networks (Cambridge, MA, and London: MIT Press, 1998), xi + 500 pp., ISBN 0-262-01167-0. Hardback £31.95. Studies in History and Philosophy of Biological and Biomedical Sciences, 32(3):557–564, September 2001. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848601000188. See [AR98].

Al-Gailani:2014:ITP

[AGD14] Salim Al-Gailani and Angela Davis. Introduction to “Transforming pregnancy since 1900”. Studies in History and Philosophy of Biological and Biomedical Sciences, 47 (part B)(??):229– 232, September 2014. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848614000818.

Aagren:2018:HVS

[Ågr18] J. Arvid Ågren. The Hamiltonian view of social evolution. Studies in History and Philosophy of Biological and Biomedical Sciences, 68–69(??):83–87, April/June 2018. CO-DEN ???? ISSN 1369-8486 (print), 1879-2499 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S1369848618300633.

Aicardi:2016:TFA

[AGS16] Christine Aicardi and Miguel García-Sancho. Towards future archives and historiographies of ‘big biology’. Studies in History and Philosophy of Biological and Biomedical Sciences, 55(??): 41–44, February 2016. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848615001375.

Anstey:2006:LB

[AH06a] Peter R. Anstey and Stephen A. Harris. Locke and botany. Studies in History and Philosophy of Biological and Biomedical Sciences, 37(2):151–171, June 2006. CODEN ???? ISSN 1369-8486 (print), 1879-2499 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S1369848606000124.

Ashcroft:2006:GDP

[AH06b] Richard E. Ashcroft and Adam M. Hedgecoe. Genetic databases and pharmacogenetics: introduction. Studies in History and Philosophy of Biological and Biomedical Sciences, 37(3):499– 502, September 2006. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S136984860600046X.

Allen-Hermanson:2017:KCE

[AH17] Sean Allen-Hermanson. Kamikazes and cultural evolution. Studies in History and Philosophy of Biological and Biomedical Sciences, 61(??):11–19, February 2017. CODEN ???? ISSN 1369-8486 (print), 1879-2499 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S1369848616301030.

Aicardi:2014:HCS

[Aic14] Christine Aicardi. Of the Helmholtz Club, South-Californian seedbed for visual and cognitive neuroscience, and its patron Francis Crick. Studies in History and Philosophy of Biological and Biomedical Sciences, 45(??):1–11, March 2014. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S136984861300174X.

Aicardi:2016:FCC

[Aic16] Christine Aicardi. Francis Crick, cross-worlds influencer: a narrative model to historicize big bioscience. Studies in History and Philosophy of Biological and Biomedical Sciences, 55(??):83–95, February 2016. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848615001119.

Albala:2012:FHC

[Alb12] Ken Albala. Food for healing: Convalescent cookery in the early modern era. Studies in History and Philosophy of Biological and Biomedical Sciences, 43(2):323–328, June 2012. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848611001166.

Allen:2001:CBC

[All01] D. E. Allen. Controlling the brambles: changing approaches to classifying a reproductively abnormal group. Studies in History and Philosophy of Biological and Biomedical Sciences, 32(2):277– 290, June 2001. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848601000097.

Allen:2005:MVO

[All05] Garland E. Allen. Mechanism, vitalism and organicism in late nineteenth and twentieth-century biology: the importance of historical context. Studies in History and Philosophy of Biological and Biomedical Sciences, 36(2):261–283, June 2005. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848605000191.

Allesch:2012:HDP

[All12] Christian G. Allesch. Hans Driesch and the problems of “normal psychology”. Rereading his Crisis in Psychology (1925). Studies in History and Philosophy of Biological and Biomedical Sciences, 43(2):455–461, June 2012. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848611001300.

Alter:2007:DLC

[Alt07] Stephen G. Alter. Darwin and the linguists: the coevolution of mind and language, Part 1. Problematic friends. Studies in History and Philosophy of Biological and Biomedical Sciences, 38(3): 573–584, September 2007. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848607000404.

Alter:2008:CPA

[Alt08a] Stephen G. Alter. “curiously parallel”: Analogies of language and race in Darwin’s Descent of man. A reply to Gregory Radick. Studies in History and Philosophy of Biological and Biomedical Sciences, 39(3):355–358, September 2008. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848608000447. See [Rad08b].

Alter:2008:DLC

[Alt08b] Stephen G. Alter. Darwin and the linguists: the coevolution of mind and language, Part 2. The language–thought relationship. Studies in History and Philosophy of Biological and Biomedical Sciences, 39(1):38–50, March 2008. CODEN ???? ISSN 1369-8486 (print), 1879-2499 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S136984860700088X.

Ambrosio:2015:PKS

[Amb15] Chiara Ambrosio. Picturing knowledge in the Sixteenth Century. Studies in History and Philosophy of Biological and Biomedical Sciences, 50(??):83–86, April 2015. CO-DEN ???? ISSN 1369-8486 (print), 1879-2499 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S1369848615000126.

Amsterdamska:2004:ADT

[Ams04] Olga Amsterdamska. Achieving disbelief: thought styles, microbial variation, and American and British epidemiology, 1900–1940. Studies in History and Philosophy of Biological and Biomedical Sciences, 35(3):483–507, September 2004. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848604000457.

Amundson:2000:ANF

[Amu00] Ron Amundson. Against normal function. Studies in History and Philosophy of Biological and Biomedical Sciences, 31(1):33–53, March 2000. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848699000333.

Amundson:2005:DE

[Amu05] Ron Amundson. Darwins for everyone. Studies in History and Philosophy of Biological and Biomedical Sciences, 36(1):209–220, March 2005. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848604000913.

Angner:2002:HHT

[Ang02] Erik Angner. The history of Hayek’s theory of cultural evolution. Studies in History and Philosophy of Biological and Biomedical Sciences, 33(4):695–718, December 2002. CODEN ???? ISSN 1369-8486 (print), 1879-2499 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S1369848602000249.

Ankeny:2020:IMD

[Ank20] Rachel A. Ankeny. The invention of the modern dog by Michael Worboys, Julie-Marie Strange, and Neil Pemberton: Learning from the concept of ‘breed’. Studies in History and Philosophy of Biological and Biomedical Sciences, 84(??):Article 101330, December 2020. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848620301412. See [Der20, WSP20].

Anonymous:1998:BRa

[Ano98a] Anonymous. Book review. Studies in History and Philosophy of Biological and Biomedical Sciences, 29(2):343–347, September 1998. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S136984869800020X.

Anonymous:1998:BRb

[Ano98b] Anonymous. Book review. Studies in History and Philosophy of Biological and Biomedical Sciences, 29(2):349–357, September 1998. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848698000211.

Anonymous:1998:BHP

[Ano98c] Anonymous. Books on history and philosophy of biological and biomedical science. Studies in History and Philosophy of Biological and Biomedical Sciences, 29(1):??, March 1998. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848698800113.

Anonymous:1998:EB

[Ano98d] Anonymous. Editorial Board. Studies in History and Philosophy of Biological and Biomedical Sciences, 29(1):??, March 1998. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848698800009.

Anonymous:1998:I

[Ano98e] Anonymous. Index. Studies in History and Philosophy of Biological and Biomedical Sciences, 29(2):??, September 1998. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848698000260.

Anonymous:1999:BRa

[Ano99a] Anonymous. Book review. Studies in History and Philosophy of Biological and Biomedical Sciences, 30(1):129–142, March 1999. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848698000326.

Anonymous:1999:BRb

[Ano99b] Anonymous. Book review. Studies in History and Philosophy of Biological and Biomedical Sciences, 30(2):237–254, June 1999. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S136984869800034X.

Anonymous:1999:BRc

[Ano99c] Anonymous. Book review. Studies in History and Philosophy of Biological and Biomedical Sciences, 30(2):255–261, June 1999. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848698000193.

Anonymous:1999:BRd

[Ano99d] Anonymous. Book review. Studies in History and Philosophy of Biological and Biomedical Sciences, 30(2):263–271, June 1999. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848698000181.

Anonymous:1999:BRe

[Ano99e] Anonymous. Book review. Studies in History and Philosophy of Biological and Biomedical Sciences, 30(4):527–535, December 1999. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848699000266.

Anonymous:1999:BRf

[Ano99f] Anonymous. Book review. Studies in History and Philosophy of Biological and Biomedical Sciences, 30(4):537–544, December 1999. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848698000338.

Anonymous:1999:BRg

[Ano99g] Anonymous. Book review. Studies in History and Philosophy of Biological and Biomedical Sciences, 30(4):545–552, December 1999. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848699000230.

Anonymous:1999:E

[Ano99h] Anonymous. Editorial. Studies in History and Philosophy of Biological and Biomedical Sciences, 30(3):315–318, September 1999. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S136984869900014X.

Anonymous:1999:I

[Ano99i] Anonymous. Index. Studies in History and Philosophy of Biological and Biomedical Sciences, 30(4):vii, December 1999. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848699000369.

Anonymous:2000:BRa

[Ano00a] Anonymous. Book review. Studies in History and Philosophy of Biological and Biomedical Sciences, 31(2):355–363, June 2000. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848699000254.

Anonymous:2000:BRb

[Ano00b] Anonymous. Book review. Studies in History and Philosophy of Biological and Biomedical Sciences, 31(4):599–614, December 2000. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848699000321.

Anonymous:2000:BRc

[Ano00c] Anonymous. Book review. Studies in History and Philosophy of Biological and Biomedical Sciences, 31(4):615–627, December 2000. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848699000357.

Anonymous:2000:BRd

[Ano00d] Anonymous. Book review. Studies in History and Philosophy of Biological and Biomedical Sciences, 31(4):659–668, December 2000. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S136984869900031X.

Anonymous:2000:Da

[Ano00e] Anonymous. Discussion. Studies in History and Philosophy of Biological and Biomedical Sciences, 31(1):1–10, March 2000. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848699000448.

Anonymous:2000:Db

[Ano00f] Anonymous. Discussion. Studies in History and Philosophy of Biological and Biomedical Sciences, 31(1):77–94, March 2000. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S136984869900028X.

Anonymous:2000:Dc

[Ano00g] Anonymous. Discussion. Studies in History and Philosophy of Biological and Biomedical Sciences, 31(1):179–191, March 2000. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848699000345.

Anonymous:2000:I

[Ano00h] Anonymous. Index. Studies in History and Philosophy of Biological and Biomedical Sciences, 31(4):??, December 2000. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848600000224.

Anonymous:2001:CAI

[Ano01a] Anonymous. 2001 contents and authors index. Studies in History and Philosophy of Biological and Biomedical Sciences, 32 (4):??, December 2001. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848601000334.

Anonymous:2001:BRa

[Ano01b] Anonymous. Book review. Studies in History and Philosophy of Biological and Biomedical Sciences, 32(1):169–177, January 2001. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848699000242.

Anonymous:2001:BRb

[Ano01c] Anonymous. Book review. Studies in History and Philosophy of Biological and Biomedical Sciences, 32(1):179–190, January 2001. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848699000278.

Anonymous:2001:BHP

[Ano01d] Anonymous. Books on history and philosophy of biological and biomedical sciences received. Studies in History and Philosophy of Biological and Biomedical Sciences, 32(4):753–756, December 2001. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848601000085.

Anonymous:2002:CAI

[Ano02a] Anonymous. 2002 calender and author index. Studies in History and Philosophy of Biological and Biomedical Sciences, 33 (4):??, December 2002. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848602000572.

Anonymous:2002:BHP

[Ano02b] Anonymous. Book in the history and philosophy of biological and biomedical sciences. Studies in History and Philosophy of Biological and Biomedical Sciences, 33(3):565–566, September 2002. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848602000171.

Anonymous:2002:EBa

[Ano02c] Anonymous. Editorial Board. Studies in History and Philosophy of Biological and Biomedical Sciences, 33(3):??, September 2002. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848602000298.

Anonymous:2002:EBb

[Ano02d] Anonymous. Editorial Board. Studies in History and Philosophy of Biological and Biomedical Sciences, 33(4):??, December 2002. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S136984860200050X.

Anonymous:2003:A

[Ano03a] Anonymous. ANNOUNCEMENT TO:34-2. Studies in History and Philosophy of Biological and Biomedical Sciences, 34(2):381– 382, June 2003. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848603000281.

Anonymous:2003:EBa

[Ano03b] Anonymous. Editorial Board. Studies in History and Philosophy of Biological and Biomedical Sciences, 34(1):??, March 2003. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848603000037.

Anonymous:2003:EBb

[Ano03c] Anonymous. Editorial board. Studies in History and Philosophy of Biological and Biomedical Sciences, 34(2):??, June 2003. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S136984860300030X.

Anonymous:2003:EBc

[Ano03d] Anonymous. EDITORIAL BOARD. Studies in History and Philosophy of Biological and Biomedical Sciences, 34(3):??, September 2003. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848603000554.

Anonymous:2003:EBd

[Ano03e] Anonymous. Editorial Board. Studies in History and Philosophy of Biological and Biomedical Sciences, 34(4):??, December 2003. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S136984860300092X.

Anonymous:2003:I

[Ano03f] Anonymous. Index. Studies in History and Philosophy of Biological and Biomedical Sciences, 34(4):??, December 2003. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848603001006.

Anonymous:2004:CAI

[Ano04a] Anonymous. 2004 contents and author index. Studies in History and Philosophy of Biological and Biomedical Sciences, 35 (4):??, December 2004. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S136984860400086X.

Anonymous:2004:A

[Ano04b] Anonymous. Announcement. Studies in History and Philosophy of Biological and Biomedical Sciences, 35(1):217, March 2004. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848603001055.

Anonymous:2004:C

[Ano04c] Anonymous. Contents. Studies in History and Philosophy of Biological and Biomedical Sciences, 35(1):119, March 2004. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848603001109.

Anonymous:2004:EBa

[Ano04d] Anonymous. Editorial Board. Studies in History and Philosophy of Biological and Biomedical Sciences, 35(1):??, March 2004. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848604000032.

Anonymous:2004:EBb

[Ano04e] Anonymous. Editorial Board. Studies in History and Philosophy of Biological and Biomedical Sciences, 35(2):??, June 2004. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848604000305.

Anonymous:2004:EBc

[Ano04f] Anonymous. Editorial Board. Studies in History and Philosophy of Biological and Biomedical Sciences, 35(3):??, September 2004. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848604000548.

Anonymous:2004:EBd

[Ano04g] Anonymous. Editorial Board. Studies in History and Philosophy of Biological and Biomedical Sciences, 35(4):??, December 2004. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848604000780.

Anonymous:2005:CAI

[Ano05a] Anonymous. 2005 contents and author index. Studies in History and Philosophy of Biological and Biomedical Sciences, 36 (4):??, December 2005. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848605000919.

Anonymous:2005:EPIa

[Ano05b] Anonymous. Editorial and publication information. Studies in History and Philosophy of Biological and Biomedical Sciences, 36 (3):??, September 2005. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848605000567.

Anonymous:2005:EPIb

[Ano05c] Anonymous. Editorial and publication information. Studies in History and Philosophy of Biological and Biomedical Sciences, 36 (4):??, December 2005. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848605000798.

Anonymous:2005:EBa

[Ano05d] Anonymous. Editorial board. Studies in History and Philosophy of Biological and Biomedical Sciences, 36(1):??, March 2005. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848605000038.

Anonymous:2005:EBb

[Ano05e] Anonymous. Editorial board. Studies in History and Philosophy of Biological and Biomedical Sciences, 36(2):??, June 2005. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848605000397.

Anonymous:2006:EPIa

[Ano06a] Anonymous. Editorial and publication information. Studies in History and Philosophy of Biological and Biomedical Sciences, 37(1):??, March 2006. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848606000033.

Anonymous:2006:EPIb

[Ano06b] Anonymous. Editorial and publication information. Studies in History and Philosophy of Biological and Biomedical Sciences, 37(2):??, June 2006. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848606000240.

Anonymous:2006:EPIc

[Ano06c] Anonymous. Editorial and publication information. Studies in History and Philosophy of Biological and Biomedical Sciences, 37 (3):??, September 2006. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848606000574.

Anonymous:2006:EPId

[Ano06d] Anonymous. Editorial and publication information. Studies in History and Philosophy of Biological and Biomedical Sciences, 37 (4):??, December 2006. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848606000781.

Anonymous:2007:BRO

[Ano07a] Anonymous. Books received to October 2006. Studies in History and Philosophy of Biological and Biomedical Sciences, 38 (1):297–301, March 2007. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848606001038.

Anonymous:2007:EPIa

[Ano07b] Anonymous. Editorial and publication information. Studies in History and Philosophy of Biological and Biomedical Sciences, 38(1):??, March 2007. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848607000039.

Anonymous:2007:EPIb

[Ano07c] Anonymous. Editorial and publication information. Studies in History and Philosophy of Biological and Biomedical Sciences, 38(2):??, June 2007. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848607000246.

Anonymous:2007:EPIc

[Ano07d] Anonymous. Editorial and publication information. Studies in History and Philosophy of Biological and Biomedical Sciences, 38 (3):??, September 2007. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848607000581.

Anonymous:2007:EPId

[Ano07e] Anonymous. Editorial and publication information. Studies in History and Philosophy of Biological and Biomedical Sciences, 38 (4):??, December 2007. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848607000696.

Anonymous:2008:EPIa

[Ano08a] Anonymous. Editorial and publication information. Studies in History and Philosophy of Biological and Biomedical Sciences, 39(1):??, March 2008. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848608000034.

Anonymous:2008:EPIb

[Ano08b] Anonymous. Editorial and publication information. Studies in History and Philosophy of Biological and Biomedical Sciences, 39(2):??, June 2008. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848608000277.

Anonymous:2008:EPIc

[Ano08c] Anonymous. Editorial and publication information. Studies in History and Philosophy of Biological and Biomedical Sciences, 39 (3):??, September 2008. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848608000630.

Anonymous:2008:EPId

[Ano08d] Anonymous. Editorial and publication information. Studies in History and Philosophy of Biological and Biomedical Sciences, 39 (4):??, December 2008. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848608000812.

Anonymous:2008:E

[Ano08e] Anonymous. Erratum. Studies in History and Philosophy of Biological and Biomedical Sciences, 39(1):51, March 2008. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848607001008.

Anonymous:2009:BRM

[Ano09a] Anonymous. Books received to March 2009. Studies in History and Philosophy of Biological and Biomedical Sciences, 40(3):239– 240, September 2009. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848609000399.

Anonymous:2009:EPIa

[Ano09b] Anonymous. Editorial and publication information. Studies in History and Philosophy of Biological and Biomedical Sciences, 40(1):??, March 2009. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S136984860900003X.

Anonymous:2009:EPIb

[Ano09c] Anonymous. Editorial and publication information. Studies in History and Philosophy of Biological and Biomedical Sciences, 40(2):??, June 2009. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848609000223.

Anonymous:2009:EPIc

[Ano09d] Anonymous. Editorial and publication information. Studies in History and Philosophy of Biological and Biomedical Sciences, 40 (3):??, September 2009. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848609000417.

Anonymous:2009:EPId

[Ano09e] Anonymous. Editorial and publication information. Studies in History and Philosophy of Biological and Biomedical Sciences, 40 (4):??, December 2009. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848609000600.

Anonymous:2009:JUT

[Ano09f] Anonymous. Journals under threat: a joint response from history of science, technology and medicine editors. Studies in History and Philosophy of Biological and Biomedical Sciences, 40(1):1–3, March 2009. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848608000885.

Anonymous:2010:EPIa

[Ano10a] Anonymous. Editorial and publication information. Studies in History and Philosophy of Biological and Biomedical Sciences, 41(1):??, March 2010. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848610000038.

Anonymous:2010:EPIb

[Ano10b] Anonymous. Editorial and publication information. Studies in History and Philosophy of Biological and Biomedical Sciences, 41(2):??, June 2010. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848610000221.

Anonymous:2010:EPIc

[Ano10c] Anonymous. Editorial and publication information. Studies in History and Philosophy of Biological and Biomedical Sciences, 41 (3):??, September 2010. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848610000506.

Anonymous:2010:EPId

[Ano10d] Anonymous. Editorial and publication information. Studies in History and Philosophy of Biological and Biomedical Sciences, 41 (4):??, December 2010. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848610000804.

Anonymous:2010:ICL

[Ano10e] Anonymous. Inside contents list. Studies in History and Philosophy of Biological and Biomedical Sciences, 41(2):??, June 2010. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848610000270.

Anonymous:2011:EPIa

[Ano11a] Anonymous. Editorial and publication information. Studies in History and Philosophy of Biological and Biomedical Sciences, 42(1):??, March 2011. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848611000033.

Anonymous:2011:EPIb

[Ano11b] Anonymous. Editorial and publication information. Studies in History and Philosophy of Biological and Biomedical Sciences, 42(2):??, June 2011. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848611000240.

Anonymous:2011:EPIc

[Ano11c] Anonymous. Editorial and publication information. Studies in History and Philosophy of Biological and Biomedical Sciences, 42 (3):??, September 2011. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848611000586.

Anonymous:2011:EPId

[Ano11d] Anonymous. Editorial and publication information. Studies in History and Philosophy of Biological and Biomedical Sciences, 42 (4):??, December 2011. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848611000975.

Anonymous:2012:EPIa

[Ano12a] Anonymous. Editorial and publication information. Studies in History and Philosophy of Biological and Biomedical Sciences, 43(1):??, March 2012. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848612000064.

Anonymous:2012:EPIb

[Ano12b] Anonymous. Editorial and publication information. Studies in History and Philosophy of Biological and Biomedical Sciences, 43(2):??, June 2012. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848612000350.

Anonymous:2012:EPIc

[Ano12c] Anonymous. Editorial and publication information. Studies in History and Philosophy of Biological and Biomedical Sciences, 43 (3):??, September 2012. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848612000854.

Anonymous:2012:EPId

[Ano12d] Anonymous. Editorial and publication information. Studies in History and Philosophy of Biological and Biomedical Sciences, 43 (4):??, December 2012. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848612001173.

Anonymous:2012:EES

[Ano12e] Anonymous. Editorial evolution at Studies C. Studies in History and Philosophy of Biological and Biomedical Sciences, 43 (3):iv, September 2012. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848612000945.

Anonymous:2013:DSM

[Ano13a] Anonymous. Diagnostic and statistical manual of mental disorders: DSM-5. American Psychiatric Association, Washington, DC, USA, fifth edition, 2013. ISBN 0-89042-554-X (hardcover), 0-89042-555-8 (paperback). xliv + 947 pp. LCCN RC455.2.C4 D54 2013.

Anonymous:2013:EPIe

[Ano13b] Anonymous. Editorial and publication information. Studies in History and Philosophy of Biological and Biomedical Sciences, 44(1):??, March 2013. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848613000071.

Anonymous:2013:EPIf

[Ano13c] Anonymous. Editorial and publication information. Studies in History and Philosophy of Biological and Biomedical Sciences, 44(2):??, June 2013. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848613000484.

Anonymous:2014:A

[Ano14a] Anonymous. Announcement. Studies in History and Philosophy of Biological and Biomedical Sciences, 47 (part A)(??): II, September 2014. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848614001186.

Anonymous:2014:EPIa

[Ano14b] Anonymous. Editorial and publication information. Studies in History and Philosophy of Biological and Biomedical Sciences, 46(??):IFC, June 2014. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848614000673.

Anonymous:2014:EPIb

[Ano14c] Anonymous. Editorial and publication information. Studies in History and Philosophy of Biological and Biomedical Sciences, 47 (part A)(??):ifc, September 2014. CO-DEN ???? ISSN 1369-8486 (print), 1879-2499 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S1369848614001101.

Anonymous:2014:EPIc

[Ano14d] Anonymous. Editorial and publication information. Studies in History and Philosophy of Biological and Biomedical Sciences, 47 (part A)(??):IFC, September 2014. CO-DEN ???? ISSN 1369-8486 (print), 1879-2499 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S1369848614001101.

Anonymous:2014:EPId

[Ano14e] Anonymous. Editorial and publication information. Studies in History and Philosophy of Biological and Biomedical Sciences, 47 (part B)(??):IFC, September 2014. CO-DEN ???? ISSN 1369-8486 (print), 1879-2499 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S1369848614001319.

Anonymous:2014:EPIe

[Ano14f] Anonymous. Editorial and publication information. Studies in History and Philosophy of Biological and Biomedical Sciences, 48 (part A)(??):IFC, December 2014. CO-DEN ???? ISSN 1369-8486 (print), 1879-2499 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S1369848614001447.

Anonymous:2014:EPIf

[Ano14g] Anonymous. Editorial and publication information. Studies in History and Philosophy of Biological and Biomedical Sciences, 48 (part B)(??):IFC, December 2014. CO-DEN ???? ISSN 1369-8486 (print), 1879-2499 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S1369848614001605.

Anonymous:2015:A

[Ano15a] Anonymous. Announcement. Studies in History and Philosophy of Biological and Biomedical Sciences, 52(??):I, August 2015. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848615000989.

Anonymous:2015:EPIa

[Ano15b] Anonymous. Editorial and publication information. Studies in History and Philosophy of Biological and Biomedical Sciences, 49 (??):IFC, February 2015. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848615000151.

Anonymous:2015:EPIb

[Ano15c] Anonymous. Editorial and publication information. Studies in History and Philosophy of Biological and Biomedical Sciences, 50(??):ifc, April 2015. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848615000436.

Anonymous:2015:EPIc

[Ano15d] Anonymous. Editorial and publication information. Studies in History and Philosophy of Biological and Biomedical Sciences, 51(??):ifc, June 2015. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848615000710.

Anonymous:2015:EPId

[Ano15e] Anonymous. Editorial and publication information. Studies in History and Philosophy of Biological and Biomedical Sciences, 52(??):ifc, August 2015. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848615000886.

Anonymous:2015:EPIe

[Ano15f] Anonymous. Editorial and publication information. Studies in History and Philosophy of Biological and Biomedical Sciences, 53(??):ifc, ???? 2015. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848615001193.

Anonymous:2015:EPIf

[Ano15g] Anonymous. Editorial and publication information. Studies in History and Philosophy of Biological and Biomedical Sciences, 54 (??):ifc, December 2015. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S136984861500148X.

Anonymous:2015:IPM

[Ano15h] Anonymous. Introduction: Philosophers meet biologists. Studies in History and Philosophy of Biological and Biomedical Sciences, 53(??):64–67, ???? 2015. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848615000631.

Anonymous:2015:P

[Ano15i] Anonymous. Preface. Studies in History and Philosophy of Biological and Biomedical Sciences, 53(??):63, ???? 2015. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848615000643.

Anonymous:2016:EPIa

[Ano16a] Anonymous. Editorial and publication information. Studies in History and Philosophy of Biological and Biomedical Sciences, 55 (??):ifc, February 2016. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S136984861600008X.

Anonymous:2016:EPIb

[Ano16b] Anonymous. Editorial and publication information. Studies in History and Philosophy of Biological and Biomedical Sciences, 56(??):ifc, April 2016. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848616000261.

Anonymous:2016:EPIc

[Ano16c] Anonymous. Editorial and publication information. Studies in History and Philosophy of Biological and Biomedical Sciences, 57(??):ifc, June 2016. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848616300267.

Anonymous:2016:EPId

[Ano16d] Anonymous. Editorial and publication information. Studies in History and Philosophy of Biological and Biomedical Sciences, 58(??):ifc, August 2016. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848616300486.

Anonymous:2016:EPIe

[Ano16e] Anonymous. Editorial and publication information. Studies in History and Philosophy of Biological and Biomedical Sciences, 59 (??):ifc, October 2016. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848616300590.

Anonymous:2016:EPIf

[Ano16f] Anonymous. Editorial and publication information. Studies in History and Philosophy of Biological and Biomedical Sciences, 60 (??):ifc, December 2016. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848616300875.

Anonymous:2017:EPIg

[Ano17a] Anonymous. Editorial and publication information. Studies in History and Philosophy of Biological and Biomedical Sciences, 61 (??):ifc, February 2017. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848617300055.

Anonymous:2017:EPIa

[Ano17b] Anonymous. Editorial and publication information. Studies in History and Philosophy of Biological and Biomedical Sciences, 62(??):ifc, April 2017. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848617300353.

Anonymous:2017:EPIb

[Ano17c] Anonymous. Editorial and publication information. Studies in History and Philosophy of Biological and Biomedical Sciences, 63(??):ifc, June 2017. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848617300882.

Anonymous:2017:EPIc

[Ano17d] Anonymous. Editorial and publication information. Studies in History and Philosophy of Biological and Biomedical Sciences, 64(??):ifc, August 2017. CODEN ???? ISSN 1369-8486 (print), 1879-2499 (electronic). URL /science/article/pii/

S1369848617301188.

Anonymous:2017:EPId

[Ano17e] Anonymous. Editorial and publication information. Studies in History and Philosophy of Biological and Biomedical Sciences, 65(??):ifc, October 2017. CODEN ???? ISSN 1369-8486 (print), 1879-2499 (electronic). URL /science/article/pii/

S1369848617301449.

Anonymous:2017:EPIe

[Ano17f] Anonymous. Editorial and publication information. Studies in History and Philosophy of Biological and Biomedical Sciences, 66 (??):ifc, December 2017. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848617302030.

Anonymous:2018:EBa

[Ano18a] Anonymous. Editorial board. Studies in History and Philosophy of Biological and Biomedical Sciences, 67(??):ii, February 2018. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848618300232.

Anonymous:2018:EBb

[Ano18b] Anonymous. Editorial Board. Studies in History and Philosophy of Biological and Biomedical Sciences, 68–69(??):1–94, April/ June 2018. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL https://www.sciencedirect.com/science/
article/pii/S1369848618300773.

Anonymous:2018:EBc

[Ano18c] Anonymous. Editorial Board. Studies in History and Philosophy of Biological and Biomedical Sciences, 70(??):1–38, August 2018. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL https://www.sciencedirect.com/science/
article/pii/S1369848618301110.

Anonymous:2018:EBd

[Ano18d] Anonymous. Editorial Board. Studies in History and Philosophy of Biological and Biomedical Sciences, 71(??):ii, October 2018. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848618301511.

Anonymous:2018:EBe

[Ano18e] Anonymous. Editorial Board. Studies in History and Philosophy of Biological and Biomedical Sciences, 72(??):ii, December 2018. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848618301651.

Anonymous:2019:A

[Ano19a] Anonymous. August 2019. Studies in History and Philosophy of Biological and Biomedical Sciences, 76(??):??, August 2019. CODEN ???? ISSN 1369-8486 (print), 1879-2499 (electronic).

Anonymous:2019:D

[Ano19b] Anonymous. December 2019. Studies in History and Philosophy of Biological and Biomedical Sciences, 78(??):??, December 2019. CODEN ???? ISSN 1369-8486 (print), 1879-2499 (electronic).

Anonymous:2019:EBa

[Ano19c] Anonymous. Editorial Board. Studies in History and Philosophy of Biological and Biomedical Sciences, 73(??):ii, February 2019. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848619300172.

Anonymous:2019:EBb

[Ano19d] Anonymous. Editorial Board. Studies in History and Philosophy of Biological and Biomedical Sciences, 74(??):ii, April 2019. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848619300469.

Anonymous:2019:EBc

[Ano19e] Anonymous. Editorial Board. Studies in History and Philosophy of Biological and Biomedical Sciences, 75(??):ii, June 2019. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848619300706.

Anonymous:2019:EBd

[Ano19f] Anonymous. Editorial Board. Studies in History and Philosophy of Biological and Biomedical Sciences, 76(??):Article 101195, August 2019. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848619301049.

Anonymous:2019:EBe

[Ano19g] Anonymous. Editorial Board. Studies in History and Philosophy of Biological and Biomedical Sciences, 77(??):Article 101211, October 2019. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848619301402.

Anonymous:2019:EBf

[Ano19h] Anonymous. Editorial Board. Studies in History and Philosophy of Biological and Biomedical Sciences, 78(??):Article 101232, December 2019. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848619301712.

Anonymous:2019:O

[Ano19i] Anonymous. October 2019. Studies in History and Philosophy of Biological and Biomedical Sciences, 77(??):??, October 2019. CODEN ???? ISSN 1369-8486 (print), 1879-2499 (electronic).

Anonymous:2019:PN

[Ano19j] Anonymous. Publisher’s note. Studies in History and Philosophy of Biological and Biomedical Sciences, 76(??):Article 101198, August 2019. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848619301074.

Anonymous:2020:Aa

[Ano20a] Anonymous. April 2020. Studies in History and Philosophy of Biological and Biomedical Sciences, 80(??):??, April 2020. CO-DEN ???? ISSN 1369-8486 (print), 1879-2499 (electronic).

Anonymous:2020:Ab

[Ano20b] Anonymous. August 2020. Studies in History and Philosophy of Biological and Biomedical Sciences, 82(??):??, August 2020. CODEN ???? ISSN 1369-8486 (print), 1879-2499 (electronic).

Anonymous:2020:D

[Ano20c] Anonymous. December 2020. Studies in History and Philosophy of Biological and Biomedical Sciences, 84(??):??, December 2020. CODEN ???? ISSN 1369-8486 (print), 1879-2499 (electronic).

Anonymous:2020:EBa

[Ano20d] Anonymous. Editorial Board. Studies in History and Philosophy of Biological and Biomedical Sciences, 79(??):Article 101258, February 2020. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848620300169.

Anonymous:2020:EBb

[Ano20e] Anonymous. Editorial Board. Studies in History and Philosophy of Biological and Biomedical Sciences, 80(??):Article 101281, April 2020. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848620300704.

Anonymous:2020:EBc

[Ano20f] Anonymous. Editorial Board. Studies in History and Philosophy of Biological and Biomedical Sciences, 81(??):Article 101299, June 2020. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848620300996.

Anonymous:2020:EBd

[Ano20g] Anonymous. Editorial Board. Studies in History and Philosophy of Biological and Biomedical Sciences, 82(??):Article 101316, August 2020. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848620301242.

Anonymous:2020:EBe

[Ano20h] Anonymous. Editorial Board. Studies in History and Philosophy of Biological and Biomedical Sciences, 83(??):Article 101344, October 2020. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848620301552.

Anonymous:2020:EBf

[Ano20i] Anonymous. Editorial Board. Studies in History and Philosophy of Biological and Biomedical Sciences, 84(??):Article 101352, December 2020. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848620301667.

Anonymous:2020:F

[Ano20j] Anonymous. February 2020. Studies in History and Philosophy of Biological and Biomedical Sciences, 79(??):??, February 2020. CODEN ???? ISSN 1369-8486 (print), 1879-2499 (electronic).

Anonymous:2020:J

[Ano20k] Anonymous. June 2020. Studies in History and Philosophy of Biological and Biomedical Sciences, 81(??):??, June 2020. CODEN ???? ISSN 1369-8486 (print), 1879-2499 (electronic).

Anonymous:2020:O

[Ano20l] Anonymous. October 2020. Studies in History and Philosophy of Biological and Biomedical Sciences, 83(??):??, October 2020. CODEN ???? ISSN 1369-8486 (print), 1879-2499 (electronic).

Anstey:2002:BSP

[Ans02] Peter R. Anstey. Boyle on seminal principles. Studies in History and Philosophy of Biological and Biomedical Sciences, 33(4):597– 630, December 2002. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848602000250.

Anderson:1998:TNO

[AR98] James A. Anderson and Edward Rosenfeld, editors. Talking nets: an oral history of neural networks. MIT Press, Cambridge, MA, USA, 1998. ISBN 0-262-01167-0. xi + 434 pp. LCCN QA76.87 .T37 1998.

Andersen:2020:UEB

[AR20] Fredrik Andersen and Elena Rocca. Underdetermination and evidence-based policy. Studies in History and Philosophy of Biological and Biomedical Sciences, 84(??):Article 101335, December 2020. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848620301461.

Ariew:2003:THG

[Ari03] Andre Ariew. The triple helix: gene, organism and environment. Studies in History and Philosophy of Biological and Biomedical Sciences, 34(4):707–712, December 2003. CODEN ???? ISSN 1369-8486 (print), 1879-2499 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S1369848603000761.

Ariew:2007:UIM

[Ari07] André Ariew. Under the influence of Malthus’s law of population growth: Darwin eschews the statistical techniques of Aldolphe Quetelet. Studies in History and Philosophy of Biological and Biomedical Sciences, 38(1):1–19, March 2007. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848606000896.

Ariew:2017:GRQ

[ARR17] André Ariew, Yasha Rohwer, and Collin Rice. Galton, reversion and the quincunx: the rise of statistical explanation. Studies in History and Philosophy of Biological and Biomedical Sciences, 66(??):63–72, December 2017. CO-DEN ???? ISSN 1369-8486 (print), 1879-2499 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S1369848617300304.

Artiga:2020:MIM

[Art20] Dr Marc Artiga. Models, information and meaning. Studies in History and Philosophy of Biological and Biomedical Sciences, 82(??):Article 101284, August 2020. CO-DEN ???? ISSN 1369-8486 (print), 1879-2499 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S1369848618301389.

Ashcroft:2002:WCE

[Ash02] Richard Ashcroft. What is clinical effectiveness? Studies in History and Philosophy of Biological and Biomedical Sciences, 33 (2):219–233, July 2002. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368102000201.

Ashcroft:2003:ATB

[Ash03] Richard E. Ashcroft. Autonomy and trust in bioethics: The Gifford Lectures, University of Edinburgh, 2001. Studies in History and Philosophy of Biological and Biomedical Sciences, 34 (2):359–366, June 2003. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848603000232.

Ashcroft:2004:BCI

[Ash04] Richard E. Ashcroft. Bioethics and conflicts of interest. Studies in History and Philosophy of Biological and Biomedical Sciences, 35(1):155–165, March 2004. CODEN ???? ISSN 1369-8486 (print), 1879-2499 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S1369848603001158.

Atalic:2012:DSR

[Ata12] Bruno Atalic. Differences and similarities in the regulation of medical practice between early modern Vienna and Osijek. Studies in History and Philosophy of Biological and Biomedical Sciences, 43(3):691–699, September 2012. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848612000337.

Aylward:2018:BRL

[Ayl18] Alex Aylward. Book review: Life Histories of Genetic Disease: Patterns and Prevention in Postwar Medical Genetics, Andrew Hogan. Johns Hopkins University Press, Baltimore, MD (2016), 280 pp. Price $40.00 hardback, ISBN: 978-1-4214-2074-5. Studies in History and Philosophy of Biological and Biomedical Sciences, 67(??):36–40, February 2018. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848617302194.

Azzolini:2010:PUA

[Azz10] Monica Azzolini. The political uses of astrology: predicting the illness and death of princes, kings and popes in the Italian Renaissance. Studies in History and Philosophy of Biological and Biomedical Sciences, 41(2):135–145, June 2010. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S136984861000018X.

Bechtel:2005:EMA

[BA05] William Bechtel and Adele Abrahamsen. Explanation: a mechanist alternative. Studies in History and Philosophy of Biological and Biomedical Sciences, 36(2):421–441, June 2005. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848605000269.

Babcock:2020:AOI

[Bab20] Gunnar Babcock. Asexual organisms, identity and vertical gene transfer. Studies in History and Philosophy of Biological and Biomedical Sciences, 81(??):Article 101265, June 2020. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848619301244.

Baca:2015:OSB

[Bac15] George Baca. Observation and “science” in British anthropology before the “Malinowskian Revolution”. Studies in History and Philosophy of Biological and Biomedical Sciences, 54(??): 81–83, December 2015. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848615001144.

Baetu:2012:GAH

[Bae12] Tudor M. Baetu. Genes after the human genome project. Studies in History and Philosophy of Biological and Biomedical Sciences, 43(1):191–201, March 2012. CODEN ???? ISSN 1369-8486 (print), 1879-2499 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S1369848611001142.

Baedke:2013:ELC

[Bae13] Jan Baedke. The epigenetic landscape in the course of time: Conrad Hal Waddington’s methodological impact on the life sciences. Studies in History and Philosophy of Biological and Biomedical Sciences, 44(4):756–773, December 2013. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848613000897.

Baetu:2014:MMS

[Bae14] Tudor M. Baetu. Models and the mosaic of scientific knowledge. The case of immunology. Studies in History and Philosophy of Biological and Biomedical Sciences, 45(??):49–56, March 2014. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848613001520.

Baetu:2020:PPB

[Bae20] Tudor M. Baetu. Pain in psychology, biology and medicine: Some implications for pain eliminativism. Studies in History and Philosophy of Biological and Biomedical Sciences, 82(??):Article 101292, August 2020. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848619301542.

Baker:2005:ASR

[Bak05] Jason M. Baker. Adaptive speciation: the role of natural selection in mechanisms of geographic and non-geographic speciation. Studies in History and Philosophy of Biological and Biomedical Sciences, 36(2):303–326, June 2005. CO-DEN ???? ISSN 1369-8486 (print), 1879-2499 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S136984860500021X.

Baker:2018:LAS

[Bak18] Tawrin Baker. Letting animals speak: Early modern scientific methods, speech, and the human/animal divide. Studies in History and Philosophy of Biological and Biomedical Sciences, 71 (??):41–45, October 2018. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848618301006.

Balinska:2000:RFN

[Bal00] Marta Aleksandra Balinska. The Rockefeller Foundation and the National Institute of Hygiene, Poland, 1918–45. Studies in History and Philosophy of Biological and Biomedical Sciences, 31(3):

419–432, September 2000. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S136984860000008X.

Bangham:2014:BGH

[Ban14] Jenny Bangham. Blood groups and human groups: Collecting and calibrating genetic data after World War Two. Studies in History and Philosophy of Biological and Biomedical Sciences, 47 (part A)(??):74–86, September 2014. CO-DEN ???? ISSN 1369-8486 (print), 1879-2499 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S136984861400048X.

Barnett:2006:EDR

[Bar06] Richard Barnett. Education or degeneration: E. Ray Lankester, H. G. Wells and The Outline of History. Studies in History and Philosophy of Biological and Biomedical Sciences, 37(2): 203–229, June 2006. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848606000148. See [Wel31, Wel56,

WPW71, W^{+}76].

Baron:2009:EVB

[Bar09] Christian Baron. Epistemic values in the Burgess Shale debate. Studies in History and Philosophy of Biological and Biomedical Sciences, 40(4):286–295, December 2009. CO-DEN ???? ISSN 1369-8486 (print), 1879-2499 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S1369848609000570.

Barham:2012:NAL

[Bar12] James Barham. Normativity, agency, and life. Studies in History and Philosophy of Biological and Biomedical Sciences, 43 (1):92–103, March 2012. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848611000446.

Barker:2013:BER

[Bar13a] Matthew J. Barker. Biological explanations, realism, ontology, and categories (reviewing J. Dupré, Processes of life: Essays in the philosophy of biology). Studies in History and Philosophy of Biological and Biomedical Sciences, 44(4):617–622, December 2013. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848613001179.

Baron:2013:HPA

[Bar13b] Christian Baron. The handicap principle and the argument of subversion from within. Studies in History and Philosophy of Biological and Biomedical Sciences, 44(3):347–355, September 2013. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848613000691.

Baranski:2015:WAG

[Bar15] Marci R. Baranski. Wide adaptation of Green Revolution wheat: International roots and the Indian context of a new plant breeding ideal, 1960–1970. Studies in History and Philosophy of Biological and Biomedical Sciences, 50(??):41–50, April 2015. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848615000060.

Barmpouti:2020:IBR

[Bar20] Alexandra Barmpouti. Issues of biopolitics of reproduction in post-war Greece. Studies in History and Philosophy of Biological and Biomedical Sciences, 83(??):Article 101276, October 2020. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848620300522.

Bassiri:2012:MTC

[Bas12] Nima Bassiri. Material translations in the Cartesian brain. Studies in History and Philosophy of Biological and Biomedical Sciences, 43(1):244–255, March 2012. CODEN ???? ISSN 1369-8486 (print), 1879-2499 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S1369848611000938.

Bateson:2001:DDD

[Bat01] Patrick Bateson. Design, development and decisions. Studies in History and Philosophy of Biological and Biomedical Sciences, 32(4):635–646, December 2001. CODEN ???? ISSN 1369-8486 (print), 1879-2499 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S1369848601000231.

Bauer:2008:MDG

[Bau08] Susanne Bauer. Mining data, gathering variables and recombining information: the flexible architecture of epidemiological studies. Studies in History and Philosophy of Biological and Biomedical Sciences, 39(4):415–428, December 2008. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848608000721.

Bauer:2014:MSP

[Bau14] Susanne Bauer. Mutations in Soviet public health science: Post-Lysenko medical genetics, 1969–1991. Studies in History and Philosophy of Biological and Biomedical Sciences, 47 (part A)(??): 163–172, September 2014. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848614000569.

Baum:2020:BRI

[Bau20] Emily Baum. Book review: The Invention of Madness by Emily Baum: Reply by the author. Studies in History and Philosophy of Biological and Biomedical Sciences, 79(??):Article 101206, February 2020. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848619300986. See [Kit20, Pol20, Edi20].

Ben-Chaim:2001:SDN

[BC01] Michael Ben-Chaim. The scientific discovery of ‘natural capital’: The production of catalytic antibodies. Studies in History and Philosophy of Biological and Biomedical Sciences, 32(3):413– 433, September 2001. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848601000127.

Bostanci:2008:IGG

[BC08] Adam Bostanci and Jane Calvert. Invisible genomes: the genomics revolution and patenting practice. Studies in History and Philosophy of Biological and Biomedical Sciences, 39(1):109–119, March 2008. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848607000933.

Baedke:2019:RNN

[BD19] Jan Baedke and Abigail Nieves Delgado. Race and nutrition in the New World: Colonial shadows in the age of epigenetics. Studies in History and Philosophy of Biological and Biomedical Sciences, 76(??):Article 101175, August 2019. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848618300463.

Bangham:2014:HHA

[BdC14] Jenny Bangham and Soraya de Chadarevian. Human heredity after 1945: Moving populations centre stage. Studies in History and Philosophy of Biological and Biomedical Sciences, 47 (part A)(??):45–49, September 2014. CO-DEN ???? ISSN 1369-8486 (print), 1879-2499 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S1369848614000454.

Brooks:2018:SLO

[BE18] Daniel S. Brooks and Markus I. Eronen. The significance of levels of organization for scientific research: a heuristic approach 1. Studies in History and Philosophy of Biological and Biomedical Sciences, 68–69(??):25–33, April/June 2018. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL https://www.sciencedirect.com/science/
article/pii/S1369848616301108.

Beach:2003:TCS

[Bea03] Jon Beach. The transition to civilization and symbolically stored genomes. Studies in History and Philosophy of Biological and Biomedical Sciences, 34(1):109–141, March 2003. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848602000687.

Beaulieu:2004:BDI

[Bea04] Anne Beaulieu. From brainbank to database: the informational turn in the study of the brain. Studies in History and Philosophy of Biological and Biomedical Sciences, 35(2):367–390, June 2004. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848604000263.

Beatty:2016:WNG

[Bea16] John Beatty. What are narratives good for? Studies in History and Philosophy of Biological and Biomedical Sciences, 58(??): 33–40, August 2016. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848616000029.

Bechtel:2010:CLO

[Bec10] William Bechtel. The cell: locus or object of inquiry? Studies in History and Philosophy of Biological and Biomedical Sciences, 41(3):172–182, September 2010. CODEN ???? ISSN 1369-8486 (print), 1879-2499 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S1369848610000403.

Becquemont:2011:SDR

[Bec11] Daniel Becquemont. Social Darwinism: from reality to myth and from myth to reality. Studies in History and Philosophy of Biological and Biomedical Sciences, 42(1):12–19, March 2011. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848610000944.

Bechtel:2013:MBC

[Bec13] William Bechtel. From molecules to behavior and the clinic: Integration in chronobiology. Studies in History and Philosophy of Biological and Biomedical Sciences, 44(4):493–502, December 2013. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848612001082.

Bechtel:2015:CME

[Bec15] William Bechtel. Can mechanistic explanation be reconciled with scale-free constitution and dynamics? Studies in History and Philosophy of Biological and Biomedical Sciences, 53(??):84–93, ???? 2015. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S136984861500059X.

Beck:2016:SMO

[Bec16] Naomi Beck. The spontaneous market order and evolution. Studies in History and Philosophy of Biological and Biomedical Sciences, 58(??):49–55, August 2016. CO-DEN ???? ISSN 1369-8486 (print), 1879-2499 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S1369848615001764.

Bednarski:2015:GSD

[Bed15] Andrew Bednarski. Global scientific dialogues: Darwin in other languages. Studies in History and Philosophy of Biological and Biomedical Sciences, 50(??):87–89, April 2015. CODEN ???? ISSN 1369-8486 (print), 1879-2499 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S1369848615000333.

Bellon:2005:QMJ

[Bel05] Richard Bellon. A question of merit: John Hutton Balfour, Joseph Hooker and the ‘concussion’ over the Edinburgh chair of botany. Studies in History and Philosophy of Biological and Biomedical Sciences, 36(1):25–54, March 2005. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848604000974.

Benninghaus:2007:GEG

[Ben07] Christina Benninghaus. Great expectations — German debates about artificial insemination in humans around 1912. Studies in History and Philosophy of Biological and Biomedical Sciences, 38 (2):374–392, June 2007. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848607000155.

Benson:2016:TLD

[Ben16] Etienne S. Benson. Trackable life: Data, sequence, and organism in movement ecology. Studies in History and Philosophy of Biological and Biomedical Sciences, 57(??):137–147, June 2016. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848616000236.

Berridge:2010:HMM

[Ber10] Virginia Berridge. History, medicine and the media. Studies in History and Philosophy of Biological and Biomedical Sciences, 41(3):304–306, September 2010. CODEN ???? ISSN 1369-8486 (print), 1879-2499 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S1369848610000464.

Berry:2014:BBP

[Ber14a] Dominic Berry. Bruno to Brünn; or the pasteurization of Mendelian genetics. Studies in History and Philosophy of Biological and Biomedical Sciences, 48 (part B)(??):280–286, December 2014. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S136984861400137X.

Berry:2014:PBI

[Ber14b] Dominic Berry. The plant breeding industry after pure line theory: Lessons from the National Institute of Agricultural Botany. Studies in History and Philosophy of Biological and Biomedical Sciences, 46(??):25–37, June 2014. CODEN ???? ISSN 1369-8486 (print), 1879-2499 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S1369848614000296.

Bertoldi:2020:AOE

[Ber20] Nicola Bertoldi. Adaptation and optimality in evolutionary biology: Historical and philosophical perspectives on the interpretations of R. A. Fisher’s “Fundamental theorem of natural selection” and the “Formal Darwinism” project. Studies in History and Philosophy of Biological and Biomedical Sciences, 81(??): Article 101285, June 2020. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848620300662.

Balz:2011:ESS

[BH11] Viola Balz and Matthias Hoheisel. East-Side story: The standardisation of psychotropic drugs at the Charité Psychiatric Clinic, 1955–1970. Studies in History and Philosophy of Biological and Biomedical Sciences, 42(4):453–466, December 2011. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848611000537.

Bergstrom:2020:ST

[BHZ20] Carl T. Bergstrom, Simon M. Huttegger, and Kevin J. S. Zollman. Signals without teleology. Studies in History and Philosophy of Biological and Biomedical Sciences, 84(??):Article 101310, December 2020. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848618301055.

Biddle:2020:ERC

[Bid20] Justin B. Biddle. Epistemic risks in cancer screening: Implications for ethics and policy. Studies in History and Philosophy of Biological and Biomedical Sciences, 79(??):Article 101200, February 2020. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848618301857.

Bilgili:2017:BTH

[Bil17] Alper Bilgili. Beating the Turkish hollow in the struggle for existence: Darwin, social Darwinism and the Turks. Studies in History and Philosophy of Biological and Biomedical Sciences, 65(??):19–25, October 2017. CODEN ???? ISSN 1369-8486 (print), 1879-2499 (electronic). URL /science/article/pii/

S1369848617301280.

Binney:2018:FHH

[Bin18a] Nicholas Binney. The function of the heart is historically contingent. Studies in History and Philosophy of Biological and Biomedical Sciences, 68–69(??):34–41, April/June 2018. CODEN ???? ISSN 1369-8486 (print), 1879-2499 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S1369848617301413.

Binney:2018:FHO

[Bin18b] Nicholas Binney. The function of the heart is not obvious. Studies in History and Philosophy of Biological and Biomedical Sciences, 68–69(??):42–55, April/June 2018. CO-DEN ???? ISSN 1369-8486 (print), 1879-2499 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S1369848617301267.

Birn:2000:WVI

[Bir00] Anne-Emanuelle Birn. Wa(i)ves of influence: Rockefeller Public Health in Mexico, 1920–50. Studies in History and Philosophy of Biological and Biomedical Sciences, 31(3):381–395, September 2000. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848600000091.

Birch:2009:ICI

[Bir09] Jonathan Birch. Irretrievably confused? Innateness in explanatory context. Studies in History and Philosophy of Biological and Biomedical Sciences, 40(4):296–301, December 2009. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848609000569.

Birch:2012:NVN

[Bir12] Jonathan Birch. The negative view of natural selection. Studies in History and Philosophy of Biological and Biomedical Sciences, 43(2):569–573, June 2012. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S136984861200026X.

Birch:2019:AD

[Bir19a] Jonathan Birch. Altruistic deception. Studies in History and Philosophy of Biological and Biomedical Sciences, 74(??):27–33, April 2019. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848618300980.

Birch:2019:IFC

[Bir19b] Jonathan Birch. Inclusive fitness as a criterion for improvement. Studies in History and Philosophy of Biological and Biomedical Sciences, 76(??):Article 101186, August 2019. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848618301274.

Burney:2013:IFC

[BKP13] Ian Burney, David A. Kirby, and Neil Pemberton. Introducing ‘forensic cultures’. Studies in History and Philosophy of Biological and Biomedical Sciences, 44(1):1–3, March 2013. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848612001069.

Buklijas:2007:SMN

[BL07] Tatjana Buklijas and Emese Lafferton. Science, medicine and nationalism in the Habsburg Empire from the 1840s to 1918. Studies in History and Philosophy of Biological and Biomedical Sciences, 38(4):679–686, December 2007. CODEN ???? ISSN 1369-8486 (print), 1879-2499 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S1369848607000489.

Blackman:2004:SLC

[Bla04] Helen Blackman. A spiritual leader? Cambridge zoology, mountaineering and the death of F. M. Balfour. Studies in History and Philosophy of Biological and Biomedical Sciences, 35(1):93–117, March 2004. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848603001092.

Blease:2010:SPP

[Ble10] Charlotte Blease. Scientific progress and the prospects for culture-bound syndromes. Studies in History and Philosophy of Biological and Biomedical Sciences, 41(4):333–339, December 2010. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848610000750.

Blount:2016:CSE

[Blo16] Zachary D. Blount. A case study in evolutionary contingency. Studies in History and Philosophy of Biological and Biomedical Sciences, 58(??):82–92, August 2016. CO-DEN ???? ISSN 1369-8486 (print), 1879-2499 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S1369848615001806.

Braun:2015:UCP

[BM15] Erez Braun and Shimon Marom. Universality, complexity and the praxis of biology: Two case studies. Studies in History and Philosophy of Biological and Biomedical Sciences, 53(??):68–72, ???? 2015. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848615000606.

Bambah-Mukku:2016:RBH

[BM16] Dhananjay Bambah-Mukku. A rush of blood to the head: the beginnings of brain imaging. Studies in History and Philosophy of Biological and Biomedical Sciences, 57(??):163–166, June 2016. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848615001867.

Baedke:2018:SET

[BM18] Jan Baedke and Siobhan F. Mc Manus. From seconds to eons: Time scales, hierarchies, and processes in evodevo. Studies in History and Philosophy of Biological and Biomedical Sciences, 72(??):38–48, December 2018. CO-DEN ???? ISSN 1369-8486 (print), 1879-2499 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S1369848617301358.

Blasimme:2013:WMC

[BMG13] Alessandro Blasimme, Paolo Maugeri, and Pierre-Luc Germain. What mechanisms can’t do: Explanatory frameworks and the function of the p53 gene in molecular oncology. Studies in History and Philosophy of Biological and Biomedical Sciences, 44(3):374– 384, September 2013. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848613000046.

Brecht:2000:DIV

[BN00] Christine Brecht and Sybilla Nikolow. Displaying the invisible: Volkskrankheiten on exhibition in Imperial Germany. Studies in History and Philosophy of Biological and Biomedical Sciences, 31(4):511–530, December 2000. CO-DEN ???? ISSN 1369-8486 (print), 1879-2499 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S1369848600000200.

Bocking:2004:EE

[Boc04] Stephen Bocking. Empires of ecology. Studies in History and Philosophy of Biological and Biomedical Sciences, 35(4):793– 801, December 2004. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848604000767.

Bogen:2001:TGH

[Bog01] James Bogen. ‘two as good as a hundred’: poorly replicated evidence in some nineteenth-century neuroscientific research. Studies in History and Philosophy of Biological and Biomedical Sciences, 32(3):491–533, September 2001. CO-DEN ???? ISSN 1369-8486 (print), 1879-2499 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S1369848601000139.

Bogen:2005:RCG

[Bog05] Jim Bogen. Regularities and causality; generalizations and causal explanations. Studies in History and Philosophy of Biological and Biomedical Sciences, 36(2):397–420, June 2005. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848605000257.

Bolduc:2012:BEE

[Bol12] Jean-Sébastien Bolduc. Behavioural ecology’s ethological roots. Studies in History and Philosophy of Biological and Biomedical Sciences, 43(3):674–683, September 2012. CO-DEN ???? ISSN 1369-8486 (print), 1879-2499 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S1369848612000271.

Bonah:2005:ESB

[Bon05] Christian Bonah. The ‘experimental stable’ of the BCG vaccine: safety, efficacy, proof, and standards, 1921–1933. Studies in History and Philosophy of Biological and Biomedical Sciences, 36(4): 696–721, December 2005. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848605000750.

Boniolo:2013:AIN

[Bon13] Giovanni Boniolo. Is an account of identity necessary for bioethics? What post-genomic biomedicine can teach us. Studies in History and Philosophy of Biological and Biomedical Sciences, 44(3):401–411, September 2013. CO-DEN ???? ISSN 1369-8486 (print), 1879-2499 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S1369848613000708.

Bonneuil:2019:SNU

[Bon19] Christophe Bonneuil. Seeing nature as a ‘universal store of genes’: How biological diversity became ‘genetic resources’, 1890–1940. Studies in History and Philosophy of Biological and Biomedical Sciences, 75(??):1–14, June 2019. CO-DEN ???? ISSN 1369-8486 (print), 1879-2499 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S1369848617301346.

Borck:2004:MBC

[Bor04a] Cornelius Borck. Message in a bottle from ‘the crisis of reality’: on Ludwik Fleck’s interventions for an open epistemology. Studies in History and Philosophy of Biological and Biomedical Sciences, 35(3):447–464, September 2004. CODEN ???? ISSN 1369-8486 (print), 1879-2499 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S1369848604000469.

Borrello:2004:SET

[Bor04b] Mark E. Borrello. The structure of evolutionary theory. Studies in History and Philosophy of Biological and Biomedical Sciences, 35(1):209–216, March 2004. CODEN ???? ISSN 1369-8486 (print), 1879-2499 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S1369848603001213.

Bordogna:2005:SPA

[Bor05] Francesca Bordogna. Scientific personae in American psychology: three case studies. Studies in History and Philosophy of Biological and Biomedical Sciences, 36(1):95–134, March 2005. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848604000950.

Borck:2016:HWM

[Bor16] Cornelius Borck. How we may think: Imaging and writing technologies across the history of the neurosciences. Studies in History and Philosophy of Biological and Biomedical Sciences, 57 (??):112–120, June 2016. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848616000248.

Botting:2004:E

[Bot04] Fred Botting. Extimatrix. Studies in History and Philosophy of Biological and Biomedical Sciences, 35(2):269–286, June 2004. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848604000214.

Boudet:2010:CAM

[Bou10] Jean-Patrice Boudet. A ‘college of astrology and medicine’? Charles V, Gervais Chrétien, and the scientific manuscripts of Maître Gervais’s College. Studies in History and Philosophy of Biological and Biomedical Sciences, 41(2):99–108, June 2010. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848610000142.

Bouchard:2011:DPM

[Bou11] Frédéric Bouchard. Darwinism without populations: a more inclusive understanding of the “Survival of the Fittest”. Studies in History and Philosophy of Biological and Biomedical Sciences, 42 (1):106–114, March 2011. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848610000956.

Boucher:2017:GSM

[Bou17] Sandy C. Boucher. Gould on species, metaphysics and macroevolution: a critical appraisal. Studies in History and Philosophy of Biological and Biomedical Sciences, 62(??):25–34, April 2017. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848617300225.

Bowler:2014:FGS

[Bow14] Peter J. Bowler. Francis Galton’s saltationism and the ambiguities of selection. Studies in History and Philosophy of Biological and Biomedical Sciences, 48 (part B)(??):272–279, December 2014. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848614001423.

Bowler:2017:APR

[Bow17] Peter J. Bowler. American palaeontology and the reception of Darwinism. Studies in History and Philosophy of Biological and Biomedical Sciences, 66(??):3–7, December 2017. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848617300511.

Boudry:2013:MMS

[BP13a] Maarten Boudry and Massimo Pigliucci. The mismeasure of machine: Synthetic biology and the trouble with engineering metaphors. Studies in History and Philosophy of Biological and Biomedical Sciences, 44(4):660–668, December 2013. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848613000812.

Burney:2013:MSC

[BP13b] Ian Burney and Neil Pemberton. Making space for criminalistics: Hans Gross and fin-de-siècle CSI. Studies in History and Philosophy of Biological and Biomedical Sciences, 44(1):16–25, March 2013. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848612000969.

Baranski:2015:ICC

[BP15] Marci Baranski and B. R. Erick Peirson. Introduction: Contexts and concepts of adaptability and plasticity in 20thcentury plant science. Studies in History and Philosophy of Biological and Biomedical Sciences, 50(??):26–28, April 2015. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848615000102.

Baylac-Paouly:2019:VDD

[BP19] Baptiste Baylac-Paouly. Vaccine development as a ‘doable problem’: the case of the meningococcal A vaccines 1962–1969. Studies in History and Philosophy of Biological and Biomedical Sciences, 74(??):7–14, April 2019. CODEN ???? ISSN 1369-8486 (print), 1879-2499 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S1369848617301735.

Borghini:2020:GFW

[BPS20] Andrea Borghini, Nicola Piras, and Beatrice Serini. A gradient framework for wild foods. Studies in History and Philosophy of Biological and Biomedical Sciences, 84(??):Article 101293, December 2020. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S136984861930202X.

Braddock:2006:PFS

[Bra06] L. E. Braddock. Psychoanalysis as functionalist social science: the legacy of Freud’s ‘Project for a scientific psychology’. Studies in History and Philosophy of Biological and Biomedical Sciences, 37(3):394–413, September 2006. CODEN ???? ISSN 1369-8486 (print), 1879-2499 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S1369848606000410.

Brassley:2007:CAN

[Bra07] Paul Brassley. Cutting across nature? The history of artificial insemination in pigs in the United Kingdom. Studies in History and Philosophy of Biological and Biomedical Sciences, 38(2):442– 461, June 2007. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848607000180.

Brauckmann:2011:APT

[Bra11a] Sabine Brauckmann. Axes, planes and tubes, or the geometry of embryogenesis. Studies in History and Philosophy of Biological and Biomedical Sciences, 42(4):381–390, December 2011. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848611000720.

Brauckmann:2011:CSE

[Bra11b] Sabine Brauckmann. Cultures of seeing embryos and cells in 3-dimensions and flatness. Studies in History and Philosophy of Biological and Biomedical Sciences, 42(4):365–367, December 2011. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848611000732.

Braunlein:2012:FBE

[Brä12] Peter J. Bräunlein. The frightening borderlands of Enlightenment: the vampire problem. Studies in History and Philosophy of Biological and Biomedical Sciences, 43(3):710–719, September 2012. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848612000313.

Brancaccio:2014:EMP

[Bra14] Maria Teresa Brancaccio. Enrico Morselli’s psychology and “spiritism” : Psychiatry, psychology and psychical research in Italy in the decades around 1900. Studies in History and Philosophy of Biological and Biomedical Sciences, 48 (part A)(??): 75–84, December 2014. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S136984861400096X.

Braun:2020:BRD

[Bra20] Lundy Braun. Book review: Difference and Disease by Suman Seth: Medicine, race, and the eighteenth-century British Empire. Cambridge University Press (2018). Studies in History and Philosophy of Biological and Biomedical Sciences, 80(??):Article 101221, April 2020. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S136984861930127X. See reply [Set20].

Boem:2016:WGL

[BRAB16] F. Boem, E. Ratti, M. Andreoletti, and G. Boniolo. Why genes are like lemons. Studies in History and Philosophy of Biological and Biomedical Sciences, 57(??):88–95, June 2016. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848616300218.

Brennan:2002:ATK

[Bre02] Andrew Brennan. Asian traditions of knowledge: the disputed questions of science, nature and ecology. Studies in History and Philosophy of Biological and Biomedical Sciences, 33(4):567– 581, December 2002. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848602000274.

Breitenbach:2005:KGF

[Bre05] Angela Breitenbach. Kant goes fishing: Kant and the right to property in environmental resources. Studies in History and Philosophy of Biological and Biomedical Sciences, 36(3):488– 512, September 2005. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848605000476.

Breitenbach:2006:MEN

[Bre06] Angela Breitenbach. Mechanical explanation of nature and its limits in Kant’s Critique of judgment. Studies in History and Philosophy of Biological and Biomedical Sciences, 37(4):694– 711, December 2006. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848606000720.

Brigandt:2013:IBP

[Bri13a] Ingo Brigandt. Integration in biology: Philosophical perspectives on the dynamics of interdisciplinarity. Studies in History and Philosophy of Biological and Biomedical Sciences, 44(4):461– 465, December 2013. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848613001246.

Brigandt:2013:SBI

[Bri13b] Ingo Brigandt. Systems biology and the integration of mechanistic explanation and mathematical explanation. Studies in History and Philosophy of Biological and Biomedical Sciences, 44(4): 477–492, December 2013. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848613000903.

Brister:2016:DCE

[Bri16] Evelyn Brister. Disciplinary capture and epistemological obstacles to interdisciplinary research: Lessons from central African conservation disputes. Studies in History and Philosophy of Biological and Biomedical Sciences, 56(??):82–91, April 2016. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848615001624.

Bright:2017:LER

[Bri17] Liam Kofi Bright. Logical empiricists on race. Studies in History and Philosophy of Biological and Biomedical Sciences, 65(??):9–18, October 2017. CODEN ???? ISSN 1369-8486 (print), 1879-2499 (electronic). URL /science/article/pii/

S1369848617301152.

Broadbent:2009:CMD

[Bro09a] Alex Broadbent. Causation and models of disease in epidemiology. Studies in History and Philosophy of Biological and Biomedical Sciences, 40(4):302–311, December 2009. CODEN ???? ISSN 1369-8486 (print), 1879-2499 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S1369848609000557.

Brosnan:2009:QIF

[Bro09b] Kevin Brosnan. Quasi-independence, fitness, and advantageousness. Studies in History and Philosophy of Biological and Biomedical Sciences, 40(3):228–234, September 2009. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848609000351.

Brosseder:2010:ASC

[Bro10] Claudia Brosseder. Astrology in seventeenth-century Peru. Studies in History and Philosophy of Biological and Biomedical Sciences, 41(2):146–157, June 2010. CODEN ???? ISSN 1369-8486 (print), 1879-2499 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S1369848610000191.

Brooks:2011:MRH

[Bro11a] Daniel R. Brooks. The mastodon in the room: how Darwinian is neo-Darwinism? Studies in History and Philosophy of Biological and Biomedical Sciences, 42(1):82–88, March 2011. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848610000968.

Brown:2011:EDM

[Bro11b] Bryson Brown. Ethics in Darwin’s melancholy vision. Studies in History and Philosophy of Biological and Biomedical Sciences, 42 (1):20–29, March 2011. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S136984861000097X.

Broadbent:2014:DTC

[Bro14] Alex Broadbent. Disease as a theoretical concept: the case of “HPV-itis”. Studies in History and Philosophy of Biological and Biomedical Sciences, 48 (part B)(??):250–257, December 2014. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848614000910.

Broadbent:2015:CPE

[Bro15a] Alex Broadbent. Causation and prediction in epidemiology: a guide to the “Methodological Revolution”. Studies in History and Philosophy of Biological and Biomedical Sciences, 54(??): 72–80, December 2015. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848615000849.

Brown:2015:CED

[Bro15b] Rachael L. Brown. A clear-eyed defense of philosophy of biology. Studies in History and Philosophy of Biological and Biomedical Sciences, 49(??):63–65, February 2015. CODEN ???? ISSN 1369-8486 (print), 1879-2499 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S1369848614001794.

Brunton:2003:SGD

[Bru03] Deborah C. Brunton. Spreading germs: disease theories and medical practice in Britain, 1865–1900. Studies in History and Philosophy of Biological and Biomedical Sciences, 34(2):367–373, June 2003. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848603000220.

Brunnander:2008:LIP

[Bru08] Björn Brunnander. Is the language of intentional psychology an efficient tool for evolutionists? Studies in History and Philosophy of Biological and Biomedical Sciences, 39(1):147–152, March 2008. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848607000969.

Bruhn:2011:LLA

[Bru11] Matthias Bruhn. Life lines: An art history of biological research around 1800. Studies in History and Philosophy of Biological and Biomedical Sciences, 42(4):368–380, December 2011. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848611000707.

Brunnander:2013:DDR

[Bru13] Björn Brunnander. Did Darwin really answer Paley’s question? Studies in History and Philosophy of Biological and Biomedical Sciences, 44(3):309–311, September 2013. CO-DEN ???? ISSN 1369-8486 (print), 1879-2499 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S1369848613000186.

Brysse:2008:WWS

[Bry08] Keynyn Brysse. From weird wonders to stem lineages: the second reclassification of the Burgess Shale fauna. Studies in History and Philosophy of Biological and Biomedical Sciences, 39(3):298– 313, September 2008. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848608000393.

Basl:2013:GNS

[BS13] John Basl and Ronald Sandler. The good of non-sentient entities: Organisms, artifacts, and synthetic biology. Studies in History and Philosophy of Biological and Biomedical Sciences, 44(4):697– 705, December 2013. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S136984861300085X.

Brzovic:2020:PFM

[BS20] Zdenka Brzović and Predrag Sustar. Postgenomics function monism. Studies in History and Philosophy of Biological and Biomedical Sciences, 80(??):Article 101243, April 2020. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848619300342.

Bud:2014:BEN

[Bud14] Robert Bud. The beer experience: Nineteenth century relations between science and praxis. Studies in History and Philosophy of Biological and Biomedical Sciences, 47 (part A)(??): 224–226, September 2014. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848614000442.

Bud:2016:RSW

[Bud16] Robert Bud. Representing scale: What should be special about the heritage of mass science? Studies in History and Philosophy of Biological and Biomedical Sciences, 55(??):117–119, February 2016. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848615001168.

Bud:2018:UC

[Bud18] Robert Bud. The unstable collection. Studies in History and Philosophy of Biological and Biomedical Sciences, 68–69(??):56– 69, April/June 2018. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S136984861830013X.

Bueno:2011:WPB

[Bue11] Otávio Bueno. When physics and biology meet: The nanoscale case. Studies in History and Philosophy of Biological and Biomedical Sciences, 42(2):180–189, June 2011. CODEN ???? ISSN 1369-8486 (print), 1879-2499 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S1369848610001184.

Buklijas:2007:SNI

[Buk07] Tatjana Buklijas. Surgery and national identity in late nineteenth-century Vienna. Studies in History and Philosophy of Biological and Biomedical Sciences, 38(4):756–774, December 2007. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848607000477.

Buklijas:2014:FGT

[Buk14] Tatjana Buklijas. Food, growth and time: Elsie Widdowson’s and Robert McCance’s research into prenatal and early postnatal growth. Studies in History and Philosophy of Biological and Biomedical Sciences, 47 (part B)(??):267–277, September 2014. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848613001751.

Burt:2006:SZM

[Bur06] Jonathan Burt. Solly Zuckerman: the making of a primatological career in Britain, 1925–1945. Studies in History and Philosophy of Biological and Biomedical Sciences, 37(2):295–310, June 2006. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848606000203.

Burnett:2010:HLA

[Bur10] Charles Burnett. Hebrew and Latin astrology in the twelfth century: the example of the location of pain. Studies in History and Philosophy of Biological and Biomedical Sciences, 41(2):70–75, June 2010. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848610000117.

Burnston:2016:DGM

[Bur16] Daniel C. Burnston. Data graphs and mechanistic explanation. Studies in History and Philosophy of Biological and Biomedical Sciences, 57(??):1–12, June 2016. CO-DEN ???? ISSN 1369-8486 (print), 1879-2499 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S1369848616000169.

Buskell:2020:SAE

[Bus20] Andrew Buskell. Synthesising arguments and the extended evolutionary synthesis. Studies in History and Philosophy of Biological and Biomedical Sciences, 80(??):Article 101244, April 2020. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848618301079.

Baravalle:2016:BBR

[BV16] Lorenzo Baravalle and Davide Vecchi. Beyond blindness: On the role of organism and environment in trial generation. Studies in History and Philosophy of Biological and Biomedical Sciences, 60(??):25–34, December 2016. CO-DEN ???? ISSN 1369-8486 (print), 1879-2499 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S1369848616300693.

Claverie:2016:GVD

[CA16] Jean-Michel Claverie and Chantal Abergel. Giant viruses: the difficult breaking of multiple epistemological barriers. Studies in History and Philosophy of Biological and Biomedical Sciences, 59 (??):89–99, October 2016. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848616300097.

Cain:2002:ECT

[Cai02] Joe Cain. Epistemic and community transition in American evolutionary studies: the ‘Committee on Common Problems of Genetics, Paleontology, and Systematics’ (1942–1949). Studies in History and Philosophy of Biological and Biomedical Sciences, 33 (2):283–313, July 2002. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368102000195.

Callebaut:2012:SPP

[Cal12a] Werner Callebaut. Scientific perspectivism: a philosopher of science’s response to the challenge of big data biology. Studies in History and Philosophy of Biological and Biomedical Sciences, 43 (1):69–80, March 2012. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848611000835.

Calvert:2012:SBS

[Cal12b] Jane Calvert. Systems biology, synthetic biology and datadriven research: a commentary on Krohs, Callebaut, and O’Malley and Soyer. Studies in History and Philosophy of Biological and Biomedical Sciences, 43(1):81–84, March 2012. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848611000847.

Caniglia:2017:HCE

[Can17] Guido Caniglia. “How complex and even perverse the real world can be”: W. D. Hamilton’s early work on social wasps (1964– 1968). Studies in History and Philosophy of Biological and Biomedical Sciences, 64(??):41–52, August 2017. CODEN ???? ISSN 1369-8486 (print), 1879-2499 (electronic). URL /science/

article/pii/S1369848616301248.

Canali:2020:MEC

[Can20] Stefano Canali. Making evidential claims in epidemiology: Three strategies for the study of the exposome. Studies in History and Philosophy of Biological and Biomedical Sciences, 82(??):Article 101248, August 2020. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848618301067.

Carson:1999:MMM

[Car99] John Carson. Minding matter/mattering mind: Knowledge and the subject in nineteenth-century psychology. Studies in History and Philosophy of Biological and Biomedical Sciences, 30(3):345– 376, September 1999. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848699000163.

Carroll:2004:NHV

[Car04] Victoria Carroll. The natural history of visiting: responses to Charles Waterton and Walton Hall. Studies in History and Philosophy of Biological and Biomedical Sciences, 35(1):31–64, March 2004. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848603001079.

Carey:2010:JAT

[Car10a] Hilary M. Carey. Judicial astrology in theory and practice in later medieval Europe. Studies in History and Philosophy of Biological and Biomedical Sciences, 41(2):90–98, June 2010. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848610000130.

Carter:2010:CTE

[Car10b] K. Codell Carter. Change of type as an explanation for the decline of therapeutic bloodletting. Studies in History and Philosophy of Biological and Biomedical Sciences, 41(1):1–11, March 2010. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848609000764.

Carson:2012:PFT

[Car12a] John Carson. Has psychology “found its true path”? Methods, objectivity, and cries of “crisis” in early twentiethcentury French psychology. Studies in History and Philosophy of Biological and Biomedical Sciences, 43(2):445–454, June 2012. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848611001294.

Carter:2012:DTB

[Car12b] K. Codell (Kay Codell) Carter. The decline of therapeutic bloodletting and the collapse of traditional medicine. Transaction Publishers, New Brunswick, NJ, USA, 2012. ISBN 1-4128-4604-8. ix + 160 pp. LCCN RM182 .C37 2012.

Carusi:2014:VVD

[Car14] Annamaria Carusi. Validation and variability: Dual challenges on the path from systems biology to systems medicine. Studies in History and Philosophy of Biological and Biomedical Sciences, 48 (part A)(??):28–37, December 2014. CO-DEN ???? ISSN 1369-8486 (print), 1879-2499 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S1369848614001265.

Cartwright:2016:CON

[Car16] Nancy Cartwright. Contingency and the order of nature. Studies in History and Philosophy of Biological and Biomedical Sciences, 58(??):56–63, August 2016. CODEN ???? ISSN 1369-8486 (print), 1879-2499 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S1369848615001818.

Castle:2003:MSA

[Cas03] David Castle. The moral significance of agricultural biotechnology. Studies in History and Philosophy of Biological and Biomedical Sciences, 34(4):713–722, December 2003. CO-DEN ???? ISSN 1369-8486 (print), 1879-2499 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S1369848603000773.

Cassier:2005:ACP

[Cas05] Maurice Cassier. Appropriation and commercialization of the Pasteur anthrax vaccine. Studies in History and Philosophy of Biological and Biomedical Sciences, 36(4):722–742, December 2005. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848605000762.

Cavallaro:2004:BVC

[Cav04] Dani Cavallaro. The brain in a vat in cyberpunk: the persistence of the flesh. Studies in History and Philosophy of Biological and Biomedical Sciences, 35(2):287–305, June 2004. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848604000226.

Cooper:2018:MHD

[CB18] Rachel Cooper and Roger Blashfield. The myth of Hempel and the DSM-III. Studies in History and Philosophy of Biological and Biomedical Sciences, 70(??):1–9, August 2018. CODEN ???? ISSN 1369-8486 (print), 1879-2499 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S1369848617301073.

Cuevas-Badallo:2011:FAB

[CBV11] Ana Cuevas-Badallo and Pieter E. Vermaas. A functional abc for biotechnology and the dissemination of its progeny. Studies in History and Philosophy of Biological and Biomedical Sciences, 42 (2):261–269, June 2011. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848610001287.

Capocci:2002:ABT

[CC02] Mauro Capocci and Gilberto Corbellini. Adriano Buzzati-Traverso and the foundation of the International Laboratory of Genetics and Biophysics in Naples (1962–1969). Studies in History and Philosophy of Biological and Biomedical Sciences, 33(3): 489–513, September 2002. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848602000079.

Carel:2010:ICB

[CC10] Havi Carel and Rachel Cooper. Introduction: culturebound syndromes. Studies in History and Philosophy of Biological and Biomedical Sciences, 41(4):307–308, December 2010. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S136984861000066X.

Clavien:2016:EUF

[CC16] Christine Clavien and Michel Chapuisat. The evolution of utility functions and psychological altruism. Studies in History and Philosophy of Biological and Biomedical Sciences, 56(??):24–31, April 2016. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848615001582.

Craver:2005:I

[CD05] Carl F. Craver and Lindley Darden. Introduction. Studies in History and Philosophy of Biological and Biomedical Sciences, 36 (2):233–244, June 2005. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848605000178.

Crowe:2015:DDB

[CDA^{+}15] Nathan Crowe, Michael R. Dietrich, Beverly S. Alomepe, Amelia F. Antrim, Bay Lauris ByrneSim, and Yi He. The diversification of developmental biology. Studies in History and Philosophy of Biological and Biomedical Sciences, 53(??):1–15, ???? 2015. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848615000667.

Calvert:2011:CLD

[CF11] Jane Calvert and Joan H. Fujimura. Calculating life? Duelling discourses in interdisciplinary systems biology. Studies in History and Philosophy of Biological and Biomedical Sciences, 42(2):155– 163, June 2011. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848610001159.

Chahrour:2007:CMA

[Cha07] Marcel Chahrour. ‘A civilizing mission’? Austrian medicine and the reform of medical structures in the Ottoman Empire, 1838–1850. Studies in History and Philosophy of Biological and Biomedical Sciences, 38(4):687–705, December 2007. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848607000490.

Charbonneau:2013:CLM

[Cha13] Mathieu Charbonneau. The cognitive life of mechanical molecular models. Studies in History and Philosophy of Biological and Biomedical Sciences, 44(4):585–594, December 2013. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848613000952.

Charnley:2016:PPH

[Cha16] Berris Charnley. Plasmids, patents and the historian. Studies in History and Philosophy of Biological and Biomedical Sciences, 60(??):109–113, December 2016. CODEN ???? ISSN 1369-8486 (print), 1879-2499 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S1369848616300462.

Chimisso:2003:TPN

[Chi03] Cristina Chimisso. The tribunal of philosophy and its norms: history and philosophy in Georges Canguilhem’s historical epistemology. Studies in History and Philosophy of Biological and Biomedical Sciences, 34(2):297–327, June 2003. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S136984860300027X.

Chiang:2009:RSH

[Chi09] Howard H. Chiang. Rethinking ‘style’ for historians and philosophers of science: converging lessons from sexuality, translation, and East Asian studies. Studies in History and Philosophy of Biological and Biomedical Sciences, 40(2):109–118, June 2009. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S136984860900017X.

Chirimuuta:2019:SCH

[Chi19] M. Chirimuuta. Synthesis of contraries: Hughlings Jackson on sensory-motor representation in the brain. Studies in History and Philosophy of Biological and Biomedical Sciences, 75(??):34–44, June 2019. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848618300359.

Christensen:2012:NSN

[Chr12] Wayne Christensen. Natural sources of normativity. Studies in History and Philosophy of Biological and Biomedical Sciences, 43 (1):104–112, March 2012. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848611000458.

Chung:2003:OTP

[Chu03] Carl Chung. On the origin of the typological/population distinction in Ernst Mayr’s changing views of species, 1942–1959.

Studies in History and Philosophy of Biological and Biomedical Sciences, 34(2):277–296, June 2003. CODEN ???? ISSN 1369-8486 (print), 1879-2499 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S1369848603000268.

Cambrosio:2000:LPT

[CK00] Alberto Cambrosio and Peter Keating. Of lymphocytes and pixels: The techno-visual production of cell populations. Studies in History and Philosophy of Biological and Biomedical Sciences, 31 (2):233–270, June 2000. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848699000370.

Clark:1998:CBE

[Cla98] J. F. M. Clark. ‘the complete biography of every animal’: ants, bees, and humanity in nineteenth-century England. Studies in History and Philosophy of Biological and Biomedical Sciences, 29(2):249–267, September 1998. CO-DEN ???? ISSN 1369-8486 (print), 1879-2499 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S1369848698000132.

Clarke:2006:ASD

[Cla06] Ellen Clarke. Anarchy, socialism and a Darwinian left. Studies in History and Philosophy of Biological and Biomedical Sciences, 37 (1):136–150, March 2006. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848605001019.

Clarke:2007:RRS

[Cla07a] Adele E. Clarke. Reflections on the reproductive sciences in agriculture in the UK and US, ca. 1900–2000+. Studies in History and Philosophy of Biological and Biomedical Sciences, 38(2):316– 339, June 2007. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S136984860700012X.

Clarke:2007:HTS

[Cla07b] John Clarke. The history of three scientific societies: the Society for the Study of Fertility (now the Society for Reproduction and Fertility) (Britain), the Société Française pour l’Étude de la Fertilité, and the Society for the Study of Reproduction (USA). Studies in History and Philosophy of Biological and Biomedical Sciences, 38(2):340–357, June 2007. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848607000131.

Claveau:2012:RWT

[Cla12] François Claveau. The Russo–Williamson Theses in the social sciences: Causal inference drawing on two types of evidence. Studies in History and Philosophy of Biological and Biomedical Sciences, 43(4):806–813, December 2012. CODEN ???? ISSN 1369-8486 (print), 1879-2499 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S136984861200060X.

Clarke:2014:MSF

[Cla14a] Brendan Clarke. Making sense of failure. Studies in History and Philosophy of Biological and Biomedical Sciences, 48 (part A)(??):122–125, December 2014. CODEN ???? ISSN 1369-8486 (print), 1879-2499 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S1369848614001228.

Clarke:2014:MMB

[Cla14b] Brendan Clarke. Mapping the methodologies of Burkitt lymphoma. Studies in History and Philosophy of Biological and Biomedical Sciences, 48 (part B)(??):210–217, December 2014. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848614001216.

Clatterbuck:2016:DHM

[Cla16] Hayley Clatterbuck. Darwin, Hume, Morgan, and the verae causae of psychology. Studies in History and Philosophy of Biological and Biomedical Sciences, 60(??):1–14, December 2016. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848616300668.

Cleland:2007:EIS

[Cle07] Carol E. Cleland. Epistemological issues in the study of microbial life: alternative terran biospheres? Studies in History and Philosophy of Biological and Biomedical Sciences, 38(4):847–861, December 2007. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848607000519.

Cleminson:2008:ESA

[Cle08] Richard Cleminson. Eugenics without the state: anarchism in Catalonia, 1900–1937. Studies in History and Philosophy of Biological and Biomedical Sciences, 39(2):232–239, June 2008. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848608000204.

Clericuzio:2012:CMT

[Cle12] Antonio Clericuzio. Chemical and mechanical theories of digestion in early modern medicine. Studies in History and Philosophy of Biological and Biomedical Sciences, 43(2):329–337, June 2012. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848611001178.

Clough:2002:WMP

[Clo02] Sharyn Clough. What is menstruation for? On the projectibility of functional predicates in menstruation research. Studies in History and Philosophy of Biological and Biomedical Sciences, 33(4): 719–732, December 2002. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848602000213.

Cobb:2015:ERF

[Cob15] Matthew Cobb. Essay review: The forgotten man of DNA. The Man in the Monkeynut Coat: William Astbury and the Forgotten Road to the Double-Helix, Kersten T. Hall. Oxford University Press, Oxford (2014). 256 pp. Price £18.99, hardback, ISBN 978-0-19-870459-1. Studies in History and Philosophy of Biological and Biomedical Sciences, 51(??):67–69, June 2015. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848615000576.

Coen:2012:CCP

[Coe12] Enrico Coen. Cells to civilizations: the principles of change that shape life. Princeton University Press, Princeton, NJ, USA, 2012. ISBN 0-691-14967-4 (hardcover), 1-4008-4165-8 (e-book). LCCN ????

Cohen:2006:KEM

[Coh06] Alix A. Cohen. Kant on epigenesis, monogenesis and human nature: The biological premises of anthropology. Studies in History and Philosophy of Biological and Biomedical Sciences, 37(4): 675–693, December 2006. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848606000719.

Cohen:2017:HNI

[Coh17] Claudine Cohen. “How nationality influences Opinion”: Darwinism and palaeontology in France (1859-1914). Studies in History and Philosophy of Biological and Biomedical Sciences, 66(??): 8–17, December 2017. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848617300717.

Cole:2013:FCE

[Col13] Simon A. Cole. Forensic culture as epistemic culture: the sociology of forensic science. Studies in History and Philosophy of Biological and Biomedical Sciences, 44(1):36–46, March 2013. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848612000970.

Colombo:2014:DBR

[Col14] Matteo Colombo. Deep and beautiful. The reward prediction error hypothesis of dopamine. Studies in History and Philosophy of Biological and Biomedical Sciences, 45(??):57–67, March 2014. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848613001313.

Collins:2015:CTE

[Col15] James P. Collins. Commentary: Tempo of evolutionary change in ecological systems. Studies in History and Philosophy of Biological and Biomedical Sciences, 50(??):80–82, April 2015. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848615000096.

Comfort:2009:PMO

[Com09] Nathaniel Comfort. The prisoner as model organism: malaria research at Stateville Penitentiary. Studies in History and Philosophy of Biological and Biomedical Sciences, 40(3):190–203, September 2009. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848609000363.

Conix:2019:RPC

[Con19] Stijn Conix. Radical pluralism, classificatory norms and the legitimacy of species classifications. Studies in History and Philosophy of Biological and Biomedical Sciences, 73(??):27–34, February 2019. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848617301681.

Cooper:2002:D

[Coo02] Rachel Cooper. Disease. Studies in History and Philosophy of Biological and Biomedical Sciences, 33(2):263–282, July 2002. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368102000183.

Cooper:2010:CBS

[Coo10] Rachel Cooper. Are culture-bound syndromes as real as universally-occurring disorders? Studies in History and Philosophy of Biological and Biomedical Sciences, 41(4):325–332, December 2010. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848610000658.

Cook:2012:CTD

[Coo12] William J. Cook. The correspondence of Thomas Dale (1700– 1750): Botany in the transatlantic Republic of Letters. Studies in History and Philosophy of Biological and Biomedical Sciences, 43 (1):232–243, March 2012. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848611001129.

Cooper:2015:WBS

[Coo15] Rachel Cooper. Why is the Diagnostic and Statistical Manual of Mental Disorders so hard to revise? Path-dependence and “lock-in” in classification. Studies in History and Philosophy of Biological and Biomedical Sciences, 51(??):1–10, June 2015. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848615000400.

Cooper:2019:LNP

[Coo19] Andrew Cooper. Living natural products in Kant’s physical geography. Studies in History and Philosophy of Biological and Biomedical Sciences, 78(??):Article 101191, December 2019. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848619300299.

Cowie:2009:PVS

[Cow09] Helen Cowie. Peripheral vision: science and creole patriotism in eighteenth-century Spanish America. Studies in History and Philosophy of Biological and Biomedical Sciences, 40(3):143– 155, September 2009. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848609000375.

Cramer:2004:EOL

[Cra04] J. S. Cramer. The early origins of the logit model. Studies in History and Philosophy of Biological and Biomedical Sciences, 35(4): 613–626, December 2004. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848604000676.

Craver:2005:BRM

[Cra05] Carl F. Craver. Beyond reduction: mechanisms, multifield integration and the unity of neuroscience. Studies in History and Philosophy of Biological and Biomedical Sciences, 36(2):373–395, June 2005. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848605000245.

Crawford:2018:EHB

[Cra18] Matthew James Crawford. Escaping the historiographical blackmail of modernity: the history of nature and knowledge in Tokugawa Japan. Studies in History and Philosophy of Biological and Biomedical Sciences, 70(??):29–32, August 2018. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL https://www.sciencedirect.com/science/
article/pii/S1369848618300682.

Creager:1999:WBT

[Cre99] Angela N. H. Creager. ‘what blood told Dr Cohn’: World War II, plasma fractionation, and the growth of human blood research. Studies in History and Philosophy of Biological and Biomedical Sciences, 30(3):377–405, September 1999. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848699000175.

Creager:2002:TPC

[Cre02] Angela N. H. Creager. Tracing the politics of changing postwar research practices: the export of ‘American’ radioisotopes to European biologists. Studies in History and Philosophy of Biological and Biomedical Sciences, 33(3):367–388, September 2002. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848602000109.

Creager:2007:ASO

[Cre07] Angela N. H. Creager. Adaptation or selection? Old issues and new stakes in the postwar debates over bacterial drug resistance. Studies in History and Philosophy of Biological and Biomedical Sciences, 38(1):159–190, March 2007. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848606000963.

Cresto:2008:SBE

[Cre08] Eleonora Cresto. In search of the best explanation about the nature of the gene: Avery on pneumococcal transformation. Studies in History and Philosophy of Biological and Biomedical Sciences, 39(1):65–79, March 2008. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848607000908.

Creager:2009:PPG

[Cre09] Angela N. H. Creager. Phosphorus-32 in the Phage Group: radioisotopes as historical tracers of molecular biology. Studies in History and Philosophy of Biological and Biomedical Sciences, 40 (1):29–42, March 2009. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848608000939.

Creager:2014:HEA

[Cre14] Angela N. H. Creager. “Happily ever after” for cancer viruses? Studies in History and Philosophy of Biological and Biomedical Sciences, 48 (part B)(??):260–262, December 2014. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848614001381.

Crozier:2008:HEE

[Cro08] Ivan Crozier. Havelock Ellis, eugenicist. Studies in History and Philosophy of Biological and Biomedical Sciences, 39(2):187–194, June 2008. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848608000150.

Cunningham:2002:PSR

[Cun02] Andrew Cunningham. The pen and the sword: recovering the disciplinary identity of physiology and anatomy before 1800: I: Old physiology — the pen. Studies in History and Philosophy of Biological and Biomedical Sciences, 33(4):631–665, December 2002. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848602000237.

Cunningham:2003:PSR

[Cun03] Andrew Cunningham. The pen and the sword: recovering the disciplinary identity of physiology and anatomy before 1800: II: Old anatomy — the sword. Studies in History and Philosophy of Biological and Biomedical Sciences, 34(1):51–76, March 2003. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848602000699.

Curry:2017:EEP

[Cur17] Helen Anne Curry. Extension and experiment: the politics of modern agricultural science. Studies in History and Philosophy of Biological and Biomedical Sciences, 63(??):80–84, June 2017. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848617300262.

Corrigan:2006:PBP

[CWJ06] Oonagh P. Corrigan and Bryn Williams-Jones. Pharmacogenetics: the bioethical problem of DNA investment banking. Studies in History and Philosophy of Biological and Biomedical Sciences, 37(3):550–565, September 2006. CO-DEN ???? ISSN 1369-8486 (print), 1879-2499 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S1369848606000501.

Dacome:2012:BAP

[Dac12] Lucia Dacome. Balancing acts: Picturing perspiration in the long eighteenth century. Studies in History and Philosophy of Biological and Biomedical Sciences, 43(2):379–391, June 2012. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848611001221.

Dietrich:2020:HCY

[DAC^{+}20] Michael R. Dietrich, Rachel A. Ankeny, Nathan Crowe, Sara Green, and Sabina Leonelli. How to choose your research organism. Studies in History and Philosophy of Biological and Biomedical Sciences, 80(??):Article 101227, April 2020. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848619301165.

Dagg:2017:HCR

[Dag17] Joachim L. Dagg. How counterfactuals of Red-Queen theory shed light on science and its historiography. Studies in History and Philosophy of Biological and Biomedical Sciences, 64 (??):53–64, August 2017. CODEN ???? ISSN 1369-8486 (print), 1879-2499 (electronic). URL /science/article/pii/

S1369848616300978.

Dagg:2019:MMC

[Dag19] Joachim L. Dagg. Motives and merits of counterfactual histories of science. Studies in History and Philosophy of Biological and Biomedical Sciences, 73(??):19–26, February 2019. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848618300049.

Darwin:1859:OSM

[Dar59] Charles Darwin. On the origin of species by means of natural selection, or preservation of favoured races in the struggle for life. John Murray, London, UK, 1859. ix + 1 + 502 + 32 pp.

Darwin:1990:OSM

[Dar90] Charles Darwin. On the origin of the species: by means of natural selection: or the preservation of favoured races in the struggle for life. Folio Society, London, UK, 1990. ???? pp.

Darwin:1998:OS

[Dar98] Charles Darwin. On the Origin of Species, volume 1228. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782,

USA, 1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext98/otoof10.zip.

Darden:2005:RAF

[Dar05] Lindley Darden. Relations among fields: Mendelian, cytological and molecular mechanisms. Studies in History and Philosophy of Biological and Biomedical Sciences, 36(2):349–371, June 2005. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848605000233.

Darwin:2008:OSM

[Dar08] Charles Darwin. On the origin of species: by means of natural selection. Arcturus, London, UK, 2008. ISBN 1-84193-989-7. 384 pp. LCCN 30.0/483; alnynat. With a preface by George Davidson.

Davis:1999:QAR

[Dav99] Claire J. Davis. The question of abortion in revolutionary Russia, 1905–1920. Studies in History and Philosophy of Biological and Biomedical Sciences, 30(1):45–67, March 1999. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848698000144.

Davis:2004:DTB

[Dav04] Sarah Davis. Darwin, Tegetmeier and the bees. Studies in History and Philosophy of Biological and Biomedical Sciences, 35 (1):65–92, March 2004. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848603001080.

Davis:2011:MNH

[Dav11] Sophia Davis. Militarised natural history: Tales of the avocet’s return to postwar Britain. Studies in History and Philosophy of Biological and Biomedical Sciences, 42(2):226–232, June 2011. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848610001202.

Davis:2014:WWG

[Dav14a] Angela Davis. Wartime women giving birth: Narratives of pregnancy and childbirth, Britain c. 1939–1960. Studies in History and Philosophy of Biological and Biomedical Sciences, 47 (part B)(??):257–266, September 2014. CO-DEN ???? ISSN 1369-8486 (print), 1879-2499 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S1369848613001696.

Davis:2014:BNH

[Dav14b] Frederick R. Davis. Biography, natural history and early America. Studies in History and Philosophy of Biological and Biomedical Sciences, 46(??):121–124, June 2014. CO-DEN ???? ISSN 1369-8486 (print), 1879-2499 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S1369848614000077.

Dawson:2014:DD

[Daw14] Gowan Dawson. Darwin decentred. Studies in History and Philosophy of Biological and Biomedical Sciences, 46(??):93–96, June 2014. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848614000041.

deChadarevian:1998:WPC

[dC98] Soraya de Chadarevian. Of worms and programmes: Caenorhabditis elegans and the study of development. Studies in History and Philosophy of Biological and Biomedical Sciences, 29 (1):81–105, March 1998. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848698000041.

daCosta:2001:NHF

[dC01] Palmira Fontes da Costa. The natural history files. Studies in History and Philosophy of Biological and Biomedical Sciences, 32(3):583–587, September 2001. CODEN ???? ISSN 1369-8486 (print), 1879-2499 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S136984860100022X.

Darden:2002:SID

[DC02a] Lindley Darden and Carl Craver. Strategies in the interfield discovery of the mechanism of protein synthesis. Studies in History and Philosophy of Biological and Biomedical Sciences, 33 (1):1–28, March 2002. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848601000218.

deChadarevian:2002:RLM

[dC02b] Soraya de Chadarevian. Reconstructing life. molecular biology in postwar Britain. Studies in History and Philosophy of Biological and Biomedical Sciences, 33(3):431–448, September 2002. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848602000080.

deChadarevian:2009:ISB

[dC09a] Soraya de Chadarevian. Interview with Sydney Brenner. Studies in History and Philosophy of Biological and Biomedical Sciences, 40(1):65–71, March 2009. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848608000964.

deChadarevian:2009:MVB

[dC09b] Soraya de Chadarevian. Microstudies versus big picture accounts? Studies in History and Philosophy of Biological and Biomedical Sciences, 40(1):13–19, March 2009. CODEN ???? ISSN 1369-8486 (print), 1879-2499 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S1369848608000915.

deChadarevian:2014:CSH

[dC14] Soraya de Chadarevian. Chromosome surveys of human populations: Between epidemiology and anthropology. Studies in History and Philosophy of Biological and Biomedical Sciences, 47 (part A)(??):87–96, September 2014. CO-DEN ???? ISSN 1369-8486 (print), 1879-2499 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S1369848614000491.

deChadarevian:2016:FHR

[dC16] Soraya de Chadarevian. The future historian: Reflections on the archives of contemporary sciences. Studies in History and Philosophy of Biological and Biomedical Sciences, 55(??):54–60, February 2016. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848615001120.

deChadarevian:2009:I

[dCR09] Soraya de Chadarevian and Hans-Jörg Rheinberger. Introduction. Studies in History and Philosophy of Biological and Biomedical Sciences, 40(1):4–5, March 2009. CO-DEN ???? ISSN 1369-8486 (print), 1879-2499 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S1369848608000897.

deChadarevian:2002:MBP

[dCS02] Soraya de Chadarevian and Bruno Strasser. Molecular biology in postwar Europe: towards a ‘glocal’ picture. Studies in History and Philosophy of Biological and Biomedical Sciences, 33(3):361– 365, September 2002. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848602000092.

DeBlock:2008:WMD

[De 08a] Andreas De Block. Why mental disorders are just mental dysfunctions (and nothing more): some Darwinian arguments. Studies in History and Philosophy of Biological and Biomedical Sciences, 39(3):338–346, September 2008. CODEN ???? ISSN 1369-8486 (print), 1879-2499 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S1369848608000423.

DeVreese:2008:CMU

[De 08b] Leen De Vreese. Causal (mis)understanding and the search for scientific explanations: a case study from the history of medicine. Studies in History and Philosophy of Biological and Biomedical Sciences, 39(1):14–24, March 2008.

CODEN ???? ISSN 1369-8486 (print), 1879-2499 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S1369848607000866.

DeGrandis:2016:PIA

[De 16] Giovanni De Grandis. Practical integration: the art of balancing values, institutions and knowledge — lessons from the History of British Public Health and Town Planning. Studies in History and Philosophy of Biological and Biomedical Sciences, 56(??):92–105, April 2016. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848615001429.

Cesare:2019:DOT

[De 19] Silvia De Cesare. Disentangling organic and technological progress: an epistemological clarification introducing a key distinction between two levels of axiology. Studies in History and Philosophy of Biological and Biomedical Sciences, 73(??):44–53, February 2019. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848617301917.

DeGrandis:2016:IGC

[DE16] Giovanni De Grandis and Sophia Efstathiou. Introduction — Grand Challenges and small steps. Studies in History and Philosophy of Biological and Biomedical Sciences, 56(??):39–47, April 2016. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848615001703.

Dean:2010:MWW

[Dea10] Sheila Ann Dean. The man who would be king of botanical classification: Imperial nature: Joseph Hooker and the practices of Victorian science Jim Endersby; University of Chicago Press, Chicago, 2008, pp. 400, Price £18.00 US$35.00 hardback, ISBN 0-226-20791-9. Studies in History and Philosophy of Biological and Biomedical Sciences, 41(3):300–303, September 2010. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848610000476. See [End08].

Deichmann:2002:EIS

[Dei02] Ute Deichmann. Emigration, isolation and the slow start of molecular biology in Germany. Studies in History and Philosophy of Biological and Biomedical Sciences, 33(3):449–471, September 2002. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848602000110.

Deichmann:2016:WEV

[Dei16] Ute Deichmann. Why epigenetics is not a vindication of Lamarckism — and why that matters. Studies in History and Philosophy of Biological and Biomedical Sciences, 57(??):80–82, June 2016. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848616300206.

Delisle:2000:BCL

[Del00] Richard G. Delisle. The biology/culture link in human evolution, 1750–1950: the problem of integration in science. Studies in History and Philosophy of Biological and Biomedical Sciences, 31(4): 531–556, December 2000. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848600000029.

Delord:2007:NE

[Del07] Julien Delord. The nature of extinction. Studies in History and Philosophy of Biological and Biomedical Sciences, 38(3):656– 667, September 2007. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848607000350.

Delisle:2009:UFN

[Del09] Richard G. Delisle. The uncertain foundation of neo-darwinism: metaphysical and epistemological pluralism in the evolutionary synthesis. Studies in History and Philosophy of Biological and Biomedical Sciences, 40(2):119–132, June 2009. CODEN ???? ISSN 1369-8486 (print), 1879-2499 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S1369848609000181.

Delisle:2011:FCC

[Del11a] Richard G. Delisle. Foreword: Celebrating Charles Darwin in disagreement. Studies in History and Philosophy of Biological and Biomedical Sciences, 42(1):1, March 2011. CODEN ???? ISSN 1369-8486 (print), 1879-2499 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S1369848610001081.

Delisle:2011:WWR

[Del11b] Richard G. Delisle. What was really synthesized during the evolutionary synthesis? A historiographic proposal. Studies in History and Philosophy of Biological and Biomedical Sciences, 42 (1):50–59, March 2011. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848610000981.

Delorme:2014:PPP

[Del14] Shannon Delorme. Physiology or psychic powers? William Carpenter and the debate over spiritualism in Victorian Britain. Studies in History and Philosophy of Biological and Biomedical Sciences, 48 (part A)(??):57–66, December 2014. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848614000843.

Dennett:2004:EPC

[Den04] Daniel C. Dennett. An evolutionary perspective on cognition: through a glass lightly. Studies in History and Philosophy of Biological and Biomedical Sciences, 35(4):721–727, December 2004. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S136984860400072X.

Depew:2008:CEB

[Dep08] David J. Depew. Consequence etiology and biological teleology in Aristotle and Darwin. Studies in History and Philosophy of Biological and Biomedical Sciences, 39(4):379–390, December 2008. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848608000733.

Depew:2011:APF

[Dep11] David J. Depew. Adaptation as process: the future of Darwinism and the legacy of Theodosius Dobzhansky. Studies in History and Philosophy of Biological and Biomedical Sciences, 42(1):89–98, March 2011. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848610000993.

Derry:2020:IMD

[Der20] Margaret Derry. The invention of the modern dog by Michael Worboys, Julie-Marie strange, and Neil Pemberton: the commercialization of breeding for beauty. Studies in History and Philosophy of Biological and Biomedical Sciences, 84(??):Article 101332, December 2020. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S1369848620301436. See [Ank20,

WSP20].

Desmond:1998:HDD

[Des98] Adrian J. Desmond. Huxley: from devil’s disciple to evolution’s high priest. Penguin, London, UK, 1998. ISBN 0-14-017309-9. xxii + 820 pp. LCCN QH31.H9 D47 1997.

DesChene:2005:MLS

[Des05] Dennis Des Chene. Mechanisms of life in the seventeenth century: Borelli, Perrault, Régis. Studies in History and Philosophy of Biological and Biomedical Sciences, 36(2):245–260, June 2005. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S136984860500018X.

DesAutels:2016:NSM

[Des16] Lane DesAutels. Natural selection and mechanistic regularity. Studies in History and Philosophy of Biological and Biomedical Sciences, 57(??):13–23, June 2016. CO-DEN ???? ISSN 1369-8486 (print), 1879-2499 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S1369848616000182.

Desmond:2018:NSP

[Des18] Hugh Desmond. Natural selection, plasticity, and the rationale for largest-scale trends. Studies in History and Philosophy of Biological and Biomedical Sciences, 68–69(??):15–24, April/ June 2018. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL https://www.sciencedirect.com/science/
article/pii/S1369848617300961.

Dettelbach:1999:FNP

[Det99] Michael Dettelbach. The face of nature: Precise measurement, mapping, and sensibility in the work of Alexander von Humboldt. Studies in History and Philosophy of Biological and Biomedical Sciences, 30(4):473–504, December 1999. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848699000114.

Devitt:2018:HBE

[Dev18] Michael Devitt. Historical biological essentialism. Studies in History and Philosophy of Biological and Biomedical Sciences, 71 (??):1–7, October 2018. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848617301887.

Devanesan:2020:MNL

[Dev20] Arjun Devanesan. Medical nihilism: the limits of a decontextualised critique of medicine. Studies in History and Philosophy of Biological and Biomedical Sciences, 79(??):Article 101189, February 2020. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S136984861930010X.

Dupre:2016:VLP

[DG16] John Dupré and Stephan Guttinger. Viruses as living processes. Studies in History and Philosophy of Biological and Biomedical Sciences, 59(??):109–116, October 2016. CODEN ???? ISSN 1369-8486 (print), 1879-2499 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S1369848616300036.

Dietrich:2016:PP

[Die16] Michael R. Dietrich. Parsing postgenomics. Studies in History and Philosophy of Biological and Biomedical Sciences, 59(??): 158–160, October 2016. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848616300164.

DiFrisco:2017:FEP

[DiF17] James DiFrisco. Functional explanation and the problem of functional equivalence. Studies in History and Philosophy of Biological and Biomedical Sciences, 65(??):1–8, October 2017. CODEN ???? ISSN 1369-8486 (print), 1879-2499 (electronic). URL /

science/article/pii/S136984861630111X.

DiFrisco:2019:IEP

[DiF19] James DiFrisco. Interdisciplinarity, epistemic pluralism, and unificationism. Studies in History and Philosophy of Biological and Biomedical Sciences, 74(??):40–44, April 2019. CODEN ???? ISSN 1369-8486 (print), 1879-2499 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S1369848619300214.

Dilley:2013:NBM

[Dil13] Stephen Dilley. Nothing in biology makes sense except in light of theology? Studies in History and Philosophy of Biological and Biomedical Sciences, 44(4):774–786, December 2013. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848613000940.

Dimech:2017:MWN

[Dim17] Dominic K. Dimech. Modelling with words: Narrative and natural selection. Studies in History and Philosophy of Biological and Biomedical Sciences, 62(??):20–24, April 2017. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848617300213.

DiMoia:2018:BRN

[DiM18] John P. DiMoia. Book review: Naming the Local: Medicine, Language, and Identity in Korea since the Fifteenth Century, Soyoung Suh. Harvard University Press, Cambridge MA (2017), 244 pp. Price $39.95 hardcover, ISBN: 978-0-674-97696-2. Studies in History and Philosophy of Biological and Biomedical Sciences, 67(??):24–27, February 2018. CODEN ???? ISSN 1369-8486 (print), 1879-2499 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S136984861730170X.

DiMarco:2020:RPM

[DiM20] Marina DiMarco. (re)producing mtEve. Studies in History and Philosophy of Biological and Biomedical Sciences, 83(??):Article 101290, October 2020. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848620300303.

Dixon:2002:GGE

[Dix02] Thomas Dixon. The genetic gods: evolution and belief in human affairs. Studies in History and Philosophy of Biological and Biomedical Sciences, 33(2):337–359, July 2002. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368102000146.

Datteri:2012:MTP

[DL12] Edoardo Datteri and Federico Laudisa. Model testing, prediction and experimental protocols in neuroscience: a case study. Studies in History and Philosophy of Biological and Biomedical Sciences, 43(3):602–610, September 2012. CO-DEN ???? ISSN 1369-8486 (print), 1879-2499 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S1369848612000556.

DeSio:2014:CHH

[DM14] Fabio De Sio and Chantal Marazia. Clever Hans and his effects: Karl Krall and the origins of experimental parapsychology in Germany. Studies in History and Philosophy of Biological and Biomedical Sciences, 48 (part A)(??):94–102, December 2014. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848614000855.

deMelo-Martin:2003:BES

[dMM03] Inmaculada de Melo-Martín. Biological explanations and social responsibility. Studies in History and Philosophy of Biological and Biomedical Sciences, 34(2):345–358, June 2003. CODEN ???? ISSN 1369-8486 (print), 1879-2499 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S1369848602000705.

Dupre:2007:MBO

[DO07] John Dupré and Maureen A. O’Malley. Metagenomics and biological ontology. Studies in History and Philosophy of Biological and Biomedical Sciences, 38(4):834–846, December 2007. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848607000453.

Dolgin:2001:IDP

[Dol01] Janet L. Dolgin. Ideologies of discrimination: personhood and the ‘genetic group’. Studies in History and Philosophy of Biological and Biomedical Sciences, 32(4):705–721, December 2001. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848601000292.

Domondon:2006:BPB

[Dom06] Andrew T. Domondon. Bringing physics to bear on the phenomenon of life: the divergent positions of Bohr, Delbrück, and Schrödinger. Studies in History and Philosophy of Biological and Biomedical Sciences, 37(3):433–458, September 2006. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848606000434.

Donaghy:2014:TDS

[Don14] Josephine Donaghy. Temporal decomposition: a strategy for building mathematical models of complex metabolic systems. Studies in History and Philosophy of Biological and Biomedical Sciences, 48 (part A)(??):1–11, December 2014. CODEN ???? ISSN 1369-8486 (print), 1879-2499 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S1369848614000909.

Donovan:2015:PHB

[Don15] Brian M. Donovan. Putting humanity back into the teaching of human biology. Studies in History and Philosophy of Biological and Biomedical Sciences, 52(??):65–75, August 2015. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848615000138.

Donhauser:2016:TEE

[Don16] Justin Donhauser. Theoretical ecology as etiological from the start. Studies in History and Philosophy of Biological and Biomedical Sciences, 60(??):67–76, December 2016. CODEN ???? ISSN 1369-8486 (print), 1879-2499 (elec-

tronic).  URL http://www.sciencedirect.com/science/

article/pii/S1369848616300723. See reply [Sag17] and re-

sponse [Don17].

Donhauser:2017:DDT

[Don17] Justin Donhauser. Differentiating and defusing theoretical ecology’s criticisms: a rejoinder to Sagoff’s reply to Donhauser (2016). Studies in History and Philosophy of Biological and Biomedical Sciences, 63(??):70–79, June 2017. CODEN ???? ISSN 1369-8486 (print), 1879-2499 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S1369848617300481. See [Don16, Sag17].

Donohue:2020:SBB

[Don20] Christopher Donohue. Social borrowings and biological appropriations: Special issue introduction. Studies in History and Philosophy of Biological and Biomedical Sciences, 83(??):Article 101309, October 2020. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S136984862030114X.

Douglas:2013:CAL

[DPS13] Thomas Douglas, Russell Powell, and Julian Savulescu. Is the creation of artificial life morally significant? Studies in History and Philosophy of Biological and Biomedical Sciences, 44(4):688– 696, December 2013. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848613000848.

Dreger:1998:LIR

[Dre98] Alice Domurat Dreger. The limits of individuality: Ritual and sacrifice in the lives and medical treatment of conjoined twins. Studies in History and Philosophy of Biological and Biomedical Sciences, 29(1):1–29, March 1998. CODEN ???? ISSN 1369-8486 (print), 1879-2499 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S1369848698000028.

Dresow:2019:MEP

[Dre19] Max Dresow. Macroevolution evolving: Punctuated equilibria and the roots of Stephen Jay Gould’s second macroevolutionary synthesis. Studies in History and Philosophy of Biological and Biomedical Sciences, 75(??):15–23, June 2019.

CODEN ???? ISSN 1369-8486 (print), 1879-2499 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S1369848617302212.

Driscoll:2009:OBB

[Dri09] Catherine Driscoll. On our best behavior: optimality models in human behavioral ecology. Studies in History and Philosophy of Biological and Biomedical Sciences, 40(2):133–141, June 2009. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848609000193.

Drouin:2001:PUT

[Dro01] Jean-Marc Drouin. Principles and uses of taxonomy in the works of Augustin-Pyramus de Candolle. Studies in History and Philosophy of Biological and Biomedical Sciences, 32(2):255–275, June 2001. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848601000024.

Droscher:2011:CDC

[Drö11] Ariane Dröscher. Cellular dimensions and cell dynamics, or the difficulty over capturing time and space in the era of electron microscopy. Studies in History and Philosophy of Biological and Biomedical Sciences, 42(4):395–402, December 2011. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848611000689.

Drouet:2012:CRC

[Dro12] Isabelle Drouet. Causal reasoning, causal probabilities, and conceptions of causation. Studies in History and Philosophy of Biological and Biomedical Sciences, 43(4):761–768, December 2012. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848612000660.

deSouza:2014:EHP

[dSV14] Vanderlei Sebastião de Souza and Ricardo Ventura Santos. The emergence of human population genetics and narratives about the formation of the Brazilian nation (1950– 1960). Studies in History and Philosophy of Biological and Biomedical Sciences, 47 (part A)(??):97–107, September 2014. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848614000508.

Dupre:2001:DC

[Dup01] John Dupré. In defence of classification. Studies in History and Philosophy of Biological and Biomedical Sciences, 32(2):203–219, June 2001. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848601000036.

Dupre:2012:PLE

[Dup12] John Dupré. Processes of life: essays in the philosophy of biology. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2012. ISBN 0-19-969198-3. viii + 350 pp.

LCCN QH331 .D85 2012. URL http://www.loc.gov/catdir/
enhancements/fy1215/2011944132-b.html;   http://www.
loc.gov/catdir/enhancements/fy1215/2011944132-d.html;
 http://www.loc.gov/catdir/enhancements/fy1215/2011944132-
t.html.

Dussault:2018:FEN

[Dus18] Antoine C. Dussault. Functional ecology’s non-selectionist understanding of function. Studies in History and Philosophy of Biological and Biomedical Sciences, 70(??):ii, August 2018. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL https://www.sciencedirect.com/science/
article/pii/S1369848617301577.

Dutfield:2018:IAG

[Dut18] Graham Dutfield. Innovation, agrobiodiversity, and the global nature of national agriculture. Studies in History and Philosophy of Biological and Biomedical Sciences, 70(??):33–35, August 2018. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL https://www.sciencedirect.com/science/
article/pii/S1369848618300724.

Dyde:2016:WMM

[Dyd16] Sean Dyde. Where is my mind? Studies in History and Philosophy of Biological and Biomedical Sciences, 60(??):105–108, December 2016. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848616300449.

Easton:2011:CDF

[Eas11] Patricia Easton. The Cartesian doctor, François Bayle (1622– 1709), on psychosomatic explanation. Studies in History and Philosophy of Biological and Biomedical Sciences, 42(2):203–209, June 2011. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S136984861000124X.

Eddy:2006:MSN

[Edd06] M. D. Eddy. The medium of signs: nominalism, language and the philosophy of mind in the early thought of Dugald Stewart. Studies in History and Philosophy of Biological and Biomedical Sciences, 37(3):373–393, September 2006. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848606000409.

Eddy:2008:AMR

[Edd08] M. D. Eddy. ‘an adept in medicine’: the Reverend Dr William Laing, nervous complaints and the commodification of spa water. Studies in History and Philosophy of Biological and Biomedical Sciences, 39(1):1–13, March 2008. CODEN ???? ISSN 1369-8486 (print), 1879-2499 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S1369848607000854.

Edington:2020:BRI

[Edi20] Claire Edington. Book review: The Invention of Madness by Emily Baum: Modernity and madness in Republican China. Studies in History and Philosophy of Biological and Biomedical Sciences, 79(??):Article 101205, February 2020. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848619300998. See reply [Bau20].

Edmond:2013:JTC

[Edm13] Gary Edmond. Just truth? Carefully applying history, philosophy and sociology of science to the forensic use of CCTV images. Studies in History and Philosophy of Biological and Biomedical Sciences, 44(1):80–91, March 2013. CODEN ???? ISSN 1369-8486 (print), 1879-2499 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S1369848612000982.

Edwards:2012:PS

[Edw12] A. W. F. Edwards. Punnett’s square. Studies in History and Philosophy of Biological and Biomedical Sciences, 43(1):219–224, March 2012. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848611001373.

Edwards:2016:PSP

[Edw16] A. W. F. Edwards. Punnett’s square: a postscript. Studies in History and Philosophy of Biological and Biomedical Sciences, 57 (??):69–70, June 2016. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848616000157.

Efstathiou:2012:NCM

[Efs12] Sophia Efstathiou. The Nazi cosmetic: Medicine in the service of beauty. Studies in History and Philosophy of Biological and Biomedical Sciences, 43(3):634–642, September 2012. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848612000581.

Efstathiou:2016:IPG

[Efs16] Sophia Efstathiou. Is it possible to give scientific solutions to Grand Challenges? On the idea of grand challenges for life science research. Studies in History and Philosophy of Biological and Biomedical Sciences, 56(??):48–61, April 2016. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848615001594.

El-Haj:2012:GSS

[EH12] Nadia Abu El-Haj. The genealogical science: the search for Jewish origins and the politics of epistemology. Chicago studies in practices of meaning. University of Chicago Press, Chicago, IL, USA and London, UK, 2012. ISBN 0-226-20140-6 (hardcover: alkaline paper). ix + 311 pp. LCCN DS143 .A15 2012.

Eliot:2007:MMC

[Eli07] Christopher Eliot. Method and metaphysics in Clements’s and Gleason’s ecological explanations. Studies in History and Philosophy of Biological and Biomedical Sciences, 38(1):85–109, March 2007. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848606000938.

Elliott:2004:SPP

[Ell04] Carl Elliott. Six problems with pharma-funded bioethics. Studies in History and Philosophy of Biological and Biomedical Sciences, 35(1):125–129, March 2004. CODEN ???? ISSN 1369-8486 (print), 1879-2499 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S1369848603001122.

Ellenbogen:2008:AOE

[Ell08] Josh Ellenbogen. Authority, objectivity, evidence: scientific photography in Victorian Britain. Studies in History and Philosophy of Biological and Biomedical Sciences, 39(1):171–175, March 2008. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848607000994.

Elliot:2014:MAC

[Ell14] Rosemary Elliot. Miscarriage, abortion or criminal feticide: Understandings of early pregnancy loss in Britain, 1900–1950. Studies in History and Philosophy of Biological and Biomedical Sciences, 47 (part B)(??):248–256, September 2014. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848614000120.

Endersby:2001:RHE

[End01] Jim Endersby. ‘the realm of hard evidence’: novelty, persuasion and collaboration in botanical cladistics. Studies in History and Philosophy of Biological and Biomedical Sciences, 32(2):343–360, June 2001. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848601000115.

Endersby:2008:INJ

[End08] Jim Endersby. Imperial nature: Joseph Hooker and the practices of Victorian science. University of Chicago Press, Chicago, IL, USA and London, UK, 2008. ISBN 0-226-20791-9 (hardcover). xii + 429 pp. LCCN QK31.H75 E53 2008. URL http://www.

loc.gov/catdir/enhancements/fy0827/2007034919-b.html;

 http://www.loc.gov/catdir/enhancements/fy0827/2007034919-
d.html;   http://www.loc.gov/catdir/toc/ecip0725/2007034919.
html.

Endersby:2009:VRI

[End09] Jim Endersby. ‘the vagaries of a Rafinesque’: imagining and classifying American nature. Studies in History and Philosophy of Biological and Biomedical Sciences, 40(3):168–178, September 2009. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S136984860900034X.

England:2020:REN

[Eng20] Renee England. Rethinking emotion as a natural kind: Correctives from Spinoza and hierarchical homology. Studies in History and Philosophy of Biological and Biomedical Sciences, 84(??):Article 101327, December 2020. CO-DEN ???? ISSN 1369-8486 (print), 1879-2499 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S1369848620301382.

Edge:2015:IAH

[ER15] Michael D. Edge and Noah A. Rosenberg. Implications of the apportionment of human genetic diversity for the apportionment of human phenotypic diversity. Studies in History and Philosophy of Biological and Biomedical Sciences, 52(??):32–45, August 2015. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848615000023.

Ereshefsky:2001:NNI

[Ere01] Marc Ereshefsky. Names, numbers and indentations: a guide to post-linnaean taxonomy. Studies in History and Philosophy of Biological and Biomedical Sciences, 32(2):361–383, June 2001. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848601000048.

Ereshefsky:2009:DHD

[Ere09] Marc Ereshefsky. Defining ‘health’ and ‘disease’. Studies in History and Philosophy of Biological and Biomedical Sciences, 40(3): 221–227, September 2009. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848609000338.

Erlingsson:2009:CBR

[Erl09] Steindór J. Erlingsson. The costs of being a restless intellect: Julian Huxley’s popular and scientific career in the 1920s. Studies in History and Philosophy of Biological and Biomedical Sciences, 40(2):101–108, June 2009. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S136984860900020X.

Erslev:2018:BWK

[Ers18] Thomas Erslev. A brain worth keeping? Waste, value and time in contemporary brain banking. Studies in History and Philosophy of Biological and Biomedical Sciences, 67(??):16–23, February 2018. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848616301005.

Esposito:2011:UBE

[Esp11] Maurizio Esposito. Utopianism in the British evolutionary synthesis. Studies in History and Philosophy of Biological and Biomedical Sciences, 42(1):40–49, March 2011. CODEN ???? ISSN 1369-8486 (print), 1879-2499 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S1369848610001007.

Esposito:2017:EFR

[Esp17] Maurizio Esposito. Expectation and futurity: the remarkable success of genetic determinism. Studies in History and Philosophy of Biological and Biomedical Sciences, 62(??):1–9, April 2017. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848617300018.

Esposito:2018:LBB

[Esp18] Maurizio Esposito. From “life” to biology and backward: the long gestation of a scientific discipline. Studies in History and Philosophy of Biological and Biomedical Sciences, 70(??):20–28, August 2018. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL https://www.sciencedirect.com/science/
article/pii/S1369848618300712.

Etkind:2008:BEF

[Etk08] Alexander Etkind. Beyond eugenics: the forgotten scandal of hybridizing humans and apes. Studies in History and Philosophy of Biological and Biomedical Sciences, 39(2):205–210, June 2008. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848608000174.

Ewens:2019:QEN

[Ewe19] Warren J. Ewens. Quantifying evolution by natural selection. Studies in History and Philosophy of Biological and Biomedical Sciences, 76(??):Article 101174, August 2019. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848618301237.

Fabian:2010:SCR

[Fab10] Ann Fabian. The skull collectors: race, science, and America’s unburied dead. University of Chicago Press, Chicago, IL, USA and London, UK, 2010. ISBN 0-226-23348-0 (hardcover). xi + 270 pp. LCCN GN90 .F33 2010.

Fagan:2007:SHS

[Fag07] Melinda B. Fagan. The search for the hematopoietic stem cell: social interaction and epistemic success in immunology. Studies in History and Philosophy of Biological and Biomedical Sciences, 38 (1):217–237, March 2007. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848606000987.

Fagan:2009:FSC

[Fag09] Melinda B. Fagan. Fleck and the social constitution of scientific objectivity. Studies in History and Philosophy of Biological and Biomedical Sciences, 40(4):272–285, December 2009. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848609000545.

Falk:2010:WGR

[Fal10] Raphael Falk. What is a gene? — Revisited. Studies in History and Philosophy of Biological and Biomedical Sciences, 41(4):396– 406, December 2010. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848610000762.

Falk:2013:NGR

[Fal13] Raphael Falk. On the nature of the gene (reviewing P. R. Sloan, B. Fogel (Eds.), Creating a physical biology: the three-man paper and early molecular biology). Studies in History and Philosophy of Biological and Biomedical Sciences, 44(4):623–625, December 2013. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848613001180.

Fantuzzi:2017:CP

[Fan17] Giamila Fantuzzi. Cancer is a propagandist. Studies in History and Philosophy of Biological and Biomedical Sciences, 63(??):28– 31, June 2017. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848617300493.

Farrell:2006:RVA

[Far06] Robert P. Farrell. Rational versus anti-rational interpretations of science: an ape-language case-study. Studies in History and Philosophy of Biological and Biomedical Sciences, 37(1):83–100, March 2006. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848605000993.

Faye:2012:ASP

[Fay12] Cathy Faye. American social psychology: Examining the contours of the 1970s crisis. Studies in History and Philosophy of Biological and Biomedical Sciences, 43(2):514–521, June 2012. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848611001361.

Fleming:2015:WFD

[FB15] Leonore Fleming and Robert Brandon. Why flying dogs are rare: a general theory of luck in evolutionary transitions. Studies in History and Philosophy of Biological and Biomedical Sciences, 49(??):24–31, February 2015. CO-DEN ???? ISSN 1369-8486 (print), 1879-2499 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S1369848614001563.

Fuller:2015:PEM

[FBF15] Jonathan Fuller, Alex Broadbent, and Luis J. Flores. Prediction in epidemiology and medicine. Studies in History and Philosophy of Biological and Biomedical Sciences, 54(??):45–48, December 2015. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848615000977.

Fiorentino:2015:EIC

[FD15] Alexander R. Fiorentino and Olaf Dammann. Evidence, illness, and causation: an epidemiological perspective on the Russo–Williamson Thesis. Studies in History and Philosophy of Biological and Biomedical Sciences, 54(??):1–9, December 2015. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848615001387.

Ferry-Danini:2019:SPA

[FD19] Juliette Ferry-Danini. Should phenomenological approaches to illness be wary of naturalism? Studies in History and Philosophy of Biological and Biomedical Sciences, 73(??):10–18, February 2019. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848617302133.

Feest:2007:HEO

[Fee07] Uljana Feest. ‘hypotheses, everywhere only hypotheses!!’: on some contexts of Dilthey’s critique of explanatory psychology. Studies in History and Philosophy of Biological and Biomedical Sciences, 38(1):43–62, March 2007. CO-DEN ???? ISSN 1369-8486 (print), 1879-2499 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S1369848606000914.

Feest:2012:HCC

[Fee12] Uljana Feest. Husserl’s Crisis as a crisis of psychology. Studies in History and Philosophy of Biological and Biomedical Sciences, 43 (2):493–503, June 2012. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848611001348.

Feller:2009:DFD

[Fel09] David Allan Feller. Dog fight: Darwin as animal advocate in the antivivisection controversy of 1875. Studies in History and Philosophy of Biological and Biomedical Sciences, 40(4):265– 271, December 2009. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848609000533.

Fellowes:2017:PPH

[Fel17] Sam Fellowes. Putting the present in the history of autism. Studies in History and Philosophy of Biological and Biomedical Sciences, 61(??):54–58, February 2017. CODEN ???? ISSN 1369-8486 (print), 1879-2499 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S1369848616300954.

Fuller:2015:RGM

[FF15] Jonathan Fuller and Luis J. Flores. The Risk GP Model: the standard model of prediction in medicine. Studies in History and Philosophy of Biological and Biomedical Sciences, 54(??): 49–61, December 2015. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848615000862.

Fichman:2014:WTT

[Fic14] Martin Fichman. Wallace’s travels and theories in the Malay Archipelago. Studies in History and Philosophy of Biological and Biomedical Sciences, 47 (part A)(??):201–205, September 2014. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848614000387.

Finnegan:2008:AMH

[Fin08] Diarmid A. Finnegan. ‘an aid to mental health’: natural history, alienists and therapeutics in Victorian Scotland. Studies in History and Philosophy of Biological and Biomedical Sciences, 39(3): 326–337, September 2008. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848608000411.

Fiorentini:2011:IVA

[Fio11] Erna Fiorentini. Inducing visibilities: An attempt at Santiago Ramón y Cajal’s aesthetic epistemology. Studies in History and Philosophy of Biological and Biomedical Sciences, 42(4):391– 394, December 2011. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848611000719.

Fiorilli:2019:PSB

[Fio19] Olivia Fiorilli. Policing the social body: Medicine and the administration of legal gender recognition in France and Italy, an historical perspective. Studies in History and Philosophy of Biological and Biomedical Sciences, 78(??):Article 101182, December 2019. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848618300475.

Fisher:2015:JCW

[Fis15] Susie Fisher. Not just “a clever way to detect whether DNA really made RNA”^{1}: The invention of DNA–RNA hybridization and its outcome. Studies in History and Philosophy of Biological and Biomedical Sciences, 53(??):40–52, ???? 2015. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848615001004.

Furlan:2020:EEO

[FJEZM20] Violeta Furlan, N. David Jiménez-Escobar, Fernando Zamudio, and Celeste Medrano. ‘Ethnobiological equivocation’ and other misunderstandings in the interpretation of natures. Studies in History and Philosophy of Biological and Biomedical Sciences, 84(??):Article 101333, December 2020. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848620301448.

Fichman:2007:RLA

[FK07] Martin Fichman and Jennifer E. Keelan. Resister’s logic: the anti-vaccination arguments of Alfred Russel Wallace and their role in the debates over compulsory vaccination in England, 1870–1907. Studies in History and Philosophy of Biological and Biomedical Sciences, 38(3):585–607, September 2007. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848607000386.

Fangerau:2007:SEJ

[FM07] Heiner Fangerau and Irmgard Müller. Scientific exchange: Jacques Loeb (1859–1924) and Emil Godlewski (1875-1944) as representatives of a transatlantic developmental biology. Studies in History and Philosophy of Biological and Biomedical Sciences, 38(3):608–617, September 2007. CODEN ???? ISSN 1369-8486 (print), 1879-2499 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S1369848607000398.

Forber:2010:CEH

[For10] Patrick Forber. Confirmation and explaining how possible. Studies in History and Philosophy of Biological and Biomedical Sciences, 41(1):32–40, March 2010. CODEN ???? ISSN 1369-8486 (print), 1879-2499 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S1369848609000752.

Forber:2012:CER

[For12] Patrick Forber. Conjecture and explanation: a reply to Reydon. Studies in History and Philosophy of Biological and Biomedical Sciences, 43(1):298–301, March 2012. CO-DEN ???? ISSN 1369-8486 (print), 1879-2499 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S136984861100094X.

Forterre:2016:AHR

[For16] Patrick Forterre. To be or not to be alive: How recent discoveries challenge the traditional definitions of viruses and life. Studies in History and Philosophy of Biological and Biomedical Sciences, 59(??):100–108, October 2016. CODEN ???? ISSN 1369-8486 (print), 1879-2499 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S1369848616300085.

Fossheim:2019:PRH

[Fos19] Hallvard J. Fossheim. Past responsibility: History and the ethics of research on ethnic groups. Studies in History and Philosophy of Biological and Biomedical Sciences, 73(??):35–43, February 2019. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848617301899.

Franklin:2007:CPE

[Fra07] Sarah Franklin. ‘Crook’ pipettes: embryonic emigrations from agriculture to reproductive biomedicine. Studies in History and Philosophy of Biological and Biomedical Sciences, 38(2):358–373, June 2007. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848607000143.

Frank:2014:BCB

[Fra14] David Frank. Biodiversity, conservation biology, and rational choice. Studies in History and Philosophy of Biological and Biomedical Sciences, 45(??):101–104, March 2014. CODEN ???? ISSN 1369-8486 (print), 1879-2499 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S1369848613001301.

French:2011:SSP

[Fre11] Steven French. Shifting to structures in physics and biology: a prophylactic for promiscuous realism. Studies in History and Philosophy of Biological and Biomedical Sciences, 42(2):164–173, June 2011. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848610001160.

Friedrich:2013:DFS

[Fri13] Kathrin Friedrich. Digital ‘faces’ of synthetic biology. Studies in History and Philosophy of Biological and Biomedical Sciences, 44 (2):217–224, June 2013. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848613000319.

Frodeman:2016:IGC

[Fro16] Robert Frodeman. Interdisciplinarity, grand challenges, and the future of knowledge. Studies in History and Philosophy of Biological and Biomedical Sciences, 56(??):108–110, April 2016. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848615001727.

Fry:2012:SMN

[Fry12] Iris Fry. Is science metaphysically neutral? Studies in History and Philosophy of Biological and Biomedical Sciences, 43(3):665– 673, September 2012. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848612000672.

Fausto-Sterling:2000:SGP

[FS00] Anne Fausto-Sterling. The sex/gender perplex. Studies in History and Philosophy of Biological and Biomedical Sciences, 31(4): 637–646, December 2000. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848600000030.

Finn:2015:MSC

[FS15] Michael A. Finn and James F. Stark. Medical science and the Cruelty to Animals Act 1876: a re-examination of antivivisectionism in provincial Britain. Studies in History and Philosophy of Biological and Biomedical Sciences, 49(??):12–23, February 2015. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848614001575.

Fuller:2018:UEM

[Ful18] Jonathan Fuller. Universal etiology, multifactorial diseases and the constitutive model of disease classification. Studies in History and Philosophy of Biological and Biomedical Sciences, 67(??): 8–15, February 2018. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S136984861730033X.

Fuller:2020:MNJ

[Ful20] Jonathan Fuller. Medical Nihilism by Jacob Stegenga: What is the right dose? Studies in History and Philosophy of Biological and Biomedical Sciences, 81(??):Article 101270, June 2020. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848620300327. See [Sol20, Hea20, Hea20,

Gab20, Ste20].

Gabriel:2020:MNJ

[Gab20] Joseph M. Gabriel. Medical Nihilism by Jacob Stegenga: Historical scholarship and the question of effectiveness. Studies in History and Philosophy of Biological and Biomedical Sciences, 81(??):Article 101273, June 2020. CODEN ???? ISSN 1369-8486 (print), 1879-2499 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S1369848620300315.

See [Ful20, Sol20, Hea20, Ste20].

Gambarotto:2014:VFO

[Gam14] Andrea Gambarotto. Vital forces and organization: Philosophy of nature and biology in Karl Friedrich Kielmeyer. Studies in History and Philosophy of Biological and Biomedical Sciences, 48 (part A)(??):12–20, December 2014. CO-DEN ???? ISSN 1369-8486 (print), 1879-2499 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S1369848614000879.

Gambarotto:2017:KPN

[Gam17] Andrea Gambarotto. The “Kantian Principle” for natural history and its historical significance. Studies in History and Philosophy of Biological and Biomedical Sciences, 64(??):22–27, August 2017. CODEN ???? ISSN 1369-8486 (print), 1879-2499 (elec-

tronic). URL /science/article/pii/S136984861630084X.

Gamboa:2020:GAC

[Gam20] J. P. Gamboa. Goltz against cerebral localization: Methodology and experimental practices. Studies in History and Philosophy of Biological and Biomedical Sciences, 84(??):Article 101304, December 2020. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848619301657.

Gannett:2003:NGT

[Gan03] Lisa Gannett. The normal genome in twentieth-century evolutionary thought. Studies in History and Philosophy of Biological and Biomedical Sciences, 34(1):143–185, March 2003. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848602000717.

Gannett:2013:TDG

[Gan13] Lisa Gannett. Theodosius Dobzhansky and the genetic race concept. Studies in History and Philosophy of Biological and Biomedical Sciences, 44(3):250–261, September 2013. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848613000459.

Gannett:2014:BAR

[Gan14] Lisa Gannett. Biogeographical ancestry and race. Studies in History and Philosophy of Biological and Biomedical Sciences, 47 (part A)(??):173–184, September 2014. CO-DEN ???? ISSN 1369-8486 (print), 1879-2499 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S1369848614000570.

Gardner:2014:DCM

[Gar14a] Kirsten E. Gardner. Dreading cancer, minimizing risk, and preventive options. Studies in History and Philosophy of Biological and Biomedical Sciences, 46(??):100–103, June 2014. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848614000314.

Garson:2014:WVH

[Gar14b] Justin Garson. What is the value of historical fidelity in restoration? Studies in History and Philosophy of Biological and Biomedical Sciences, 45(??):97–100, March 2014. CODEN ???? ISSN 1369-8486 (print), 1879-2499 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S1369848613001283.

Garson:2016:TTP

[Gar16] Justin Garson. Two types of psychological hedonism. Studies in History and Philosophy of Biological and Biomedical Sciences, 56(??):7–14, April 2016. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848615001612.

Gaudilliere:2000:RSS

[Gau00] Jean-Paul Gaudillière. Rockefeller strategies for scientific medicine: molecular machines, viruses and vaccines. Studies in History and Philosophy of Biological and Biomedical Sciences, 31(3):491–509, September 2000. CODEN ???? ISSN 1369-8486 (print), 1879-2499 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S1369848600000170.

Gaudilliere:2001:PIB

[Gau01] Jean-Paul Gaudillière. The pharmaceutical industry in the biotech century: toward a history of science, technology and business? Studies in History and Philosophy of Biological and Biomedical Sciences, 32(1):191–201, January 2001. CODEN ???? ISSN 1369-8486 (print), 1879-2499 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S1369848600000042.

Gaudilliere:2002:IBF

[Gau02a] Jean-Paul Gaudilliére. Inventer la bioméedecine: la France, l’Amérique et la production des savoirs du vivant, 1945–1965. (French) [Inventing biomedicine: France, America, and the production of life savers (1945–1965)]. Éd. la Découverte, Paris, France, 2002. ISBN 2-7071-3607-7 (paperback). 391 + 8 pp. LCCN ????

Gaudilliere:2002:PNY

[Gau02b] Jean-Paul Gaudillière. Paris–New York roundtrip: transatlantic crossings and the reconstruction of the biological sciences in post-war France. Studies in History and Philosophy of Biological and Biomedical Sciences, 33(3):389–417, September 2002. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848602000122.

Gaudilliere:2004:GDB

[Gau04] Jean-Paul Gaudillière. Genesis and development of a biomedical object: styles of thought, styles of work and the history of the sex steroids. Studies in History and Philosophy of Biological and Biomedical Sciences, 35(3):525–543, September 2004. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848604000470.

Gaudilliere:2005:BPT

[Gau05a] Jean-Paul Gaudillière. Better prepared than synthesized: Adolf Butenandt, Schering AG and the transformation of sex steroids into drugs (1930–1946). Studies in History and Philosophy of Biological and Biomedical Sciences, 36(4):612–644, December 2005. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848605000725.

Gaudilliere:2005:IDT

[Gau05b] Jean-Paul Gaudillière. Introduction: drug trajectories. Studies in History and Philosophy of Biological and Biomedical Sciences, 36(4):603–611, December 2005. CODEN ???? ISSN 1369-8486 (print), 1879-2499 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S1369848605000713.

Gaudilliere:2007:FCI

[Gau07] Jean-Paul Gaudillière. The farm and the clinic: an inquiry into the making of our biotechnological modernity. Studies in History and Philosophy of Biological and Biomedical Sciences, 38(2):521– 529, June 2007. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848607000222.

Gaudilliere:2009:NWO

[Gau09] Jean-Paul Gaudillière. New wine in old bottles? The biotechnology problem in the history of molecular biology. Studies in History and Philosophy of Biological and Biomedical Sciences, 40 (1):20–28, March 2009. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848608000927.

Gawin:2008:SRM

[Gaw08] Magdalena Gawin. The sex reform movement and eugenics in interwar Poland. Studies in History and Philosophy of Biological and Biomedical Sciences, 39(2):181–186, June 2008. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848608000149.

Green:2017:BMP

[GB17] Sara Green and Robert Batterman. Biology meets physics: Reductionism and multi-scale modeling of morphogenesis. Studies in History and Philosophy of Biological and Biomedical Sciences, 61(??):20–34, February 2017. CODEN ???? ISSN 1369-8486 (print), 1879-2499 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S1369848616301157.

Germain:2017:EPA

[GCT17] Pierre-Luc Germain, Luca Chiapperino, and Giuseppe Testa. The European politics of animal experimentation: From Victorian Britain to ‘stop vivisection’. Studies in History and Philosophy of Biological and Biomedical Sciences, 64(??):75–87, August 2017. CODEN ???? ISSN 1369-8486 (print), 1879-2499 (elec-

tronic). URL /science/article/pii/S1369848616301091.

Gelfert:2013:SBB

[Gel13] Axel Gelfert. Synthetic biology between technoscience and thing knowledge. Studies in History and Philosophy of Biological and Biomedical Sciences, 44(2):141–149, June 2013. CODEN ???? ISSN 1369-8486 (print), 1879-2499 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S136984861300023X.

Gere:1999:BMS

[Ger99] Cathy Gere. Bones that matter: Sex determination in paleodemography 1948–1995. Studies in History and Philosophy of Biological and Biomedical Sciences, 30(4):455–471, December 1999. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848699000217.

Gere:2004:BV

[Ger04a] Cathy Gere. The brain in a vat. Studies in History and Philosophy of Biological and Biomedical Sciences, 35(2):219–225, June 2004. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848604000172.

Gere:2004:TVT

[Ger04b] Cathy Gere. Thought in a vat: thinking through Annie Cattrell. Studies in History and Philosophy of Biological and Biomedical Sciences, 35(2):415–436, June 2004. CODEN ???? ISSN 1369-8486 (print), 1879-2499 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S1369848604000287.

Gere:2004:BVGa

[Ger04c] Charlie Gere. Brains-in-vats, giant brains and world brains: the brain as metaphor in digital culture. Studies in History and Philosophy of Biological and Biomedical Sciences, 35(2):351–366, June 2004. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848604000251.

Gere:2004:TPD

[Ger04d] Charlie Gere. The technologies and politics of delusion: an interview with artist Rod Dickinson. Studies in History and Philosophy of Biological and Biomedical Sciences, 35(2):333–349, June 2004. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S136984860400024X.

Gerodetti:2008:RSM

[Ger08] Natalia Gerodetti. Rational subjects, marriage counselling and the conundrums of eugenics. Studies in History and Philosophy of Biological and Biomedical Sciences, 39(2):255–262, June 2008. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S136984860800023X.

Gerson:2013:ISI

[Ger13] Elihu M. Gerson. Integration of specialties: an institutional and organizational view. Studies in History and Philosophy of Biological and Biomedical Sciences, 44(4):515–524, December 2013. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848612001094.

Gere:2020:PPG

[Ger20a] Cathy Gere. Pain, Pleasure, and the Greater Good, by Cathy Gere: Reply by the author. Studies in History and Philosophy of Biological and Biomedical Sciences, 83(??):Article 101252, October 2020. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848619301578.

Germain:2020:BEC

[Ger20b] Pierre-Luc Germain. Beyond explanation, the cancer biology patchwork. Studies in History and Philosophy of Biological and Biomedical Sciences, 84(??):Article 101324, December 2020. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848620301357.

Gradmann:2009:LDR

[GF09] Christoph Gradmann and Elborg Forster. Laboratory disease: Robert Koch’s medical bacteriology. The Johns Hopkins University Press, Baltimore, MD, USA, 2009. ISBN 0-8018-9313-5 (hardcover), 1-71860-320-7. 318 pp. LCCN QR46 .G6713

2009. URL http://d-nb.info/1009572067/04;   http://
www.gbv.de/dms/bowker/toc/9780801893131.pdf.

Gere:2004:BVGb

[GG04] Cathy Gere and Charlie Gere. The brain in a vat: Guest Editors. Studies in History and Philosophy of Biological and Biomedical Sciences, 35(2):iii–iv, June 2004. CO-DEN ???? ISSN 1369-8486 (print), 1879-2499 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S1369848604000160.

Griesemer:2006:MML

[GG06] James R. Griesemer and Elihu M. Gerson. Of mice and men and low unit cost. Studies in History and Philosophy of Biological and Biomedical Sciences, 37(2):363–372, June 2006. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848606000173.

Giglioni:2014:DRL

[Gig14] Guido Giglioni. Death in Rome: Lancisi, Pope Clement XI, and the medicalisation of life. Studies in History and Philosophy of Biological and Biomedical Sciences, 46(??):97–99, June 2014. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848614000326.

Gildenhuys:2004:DHR

[Gil04] Peter Gildenhuys. Darwin, Herschel, and the role of analogy in Darwin’s origin. Studies in History and Philosophy of Biological and Biomedical Sciences, 35(4):593–611, December 2004. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848604000664.

Gillies:2005:HKA

[Gil05] Donald Gillies. Hempelian and Kuhnian approaches in the philosophy of medicine: the Semmelweis case. Studies in History and Philosophy of Biological and Biomedical Sciences, 36(1):159–181, March 2005. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848604000937.

Gillies:2013:WDB

[Gil13] Donald Gillies. Why did bloodletting decline? (reviewing K. C. Carter, The decline of therapeutic bloodletting and the collapse of traditional medicine). Studies in History and Philosophy of Biological and Biomedical Sciences, 44(3):433–434, September 2013. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848613000381.

Gillies:2017:EME

[Gil17] Donald Gillies. Evidence of mechanism in the evaluation of streptomycin and thalidomide. Studies in History and Philosophy of Biological and Biomedical Sciences, 66(??):55–62, December 2017. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848616301066.

Giraud:2020:MBN

[Gir20] Eva Haifa Giraud. Model Behavior by Nicole Nelson: Complex ethics. Studies in History and Philosophy of Biological and Biomedical Sciences, 82(??):Article 101267, August 2020. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848620300066.

Gissis:2008:WRR

[Gis08] Snait B. Gissis. When is ‘race’ a race? 1946–2003. Studies in History and Philosophy of Biological and Biomedical Sciences, 39(4): 437–450, December 2008. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848608000757.

Gissis:2014:CVP

[Gis14] Snait B. Gissis. The continuing vitality of the problématique of vitalism? Studies in History and Philosophy of Biological and Biomedical Sciences, 47 (part A)(??):196–200, September 2014. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848614000430.

Gissis:2017:TFC

[Gis17] Snait B. Gissis. Is time future contained in time past? Studies in History and Philosophy of Biological and Biomedical Sciences, 62(??):51–55, April 2017. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S136984861730002X.

Gissis:2020:TLE

[Gis20] Snait B. Gissis. Transfer of Lamarckisms and emerging ‘scientific’ psychologies: 19th–early 20th centuries Britain and France. Studies in History and Philosophy of Biological and Biomedical Sciences, 83(??):Article 101146, October 2020. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848617301334.

Gramelsberger:2013:PPS

[GKG13] Gabriele Gramelsberger, Tarja Knuuttila, and Axel Gelfert. Philosophical perspectives on synthetic biology. Studies in History and Philosophy of Biological and Biomedical Sciences, 44 (2):119–121, June 2013. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848613000204.

Geison:2001:VLO

[GL01] Gerald L. Geison and Manfred D. Laubichler. The varied lives of organisms: variation in the historiography of the biological sciences. Studies in History and Philosophy of Biological and Biomedical Sciences, 32(1):1–29, January 2001. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848600000236.

Guillem-Llobat:2015:SRC

[GL15] Ximo Guillem-Llobat. Santiago Ramón y Cajal and the Spanish historiography of science. Studies in History and Philosophy of Biological and Biomedical Sciences, 53(??):111–113, ???? 2015. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848615000825.

Glackin:2008:DNH

[Gla08] Shane Nicholas Glackin. Dolphin natures, human virtues: Mac-Intyre and ethical naturalism. Studies in History and Philosophy of Biological and Biomedical Sciences, 39(3):292–297, September 2008. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848608000381.

Glackin:2012:KMO

[Gla12] Shane Nicholas Glackin. Kind-making, objectivity, and political neutrality; the case of Solastalgia. Studies in History and Philosophy of Biological and Biomedical Sciences, 43(1):209–218, March 2012. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848611001397.

Glennan:2005:MM

[Gle05] Stuart Glennan. Modeling mechanisms. Studies in History and Philosophy of Biological and Biomedical Sciences, 36(2):443–464, June 2005. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848605000270.

Gliboff:2000:PDA

[Gli00] Sander Gliboff. Paley’s design argument as an inference to the best explanation, or, Dawkins’ dilemma. Studies in History and Philosophy of Biological and Biomedical Sciences, 31(4):579– 597, December 2000. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848600000194.

Gliboff:2018:CCR

[Gli18] Sander Gliboff. Cold case reopened. Studies in History and Philosophy of Biological and Biomedical Sciences, 68–69(??):70– 72, April/June 2018. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S1369848618300360.

Gonzalez-Moreno:2015:DMT

[GMST15] María González-Moreno, Cristian Saborido, and David Teira. Disease-mongering through clinical trials. Studies in History and Philosophy of Biological and Biomedical Sciences, 51(??):11–18, June 2015. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848615000382.

Golden:1996:SHW

[Gol96] Janet Lynne Golden. A social history of wet nursing in America: from breast to bottle. Cambridge history of medicine. Cambridge University Press, Cambridge, UK, 1996. ISBN 0-521-49544-X (hardcover). xiii + 215 pp. LCCN RJ216 .G63 1996. URL http:/

/www.loc.gov/catdir/description/cam027/95022149.html;
 http://www.loc.gov/catdir/toc/cam026/95022149.html.

Goldberg:2013:DBF

[Gol13] Benjamin Goldberg. A dark business, full of shadows: Analogy and theology in William Harvey. Studies in History and Philosophy of Biological and Biomedical Sciences, 44(3):419– 432, September 2013. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848613000745.

Gomez:2020:BRD

[Góm20] Pablo F. Gómez. Book review: Difference and Disease by Suman Seth: Colonial Hippocratic medicine and the history of race in the British Empire. Studies in History and Philosophy of Biological and Biomedical Sciences, 80(??):Article 101220, April 2020. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848619301268. See reply [Set20].

Gaudilliere:2004:LS

[GR04] Jean-Paul Gaudillière and Hans-Jörg Rheinberger. Life stories. Studies in History and Philosophy of Biological and Biomedical Sciences, 35(4):753–764, December 2004. CO-DEN ???? ISSN 1369-8486 (print), 1879-2499 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S1369848604000755.

Grasseni:1998:TRS

[Gra98] Cristina Grasseni. Taxidermy as rhetoric of self-making: Charles Waterton (1782–1865), wandering naturalist. Studies in History and Philosophy of Biological and Biomedical Sciences, 29(2):269– 294, September 1998. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848698000223.

Gradmann:2004:HIC

[Gra04] Christoph Gradmann. A harmony of illusions: clinical and experimental testing of Robert Koch’s tuberculin 1890– 1900. Studies in History and Philosophy of Biological and Biomedical Sciences, 35(3):465–481, September 2004. CO-DEN ???? ISSN 1369-8486 (print), 1879-2499 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S1369848604000482.

Grasseni:2007:MCE

[Gra07] Cristina Grasseni. Managing cows: an ethnography of breeding practices and uses of reproductive technology in contemporary dairy farming in Lombardy (Italy). Studies in History and Philosophy of Biological and Biomedical Sciences, 38(2):488–510, June 2007. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848607000209.

Gradmann:2010:RKI

[Gra10] Christoph Gradmann. Robert Koch and the invention of the carrier state: tropical medicine, veterinary infections and epidemiology around 1900. Studies in History and Philosophy of Biological and Biomedical Sciences, 41(3):232–240, September 2010. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S136984861000035X.

Gradmann:2012:MPH

[Gra12] Christoph Gradmann. Modernity, public health and the welfare state. Studies in History and Philosophy of Biological and Biomedical Sciences, 43(4):825–827, December 2012. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848612001112.

Gramelsberger:2013:SAS

[Gra13] Gabriele Gramelsberger. The simulation approach in synthetic biology. Studies in History and Philosophy of Biological and Biomedical Sciences, 44(2):150–157, June 2013. CODEN ???? ISSN 1369-8486 (print), 1879-2499 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S1369848613000241.

Graus:2014:HSS

[Gra14] Andrea Graus. Hypnosis in Spain (1888–1905): From spectacle to medical treatment of mediumship. Studies in History and Philosophy of Biological and Biomedical Sciences, 48 (part A)(??): 85–93, December 2014. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S136984861400082X.

Gradmann:2017:TTC

[Gra17] Christoph Gradmann. Transitions, traditions: From colonial to global health. Studies in History and Philosophy of Biological and Biomedical Sciences, 64(??):101–105, August 2017. CODEN ???? ISSN 1369-8486 (print), 1879-2499 (electronic). URL /

science/article/pii/S1369848617300821.

Grafen:2019:SWA

[Gra19] Alan Grafen. Should we ask for more than consistency of Darwinism with Mendelism? Studies in History and Philosophy of Biological and Biomedical Sciences, 78(??):Article 101224, December 2019. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848619301189.

Gregory:2001:HAW

[Gre01] Andrew Gregory. Harvey, Aristotle and the weather cycle. Studies in History and Philosophy of Biological and Biomedical Sciences, 32(1):153–168, January 2001. CO-DEN ???? ISSN 1369-8486 (print), 1879-2499 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S1369848600000248.

Green:2013:WOM

[Gre13] Sara Green. When one model is not enough: Combining epistemic tools in systems biology. Studies in History and Philosophy of Biological and Biomedical Sciences, 44(2):170–180, June 2013. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848613000265.

Green:2015:CBC

[Gre15a] Sara Green. Can biological complexity be reverse engineered? Studies in History and Philosophy of Biological and Biomedical Sciences, 53(??):73–83, ???? 2015. CO-DEN ???? ISSN 1369-8486 (print), 1879-2499 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S1369848615000618.

Greif:2015:DTR

[Gre15b] Hajo Greif. The Darwinian tension: Romantic science and the causal laws of nature. Studies in History and Philosophy of Biological and Biomedical Sciences, 53(??):53–61, ???? 2015. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848615001016.

Green:2016:EPB

[Gre16] Sara Green. Explanatory pluralism in biology. Studies in History and Philosophy of Biological and Biomedical Sciences, 59(??): 154–157, October 2016. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848616000200.

Greene:2017:FGI

[Gre17] Mott T. Greene. First get it right, then get it written. Studies in History and Philosophy of Biological and Biomedical Sciences, 64(??):97–100, August 2017. CODEN ???? ISSN 1369-8486 (print), 1879-2499 (electronic). URL /science/article/pii/

S136984861730078X.

Green:2020:ECA

[Gre20a] Sara Green. Explaining Cancer by Anya Plutynski: Cancer explained and unexplained. Studies in History and Philosophy of Biological and Biomedical Sciences, 84(??):Article 101325, December 2020. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848620301369. See [Plu20].

Greslehner:2020:SAC

[Gre20b] Gregor P. Greslehner. Not by structures alone: Can the immune system recognize microbial functions? Studies in History and Philosophy of Biological and Biomedical Sciences, 84(??):Article 101336, December 2020. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848620301473.

Griesemer:2013:IAD

[Gri13] James Griesemer. Integration of approaches in David Wake’s model-taxon research platform for evolutionary morphology. Studies in History and Philosophy of Biological and Biomedical Sciences, 44(4):525–536, December 2013. CO-DEN ???? ISSN 1369-8486 (print), 1879-2499 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S1369848613000356.

Grote:2010:SAC

[Gro10] Mathias Grote. Surfaces of action: cells and membranes in electrochemistry and the life sciences. Studies in History and Philosophy of Biological and Biomedical Sciences, 41(3):183– 193, September 2010. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848610000415.

Grosholz:2011:SPM

[Gro11] Emily Grosholz. Studying populations without molecular biology: Aster models and a new argument against reductionism. Studies in History and Philosophy of Biological and Biomedical Sciences, 42(2):246–251, June 2011. CO-DEN ???? ISSN 1369-8486 (print), 1879-2499 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S1369848610001251.

Garcia-Sancho:2011:AMM

[GS11a] Miguel García-Sancho. Academic and molecular matrices: a study of the transformations of connective tissue research at the University of Manchester (1947–1996). Studies in History and Philosophy of Biological and Biomedical Sciences, 42(2):233–245, June 2011. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848610001275.

Greek:2011:CSE

[GS11b] Ray Greek and Niall Shanks. Complex systems, evolution, and animal models. Studies in History and Philosophy of Biological and Biomedical Sciences, 42(4):542–544, December 2011. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848611000562.

Garcia-Sancho:2012:GCP

[GS12] Miguel García-Sancho. From the genetic to the computer program: the historicity of ‘data’ and ‘computation’ in the investigations on the nematode worm C. elegans (1963–1998). Studies in History and Philosophy of Biological and Biomedical Sciences, 43(1):16–28, March 2012. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848611000781.

Garcia-Sancho:2016:PHM

[GS16a] Miguel García-Sancho. The proactive historian: Methodological opportunities presented by the new archives documenting genomics. Studies in History and Philosophy of Biological and Biomedical Sciences, 55(??):70–82, February 2016. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848615001314.

Garson:2016:IBP

[GS16b] Justin Garson and Armin W. Schulz. Introduction: the biology of psychological altruism. Studies in History and Philosophy of Biological and Biomedical Sciences, 56(??):1–2, April 2016. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848615001685.

Godfrey-Smith:2020:BTW

[GS20] Peter Godfrey-Smith. In the beginning there was information? Studies in History and Philosophy of Biological and Biomedical Sciences, 80(??):Article 101239, April 2020. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848619301001.

Garcia-Sancho:2019:BMS

[GSM19] Miguel García-Sancho and Dmitriy Myelnikov. Between mice and sheep: Biotechnology, agricultural science and animal models in late-twentieth century Edinburgh. Studies in History and Philosophy of Biological and Biomedical Sciences, 75(??):24–33, June 2019. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848617301814.

Guerrero:2010:NTH

[Gue10] J. David Guerrero. On a naturalist theory of health: a critique. Studies in History and Philosophy of Biological and Biomedical Sciences, 41(3):272–278, September 2010.

CODEN ???? ISSN 1369-8486 (print), 1879-2499 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S1369848610000348.

Guerrini:2012:HNC

[Gue12] Anita Guerrini. Health, national character and the English diet in 1700. Studies in History and Philosophy of Biological and Biomedical Sciences, 43(2):349–356, June 2012. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848611001191.

Guenther:2020:PPG

[Gue20] Katja Guenther. Pain, pleasure, and the greater good by Cathy Gere: History as moral work. Studies in History and Philosophy of Biological and Biomedical Sciences, 83(??):Article 101251, October 2020. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S136984861930158X.

Gundlach:2012:BRT

[Gun12] Horst Gundlach. Bühler revisited in times of war — Peter R. Hofstätter’s The Crisis of Psychology (1941). Studies in History and Philosophy of Biological and Biomedical Sciences, 43(2):504– 513, June 2012. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S136984861100135X.

Gurjeva:1998:BRS

[Gur98] Lyuba Gurjeva. Book review: A social history of wet nursing in America: From breast to bottle: Janet Golden, (Cambridge: Cambridge University Press, 1996), 215 pp., ISBN 0-521-49544-X hardback. Studies in History and Philosophy of Biological and Biomedical Sciences, 29(1):189–199, March 1998. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848698000089. See [Gol96].

Guttinger:2013:CPA

[Güt13] Stephan Güttinger. Creating parts that allow for rational design: Synthetic biology and the problem of context-sensitivity. Studies in History and Philosophy of Biological and Biomedical Sciences, 44(2):199–207, June 2013. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848613000290.

Guttinger:2018:BRC

[Gut18] Stephan Guttinger. Book review: A Crack in Creation: The New Power to Control Evolution, Jennifer Doudna, Samuel Sternberg. Bodley Head, London (2017), 304, Price £20 hardcover, ISBN: 978-1-84792-381-3. Studies in History and Philosophy of Biological and Biomedical Sciences, 67(??):32–35, February 2018. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S136984861730184X.

Giallombardo:2019:PBV

[GvA19] F. Giallombardo and T. R. van Andel. Paolo Boccone and the visual communication of pre-linnean botany. A comparison between his Leiden herbarium, Paris autoprint and published Icones (1674). Studies in History and Philosophy of Biological and Biomedical Sciences, 74(??):15–26, April 2019. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848618300104.

Gervais:2013:IEB

[GW13] Raoul Gervais and Erik Weber. Inferential explanations in biology. Studies in History and Philosophy of Biological and Biomedical Sciences, 44(3):356–364, September 2013. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848613000915.

Harris:2009:JLS

[HA09] Stephen A. Harris and Peter R. Anstey. John Locke’s seed lists: a case study in botanical exchange. Studies in History and Philosophy of Biological and Biomedical Sciences, 40(4):256–264, December 2009. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848609000521.

Hacking:2010:PWR

[Hac10] Ian Hacking. Pathological withdrawal of refugee children seeking asylum in Sweden. Studies in History and Philosophy of Biological and Biomedical Sciences, 41(4):309–317, December 2010. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848610000634.

Hadreas:1999:INP

[Had99] Peter Hadreas. Intentionality and the neurobiology of pleasure. Studies in History and Philosophy of Biological and Biomedical Sciences, 30(2):219–236, June 1999. CODEN ???? ISSN 1369-8486 (print), 1879-2499 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S1369848698000235.

Hagen:2001:ICS

[Hag01] Joel B. Hagen. 1the introduction of computers into systematic research in the United States during the 1960s. Studies in History and Philosophy of Biological and Biomedical Sciences, 32(2):291– 314, June 2001. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S136984860100005X.

Hakosalo:2006:BUK

[Hak06] Heini Hakosalo. The brain under the knife: serial sectioning and the development of late nineteenth-century neuroanatomy. Studies in History and Philosophy of Biological and Biomedical Sciences, 37(2):172–202, June 2006. CODEN ???? ISSN 1369-8486 (print), 1879-2499 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S1369848606000136.

Hall:2008:ESS

[Hal08] Lesley A. Hall. Eugenics, sex and the state: some introductory remarks. Studies in History and Philosophy of Biological and Biomedical Sciences, 39(2):177–180, June 2008. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848608000137.

Hall:2011:WAB

[Hal11] Kersten Hall. William Astbury and the biological significance of nucleic acids, 1938–1951. Studies in History and Philosophy of Biological and Biomedical Sciences, 42(2):119–128, June 2011. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848610001111.

Hale:2014:PDM

[Hal14a] Piers J. Hale. Political descent: Malthus, mutualism, and the politics of evolution in Victorian England. University of Chicago Press, Chicago, IL, USA and London, UK, 2014. ISBN 0-226-10849-X (hardcover), 0-226-10852-X (e-book). 442 pp. LCCN JC336 .H254 2014.

Hall:2014:MMC

[Hal14b] Kersten T. Hall. The man in the monkeynut coat: William Astbury and the forgotten road to the double-helix. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2014. ISBN 0-19-870459-3. ix + 242 pp. LCCN QP624 .H348

2014; Q143.A88 H35 2014. URL http://www.loc.gov/catdir/
enhancements/fy1506/2013953775-b.html;   http://www.
loc.gov/catdir/enhancements/fy1506/2013953775-d.html;
 http://www.loc.gov/catdir/enhancements/fy1506/2013953775-
t.html.

Hall:2016:TOB

[Hal16] Kersten Hall. Thinking outside the black-box: the case of Marshall Nirenberg and Oswald Avery. Studies in History and Philosophy of Biological and Biomedical Sciences, 60(??):97–101, December 2016. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848616300395.

Halina:2018:BRO

[Hal18] Marta Halina. Book review: Other Minds: The Octopus and the Evolution of Intelligent Life, Peter Godfrey-Smith. William Collins, London (2017), pp. 255, Price £14.99 paperback, ISBN 978-0-00-822631-2. Studies in History and Philosophy of Biological and Biomedical Sciences, 67(??):28–31, February 2018. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848617301796.

Hall:2020:MBN

[Hal20] Wayne Hall. Model behaviour by Nicole Nelson: Is animal behaviour genetics a degenerating research program?

Studies in History and Philosophy of Biological and Biomedical Sciences, 82(??):Article 101268, August 2020. CO-DEN ???? ISSN 1369-8486 (print), 1879-2499 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S1369848620300078.

Hamlin:2013:FCH

[Ham13] Christopher Hamlin. Forensic cultures in historical perspective: Technologies of witness, testimony, judgment (and justice?). Studies in History and Philosophy of Biological and Biomedical Sciences, 44(1):4–15, March 2013. CO-DEN ???? ISSN 1369-8486 (print), 1879-2499 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S1369848612000994.

Hansson:2006:CEC

[Han06] Mats G. Hansson. Combining efficiency and concerns about integrity when using human biobanks. Studies in History and Philosophy of Biological and Biomedical Sciences, 37(3):520– 532, September 2006. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848606000483.

Hanna:2014:KAM

[Han14] Robert Hanna. Kant’s anti-mechanism and Kantian antimechanism. Studies in History and Philosophy of Biological and Biomedical Sciences, 46(??):112–116, June 2014. CODEN ???? ISSN 1369-8486 (print), 1879-2499 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S1369848614000107.

Hardcastle:1999:TSG

[Har99a] Gary Hardcastle. Are there scientific goals? Studies in History and Philosophy of Biological and Biomedical Sciences, 30(3):297– 311, September 1999. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848699000187.

Hardcastle:1999:WWD

[Har99b]  Valerie Gray Hardcastle.  What we don’t know about

brains. Studies in History and Philosophy of Biological and Biomedical Sciences, 30(1):69–89, March 1999. CO-DEN ???? ISSN 1369-8486 (print), 1879-2499 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S1369848698000272.

Harman:2013:ERS

[Har13a] Oren Harman. Essay review: Shakespeare among the ants: The social conquest of earth, Edward O. Wilson. Liveright Publishing Corporation, New York (2012). 330 pp., Price US $27.95 Hardcover, ISBN: 978-0-87140-413-8. Studies in History and Philosophy of Biological and Biomedical Sciences, 44(1):114–118, March 2013. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S136984861200129X.

Harman:2013:UMJ

[Har13b] Oren Harman. Unformed minds: Juveniles, neuroscience, and the law. Studies in History and Philosophy of Biological and Biomedical Sciences, 44(3):455–459, September 2013. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848613000411.

Harbecke:2015:RTM

[Har15] Jens Harbecke. The regularity theory of mechanistic constitution and a methodology for constitutive inference. Studies in History and Philosophy of Biological and Biomedical Sciences, 54(??): 10–19, December 2015. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848615001302.

Harrison:2016:WWH

[Har16] Peter Harrison. What was historical about natural history? Contingency and explanation in the science of living things. Studies in History and Philosophy of Biological and Biomedical Sciences, 58(??):8–16, August 2016. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848615001855.

Hardesty:2018:MAA

[Har18] Rebecca A. Hardesty. Much ado about mice: Standard-setting in model organism research. Studies in History and Philosophy of Biological and Biomedical Sciences, 68–69(??):1–14, April/ June 2018. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL https://www.sciencedirect.com/science/
article/pii/S1369848617301279.

Haselager:1997:CSF

[Has97] W. F. G. Haselager. Cognitive science and folk psychology: the right frame of mind. Sage Publications, London, UK, 1997. ISBN 0-7619-5425-2, 0-7619-5426-0 (paperback). viii + 165 pp. LCCN BF311 .H38 1997. URL http://www.loc.gov/

catdir/enhancements/fy1006/98150593-d.html;   http://
www.loc.gov/catdir/enhancements/fy1006/98150593-t.html.

Hattiangadi:2000:CE

[Hat00] Anandi Hattiangadi. Credibility@feminist.epistemology. Studies in History and Philosophy of Biological and Biomedical Sciences, 31(4):647–657, December 2000. CODEN ???? ISSN 1369-8486 (print), 1879-2499 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S1369848600000054.

Hatfield:2012:KKC

[Hat12] Gary Hatfield. Koffka, Köhler, and the “crisis” in psychology. Studies in History and Philosophy of Biological and Biomedical Sciences, 43(2):483–492, June 2012. CO-DEN ???? ISSN 1369-8486 (print), 1879-2499 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S1369848611001312.

Haufe:2012:DL

[Hau12] Chris Haufe. Darwin’s laws. Studies in History and Philosophy of Biological and Biomedical Sciences, 43(1):269–280, March 2012. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848611000872.

Hawthorne:2010:EVH

[Haw10] Susan Hawthorne. Embedding values: how science and society jointly valence a concept-the case of ADHD. Studies in History and Philosophy of Biological and Biomedical Sciences, 41 (1):21–31, March 2010. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848609000740.

Hayton:2010:IDA

[Hay10] Darin Hayton. Instruments and demonstrations in the astrological curriculum: evidence from the University of Vienna, 1500–1530. Studies in History and Philosophy of Biological and Biomedical Sciences, 41(2):125–134, June 2010. CODEN ???? ISSN 1369-8486 (print), 1879-2499 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S1369848610000178.

Healy:2020:MNJ

[Hea20] David Healy. Medical Nihilism by Jacob Stegenga: Is operationalism the answer to nihilism? Studies in History and Philosophy of Biological and Biomedical Sciences, 81(??):Article 101272, June 2020. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848620300480. See [Ful20, Sol20, Gab20,

Ste20].

Hedgecoe:2006:CEP

[Hed06] Adam M. Hedgecoe. Context, ethics and pharmacogenetics. Studies in History and Philosophy of Biological and Biomedical Sciences, 37(3):566–582, September 2006. CO-DEN ???? ISSN 1369-8486 (print), 1879-2499 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S1369848606000513.

Hedfors:2007:MEW

[Hed07a] Eva Hedfors. Medical ethics in the wake of the Holocaust: departing from a postwar paper by Ludwik Fleck. Studies in History and Philosophy of Biological and Biomedical Sciences, 38(3): 642–655, September 2007. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848607000374.

Hedfors:2007:RST

[Hed07b] Eva Hedfors. The reading of scientific texts: questions on interpretation and evaluation, with special reference to the scientific writings of Ludwik Fleck. Studies in History and Philosophy of Biological and Biomedical Sciences, 38(1):136–158, March 2007. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848606000951.

Hellstrom:2013:GDG

[Hel13] Petter Hellström. Genetic diaspora, genetic return (reviewing N. A. El-Haj, The genealogical science: the search for Jewish origins and the politics of epistemology). Studies in History and Philosophy of Biological and Biomedical Sciences, 44(3):439– 442, September 2013. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848613000757.

Henckes:2011:RCN

[Hen11] Nicolas Henckes. Reshaping chronicity: neuroleptics and changing meanings of therapy in French psychiatry, 1950– 1975. Studies in History and Philosophy of Biological and Biomedical Sciences, 42(4):434–442, December 2011. CO-DEN ???? ISSN 1369-8486 (print), 1879-2499 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S1369848611000409.

Herzberg:2011:BCS

[Her11] David Herzberg. Blockbusters and controlled substances: Miltown, Quaalude, and consumer demand for drugs in postwar America. Studies in History and Philosophy of Biological and Biomedical Sciences, 42(4):415–426, December 2011. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848611000410.

Hesketh:2015:GDW

[Hes15] Ian Hesketh. A good Darwinian? Winwood Reade and the making of a late Victorian evolutionary epic. Studies in History and Philosophy of Biological and Biomedical Sciences, 51(??):44–52, June 2015. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848615000278.

Hesketh:2016:CHC

[Hes16] Ian Hesketh. Counterfactuals and history: Contingency and convergence in histories of science and life. Studies in History and Philosophy of Biological and Biomedical Sciences, 58(??):41–48, August 2016. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848615001880.

Heyck:2015:LIB

[Hey15] Hunter Heyck. Leviathan and the ink blot: the politics of the mind and its sciences in Cold War America. Studies in History and Philosophy of Biological and Biomedical Sciences, 53(??): 114–117, ???? 2015. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848615001028.

Harrison:2016:IEH

[HH16] Peter Harrison and Ian Hesketh. Introduction: Evolution and historical explanation. Studies in History and Philosophy of Biological and Biomedical Sciences, 58(??):1–7, August 2016. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848615001831.

Hicks:2015:EDG

[Hic15] Daniel J. Hicks. Epistemological depth in a GM crops controversy. Studies in History and Philosophy of Biological and Biomedical Sciences, 50(??):1–12, April 2015. CO-DEN ???? ISSN 1369-8486 (print), 1879-2499 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S1369848615000321.

Huber:2013:MME

[HK13] Lara Huber and Lara K. Keuck. Mutant mice: Experimental organisms as materialised models in biomedicine. Studies in History and Philosophy of Biological and Biomedical Sciences, 44(3): 385–391, September 2013. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848613000150.

Hochman:2013:RDH

[Hoc13] Adam Hochman. Racial discrimination: How not to do it. Studies in History and Philosophy of Biological and Biomedical Sciences, 44(3):278–286, September 2013. CO-DEN ???? ISSN 1369-8486 (print), 1879-2499 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S1369848613000174.

Hochman:2014:URN

[Hoc14] Adam Hochman. Unnaturalised racial naturalism. Studies in History and Philosophy of Biological and Biomedical Sciences, 46(??):79–87, June 2014. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848614000429.

Hochman:2015:VNN

[Hoc15] Adam Hochman. Of Vikings and Nazis: Norwegian contributions to the rise and the fall of the idea of a superior Aryan race. Studies in History and Philosophy of Biological and Biomedical Sciences, 54(??):84–88, December 2015. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848615001296.

Hochman:2016:RDP

[Hoc16] Adam Hochman. Race: Deflate or pop? Studies in History and Philosophy of Biological and Biomedical Sciences, 57(??):60–68, June 2016. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848616300140.

Hodge:2011:DAM

[Hod11] Jonathan Hodge. Darwinism after Mendelism: the case of Sewall Wright’s intellectual synthesis in his shifting balance theory of evolution (1931). Studies in History and Philosophy of Biological and Biomedical Sciences, 42(1):30–39, March 2011. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848610001019.

Hogan:2016:MMU

[Hog16] Andrew J. Hogan. Making the most of uncertainty: Treasuring exceptions in prenatal diagnosis. Studies in History and Philosophy of Biological and Biomedical Sciences, 57(??):24–33, June 2016. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S136984861600039X.

Hogan:2018:PDC

[Hog18] Andrew J. Hogan. Prenatal diagnosis in context. Studies in History and Philosophy of Biological and Biomedical Sciences, 72(??):55–58, December 2018. CODEN ???? ISSN 1369-8486 (print), 1879-2499 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S1369848618301456.

Holcomb:1993:SSS

[Hol93] Harmon R. Holcomb. Sociobiology, sex, and science. SUNY series in philosophy and biology. State University of New York Press, Albany, NY, USA, 1993. ISBN 0-7914-1259-8, 0-7914-1260-1 (paperback). x + 447 pp. LCCN GN365.9 .H65 1993.

Holm:2013:OAP

[Hol13] Sune Holm. Organism and artifact: Proper functions in Paley organisms. Studies in History and Philosophy of Biological and Biomedical Sciences, 44(4):706–713, December 2013. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848613000861.

Holm:2016:BBB

[Hol16] Sune Holm. Bridging bioethics and biology. Studies in History and Philosophy of Biological and Biomedical Sciences, 55(??): 133–136, February 2016. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848615001661.

Holmes:2017:WTC

[Hol17] Tarquin Holmes. The wild type as concept and in experimental practice: a history of its role in classical genetics and evolutionary theory. Studies in History and Philosophy of Biological and Biomedical Sciences, 63(??):15–27, June 2017. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848617300456.

Holmes:2019:INA

[Hol19] Matthew Holmes. Imitating nature: Analogy and experiment in D’Arcy Thompson’s Science of Form. Studies in History and Philosophy of Biological and Biomedical Sciences, 78(??):Article 101181, December 2019. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848618301687.

Homei:2008:SMM

[Hom08] Aya Homei. Specialization and medical mycology in the US, Britain and Japan. Studies in History and Philosophy of Biological and Biomedical Sciences, 39(1):80–92, March 2008. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S136984860700091X.

Honenberger:2015:GHT

[Hon15] Phillip Honenberger. Grene and Hull on types and typological thinking in biology. Studies in History and Philosophy of Biological and Biomedical Sciences, 50(??):13–25, April 2015. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848615000291.

Hoover:2012:CSH

[Hoo12] Kevin D. Hoover. Causal structure and hierarchies of models. Studies in History and Philosophy of Biological and Biomedical Sciences, 43(4):778–786, December 2012. CO-DEN ???? ISSN 1369-8486 (print), 1879-2499 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S1369848612000635.

Horder:2008:GA

[Hor08] Tim. J. Horder. A growing alienation? Studies in History and Philosophy of Biological and Biomedical Sciences, 39(3):374– 377, September 2008. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848608000472.

Horwitz:2011:NPN

[Hor11] Allan V. Horwitz. Naming the problem that has no name: creating targets for standardized drugs. Studies in History and Philosophy of Biological and Biomedical Sciences, 42(4):427–433, December 2011. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848611000380.

Hourdequin:2005:TTP

[Hou05] Marion Hourdequin. Theories as tools: a pluralistic approach to ecological modeling. Studies in History and Philosophy of Biological and Biomedical Sciences, 36(3):594–601, September 2005. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848605000543.

Holm:2013:OMA

[HP13] Sune Holm and Russell Powell. Organism, machine, artifact: the conceptual and normative challenges of synthetic biology. Studies in History and Philosophy of Biological and Biomedical Sciences, 44(4):627–631, December 2013. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848613000770.

Huzair:2017:BTV

[HS17] Farah Huzair and Steve Sturdy. Biotechnology and the transformation of vaccine innovation: the case of the hepatitis B vaccines 1968–2000. Studies in History and Philosophy of Biological and Biomedical Sciences, 64(??):11–21, August 2017. CODEN ???? ISSN 1369-8486 (print), 1879-2499 (electronic). URL /science/

article/pii/S1369848617300857.

Hull:2001:RTB

[Hul01] David L. Hull. The role of theories in biological systematics. Studies in History and Philosophy of Biological and Biomedical Sciences, 32(2):221–238, June 2001. CO-DEN ???? ISSN 1369-8486 (print), 1879-2499 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S1369848601000061.

Hull:2011:DD

[Hul11] David L. Hull. Defining Darwinism. Studies in History and Philosophy of Biological and Biomedical Sciences, 42(1):2–4, March 2011. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848610001020.

Huneman:2006:NPC

[Hun06] Philippe Huneman. Naturalising purpose: From comparative anatomy to the ‘adventure of reason’. Studies in History and Philosophy of Biological and Biomedical Sciences, 37(4):649– 674, December 2006. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848606000707.

Huneman:2013:ASV

[Hun13] Philippe Huneman. Assessing statistical views of natural selection: Room for non-local causation? Studies in History and Philosophy of Biological and Biomedical Sciences, 44(4):604– 612, December 2013. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848613001192.

Huneman:2019:RDT

[Hun19] Philippe Huneman. Revisiting Darwinian teleology: a case for inclusive fitness as design explanation. Studies in History and Philosophy of Biological and Biomedical Sciences, 76(??):Article 101188, August 2019. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848618301201.

Hutton:2012:CPM

[Hut12] Joseph Hutton. Composite paradigms in medicine: Analysing Gillies’ claim of reclassification of disease without paradigm shift in the case of Helicobacter pylori. Studies in History and Philosophy of Biological and Biomedical Sciences, 43(3):643– 654, September 2012. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848612000593.

Hyman:2012:VCA

[Hym12] Ludmila Hyman. Vygotsky’s Crisis: Argument, context, relevance. Studies in History and Philosophy of Biological and Biomedical Sciences, 43(2):473–482, June 2012. CO-DEN ???? ISSN 1369-8486 (print), 1879-2499 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S1369848611001336.

Iida:2020:BRP

[Iid20] Kaori Iida. Book review: Peaceful atoms in Japan: Radioisotopes as shared technical and sociopolitical resources for the Atomic Bomb Casualty Commission and the Japanese scientific community in the 1950s. Studies in History and Philosophy of Biological and Biomedical Sciences, 80(??):Article 101240, April 2020. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848618301699.

Ingold:1999:THL

[Ing99] Tim Ingold. ‘tools for the hand, language for the face’: An appreciation of Leroi-Gourhan’s gesture and speech. Studies in History and Philosophy of Biological and Biomedical Sciences, 30(4): 411–453, December 1999. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848699000229.

Inkpen:2016:LHH

[Ink16] S. Andrew Inkpen. Like Hercules and the Hydra: Tradeoffs and strategies in ecological model-building and experimental design. Studies in History and Philosophy of Biological and Biomedical Sciences, 57(??):34–43, June 2016. CODEN ???? ISSN 1369-8486 (print), 1879-2499 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S1369848616000388.

Ishida:2020:OST

[IR20] Yoichi Ishida and Alirio Rosales. The origins of the stochastic theory of population genetics: the Wright–Fisher model. Studies in History and Philosophy of Biological and Biomedical Sciences, 79(??):Article 101226, February 2020. CO-DEN ???? ISSN 1369-8486 (print), 1879-2499 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S1369848619300639.

Illari:2012:I

[IRR12] Phyllis Illari, Julian Reiss, and Federica Russo. Introduction. Studies in History and Philosophy of Biological and Biomedical Sciences, 43(4):758–760, December 2012. CO-DEN ???? ISSN 1369-8486 (print), 1879-2499 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S1369848612000957.

Ishida:2017:SWS

[Ish17] Yoichi Ishida. Sewall Wright, shifting balance theory, and the hardening of the modern synthesis. Studies in History and Philosophy of Biological and Biomedical Sciences, 61(??):1–10, February 2017. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848616300991. See [Ste15b].

Istvan:2013:GTP

[Ist13] M. A. Istvan, Jr. Gould talking past Dawkins on the unit of selection issue. Studies in History and Philosophy of Biological and Biomedical Sciences, 44(3):327–335, September 2013. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848613000885.

Illari:2010:FOC

[IW10] Phyllis McKay Illari and Jon Williamson. Function and organization: comparing the mechanisms of protein synthesis and natural selection. Studies in History and Philosophy of Biological and Biomedical Sciences, 41(3):279–291, September 2010. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848610000336.

Jimenez-Alonso:2008:ESP

[JA08] Belén Jiménez-Alonso. Eugenics, sexual pedagogy and social change: constructing the responsible subject of governmentality in the Spanish Second Republic. Studies in History and Philosophy of Biological and Biomedical Sciences, 39(2):247–254, June 2008. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848608000228.

Jackson:2003:AH

[Jac03a] Mark Jackson. Allergy and history. Studies in History and Philosophy of Biological and Biomedical Sciences, 34(3):383– 398, September 2003. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848603000475.

Jackson:2003:JFH

[Jac03b] Mark Jackson. John Freeman, hay fever and the origins of clinical allergy in Britain, 1900–1950. Studies in History and Philosophy of Biological and Biomedical Sciences, 34(3):473–490, September 2003. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848603000505.

Jamieson:2010:IRA

[Jam10] Michelle Jamieson. Imagining ‘reactivity’: allergy within the history of immunology. Studies in History and Philosophy of Biological and Biomedical Sciences, 41(4):356–366, December 2010. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848610000786.

Jeffares:2008:TTR

[Jef08] Ben Jeffares. Testing times: regularities in the historical sciences. Studies in History and Philosophy of Biological and Biomedical Sciences, 39(4):469–475, December 2008. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848608000769.

Jermyn:2013:LST

[Jer13] Deborah Jermyn. Labs and slabs: Television crime drama and the quest for forensic realism. Studies in History and Philosophy of Biological and Biomedical Sciences, 44(1):103–109, March 2013. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848612001008.

Jiang:2017:ONS

[Jia17] Lijing Jiang. The old, the new and the state in the making of modern East Asian medicine. Studies in History and Philosophy of Biological and Biomedical Sciences, 64(??):88–91, August 2017. CODEN ???? ISSN 1369-8486 (print), 1879-2499 (elec-

tronic). URL /science/article/pii/S1369848617300432.

Johnson:2001:DBI

[Joh01] Martin H. Johnson. The developmental basis of identity. Studies in History and Philosophy of Biological and Biomedical Sciences, 32(4):601–617, December 2001. CO-DEN ???? ISSN 1369-8486 (print), 1879-2499 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S1369848601000243.

John:2004:B

[Joh04] Stephen John. Bioethics. Studies in History and Philosophy of Biological and Biomedical Sciences, 35(1):177–184, March 2004. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848603001171.

John:2012:NGP

[Joh12] S. D. John. No genes, please: we’re British. Studies in History and Philosophy of Biological and Biomedical Sciences, 43(4):828– 830, December 2012. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848612001124.

Jones:2018:ADH

[Jon18] Elizabeth D. Jones. Ancient DNA: a history of the science before Jurassic Park. Studies in History and Philosophy of Biological and Biomedical Sciences, 68–69(??):ii, April/ June 2018. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL https://www.sciencedirect.com/science/
article/pii/S1369848617300833.

Jukola:2017:IOJ

[Juk17] Saana Jukola. On ideals of objectivity, judgments, and bias in medical research — a comment on Stegenga. Studies in History and Philosophy of Biological and Biomedical Sciences, 62(??):35– 41, April 2017. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848616300966.

Jukola:2019:ESN

[Juk19] Saana Jukola. On the evidentiary standards for nutrition advice. Studies in History and Philosophy of Biological and Biomedical Sciences, 73(??):1–9, February 2019. CODEN ???? ISSN 1369-8486 (print), 1879-2499 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S1369848617301942.

Juzda:2009:SSS

[Juz09] Elise Juzda. Skulls, science, and the spoils of war: craniological studies at the United States Army Medical Museum, 1868–1900. Studies in History and Philosophy of Biological and Biomedical Sciences, 40(3):156–167, September 2009. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848609000387.

Kraft:2003:ETD

[KA03] Alison Kraft and Samuel J. M. M. Alberti. ‘equal though different’: laboratories, museums and the institutional development of biology in late-Victorian Northern England. Studies in History and Philosophy of Biological and Biomedical Sciences, 34 (2):203–236, June 2003. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848603000256.

Kaitaro:2004:BMI

[Kai04] Timo Kaitaro. Brain-mind identities in dualism and materialism: a historical perspective. Studies in History and Philosophy of Biological and Biomedical Sciences, 35(4):627–645, December 2004. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848604000688.

Kaiser:2018:EPH

[Kai18] Marie I. Kaiser. ENCODE and the parts of the human genome. Studies in History and Philosophy of Biological and Biomedical Sciences, 72(??):28–37, December 2018. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848617301760.

Kaplan:2013:ALC

[Kap13] Jonathan Michael Kaplan. Adaptive landscapes: Concepts, tools and metaphors (reviewing E. I. Svensson and R. Calsbeek (Eds.), The adaptive landscape in evolutionary biology). Studies in History and Philosophy of Biological and Biomedical Sciences, 44(4): 613–616, December 2013. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848613001155.

Karfakis:2018:BCM

[Kar18] Nikos Karfakis. The biopolitics of CFS/ME. Studies in History and Philosophy of Biological and Biomedical Sciences, 70(??): 10–19, August 2018. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S1369848617300705.

Kassell:2010:SSS

[Kas10] Lauren Kassell. Stars, spirits, signs: towards a history of astrology 1100–1800. Studies in History and Philosophy of Biological and Biomedical Sciences, 41(2):67–69, June 2010. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848610000105.

Kastenhofer:2013:TSS

[Kas13] Karen Kastenhofer. Two sides of the same coin? The (techno)epistemic cultures of systems and synthetic biology. Studies in History and Philosophy of Biological and Biomedical Sciences, 44(2):130–140, June 2013. CODEN ???? ISSN 1369-8486 (print), 1879-2499 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S1369848613000228.

Keating:2012:TMN

[KC12] Peter Keating and Alberto Cambrosio. Too many numbers: Microarrays in clinical cancer research. Studies in History and Philosophy of Biological and Biomedical Sciences, 43(1):37–51, March 2012. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S136984861100080X.

Keating:2016:TNB

[KCN16] Peter Keating, Alberto Cambrosio, and Nicole C. Nelson. “triple negative breast cancer”: Translational research and the (re)assembling of diseases in post-genomic medicine. Studies in History and Philosophy of Biological and Biomedical Sciences, 59 (??):20–34, October 2016. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848616300401.

Keirns:2003:BTN

[Kei03] Carla C. Keirns. Better than nature: the changing treatment of asthma and hay fever in the United States, 1910– 1945. Studies in History and Philosophy of Biological and Biomedical Sciences, 34(3):511–531, September 2003. CO-DEN ???? ISSN 1369-8486 (print), 1879-2499 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S1369848603000487.

Keller:2011:TSI

[Kel11a] Evelyn Fox Keller. Towards a science of informed matter. Studies in History and Philosophy of Biological and Biomedical Sciences, 42(2):174–179, June 2011. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848610001172.

Kelly:2011:ADS

[Kel11b] Laura Kelly. Anatomy dissections and student experience at Irish universities, c.1900s–1960s. Studies in History and Philosophy of Biological and Biomedical Sciences, 42(4):467–474, December 2011. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848611000756.

Keller:2017:CST

[Kel17] Evelyn Fox Keller. Climate science, truth, and democracy. Studies in History and Philosophy of Biological and Biomedical Sciences, 64(??):106–122, August 2017. CODEN ???? ISSN 1369-8486 (print), 1879-2499 (electronic). URL /science/article/

pii/S136984861730105X.

Kendig:2020:OVA

[Ken20] Catherine Kendig. Ontology and values anchor indigenous and grey nomenclatures: a case study in lichen naming practices among the Samí, Sherpa, Scots, and Okanagan. Studies in History and Philosophy of Biological and Biomedical Sciences, 84(??):Article 101340, December 2020. CODEN ???? ISSN 1369-8486 (print), 1879-2499 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S1369848620301515.

Kerasidou:2016:HES

[Ker16] Angeliki Kerasidou. Human embryonic stem cell research: Middle-ground positions and moral compromise. Studies in History and Philosophy of Biological and Biomedical Sciences, 57 (??):167–169, June 2016. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848616000170.

Keyes:1999:PCCa

[Key99a] Martha E. Keyes. The prion challenge to the ‘central dogma’ of molecular biology, 1965–1991: Part I: Prelude to prions. Studies in History and Philosophy of Biological and Biomedical Sciences, 30(1):1–19, March 1999. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848698000284.

Keyes:1999:PCCb

[Key99b] Martha E. Keyes. The prion challenge to the ‘central dogma’ of molecular biology, 1965–1991: Part II: The problem with prions. Studies in History and Philosophy of Biological and Biomedical Sciences, 30(2):181–218, June 1999. CODEN ???? ISSN 1369-8486 (print), 1879-2499 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S1369848698000296.

Kater:2003:HCE

[KHDW03] Loes Kater, Rob Houtepen, Raymond De Vries, and Guy Widdershoven. Health care ethics and health law in the Dutch discussion on end-of-life decisions: a historical analysis of the dynamics and development of both disciplines. Studies in History and Philosophy of Biological and Biomedical Sciences, 34(4):669– 684, December 2003. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848603000700.

Kidd:2013:PCS

[Kid13] Ian James Kidd. A pluralist challenge to “integrative medicine”: Feyerabend and Popper on the cognitive value of alternative medicine. Studies in History and Philosophy of Biological and Biomedical Sciences, 44(3):392–400, September 2013. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848613000733.

Kidd:2014:WSW

[Kid14] Ian James Kidd. Was Sir William Crookes epistemically virtuous? Studies in History and Philosophy of Biological and Biomedical Sciences, 48 (part A)(??):67–74, December 2014. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848614000788.

Kidd:2017:PIP

[Kid17] Ian James Kidd. Phenomenology of illness, philosophy, and life. Studies in History and Philosophy of Biological and Biomedical Sciences, 62(??):56–60, April 2017. CO-DEN ???? ISSN 1369-8486 (print), 1879-2499 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S1369848617300183.

Kinzel:2016:CCC

[Kin16] Katherina Kinzel. Counterfactuals, causes and contingency in the history of science. Studies in History and Philosophy of Biological and Biomedical Sciences, 60(??):92–96, December 2016. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848616300231.

Kirk:2008:WSG

[Kir08] Robert G. W. Kirk. ‘wanted — standard guinea pigs’: standardisation and the experimental animal market in Britain ca. 1919–1947. Studies in History and Philosophy of Biological and Biomedical Sciences, 39(3):280–291, September 2008. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S136984860800037X.

Kirby:2013:FFS

[Kir13] David A. Kirby. Forensic fictions: Science, television production, and modern storytelling. Studies in History and Philosophy of Biological and Biomedical Sciences, 44(1):92–102, March 2013. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S136984861200101X.

Kitanaka:2020:BRI

[Kit20] Junko Kitanaka. Book review: The Invention of Madness by Emily Baum: Recovering incommensurability: Theorizing psychiatry in Asia. Studies in History and Philosophy of Biological and Biomedical Sciences, 79(??):Article 101204, February 2020. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848619301177. See reply [Bau20].

Knuuttila:2013:BST

[KL13] Tarja Knuuttila and Andrea Loettgers. Basic science through engineering? Synthetic modeling and the idea of biologyinspired engineering. Studies in History and Philosophy of Biological and Biomedical Sciences, 44(2):158–169, June 2013. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848613000253.

Kyllingstad:2015:EPA

[KL15] Jon Røyne Kyllingstad and Ageliki Lefkaditou. Eugenics and physical anthropology in Hungary and Greece. Studies in History and Philosophy of Biological and Biomedical Sciences, 49(??): 70–74, February 2015. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848614001770.

Klasios:2014:EPH

[Kla14] John Klasios. The evolutionary psychology of human mating: a response to Buller’s critique. Studies in History and Philosophy of Biological and Biomedical Sciences, 47 (part A)(??):1– 11, September 2014. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848614000417.

Klein:2003:EHH

[Kle03] Ursula Klein. Experimental history and Herman Boerhaave’s chemistry of plants. Studies in History and Philosophy of Biological and Biomedical Sciences, 34(4):533–567, December 2003. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848603000712.

Klein:2016:CCG

[Kle16] Julie Thompson Klein. Conceptual clarification for Grand Challenges. Studies in History and Philosophy of Biological and Biomedical Sciences, 56(??):106–107, April 2016. CODEN ???? ISSN 1369-8486 (print), 1879-2499 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S1369848615001715.

Kleinman:2018:GEB

[Kle18] Kim Kleinman. Genera, evolution, and botanists in 1940: Edgar Anderson’s “Survey of Modern Opinion”. Studies in History and Philosophy of Biological and Biomedical Sciences, 67(??): 1–7, February 2018. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848616301200.

Knoeff:2015:TAH

[Kno15] Rina Knoeff. Touching anatomy: On the handling of preparations in the anatomical cabinets of Frederik Ruysch (1638–1731). Studies in History and Philosophy of Biological and Biomedical Sciences, 49(??):32–44, February 2015. CODEN ???? ISSN 1369-8486 (print), 1879-2499 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S1369848614001757.

Kosso:2001:ES

[Kos01] Peter Kosso. The epidemiology of science. Studies in History and Philosophy of Biological and Biomedical Sciences, 32(3):575– 581, September 2001. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S136984860100019X.

Kostyrka:2016:WRV

[Kos16] Gladys Kostyrka. What roles for viruses in origin of life scenarios? Studies in History and Philosophy of Biological and Biomedical Sciences, 59(??):135–144, October 2016. CODEN ???? ISSN 1369-8486 (print), 1879-2499 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S1369848616300073.

Kovaka:2017:DRP

[Kov17] Karen Kovaka. Different research projects require their own individuality concepts. Studies in History and Philosophy of Biological and Biomedical Sciences, 61(??):50–53, February 2017. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848616300577.

Kaplan:2015:GMR

[KPB15] Jonathan Michael Kaplan, Massimo Pigliucci, and Joshua Alexander Banta. Gould on Morton, redux: What can the debate reveal about the limits of data? Studies in History and Philosophy of Biological and Biomedical Sciences, 52(??):22–31, August 2015. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848615000035.

Kogge:2013:SBA

[KR13] Werner Kogge and Michael Richter. Synthetic biology and its alternatives. Descartes, Kant and the idea of engineering biological machines. Studies in History and Philosophy of Biological and Biomedical Sciences, 44(2):181–189, June 2013. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848613000277.

Krasz:2012:QVP

[Krá12] Lilla Krász. Quackery versus professionalism? Characters, places and media of medical knowledge in eighteenthcentury Hungary. Studies in History and Philosophy of Biological and Biomedical Sciences, 43(3):700–709, September 2012. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848612000301.

Krasnodebski:2014:CCF

[Kra14] Marcin Krasnodebski. Constructing creationists: French and British narratives and policies in the wake of the resurgence of anti-evolution movements. Studies in History and Philosophy of Biological and Biomedical Sciences, 47 (part A)(??):35– 44, September 2014. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848614000375.

Krementsov:2009:YHI

[Kre09] Nikolai Krementsov. Off with your heads: isolated organs in early Soviet science and fiction. Studies in History and Philosophy of Biological and Biomedical Sciences, 40(2):87–100, June 2009. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848609000156.

Krige:2002:BED

[Kri02] John Krige. The birth of EMBO and the difficult road to EMBL. Studies in History and Philosophy of Biological and Biomedical Sciences, 33(3):547–564, September 2002. CODEN ???? ISSN 1369-8486 (print), 1879-2499 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S1369848602000134.

Kroker:1999:IOA

[Kro99] Kenton Kroker. Immunity and its other: The anaphylactic selves of Charles Richet. Studies in History and Philosophy of Biological and Biomedical Sciences, 30(3):273–296, September 1999. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848699000126.

Kroker:2003:PIA

[Kro03] Kenton Kroker. The progress of introspection in America, 1896–1938. Studies in History and Philosophy of Biological and Biomedical Sciences, 34(1):77–108, March 2003. CODEN ???? ISSN 1369-8486 (print), 1879-2499 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S1369848602000729.

Krohs:2012:CE

[Kro12] Ulrich Krohs. Convenience experimentation. Studies in History and Philosophy of Biological and Biomedical Sciences, 43 (1):52–57, March 2012. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848611000811.

Krohs:2015:CFE

[Kro15] Ulrich Krohs. Can functionality in evolving networks be explained reductively? Studies in History and Philosophy of Biological and Biomedical Sciences, 53(??):94–101, ???? 2015. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S136984861500062X.

Koonin:2016:VAR

[KS16] Eugene V. Koonin and Petro Starokadomskyy. Are viruses alive? The replicator paradigm sheds decisive light on an old but misguided question. Studies in History and Philosophy of Biological and Biomedical Sciences, 59(??):125–134, October 2016. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848616300103.

Kuklick:2012:FF

[Kuk12] Henrika Kuklick. Family feud? Studies in History and Philosophy of Biological and Biomedical Sciences, 43(2):574–577, June 2012. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848612000027.

Knaapen:2008:BSP

[KW08] Loes Knaapen and George Weisz. The biomedical standardization of premenstrual syndrome. Studies in History and Philosophy of Biological and Biomedical Sciences, 39(1):120–134, March 2008. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848607000945.

Leonelli:2012:RTO

[LA12] Sabina Leonelli and Rachel A. Ankeny. Re-thinking organisms: The impact of databases on model organism biology. Studies in History and Philosophy of Biological and Biomedical Sciences, 43 (1):29–36, March 2012. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848611000793.

LaBonte:2014:AFV

[LaB14] Michelle L. LaBonte. Anticoagulant factor V: Factors affecting the integration of novel scientific discoveries into the broader framework. Studies in History and Philosophy of Biological and Biomedical Sciences, 47 (part A)(??):23–34, September 2014. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848614000363.

Lafferton:2007:MMF

[Laf07] Emese Lafferton. The Magyar moustache: the faces of Hungarian state formation, 1867–1918. Studies in History and Philosophy of Biological and Biomedical Sciences, 38(4):706–732, December 2007. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848607000507.

Landecker:2002:NTB

[Lan02] Hannah Landecker. New times for biology: nerve cultures and the advent of cellular life in vitro. Studies in History and Philosophy of Biological and Biomedical Sciences, 33(4):667–694, December 2002. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848602000262.

Lans:2008:MBM

[Lan08] Cheryl Lans. Man better man: the politics of disappearance. Studies in History and Philosophy of Biological and Biomedical Sciences, 39(4):429–436, December 2008. CODEN ???? ISSN 1369-8486 (print), 1879-2499 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S1369848608000770.

Landecker:2016:IWI

[Lan16] Hannah Landecker. It is what it eats: Chemically defined media and the history of surrounds. Studies in History and Philosophy of Biological and Biomedical Sciences, 57(??):148–160, June 2016. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848616000224.

LaPorte:2007:DS

[LaP07] Joseph LaPorte. In defense of species. Studies in History and Philosophy of Biological and Biomedical Sciences, 38(1):255–269, March 2007. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848606001002.

Laveaga:2005:UTS

[Lav05] Gabriela Soto Laveaga. Uncommon trajectories: steroid hormones, Mexican peasants, and the search for a wild yam. Studies in History and Philosophy of Biological and Biomedical Sciences, 36(4):743–760, December 2005. CO-DEN ???? ISSN 1369-8486 (print), 1879-2499 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S1369848605000774.

Lawrence:2012:PMP

[Law12] Natalie Lawrence. The Prime Minister and the platypus: a paradox goes to war. Studies in History and Philosophy of Biological and Biomedical Sciences, 43(1):290–297, March 2012. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848611000768.

Lawrence:2014:PWO

[Law14] Natalie Lawrence. Plumed wonders and ornithological passions. Studies in History and Philosophy of Biological and Biomedical Sciences, 47 (part A)(??):206–209, September 2014. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848614000399.

Lopez-Beltran:2006:SSH

[LB06] Carlos López-Beltrán. Storytelling, statistics and hereditary thought: the narrative support of early statistics. Studies in History and Philosophy of Biological and Biomedical Sciences, 37(1):41–58, March 2006. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S136984860500097X.

Lorusso:2008:CHB

[LB08] Ludovica Lorusso and Giovanni Boniolo. Clustering humans: on biological boundaries. Studies in History and Philosophy of Biological and Biomedical Sciences, 39(1):163–170, March 2008. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848607000982.

Lorusso:2015:RRS

[LB15] Ludovica Lorusso and Fabio Bacchini. A reconsideration of the role of self-identified races in epidemiology and biomedical research. Studies in History and Philosophy of Biological and Biomedical Sciences, 52(??):56–64, August 2015. CODEN ???? ISSN 1369-8486 (print), 1879-2499 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S1369848615000345.

Leeming:2018:SCD

[LB18] William Leeming and Ana Barahona. Synthesis, convergence, and differences in the entangled histories of cytogenetics in medicine: a comparative study of Canada and Mexico. Studies in History and Philosophy of Biological and Biomedical Sciences, 71 (??):8–16, October 2018. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848617300651.

Morvan:2018:SBS

[Le 18] Pierre Le Morvan. Searle on the biology of seeing. Studies in History and Philosophy of Biological and Biomedical Sciences, 71 (??):26–31, October 2018. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848617302273.

Lean:2019:CAC

[Lea19] Oliver M. Lean. Chemical arbitrariness and the causal role of molecular adapters. Studies in History and Philosophy of Biological and Biomedical Sciences, 78(??):Article 101180, December 2019. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848618301808.

Leeming:2005:IAH

[Lee05] William Leeming. Ideas about heredity, genetics, and ‘medical genetics’ in Britain, 1900–1982. Studies in History and Philosophy of Biological and Biomedical Sciences, 36(3):538– 558, September 2005. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S136984860500049X.

Leeming:2010:TSS

[Lee10] William Leeming. Tracing the shifting sands of ‘medical genetics’: what’s in a name? Studies in History and Philosophy of Biological and Biomedical Sciences, 41(1):50–60, March 2010. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848609000727.

Lee:2015:USM

[Lee15] Victoria Lee. Unraveling the search for microbial control in twentieth-century pandemics. Studies in History and Philosophy of Biological and Biomedical Sciences, 53(??):122–125, ???? 2015. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848615001089.

Lefkaditou:2017:PRP

[Lef17] Ageliki Lefkaditou. Practicing race and photography. Studies in History and Philosophy of Biological and Biomedical Sciences, 64(??):92–96, August 2017. CODEN ???? ISSN 1369-8486 (print), 1879-2499 (electronic). URL /science/article/pii/

S1369848617300778.

Leland:2020:KOR

[Lel20] Patrick R. Leland. Kant, organisms, and representation. Studies in History and Philosophy of Biological and Biomedical Sciences, 79(??):Article 101223, February 2020. CO-DEN ???? ISSN 1369-8486 (print), 1879-2499 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S1369848617301140.

Lemeire:2016:BRD

[Lem16] Olivier Lemeire. Beyond the realism debate: the metaphysics of ‘racial’ distinctions. Studies in History and Philosophy of Biological and Biomedical Sciences, 59(??):47–56, October 2016. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S136984861630053X.

Lemoine:2017:AEP

[Lem17] Maël Lemoine. Animal extrapolation in preclinical studies: an analysis of the tragic case of TGN1412. Studies in History and Philosophy of Biological and Biomedical Sciences, 61(??):35–45, February 2017. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848616301224.

Leonelli:2012:IMS

[Leo12] S. Leonelli. Introduction: Making sense of data-driven research in the biological and biomedical sciences. Studies in History and Philosophy of Biological and Biomedical Sciences, 43(1):1–3, March 2012. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S136984861100077X.

Leonelli:2013:IDA

[Leo13] Sabina Leonelli. Integrating data to acquire new knowledge: Three modes of integration in plant science. Studies in History and Philosophy of Biological and Biomedical Sciences, 44(4): 503–514, December 2013. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848613000344.

Levy:2006:WEW

[Lev06] Neil Levy. What evolves when morality evolves? Studies in History and Philosophy of Biological and Biomedical Sciences, 37(3):

612–620, September 2006. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848606000549.

Lewens:2000:FTA

[Lew00] Tim Lewens. Function talk and the artefact model. Studies in History and Philosophy of Biological and Biomedical Sciences, 31 (1):95–111, March 2000. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848699000400.

Lewens:2002:DAO

[Lew02] Tim Lewens. Development aid: on ontogeny and ethics. Studies in History and Philosophy of Biological and Biomedical Sciences, 33(2):195–217, July 2002. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0039368102000171.

Lewens:2004:CEE

[Lew04a] Tim Lewens. The commercial exploitation of ethics. Studies in History and Philosophy of Biological and Biomedical Sciences, 35 (1):145–153, March 2004. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848603001146.

Lewens:2004:SWB

[Lew04b] Tim Lewens. Is something wrong with bioethics? Studies in History and Philosophy of Biological and Biomedical Sciences, 35 (1):121–123, March 2004. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848603001110.

Lewens:2012:SEE

[Lew12] Tim Lewens. Species, essence and explanation. Studies in History and Philosophy of Biological and Biomedical Sciences, 43(4):751– 757, December 2012. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848612001070.

Lewens:2013:BBS

[Lew13] Tim Lewens. From bricolage to BioBricks^{T M}: Synthetic biology and rational design. Studies in History and Philosophy of Biological and Biomedical Sciences, 44(4):641–648, December 2013. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848613000794.

Lewens:2019:NPB

[Lew19] Tim Lewens. Neo-Paleyan biology. Studies in History and Philosophy of Biological and Biomedical Sciences, 76(??):Article 101185, August 2019. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848618301353.

Lewens:2020:PPG

[Lew20] Tim Lewens. Pain, Pleasure and the Greater Good by Cathy Gere: the abhorrent consequences of consequentialism. Studies in History and Philosophy of Biological and Biomedical Sciences, 83(??):Article 101250, October 2020. CODEN ???? ISSN 1369-8486 (print), 1879-2499 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S1369848619302109.

Lachapelle:2010:HZW

[LH10] Sofie Lachapelle and Jenna Healey. On Hans, Zou and the others: wonder animals and the question of animal intelligence in early twentieth-century France. Studies in History and Philosophy of Biological and Biomedical Sciences, 41(1):12–20, March 2010. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848609000703.

Lightman:2016:HVS

[Lig16] Bernard Lightman. The “history” of Victorian scientific naturalism: Huxley, Spencer and the “end” of natural history. Studies in History and Philosophy of Biological and Biomedical Sciences, 58 (??):17–23, August 2016. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848615001776.

Lightman:2020:IMD

[Lig20] Bernard Lightman. The Invention of the Modern Dog by Michael Worboys, Julie-Marie Strange, and Neil Pemberton: (Breeding Spectacle). Studies in History and Philosophy of Biological and Biomedical Sciences, 84(??):Article 101329, December 2020. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848620301400.

Lillehammer:2001:GE

[Lil01] Hallvard Lillehammer. From genes to eugenics. Studies in History and Philosophy of Biological and Biomedical Sciences, 32(4):589– 600, December 2001. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848601000255.

Lillehammer:2004:JEA

[Lil04a] Hallvard Lillehammer. Jamieson on the ethics of animals and the environment. Studies in History and Philosophy of Biological and Biomedical Sciences, 35(4):743–751, December 2004. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848604000743.

Lillehammer:2004:WNB

[Lil04b] Hallvard Lillehammer. Who needs bioethicists? Studies in History and Philosophy of Biological and Biomedical Sciences, 35 (1):131–144, March 2004. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848603001134.

Lindenmann:2001:SSF

[Lin01] Jean Lindenmann. Siegel, Schaudinn, Fleck and the etiology of syphilis. Studies in History and Philosophy of Biological and Biomedical Sciences, 32(3):435–455, September 2001. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848601000140. See rejoinder [vdB02] and

reply [Lin02].

Lindenmann:2002:SSF

[Lin02] Jean Lindenmann. Siegel, Schaudinn, Fleck and the etiology of syphilis: a response to Henk van den Belt. Studies in History and Philosophy of Biological and Biomedical Sciences, 33(4):751– 752, December 2002. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S1369848602000183. See [Lin01,

vdB02].

Lindee:2014:SHG

[Lin14] Susan Lindee. Scaling up: Human genetics as a Cold War network. Studies in History and Philosophy of Biological and Biomedical Sciences, 47 (part A)(??):185–190, September 2014. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848614000582.

Lindee:2016:HGA

[Lin16] Susan Lindee. Human genetics after the bomb: Archives, clinics, proving grounds and board rooms. Studies in History and Philosophy of Biological and Biomedical Sciences, 55(??):45–53, February 2016. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S136984861500117X.

Lipphardt:2014:GDP

[Lip14] Veronika Lipphardt. “Geographical Distribution Patterns of Various Genes”: Genetic studies of human variation after 1945. Studies in History and Philosophy of Biological and Biomedical Sciences, 47 (part A)(??):50–61, September 2014. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848614000466.

Liu:2017:SB

[Liu17] Daniel Liu. This is the synthetic biology that is. Studies in History and Philosophy of Biological and Biomedical Sciences, 63(??):89–93, June 2017. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848617300274.

Livingstone:2004:PSS

[Liv04] David N. Livingstone. Public spectacle and scientific theory: William Robertson Smith and the reading of evolution in Victorian Scotland. Studies in History and Philosophy of Biological and Biomedical Sciences, 35(1):1–29, March 2004. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848603001067.

Love:2013:DII

[LL13] Alan C. Love and Gary L. Lugar. Dimensions of integration in interdisciplinary explanations of the origin of evolutionary novelty. Studies in History and Philosophy of Biological and Biomedical Sciences, 44(4):537–550, December 2013. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848613001234.

Leychkis:2009:WS

[LMR09] Yan Leychkis, Stephen R. Munzer, and Jessica L. Richardson. What is stemness? Studies in History and Philosophy of Biological and Biomedical Sciences, 40(4):312–320, December 2009. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S136984860900051X.

LoTufo:2003:INS

[LoT03] Ilaria LoTufo. Images of the natural (and social) universe in Rétif De La Bretonne’s La découverte australe. Studies in History and Philosophy of Biological and Biomedical Sciences, 34 (1):1–50, March 2003. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848602000730. See [R81, ?].

Loughlin:2003:SSE

[Lou03] Kelly Loughlin. Science on stage: expert advice as public drama. Studies in History and Philosophy of Biological and Biomedical Sciences, 34(2):375–380, June 2003. CO-DEN ???? ISSN 1369-8486 (print), 1879-2499 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S1369848603000219.

Louis:2016:CCH

[Lou16] Ard A. Louis. Contingency, convergence and hyper-astronomical numbers in biological evolution. Studies in History and Philosophy of Biological and Biomedical Sciences, 58(??):107–116, August 2016. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848615001879.

Lowy:2002:EP

[Löw02] Ilana Löwy. Epidemics and populations. Studies in History and Philosophy of Biological and Biomedical Sciences, 33(1):187–194, March 2002. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848601000371.

Lowy:2003:GPD

[Löw03] Ilana Löwy. On guinea pigs, dogs and men: anaphylaxis and the study of biological individuality, 1902–1939. Studies in History and Philosophy of Biological and Biomedical Sciences, 34(3):399– 423, September 2003. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848603000530.

Lowy:2004:RCO

[Löw04a] Ilana Löwy. ‘A river that is cutting its own bed’: the serology of syphilis between laboratory, society and the law. Studies in History and Philosophy of Biological and Biomedical Sciences, 35(3): 509–524, September 2004. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848604000500.

Lowy:2004:ILF

[Löw04b] Ilana Löwy. Introduction: Ludwik Fleck’s epistemology of medicine and biomedical sciences. Studies in History and Philosophy of Biological and Biomedical Sciences, 35(3):437–445, September 2004. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848604000494.

Lowy:2005:BCD

[Löw05] Ilana Löwy. Biotherapies of chronic diseases in the inter-war period: from Witte’s peptone to Penicillium extract. Studies in History and Philosophy of Biological and Biomedical Sciences, 36(4):675–695, December 2005. CODEN ???? ISSN 1369-8486 (print), 1879-2499 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S1369848605000749.

Lowy:2010:RR

[Löw10] Ilana Löwy. Reproductive revolutions. Studies in History and Philosophy of Biological and Biomedical Sciences, 41(4):422– 424, December 2010. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848610000713.

Lowy:2012:DPB

[Löw12] Ilana Löwy. Defusing the population bomb in the 1950s: Foam tablets in India. Studies in History and Philosophy of Biological and Biomedical Sciences, 43(3):583–593, September 2012. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848612000465.

Lowy:2014:HGC

[Löw14a] Ilana Löwy. How genetics came to the unborn: 1960– 2000. Studies in History and Philosophy of Biological and Biomedical Sciences, 47 (part A)(??):154–162, September 2014. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848614000557.

Lowy:2014:PDI

[Löw14b] Ilana Löwy. Prenatal diagnosis: the irresistible rise of the ‘visible fetus’. Studies in History and Philosophy of Biological and Biomedical Sciences, 47 (part B)(??):290–299, September 2014. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848613001775.

Lowe:2016:NDE

[Low16] James W. E. Lowe. Normal development and experimental embryology: Edmund beecher Wilson and amphioxus. Studies in History and Philosophy of Biological and Biomedical Sciences, 57(??):44–59, June 2016. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848616300152.

Lowe:2018:STT

[Low18] James W. E. Lowe. Sequencing through thick and thin: Historiographical and philosophical implications. Studies in History and Philosophy of Biological and Biomedical Sciences, 72(??): 10–27, December 2018. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848617301802.

Ling:2006:PUI

[LR06] Tom Ling and Ann Raven. Pharmacogenetics and uncertainty: implications for policy makers. Studies in History and Philosophy of Biological and Biomedical Sciences, 37(3):533–549, September 2006. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848606000495.

Ludwig:2014:HRP

[Lud14] David Ludwig. Hysteria, race, and phlogiston. A model of ontological elimination in the human sciences. Studies in History and Philosophy of Biological and Biomedical Sciences, 45(??):68–77, March 2014. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848613001325.

Luque:2016:PSW

[Luq16] Victor J. Luque. The Principle of Stasis: Why drift is not a Zero-Cause Law. Studies in History and Philosophy of Biological and Biomedical Sciences, 57(??):71–79, June 2016. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848616300176.

Lyman:2019:MGC

[Lym19] R. Lee Lyman. Misunderstanding graphs: the confusion of biological clade diversity diagrams and archaeological frequency seriation diagrams. Studies in History and Philosophy of Biological and Biomedical Sciences, 77(??):Article 101178, October 2019. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848618301766.

Lynch:2013:STF

[Lyn13] Michael Lynch. Science, truth, and forensic cultures: the exceptional legal status of DNA evidence. Studies in History and Philosophy of Biological and Biomedical Sciences, 44(1):60–70, March 2013. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848612001021.

Lyon:2007:QCL

[Lyo07] Pamela Lyon. From quorum to cooperation: lessons from bacterial sociality for evolutionary theory. Studies in History and Philosophy of Biological and Biomedical Sciences, 38(4):820– 833, December 2007. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848607000520.

Lowy:2000:MSI

[LZ00] Ilana Löwy and Patrick Zylberman. Medicine as a social instrument: Rockefeller Foundation, 1913–45. Studies in History and Philosophy of Biological and Biomedical Sciences, 31(3):365– 379, September 2000. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S136984860000011X.

Macdonald:2002:WMY

[Mac02] Helen Macdonald. ‘what makes you a scientist is the way you look at things’: ornithology and the observer 1930–1955. Studies in History and Philosophy of Biological and Biomedical Sciences, 33(1):53–77, March 2002. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848601000346.

Machery:2007:YPC

[Mac07] Edouard Machery. 100 years of psychology of concepts: the theoretical notion of concept and its operationalization. Studies in History and Philosophy of Biological and Biomedical Sciences, 38 (1):63–84, March 2007. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848606000926.

Mackenzie:2013:SBT

[Mac13] Adrian Mackenzie. Synthetic biology and the technicity of biofuels. Studies in History and Philosophy of Biological and Biomedical Sciences, 44(2):190–198, June 2013. CODEN ???? ISSN 1369-8486 (print), 1879-2499 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S1369848613000289.

MacCord:2015:WVE

[Mac15] Kate MacCord. Whose view of embryos? Studies in History and Philosophy of Biological and Biomedical Sciences, 53(??):

103–106, ???? 2015. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848615001090.

Maerker:2012:FAM

[Mae12] Anna Maerker. Florentine anatomical models and the challenge of medical authority in late-eighteenth-century Vienna. Studies in History and Philosophy of Biological and Biomedical Sciences, 43(3):730–740, September 2012. CO-DEN ???? ISSN 1369-8486 (print), 1879-2499 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S1369848612000295.

Magnus:2005:HRE

[Mag05] P. D. Magnus. Hormone research as an exemplar of underdetermination. Studies in History and Philosophy of Biological and Biomedical Sciences, 36(3):559–567, September 2005. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848605000506.

Magnus:2014:ECC

[Mag14] P. D. Magnus. Epistemic categories and causal kinds. Studies in History and Philosophy of Biological and Biomedical Sciences, 48 (part B)(??):263–266, December 2014. CO-DEN ???? ISSN 1369-8486 (print), 1879-2499 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S1369848614001411.

Maienschein:2014:SVT

[Mai14] Jane Maienschein. A surgeon’s view of transplantation. Studies in History and Philosophy of Biological and Biomedical Sciences, 46(??):104–106, June 2014. CODEN ???? ISSN 1369-8486 (print), 1879-2499 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S1369848614000089.

Maienschein:2016:EMS

[Mai16] Jane Maienschein. Embryos, microscopes, and society. Studies in History and Philosophy of Biological and Biomedical Sciences, 57 (??):129–136, June 2016. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848616000212.

Mameli:2006:NEB

[Mam06] Matteo Mameli. Norms for emotions: biological functions and representational contents. Studies in History and Philosophy of Biological and Biomedical Sciences, 37(1):101–121, March 2006. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S136984860500052X.

Manson:2004:BVN

[Man04] Neil C. Manson. Brains, vats, and neurally-controlled animats. Studies in History and Philosophy of Biological and Biomedical Sciences, 35(2):249–268, June 2004. CO-DEN ???? ISSN 1369-8486 (print), 1879-2499 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S1369848604000202.

Manus:2012:DME

[Man12] Fabrizzio Mc Manus. Development and mechanistic explanation. Studies in History and Philosophy of Biological and Biomedical Sciences, 43(2):532–541, June 2012. CO-DEN ???? ISSN 1369-8486 (print), 1879-2499 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S1369848611001415.

Manias:2015:PCP

[Man15] Chris Manias. The problematic construction of ‘Palaeolithic Man’: the Old Stone Age and the difficulties of the comparative method, 1859–1914. Studies in History and Philosophy of Biological and Biomedical Sciences, 51(??):32–43, June 2015. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S136984861500028X.

Manias:2017:PLH

[Man17] Chris Manias. Progress in life’s history: Linking Darwinism and palaeontology in Britain, 1860–1914. Studies in History and Philosophy of Biological and Biomedical Sciences, 66(??):18–26, December 2017. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848617300602.

Marshall:1998:PNH

[Mar98] Alan Marshall. A postmodern natural history of the world: eviscerating the GUTs from ecology and environmentalism. Studies in History and Philosophy of Biological and Biomedical Sciences, 29(1):137–164, March 1998. CODEN ???? ISSN 1369-8486 (print), 1879-2499 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S1369848698000077.

Marks:2016:SRR

[Mar16] Jonathan Marks. Solving the riddle of race. Studies in History and Philosophy of Biological and Biomedical Sciences, 59(??): 161–164, October 2016. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848616300188.

Martens:2019:HMC

[Mar19a] Johannes Martens. Hamilton meets causal decision theory. Studies in History and Philosophy of Biological and Biomedical Sciences, 77(??):Article 101187, October 2019. CODEN ???? ISSN 1369-8486 (print), 1879-2499 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S1369848618301304.

Martinez:2019:DC

[Mar19b] Manolo Martínez. Deception as cooperation. Studies in History and Philosophy of Biological and Biomedical Sciences, 77 (??):Article 101184, October 2019. CODEN ???? ISSN 1369-8486 (print), 1879-2499 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S1369848618301602.

Marlor:2020:EKP

[Mar20] Chantelle Marlor. Explaining knowledge pluralisms; the intertwining of culture and materiality. Studies in History and Philosophy of Biological and Biomedical Sciences, 84(??):Article 101339, December 2020. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848620301503.

Matthews:2016:CGB

[Mat16] Lucas J. Matthews. On closing the gap between philosophical concepts and their usage in scientific practice: a lesson from the debate about natural selection as mechanism. Studies in History and Philosophy of Biological and Biomedical Sciences, 55(??): 21–28, February 2016. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848615001739.

Mathew:2017:IHO

[Mat17] John Mathew. The implications of human and other animal displays in U.S. based museums. Studies in History and Philosophy of Biological and Biomedical Sciences, 66(??):87–93, December 2017. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848617301620.

Maund:2000:PFA

[Mau00] Barry Maund. Proper functions and Aristotelian functions in biology. Studies in History and Philosophy of Biological and Biomedical Sciences, 31(1):155–178, March 2000. CODEN ???? ISSN 1369-8486 (print), 1879-2499 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S1369848699000382.

Maung:2016:DCE

[Mau16a] Hane Htut Maung. Diagnosis and causal explanation in psychiatry. Studies in History and Philosophy of Biological and Biomedical Sciences, 60(??):15–24, December 2016. CODEN ???? ISSN 1369-8486 (print), 1879-2499 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S136984861630067X.

Maung:2016:WDP

[Mau16b] Hane Htut Maung. To what do psychiatric diagnoses refer? a two-dimensional semantic analysis of diagnostic terms. Studies in History and Philosophy of Biological and Biomedical Sciences, 55 (??):1–10, February 2016. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848615001399.

Maung:2020:PIS

[Mau20] Hane Htut Maung. Pluralism and incommensurability in suicide research. Studies in History and Philosophy of Biological and Biomedical Sciences, 80(??):Article 101247, April 2020. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848619300834.

Mazzarello:2016:VGN

[Maz16] Paolo Mazzarello. A Vesalian guide to neuroscience. Studies in History and Philosophy of Biological and Biomedical Sciences, 55(??):121–123, February 2016. CODEN ???? ISSN 1369-8486 (print), 1879-2499 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S136984861500165X.

McGhee:2016:CED

[McG16] George R. McGhee, Jr. Can evolution be directional without being teleological? Studies in History and Philosophy of Biological and Biomedical Sciences, 58(??):93–99, August 2016. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S136984861500179X.

McLean:2001:GGG

[McL01] Sheila A. M. McLean. The gene genie: good fairy or wicked witch? Studies in History and Philosophy of Biological and Biomedical Sciences, 32(4):723–739, December 2001. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848601000309.

McLeish:2007:R

[McL07] Christina McLeish. Am I a rodent? Studies in History and Philosophy of Biological and Biomedical Sciences, 38(3):668– 677, September 2007. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848607000362.

McLaughlin:2008:RPN

[McL08] Peter McLaughlin. Reverend Paley’s naturalist revival. Studies in History and Philosophy of Biological and Biomedical Sciences, 39(1):25–37, March 2008. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848607000878.

McLeish:2011:PMB

[McL11] Tom McLeish. Physics met biology, and the consequence was …. Studies in History and Philosophy of Biological and Biomedical Sciences, 42(2):190–192, June 2011. CO-DEN ???? ISSN 1369-8486 (print), 1879-2499 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S1369848610001196.

McLoone:2013:SET

[McL13] Brian McLoone. Selection explanations of token traits. Studies in History and Philosophy of Biological and Biomedical Sciences, 44(3):342–346, September 2013. CODEN ???? ISSN 1369-8486 (print), 1879-2499 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S136984861300068X.

McLoone:2020:POP

[McL20] Brian McLoone. Population and organismal perspectives on trait origins. Studies in History and Philosophy of Biological and Biomedical Sciences, 83(??):Article 101288, October 2020. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848619302031.

McMillan:2004:CMB

[McM04] John McMillan. Is corporate money bad for bioethics? Studies in History and Philosophy of Biological and Biomedical Sciences, 35 (1):167–175, March 2004. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S136984860300116X.

McMillan:2007:RIE

[McM07] John McMillan. The return of the inseminator: Eutelegenesis in past and contemporary reproductive ethics. Studies in History and Philosophy of Biological and Biomedical Sciences, 38(2):393– 410, June 2007. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848607000167.

McMahon:2020:RRG

[McM20] Richard McMahon. Resurecting raciology? Genetic ethnology and pre-1945 anthropological race classification. Studies in History and Philosophy of Biological and Biomedical Sciences, 83(??):Article 101242, October 2020. CO-DEN ???? ISSN 1369-8486 (print), 1879-2499 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S1369848617301231.

McShea:2013:MW

[McS13] Daniel W. McShea. Machine wanting. Studies in History and Philosophy of Biological and Biomedical Sciences, 44(4):679– 687, December 2013. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848613000836.

McShea:2016:FPB

[McS16] Daniel W. McShea. Freedom and purpose in biology. Studies in History and Philosophy of Biological and Biomedical Sciences, 58 (??):64–72, August 2016. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848615001752.

Mebius:2014:WMS

[Meb14] Alexander Mebius. A weakened mechanism is still a mechanism: On the causal role of absences in mechanistic explanation. Studies in History and Philosophy of Biological and Biomedical Sciences, 45(??):43–48, March 2014. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848613001507.

Megill:2016:TPE

[Meg16] Allan Megill. Theological presuppositions of the evolutionary epic: From Robert Chambers to E. O. Wilson. Studies in History and Philosophy of Biological and Biomedical Sciences, 58(??): 24–32, August 2016. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848615001788.

Meloni:2017:DLD

[Mel17] Maurizio Meloni. Disentangling life: Darwin, selectionism, and the postgenomic return of the environment. Studies in History and Philosophy of Biological and Biomedical Sciences, 62(??):10– 19, April 2017. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848617300171.

Menzies:2012:CSM

[Men12] Peter Menzies. The causal structure of mechanisms. Studies in History and Philosophy of Biological and Biomedical Sciences, 43(4):796–805, December 2012. CODEN ???? ISSN 1369-8486 (print), 1879-2499 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S1369848612000647.

Menon:2017:GPR

[Men17] Minakshi Menon. Grains of paradise and reading against the grain: Telling stories about science in the Global South. Studies in History and Philosophy of Biological and Biomedical Sciences, 66(??):83–86, December 2017. CODEN ???? ISSN 1369-8486 (print), 1879-2499 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S1369848617301395.

Mertens:2014:FAP

[Mer14] Rebecca Mertens. A functional analysis in practice? Studies in History and Philosophy of Biological and Biomedical Sciences, 47 (part A)(??):210–212, September 2014. CO-DEN ???? ISSN 1369-8486 (print), 1879-2499 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S1369848614000351.

Methot:2014:EEM

[Mét14] Pierre-Olivier Méthot. Empirical evolutionary medicine. Studies in History and Philosophy of Biological and Biomedical Sciences, 47 (part A)(??):213–217, September 2014. CO-DEN ???? ISSN 1369-8486 (print), 1879-2499 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S136984861400034X.

Methot:2016:WHV

[Mét16] Pierre-Olivier Méthot. Writing the history of virology in the twentieth century: Discovery, disciplines, and conceptual change. Studies in History and Philosophy of Biological and Biomedical Sciences, 59(??):145–153, October 2016. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848616300061.

Meunier:2012:SDM

[Meu12] Robert Meunier. Stages in the development of a model organism as a platform for mechanistic models in developmental biology: Zebrafish, 1970–2000. Studies in History and Philosophy of Biological and Biomedical Sciences, 43(2):522–531, June 2012. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848611001403.

Martin:2014:STD

[MH14] Aryn Martin and Kelly Holloway. ‘something there is that doesn’t love a wall’: Histories of the placental barrier. Studies in History and Philosophy of Biological and Biomedical Sciences, 47 (part B)(??):300–310, September 2014. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848613001702.

Masterton:2010:SMS

[MHH10] Malin Masterton, Mats G. Hansson, and Anna T. Höglund. In search of the missing subject: narrative identity and posthumous wronging. Studies in History and Philosophy of Biological and Biomedical Sciences, 41(4):340–346, December 2010. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848610000725.

Michell:2006:PIM

[Mic06] Joel Michell. Psychophysics, intensive magnitudes, and the psychometricians’ fallacy. Studies in History and Philosophy of Biological and Biomedical Sciences, 37(3):414–432, September 2006. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848606000422.

Michael:2008:VFD

[Mic08] Michael Michael. On the validity of Freud’s dream interpretations. Studies in History and Philosophy of Biological and Biomedical Sciences, 39(1):52–64, March 2008. CODEN ???? ISSN 1369-8486 (print), 1879-2499 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S1369848607000891.

Michelini:2012:HNN

[Mic12] Francesca Michelini. Hegel’s notion of natural purpose. Studies in History and Philosophy of Biological and Biomedical Sciences, 43 (1):133–139, March 2012. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848611000483.

Miller:2006:YTP

[Mil06] Fiona Alice Miller. ‘your true and proper gender’: the Barr body as a good enough science of sex. Studies in History and Philosophy of Biological and Biomedical Sciences, 37(3):459– 483, September 2006. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848606000446.

Millstein:2012:DER

[Mil12] Roberta L. Millstein. Darwin’s explanation of races by means of sexual selection. Studies in History and Philosophy of Biological and Biomedical Sciences, 43(3):627–633, September 2012. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S136984861200057X.

Millstein:2015:TAP

[Mil15]  Roberta L. Millstein.  Thinking about populations and

races in time. Studies in History and Philosophy of Biological and Biomedical Sciences, 52(??):5–11, August 2015. CODEN ???? ISSN 1369-8486 (print), 1879-2499 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S136984861500031X.

Milne:2016:PMB

[Mil16] Anne Milne. The pollen of metaphor: Box, cage, and trap as containment in the eighteenth century. Studies in History and Philosophy of Biological and Biomedical Sciences, 57(??):121–128, June 2016. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848616000376.

Mills:2019:PCC

[Mil19] James H. Mills. Patients, carers and consumers: Agency and the history of pharmaceuticals. Studies in History and Philosophy of Biological and Biomedical Sciences, 76(??):Article 101172, August 2019. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848619300263.

Mills:2020:PBC

[Mil20] James Mills. Pandora’s box closed: the Royal Air Force Institute of Aviation Medicine and Nazi medical experiments on human beings during World War II. Studies in History and Philosophy of Biological and Biomedical Sciences, 79(??):Article 101190, February 2020. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848619300317.

Mitman:2003:NHC

[Mit03] Gregg Mitman. Natural history and the clinic: the regional ecology of allergy in America. Studies in History and Philosophy of Biological and Biomedical Sciences, 34(3):491–510, September 2003. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848603000499.

Mitchell:2014:CWI

[Mit14] Benjamin David Mitchell. Capturing the will: Imposture, delusion, and exposure in Alfred Russel Wallace’s defence of spirit photography. Studies in History and Philosophy of Biological and Biomedical Sciences, 46(??):15–24, June 2014. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848614000119.

Chang:2010:CPE

[mKC10] Ku ming (Kevin) Chang. Collaborative production and experimental labor: two models of dissertation authorship in the eighteenth century. Studies in History and Philosophy of Biological and Biomedical Sciences, 41(4):347–355, December 2010. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848610000683.

Muszynski:2020:BBP

[MM20] Eric Muszynski and Christophe Malaterre. Best behaviour: a proposal for a non-binary conceptualization of behaviour in biology. Studies in History and Philosophy of Biological and Biomedical Sciences, 79(??):Article 101222, February 2020. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848619300767.

Moss:2012:NNN

[MN12] Lenny Moss and Daniel J. Nicholson. On nature and normativity: Normativity, teleology, and mechanism in biological explanation. Studies in History and Philosophy of Biological and Biomedical Sciences, 43(1):88–91, March 2012. CODEN ???? ISSN 1369-8486 (print), 1879-2499 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S1369848611000434.

MacLeod:2013:CSE

[MN13] Miles MacLeod and Nancy J. Nersessian. Coupling simulation and experiment: the bimodal strategy in integrative systems biology. Studies in History and Philosophy of Biological and Biomedical Sciences, 44(4):572–584, December 2013. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848613000964.

MacLeod:2015:MSL

[MN15] Miles MacLeod and Nancy J. Nersessian. Modeling systemslevel dynamics: Understanding without mechanistic explanation in integrative systems biology. Studies in History and Philosophy of Biological and Biomedical Sciences, 49(??):1–11, February 2015. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S136984861400154X.

MacLeod:2019:MMC

[MN19] Miles MacLeod and Nancy J. Nersessian. Mesoscopic modeling as a cognitive strategy for handling complex biological systems. Studies in History and Philosophy of Biological and Biomedical Sciences, 78(??):Article 101201, December 2019. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848618300839.

Montgomery:2016:WDA

[Mon16] Georgina M. Montgomery. Why did attachment stick? Studies in History and Philosophy of Biological and Biomedical Sciences, 60(??):102–104, December 2016. CODEN ???? ISSN 1369-8486 (print), 1879-2499 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S1369848616300425.

Monteiro:2018:SDC

[Mon18] Filipe Pinto Monteiro. The “sick dancers”: the construction of medical knowledge about the “epidemic of dance” in Itapagipe, Salvador, Bahia (1882–1901). Studies in History and Philosophy of Biological and Biomedical Sciences, 71(??):32–40, October 2018. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848617301759.

Moore:2007:RFF

[Moo07] James Moore. R. A. Fisher: a faith fit for eugenics. Studies in History and Philosophy of Biological and Biomedical Sciences, 38 (1):110–135, March 2007. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S136984860600094X.

Morgan:2009:MDB

[Mor09] Gregory J. Morgan. The many dimensions of biodiversity. Studies in History and Philosophy of Biological and Biomedical Sciences, 40(3):235–238, September 2009. CO-DEN ???? ISSN 1369-8486 (print), 1879-2499 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S1369848609000296.

Morange:2011:ROI

[Mor11a] Michel Morange. Recent opportunities for an increasing role for physical explanations in biology. Studies in History and Philosophy of Biological and Biomedical Sciences, 42(2):139–144, June 2011. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848610001135.

Morange:2011:WWR

[Mor11b] Michel Morange. What will result from the interaction between functional and evolutionary biology? Studies in History and Philosophy of Biological and Biomedical Sciences, 42(1):69–74, March 2011. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848610001032.

Morange:2012:WMN

[Mor12] Michel Morange. What might be a new “view of evolution’? Studies in History and Philosophy of Biological and Biomedical Sciences, 43(2):578–581, June 2012. CO-DEN ???? ISSN 1369-8486 (print), 1879-2499 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S1369848612000039.

Morgan:2014:LGS

[Mor14] Gregory J. Morgan. Ludwik Gross, Sarah Stewart, and the 1950s discoveries of Gross murine leukemia virus and polyoma virus. Studies in History and Philosophy of Biological and Biomedical Sciences, 48 (part B)(??):200–209, December 2014. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848614000946.

Morgan:2016:WVS

[Mor16a] Gregory J. Morgan. What is a virus species? Radical pluralism in viral taxonomy. Studies in History and Philosophy of Biological and Biomedical Sciences, 59(??):64–70, October 2016. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848616300048.

Morris:2016:IAA

[Mor16b] Simon Conway Morris. It all adds up … or does it? Numbers, mathematics and purpose. Studies in History and Philosophy of Biological and Biomedical Sciences, 58(??):117–122, August 2016. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848615001843.

Moreira:2019:AMA

[Mor19] Tiago Moreira. Anticipatory measure: Alex Comfort, experimental gerontology and the measurement of senescence. Studies in History and Philosophy of Biological and Biomedical Sciences, 77(??):Article 101179, October 2019. CO-DEN ???? ISSN 1369-8486 (print), 1879-2499 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S1369848618301316.

Moss:2012:PMP

[Mos12] Lenny Moss. Is the philosophy of mechanism philosophy enough? Studies in History and Philosophy of Biological and Biomedical Sciences, 43(1):164–172, March 2012.

CODEN ???? ISSN 1369-8486 (print), 1879-2499 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S1369848611000513.

Mottier:2008:EPS

[Mot08] Véronique Mottier. Eugenics, politics and the state: social democracy and the Swiss ‘gardening state’. Studies in History and Philosophy of Biological and Biomedical Sciences, 39(2):263– 269, June 2008. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848608000241.

Moxham:2018:WSW

[Mox18] Noah Moxham. Where to start and where to end up: Early modern knowledge-making from wish-list to notebook to archive. Studies in History and Philosophy of Biological and Biomedical Sciences, 68–69(??):78–82, April/June 2018. CODEN ???? ISSN 1369-8486 (print), 1879-2499 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S1369848618300566.

Moreira:2011:PLL

[MP11] Tiago Moreira and Paolo Palladino. ‘population laboratories’ or ‘laboratory populations’? Making sense of the Baltimore Longitudinal Study of Aging, 1965–1987. Studies in History and Philosophy of Biological and Biomedical Sciences, 42(3):317– 327, September 2011. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848611000379.

Muller-Strahl:2014:MMM

[MS14] Gerhard Müller-Strahl. Matter, metaphors, and mechanisms: Rethinking cell theories. Studies in History and Philosophy of Biological and Biomedical Sciences, 48 (part B)(??):130–150, December 2014. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848614000764.

Moll:2016:ACC

[MSM16] Remington J. Moll, Daniel Steel, and Robert A. Montgomery. AIC and the challenge of complexity: a case study from ecology. Studies in History and Philosophy of Biological and Biomedical Sciences, 60(??):35–43, December 2016. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848616300711.

Matthews:2017:FAT

[MT17] Lucas J. Matthews and Eric Turkheimer. Flynn, the age-table method, and a metatheory of intelligence. Studies in History and Philosophy of Biological and Biomedical Sciences, 65 (??):35–40, October 2017. CODEN ???? ISSN 1369-8486 (print), 1879-2499 (electronic). URL /science/article/pii/

S136984861730047X.

Mulberger:2012:WCF

[Mül12] Annette Mülberger. Wundt contested: the first crisis declaration in psychology. Studies in History and Philosophy of Biological and Biomedical Sciences, 43(2):434–444, June 2012. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848611001282.

Munns:2015:PPP

[Mun15] David P. D. Munns. The phytotronist and the phenotype: Plant physiology, Big Science, and a Cold War biology of the whole plant. Studies in History and Philosophy of Biological and Biomedical Sciences, 50(??):29–40, April 2015. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848615000047.

Murphy:2001:BRF

[Mur01] Dominic Murphy. Book review: Folk psychology meets the frame problem: W. F. G. Haselager, Cognitive Science and Folk Psychology (London: Sage Publications, 1997), x + 165 pp. ISBN 0-7619-5425-2 Hardback £55.00; ISBN 0-7619-5426-0 Paperback £17.99. Studies in History and Philosophy of Biological and Biomedical Sciences, 32(3):565–573, September 2001. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848601000206. See [Has97].

Muller-Wille:2005:EMS

[MW05] Staffan Müller-Wille. Early Mendelism and the subversion of taxonomy: epistemological obstacles as institutions. Studies in History and Philosophy of Biological and Biomedical Sciences, 36(3):

465–487, September 2005. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848605000464.

Muller-Wille:2007:CCT

[MW07a] Staffan Müller-Wille. Collection and collation: theory and practice of Linnaean botany. Studies in History and Philosophy of Biological and Biomedical Sciences, 38(3):541–562, September 2007. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848607000428.

Muller-Wille:2007:HPC

[MW07b] Staffan Müller-Wille. Hybrids, pure cultures, and pure lines: from nineteenth-century biology to twentieth-century genetics. Studies in History and Philosophy of Biological and Biomedical Sciences, 38(4):796–806, December 2007. CODEN ???? ISSN 1369-8486 (print), 1879-2499 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S1369848607000556.

Muller-Wille:2010:CTS

[MW10] Staffan Müller-Wille. Cell theory, specificity, and reproduction, 1837–1870. Studies in History and Philosophy of Biological and Biomedical Sciences, 41(3):225–231, September 2010. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848610000427.

Muller-Wille:2012:NHI

[MWC12] Staffan Müller-Wille and Isabelle Charmantier. Natural history and information overload: The case of Linnaeus. Studies in History and Philosophy of Biological and Biomedical Sciences, 43 (1):4–15, March 2012. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848611001130.

Muller-Wille:2007:TCL

[MWR07] Staffan Müller-Wille and Karen Reeds. A translation of Carl Linnaeus’s Introduction to Genera plantarum (1737). Studies in History and Philosophy of Biological and Biomedical Sciences, 38(3): 563–572, September 2007. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848607000349. See [vL37].

Myelnikov:2016:MTR

[Mye16] Dmitriy Myelnikov. Metaphors and tracers: Radioactivity in twentieth-century biology. Studies in History and Philosophy of Biological and Biomedical Sciences, 55(??):124–127, February 2016. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848615001405.

Murard:2000:SFH

[MZ00] Lion Murard and Patrick Zylberman. Seeds for French health care: did the Rockefeller Foundation plant the seeds between the two World Wars? Studies in History and Philosophy of Biological and Biomedical Sciences, 31(3):463–475, September 2000. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848600000169.

Nanay:2010:NSL

[Nan10] Bence Nanay. Natural selection and the limitations of environmental resources. Studies in History and Philosophy of Biological and Biomedical Sciences, 41(4):418–419, December 2010. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848610000695. See reply [Ste10a].

Nash:1999:FNI

[Nas99] Jessica Nash. Freaks of nature: images of Barbara Mc-Clintock. Studies in History and Philosophy of Biological and Biomedical Sciences, 30(1):21–43, March 1999. CO-DEN ???? ISSN 1369-8486 (print), 1879-2499 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S1369848698000156.

Nederbragt:2003:SIR

[Ned03] Hubertus Nederbragt. Strategies to improve the reliability of a theory: the experiment of bacterial invasion into cultured epithelial cells. Studies in History and Philosophy of Biological and Biomedical Sciences, 34(4):593–614, December 2003. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848603000724.

Nederbragt:2010:PPP

[Ned10] Hubertus Nederbragt. Protocol, pattern and paper: interactive stabilization of immunohistochemical knowledge. Studies in History and Philosophy of Biological and Biomedical Sciences, 41(4): 386–395, December 2010. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848610000646.

Nelson:2020:MBN

[Nel20] Nicole C. Nelson. Model Behavior by Nicole Nelson: Reply by the author. Studies in History and Philosophy of Biological and Biomedical Sciences, 82(??):Article 101269, August 2020. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848620300340.

Nickelsen:2011:PFC

[NG11] Kärin Nickelsen and Gerd Graßhoff. In pursuit of formaldehyde: Causally explanatory models and falsification. Studies in History and Philosophy of Biological and Biomedical Sciences, 42(3):297– 305, September 2011. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848611000227.

Nickelsen:2006:DBN

[Nic06] Kärin Nickelsen. Draughtsmen, botanists and nature: constructing eighteenth-century botanical illustrations. Studies in History and Philosophy of Biological and Biomedical Sciences, 37 (1):1–25, March 2006. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848605000956.

Nickelsen:2009:CSM

[Nic09] Kärin Nickelsen. The construction of a scientific model: Otto Warburg and the building block strategy. Studies in History and Philosophy of Biological and Biomedical Sciences, 40(2):73–86, June 2009. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848609000168.

Nicholson:2010:BAC

[Nic10]  Daniel J. Nicholson.  Biological atomism and cell the-

ory. Studies in History and Philosophy of Biological and Biomedical Sciences, 41(3):202–211, September 2010. CO-DEN ???? ISSN 1369-8486 (print), 1879-2499 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S1369848610000439.

Nicholson:2012:CMB

[Nic12] Daniel J. Nicholson. The concept of mechanism in biology. Studies in History and Philosophy of Biological and Biomedical Sciences, 43(1):152–163, March 2012. CO-DEN ???? ISSN 1369-8486 (print), 1879-2499 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S1369848611000501.

Nicholson:2013:OM

[Nic13] Daniel J. Nicholson. Organisms / = machines. Studies in History and Philosophy of Biological and Biomedical Sciences, 44(4):669– 678, December 2013. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848613000824.

Nicholson:2014:MCO

[Nic14] Daniel J. Nicholson. The machine conception of the organism in development and evolution: a critical analysis. Studies in History and Philosophy of Biological and Biomedical Sciences, 48 (part B)(??):162–174, December 2014. CODEN ???? ISSN 1369-8486 (print), 1879-2499 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S1369848614001083.

Nicoglou:2015:EPP

[Nic15] Antonine Nicoglou. The evolution of phenotypic plasticity: Genealogy of a debate in genetics. Studies in History and Philosophy of Biological and Biomedical Sciences, 50(??):67–76, April 2015. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848615000059.

Nilsen:2015:BLF

[Nil15] Håvard Friis Nilsen. The biologist of love and fear. Studies in History and Philosophy of Biological and Biomedical Sciences, 54(??):89–93, December 2015. CODEN ???? ISSN 1369-8486 (print), 1879-2499 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S1369848615001363.

Nicoglou:2017:EWB

[NM17] Antonine Nicoglou and Francesca Merlin. Epigenetics: a way to bridge the gap between biological fields. Studies in History and Philosophy of Biological and Biomedical Sciences, 66(??): 73–82, December 2017. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848617300444.

Noakes:2014:HTC

[Noa14] Richard Noakes. Haunted thoughts of the careful experimentalist: Psychical research and the troubles of experimental physics. Studies in History and Philosophy of Biological and Biomedical Sciences, 48 (part A)(??):46–56, December 2014. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848614000831.

Nordlund:2007:HLB

[Nor07] Christer Nordlund. Hormones for life? Behind the rise and fall of a hormone remedy (Gonadex) against sterility in the Swedish welfare state. Studies in History and Philosophy of Biological and Biomedical Sciences, 38(1):191–216, March 2007. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848606000975.

Nachtomy:2007:GEC

[NSY07] Ohad Nachtomy, Ayelet Shavit, and Zohar Yakhini. Gene expression and the concept of the phenotype. Studies in History and Philosophy of Biological and Biomedical Sciences, 38(1):238–254, March 2007. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848606000999.

OMalley:2005:PCE

[OB05] Maureen A. O’Malley and Yan Boucher. Paradigm change in evolutionary microbiology. Studies in History and Philosophy of Biological and Biomedical Sciences, 36(1):183–208, March 2005. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848604000925.

Okasha:2012:ERD

[OB12] Samir Okasha and K. G. Binmore, editors. Evolution and rationality: decisions, co-operation and strategic behaviour. Cambridge University Press, Cambridge, UK, 2012. ISBN 1-107-00499-3 (hardcover), 1-139-51851-8 (ebook), 0-511-79260-3 (e-book), 1-139-51665-5. x + 281 pp.

LCCN HM495 .E96 2012. URL http://assets.cambridge.
org/97811070/04993/cover/9781107004993.jpg;   http://
catdir.loc.gov/catdir/enhancements/fy1212/2012019850-
b.html;   http://catdir.loc.gov/catdir/enhancements/
fy1212/2012019850-d.html;   http://catdir.loc.gov/catdir/
enhancements/fy1212/2012019850-t.html.

ORourke:2016:NCD

[OCG16] Michael O’Rourke, Stephen Crowley, and Chad Gonnerman. On the nature of cross-disciplinary integration: a philosophical framework. Studies in History and Philosophy of Biological and Biomedical Sciences, 56(??):62–70, April 2016. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848615001417.

OMalley:2007:TPM

[OD07] Maureen A. O’Malley and John Dupré. Towards a philosophy of microbiology. Studies in History and Philosophy of Biological and Biomedical Sciences, 38(4):775–779, December 2007. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848607000465.

Odenbaugh:2014:EPE

[Ode14] Jay Odenbaugh. Environmental philosophy 2.0: Ethics and conservation biology for the 21st century. Studies in History and Philosophy of Biological and Biomedical Sciences, 45(??):92–96, March 2014. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848613001519.

OMalley:2010:GGR

[OEB10] Maureen A. O’Malley, Kevin C. Elliott, and Richard M. Burian. From genetic to genomic regulation: iterativity in microRNA research. Studies in History and Philosophy of Biological and Biomedical Sciences, 41(4):407–417, December 2010. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848610000737.

Olszynko-Gryn:2014:DPT

[OG14] Jesse Olszynko-Gryn. The demand for pregnancy testing: the Aschheim–Zondek reaction, diagnostic versatility, and laboratory services in 1930s Britain. Studies in History and Philosophy of Biological and Biomedical Sciences, 47 (part B)(??):233– 247, September 2014. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848613001763.

Okasha:2003:BRC

[Oka03] Samir Okasha. Book review: Could religion be a group-level adaptation of Homo sapiens?: Darwin’s cathedral: evolution, religion and the nature of society, David Sloan Wilson; University of Chicago Press, 2002, pp. v + 268, Price $25 hardback, ISBN 0-226-90134-3. Studies in History and Philosophy of Biological and Biomedical Sciences, 34(4):699–705, December 2003. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848603000785. See [Wil02b].

Olby:2000:WA

[Olb00] Robert Olby. Whetting the appetite. Studies in History and Philosophy of Biological and Biomedical Sciences, 31(4):629– 636, December 2000. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848600000066.

Olby:2015:WAF

[Olb15] Robert Olby. What’s all this fuss about the gene? Studies in History and Philosophy of Biological and Biomedical Sciences, 49(??):66–69, February 2015. CODEN ???? ISSN 1369-8486 (print), 1879-2499 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S1369848614001587.

Olszewski:2011:DAB

[Ols11] Margaret Maria Olszewski. Dr. Auzoux’s botanical teaching models and medical education at the universities of Glasgow and Aberdeen. Studies in History and Philosophy of Biological and Biomedical Sciences, 42(3):285–296, September 2011. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848611000161.

OMalley:2008:EEE

[O’M08] Maureen A. O’Malley. ‘everything is everywhere: but the environment selects’: ubiquitous distribution and ecological determinism in microbial biogeography. Studies in History and Philosophy of Biological and Biomedical Sciences, 39(3):314– 325, September 2008. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S136984860800040X.

OMalley:2010:FEC

[O’M10] Maureen A. O’Malley. The first eukaryote cell: an unfinished history of contestation. Studies in History and Philosophy of Biological and Biomedical Sciences, 41(3):212–224, September 2010. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848610000440.

OMalley:2013:WIF

[O’M13] Maureen A. O’Malley. When integration fails: Prokaryote phylogeny and the tree of life. Studies in History and Philosophy of Biological and Biomedical Sciences, 44(4):551–562, December 2013. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848612001100.

OMalley:2014:EPS

[O’M14] Maureen A. O’Malley. Exemplary philosophy of science: How to do it. Studies in History and Philosophy of Biological and Biomedical Sciences, 45(??):149–152, March 2014. CODEN ???? ISSN 1369-8486 (print), 1879-2499 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S1369848613001295.

OMalley:2016:EV

[O’M16] Maureen A. O’Malley. The ecological virus. Studies in History and Philosophy of Biological and Biomedical Sciences, 59(??): 71–79, October 2016. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S136984861630005X.

OMalley:2010:CNC

[OMW10] Maureen A. O’Malley and Staffan Müller-Wille. The cell as nexus: connections between the history, philosophy and science of cell biology. Studies in History and Philosophy of Biological and Biomedical Sciences, 41(3):169–171, September 2010. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848610000397.

ONeill:2001:ICG

[O’N01]  Onora O’Neill.  Informed consent and genetic informa-

tion. Studies in History and Philosophy of Biological and Biomedical Sciences, 32(4):689–704, December 2001. CO-DEN ???? ISSN 1369-8486 (print), 1879-2499 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S1369848601000267.

Onaga:2014:RWF

[Ona14] Lisa A. Onaga. Ray Wu as fifth business: Deconstructing collective memory in the history of DNA sequencing. Studies in History and Philosophy of Biological and Biomedical Sciences, 46(??):1–14, June 2014. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S136984861400003X.

Oftedal:2013:SBG

[OP13] Gry Oftedal and Veli-Pekka Parkkinen. Synthetic biology and genetic causation. Studies in History and Philosophy of Biological and Biomedical Sciences, 44(2):208–216, June 2013. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848613000307.

Orland:2012:FMN

[Orl12] Barbara Orland. The fluid mechanics of nutrition: Herman Boerhaave’s synthesis of seventeenth-century circulation physiology. Studies in History and Philosophy of Biological and Biomedical Sciences, 43(2):357–369, June 2012. CODEN ???? ISSN 1369-8486 (print), 1879-2499 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S1369848611001208.

OMalley:2012:RIM

[OS12a] Maureen A. O’Malley and Orkun S. Soyer. The roles of integration in molecular systems biology. Studies in History and Philosophy of Biological and Biomedical Sciences, 43(1):58–68, March 2012. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848611000823.

Orland:2012:I

[OS12b] Barbara Orland and E. C. Spary. Introduction. Studies in History and Philosophy of Biological and Biomedical Sciences, 43(2):317– 322, June 2012. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848611001154.

Overy:2008:ESS

[Ove08] Richard Overy. Eugenics, sex and the state: an afterword. Studies in History and Philosophy of Biological and Biomedical Sciences, 39(2):270–272, June 2008. CO-DEN ???? ISSN 1369-8486 (print), 1879-2499 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S1369848608000253.

Pallo:2000:RCS

[Pal00] Gábor Palló. Rescue and cordon sanitaire: The Rockefeller Foundation in Hungarian public health. Studies in History and Philosophy of Biological and Biomedical Sciences, 31(3):433– 445, September 2000. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848600000133.

Pannese:2012:GMT

[Pan12] Alessia Pannese. A gray matter of taste: Sound perception, music cognition, and Baumgarten’s aesthetics. Studies in History and Philosophy of Biological and Biomedical Sciences, 43(3):594– 601, September 2012. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848612000453.

Pantazakos:2019:TWT

[Pan19] Themistoklis Pantazakos. Treatment for whom? Towards a phenomenological resolution of controversy within autism treatment. Studies in History and Philosophy of Biological and Biomedical Sciences, 77(??):Article 101176, October 2019. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848618301018.

Parusnikova:2002:IMP

[Par02] Zuzana Parusnikova. Integrative medicine: partnership or control? Studies in History and Philosophy of Biological and Biomedical Sciences, 33(1):169–186, March 2002. CODEN ???? ISSN 1369-8486 (print), 1879-2499 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S1369848601000358.

Parnes:2003:TWA

[Par03] Ohad Parnes. ‘trouble from within’: allergy, autoimmunity, and pathology in the first half of the twentieth century. Studies in History and Philosophy of Biological and Biomedical Sciences, 34(3): 425–454, September 2003. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848603000529.

Parry:2004:TIB

[Par04] Bronwyn Parry. Technologies of immortality: the brain on ice. Studies in History and Philosophy of Biological and Biomedical Sciences, 35(2):391–413, June 2004. CO-DEN ???? ISSN 1369-8486 (print), 1879-2499 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S1369848604000275.

Parke:2014:FMW

[Par14] Emily C. Parke. Flies from meat and wasps from trees: Reevaluating Francesco Redi’s spontaneous generation experiments. Studies in History and Philosophy of Biological and Biomedical Sciences, 45(??):34–42, March 2014. CO-DEN ???? ISSN 1369-8486 (print), 1879-2499 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S1369848614000028.

Parkkinen:2016:REM

[Par16] Veli-Pekka Parkkinen. Robustness and evidence of mechanisms in early experimental atherosclerosis research. Studies in History and Philosophy of Biological and Biomedical Sciences, 60(??): 44–55, December 2016. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S136984861630070X.

Paslaru:2014:MAT

[Pâs14] Viorel Pâslaru. The mechanistic approach of The Theory of Island Biogeography and its current relevance. Studies in History and Philosophy of Biological and Biomedical Sciences, 45(??): 22–33, March 2014. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848613001738.

Paternotte:2020:SEI

[Pat20] Cédric Paternotte. Social evolution and the individual-as-maximising-agent analogy. Studies in History and Philosophy of Biological and Biomedical Sciences, 79(??):Article 101225, February 2020. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S136984861930113X.

Paul:2008:JSM

[PD08] Diane B. Paul and Benjamin Day. John Stuart Mill, innate differences, and the regulation of reproduction. Studies in History and Philosophy of Biological and Biomedical Sciences, 39(2):222–231, June 2008. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848608000198.

Powell:2009:MSI

[PD09] Alexander Powell and John Dupré. From molecules to systems: the importance of looking both ways. Studies in History and Philosophy of Biological and Biomedical Sciences, 40(1):54–64, March 2009. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848608000952.

Pearson:2007:HEU

[Pea07] Christopher H. Pearson. Is heritability explanatorily useful? Studies in History and Philosophy of Biological and Biomedical Sciences, 38(1):270–288, March 2007. CO-DEN ???? ISSN 1369-8486 (print), 1879-2499 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S1369848606001026.

Pearce:2010:CEH

[Pea10] Trevor Pearce. From ‘circumstances’ to ‘environment’: Herbert Spencer and the origins of the idea of organism-environment interaction. Studies in History and Philosophy of Biological and Biomedical Sciences, 41(3):241–252, September 2010. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848610000373.

Peacock:2011:TFE

[Pea11] Kent A. Peacock. The three faces of ecological fitness. Studies in History and Philosophy of Biological and Biomedical Sciences, 42 (1):99–105, March 2011. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848610001044.

Pearce:2012:PBT

[Pea12] Trevor Pearce. Philosophy of biology in the twenty-first century. Studies in History and Philosophy of Biological and Biomedical Sciences, 43(1):312–315, March 2012. CO-DEN ???? ISSN 1369-8486 (print), 1879-2499 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S1369848611000926.

Pedroso:2012:EHB

[Ped12] Makmiller Pedroso. Essentialism, history, and biological taxa. Studies in History and Philosophy of Biological and Biomedical Sciences, 43(1):182–190, March 2012. CODEN ???? ISSN 1369-8486 (print), 1879-2499 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S1369848611000951.

Pedroso:2015:SSU

[Ped15] Makmiller Pedroso. Starting small: Using little microbes to tackle big philosophical problems. Studies in History and Philosophy of Biological and Biomedical Sciences, 53(??):126–128, ???? 2015. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848615000679.

Peirson:2015:PSY

[Pei15] B. R. Erick Peirson. Plasticity, stability, and yield: the origins of Anthony David Bradshaw’s model of adaptive phenotypic plasticity. Studies in History and Philosophy of Biological and Biomedical Sciences, 50(??):51–66, April 2015. CODEN ???? ISSN 1369-8486 (print), 1879-2499 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S1369848615000072.

Pence:2011:DOW

[Pen11] Charles H. Pence. “describing our whole experience”: The statistical philosophies of W. F. R. Weldon and Karl Pearson. Studies in History and Philosophy of Biological and Biomedical Sciences, 42(4):475–485, December 2011. CO-DEN ???? ISSN 1369-8486 (print), 1879-2499 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S1369848611000744.

Pennell:2012:MGI

[Pen12] Sara Pennell. ‘A matter of so great importance to my health’: Alimentary knowledge in practice. Studies in History and Philosophy of Biological and Biomedical Sciences, 43(2):418–424, June 2012. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848611001257.

Pence:2015:MCC

[Pen15] Charles H. Pence. The many chances of Charles Darwin. Studies in History and Philosophy of Biological and Biomedical Sciences, 53(??):107–110, ???? 2015. CO-DEN ???? ISSN 1369-8486 (print), 1879-2499 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S1369848615001065.

Pence:2020:HCP

[Pen20] David Evan Pence. How comparative psychology lost its soul: Psychical research and the new science of animal behavior. Studies in History and Philosophy of Biological and Biomedical Sciences, 82(??):Article 101275, August 2020. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848619300305.

Persson:2009:SMM

[Per09]  Johannes Persson.  Semmelweis’s methodology from the

modern stand-point: intervention studies and causal ontology. Studies in History and Philosophy of Biological and Biomedical Sciences, 40(3):204–209, September 2009. CO-DEN ???? ISSN 1369-8486 (print), 1879-2499 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S1369848609000314.

Peres:2016:SGP

[Per16] Sara Peres. Saving the gene pool for the future: Seed banks as archives. Studies in History and Philosophy of Biological and Biomedical Sciences, 55(??):96–104, February 2016. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848615001284.

Pettitt:2002:KNV

[Pet02] Clare Pettitt. Kindred nature: Victorian and Edwardian women embrace the living world. Studies in History and Philosophy of Biological and Biomedical Sciences, 33(2):325–335, July 2002. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368102000158.

Pieters:1998:MDB

[Pie98] Toine Pieters. Managing differences in biomedical research: The case of standardizing interferons. Studies in History and Philosophy of Biological and Biomedical Sciences, 29(1):31–79, March 1998. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S136984869800003X.

Pietsch:2012:TLV

[Pie12] Theodore W. Pietsch. Trees of life: a visual history of evolution. The Johns Hopkins University Press, Baltimore, MD, USA, 2012. ISBN 1-4214-0479-6 (hardcover). xi + 358 pp. LCCN QH361 .P53 2012.

Pierotti:2020:HLB

[Pie20] Raymond Pierotti. Historical links between ethnobiology and evolution: Conflicts and possible resolutions. Studies in History and Philosophy of Biological and Biomedical Sciences, 81(??): Article 101277, June 2020. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848619301955.

Pigliucci:2013:WWM

[Pig13] Massimo Pigliucci. What are we to make of the concept of race?: Thoughts of a philosopher-scientist. Studies in History and Philosophy of Biological and Biomedical Sciences, 44(3):272– 277, September 2013. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848613000447.

Pilpel:2007:SER

[Pil07] Avital Pilpel. Statistics is not enough: revisiting Ronald A. Fisher’s critique (1936) of Mendel’s experimental results (1866). Studies in History and Philosophy of Biological and Biomedical Sciences, 38(3):618–626, September 2007. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848607000416.

Pradeu:2016:UVP

[PKD16] Thomas Pradeu, Gladys Kostyrka, and John Dupré. Understanding viruses: Philosophical investigations. Studies in History and Philosophy of Biological and Biomedical Sciences, 59(??):57–63, October 2016. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848616300024.

Planer:2015:GCP

[Pla15] Ronald J. Planer. Gene-concept pluralism, causal specificity, and information. Studies in History and Philosophy of Biological and Biomedical Sciences, 53(??):129–133, ???? 2015. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848615000801.

Plutynski:2006:WWF

[Plu06] Anya Plutynski. What was Fisher’s fundamental theorem of natural selection and what was it for? Studies in History and Philosophy of Biological and Biomedical Sciences, 37(1):59–82, March 2006. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848605000981.

Plutynski:2007:PGW

[Plu07] Anya Plutynski. A philosopher goes wild. Studies in History and Philosophy of Biological and Biomedical Sciences, 38(1):289–296, March 2007. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848606001014.

Plutynski:2013:CGI

[Plu13] Anya Plutynski. Cancer and the goals of integration. Studies in History and Philosophy of Biological and Biomedical Sciences, 44(4):466–476, December 2013. CODEN ???? ISSN 1369-8486 (print), 1879-2499 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S1369848613000332.

Plutynski:2014:PE

[Plu14] Anya Plutynski. Philosophy of epidemiology. Studies in History and Philosophy of Biological and Biomedical Sciences, 46(??): 107–111, June 2014. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848614000053.

Plutynski:2020:ECA

[Plu20] Anya Plutynski. Explaining Cancer by Anya Plutynski: Reply by the author. Studies in History and Philosophy of Biological and Biomedical Sciences, 84(??):Article 101326, December 2020. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848620301370. See [Gre20a].

Peyrieras:2002:SLP

[PM02] Nadine Peyrieras and Michel Morange. The study of lysogeny at the Pasteur Institute (1950–1960): an epistemologically open system. Studies in History and Philosophy of Biological and Biomedical Sciences, 33(3):419–430, September 2002. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848602000146.

Pieters:2011:ICB

[PM11] Toine Pieters and Benoît Majerus. The introduction of chlorpromazine in Belgium and the Netherlands (1951–1968); tango between old and new treatment features. Studies in History and Philosophy of Biological and Biomedical Sciences, 42(4):443– 452, December 2011. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848611000392.

Podolsky:2015:GYG

[Pod15] Scott H. Podolsky. Generic, yet not generic. Studies in History and Philosophy of Biological and Biomedical Sciences, 50(??):90– 93, April 2015. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848615000369.

Podgorny:2017:MAI

[Pod17] Irina Podgorny. Manifest ambiguity: Intermediate forms, variation, and mammal paleontology in Argentina, 1830– 1880. Studies in History and Philosophy of Biological and Biomedical Sciences, 66(??):27–36, December 2017. CO-DEN ???? ISSN 1369-8486 (print), 1879-2499 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S1369848617300742.

Polge:2007:WAR

[Pol07] Chris Polge. The work of the Animal Research Station, Cambridge. Studies in History and Philosophy of Biological and Biomedical Sciences, 38(2):511–520, June 2007. CODEN ???? ISSN 1369-8486 (print), 1879-2499 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S1369848607000210.

Polonioli:2015:SAA

[Pol15] Andrea Polonioli. Stanovich’s arguments against the “adaptive rationality” project: an assessment. Studies in History and Philosophy of Biological and Biomedical Sciences, 49(??):55–62, February 2015. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848614001782.

Pols:2020:BRI

[Pol20] Hans Pols. Book review: The Invention of Madness by Emily Baum: What can historians of psychiatry learn from China? Studies in History and Philosophy of Biological and Biomedical Sciences, 79(??):Article 101203, February 2020. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848619301207. See reply [Bau20].

Poon:2011:IRC

[Poo11] Wilson C. K. Poon. Interdisciplinary reflections: The case of physics and biology. Studies in History and Philosophy of Biological and Biomedical Sciences, 42(2):115–118, June 2011. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S136984861000110X.

Popa:2020:MHN

[Pop20] Elena Popa. Mental health, normativity, and local knowledge in global perspective. Studies in History and Philosophy of Biological and Biomedical Sciences, 84(??):Article 101334, December 2020. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S136984862030145X.

Porter:2014:CCB

[Por14] Theodore M. Porter. The curious case of blending inheritance. Studies in History and Philosophy of Biological and Biomedical Sciences, 46(??):125–132, June 2014. CODEN ???? ISSN 1369-8486 (print), 1879-2499 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S1369848614000260.

Potochnik:2012:MSE

[Pot12] Angela Potochnik. Modeling social and evolutionary games. Studies in History and Philosophy of Biological and Biomedical Sciences, 43(1):202–208, March 2012. CODEN ???? ISSN 1369-8486 (print), 1879-2499 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S1369848611001385.

Powers:2013:FEM

[Pow13a] Jack Powers. Finding Ernst Mayr’s Plato. Studies in History and Philosophy of Biological and Biomedical Sciences, 44(4):714– 723, December 2013. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848613001222.

Powers:2013:CLR

[Pow13b] Simon T. Powers. The circle of life (reviewing E. Coen, Cells to Civilizations: the principles of change that shape life). Studies in History and Philosophy of Biological and Biomedical Sciences, 44(3):447–450, September 2013. CO-DEN ???? ISSN 1369-8486 (print), 1879-2499 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S1369848613000721.

Pradeu:2016:MVH

[Pra16] Thomas Pradeu. Mutualistic viruses and the heteronomy of life. Studies in History and Philosophy of Biological and Biomedical Sciences, 59(??):80–88, October 2016. CO-DEN ???? ISSN 1369-8486 (print), 1879-2499 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S1369848616300012.

Prebble:2010:DOP

[Pre10] John N. Prebble. The discovery of oxidative phosphorylation: a conceptual off-shoot from the study of glycolysis. Studies in History and Philosophy of Biological and Biomedical Sciences, 41(3): 253–262, September 2010. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848610000488.

Preston:2013:SBR

[Pre13] Beth Preston. Synthetic biology as red herring. Studies in History and Philosophy of Biological and Biomedical Sciences, 44(4):649– 659, December 2013. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848613000800.

Prince:2012:SKC

[Pri12] Ruth J. Prince. Science, knowledge and colonial rule in Africa. Studies in History and Philosophy of Biological and Biomedical Sciences, 43(4):821–824, December 2012. CO-DEN ???? ISSN 1369-8486 (print), 1879-2499 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S1369848612000702.

Principe:2013:SA

[Pri13] Lawrence Principe. The secrets of alchemy. Synthesis. University of Chicago Press, Chicago, IL, USA and London, UK, 2013. ISBN 0-226-68295-1 (hardcover), 0-226-92378-9 (e-book), 0-226-10379-X. v + 281 + 8 pp. LCCN QD13 .P75 2013.

Price:2014:TPE

[Pri14] Katy Price. Testimonies of precognition and encounters with psychiatry in letters to J. B. Priestley. Studies in History and Philosophy of Biological and Biomedical Sciences, 48 (part A)(??): 103–111, December 2014. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848614000867.

Priest:2017:FCQ

[Pri17] Greg Priest. Framing causal questions about the past: the Cambrian explosion as case study. Studies in History and Philosophy of Biological and Biomedical Sciences, 63(??):55–63, June 2017. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848617300626.

Pieters:2011:SPD

[PS11] Toine Pieters and Stephen Snelders. Standardizing psychotropic drugs and drug practices in the twentieth century: paradox of order and disorder. Studies in History and Philosophy of Biological and Biomedical Sciences, 42(4):412–414, December 2011. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848611000422.

Prainsack:2013:PUP

[PT13] Barbara Prainsack and Victor Toom. Performing the Union: the Prüm Decision and the European dream. Studies in History and Philosophy of Biological and Biomedical Sciences, 44(1):71–79, March 2013. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848612001033.

Pugliano:2009:NCB

[Pug09] Valentina Pugliano. Non-colonial botany or, the late rise of local knowledge? Studies in History and Philosophy of Biological and Biomedical Sciences, 40(4):321–328, December 2009. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848609000582.

Quarfood:2006:KBT

[Qua06] Marcel Quarfood. Kant on biological teleology: Towards a two-level interpretation. Studies in History and Philosophy of Biological and Biomedical Sciences, 37(4):735–747, December 2006. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848606000744.

Quirke:2004:FBM

[Qui04] Viviane Quirke. French biomedicine in the mirror of America: Inventer la biomédecine: la France, l’Amérique et la production des savoirs du vivant (1945–1965), Jean-Paul Gaudillière, La Découverte & Syros, Paris, 2002, pp. 392, Price F 219,75 paperback, ISBN 2-7071-3607-7. Studies in History and Philosophy of Biological and Biomedical Sciences, 35(4):765–776, December 2004. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848603001237. See [Gau02a].

Quirke:2005:MBC

[Qui05] Viviane Quirke. Making British cortisone: Glaxo and the development of corticosteroids in Britain in the 1950s– 1960s. Studies in History and Philosophy of Biological and Biomedical Sciences, 36(4):645–674, December 2005. CO-DEN ???? ISSN 1369-8486 (print), 1879-2499 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S1369848605000737.

Quinn:2016:WWP

[Qui16] Aleta Quinn. William Whewell’s philosophy of architecture and the historicization of biology. Studies in History and Philosophy of Biological and Biomedical Sciences, 59(??):11–19, October 2016. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848616300413.

Quinn:2017:WCC

[Qui17] Aleta Quinn. Whewell on classification and consilience. Studies in History and Philosophy of Biological and Biomedical Sciences, 64(??):65–74, August 2017. CODEN ???? ISSN 1369-8486 (print), 1879-2499 (electronic). URL /science/article/pii/

S1369848617300663.

RetifdeLaBretonne:1781:DAP

[R81] Nicolas-Edme Rétif de La Bretonne. La découverte australe par un homme volant, ou Le dédale français nouvelle très philosophique, suivie de la Lettre d’un signé. … ????, Leipzig, Germany, 1781. ???? pp. Four volumes.

RetifdeLaBretonne:1977:DAP

[R77] Nicolas-Edme Rétif de La Bretonne. La découverte australe par un homme-volant; ou, Le dédale français: nouvelle philosophique. Bibliothèque des utopies. France Adel, Paris, 1977. ISBN 2-85965-001-6. 260 pp.

Racine:2016:MHD

[Rac16] Valérie Racine. The mechanistic-holistic divide revisited: the case of the lac operon. Studies in History and Philosophy of Biological and Biomedical Sciences, 59(??):1–10, October 2016. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848616300243.

Rader:1999:MMG

[Rad99] Karen A. Rader. Of mice, medicine, and genetics: C. C. Little’s creation of the inbred laboratory mouse, 1909– 1918. Studies in History and Philosophy of Biological and Biomedical Sciences, 30(3):319–343, September 1999. CO-DEN ???? ISSN 1369-8486 (print), 1879-2499 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S1369848699000151.

Radick:2000:LBF

[Rad00] Gregory Radick. Language, brain function, and human origins in the Victorian debates on evolution. Studies in History and Philosophy of Biological and Biomedical Sciences, 31(1):55–75, March 2000. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S136984869900045X.

Radick:2001:CKE

[Rad01] Gregory Radick. A critique of Kitcher on eugenic reasoning. Studies in History and Philosophy of Biological and Biomedical Sciences, 32(4):741–751, December 2001. CO-DEN ???? ISSN 1369-8486 (print), 1879-2499 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S1369848601000310.

Radick:2003:MME

[Rad03] Gregory Radick. Mystery of mysteries: is evolution a social construction? Studies in History and Philosophy of Biological and Biomedical Sciences, 34(1):187–200, March 2003. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848602000754.

Radick:2006:WNV

[Rad06] Gregory Radick. What’s in a name? The vervet predator calls and the limits of the Washburnian synthesis. Studies in History and Philosophy of Biological and Biomedical Sciences, 37(2):334– 362, June 2006. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848606000227.

Rader:2008:WHG

[Rad08a] Karen A. Rader. Whose history is a guinea pig’s history? Studies in History and Philosophy of Biological and Biomedical Sciences, 39(3):371–373, September 2008. CODEN ???? ISSN 1369-8486 (print), 1879-2499 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S1369848608000460.

Radick:2008:RLD

[Rad08b] Gregory Radick. Race and language in the Darwinian tradition (and what Darwin’s language–species parallels have to do with it). Studies in History and Philosophy of Biological and Biomedical Sciences, 39(3):359–370, September 2008. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848608000459. See reply [?].

Radick:2011:PGS

[Rad11]  Gregory Radick.  Physics in the Galtonian sciences of

heredity. Studies in History and Philosophy of Biological and Biomedical Sciences, 42(2):129–138, June 2011. CO-DEN ???? ISSN 1369-8486 (print), 1879-2499 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S1369848610001123.

Radick:2013:BD

[Rad13] Gregory Radick. Biomachine dreams. Studies in History and Philosophy of Biological and Biomedical Sciences, 44(4):790– 792, December 2013. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848613001532.

Radin:2014:UES

[Rad14] Joanna Radin. Unfolding epidemiological stories: How the WHO made frozen blood into a flexible resource for the future. Studies in History and Philosophy of Biological and Biomedical Sciences, 47 (part A)(??):62–73, September 2014. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848614000478.

Ralley:2010:SDB

[Ral10] Robert Ralley. Stars, demons and the body in fifteenthcentury England. Studies in History and Philosophy of Biological and Biomedical Sciences, 41(2):109–116, June 2010. CODEN ???? ISSN 1369-8486 (print), 1879-2499 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S1369848610000154.

Ramsey:2006:BF

[Ram06] Grant Ramsey. Block fitness. Studies in History and Philosophy of Biological and Biomedical Sciences, 37(3):484–498, September 2006. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848606000458.

Ramsden:2008:END

[Ram08] Edmund Ramsden. Eugenics from the New Deal to the Great Society: genetics, demography and population quality. Studies in History and Philosophy of Biological and Biomedical Sciences, 39(4):391–406, December 2008. CO-DEN ???? ISSN 1369-8486 (print), 1879-2499 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S1369848608000782.

Ramsden:2014:SMP

[Ram14] Edmund Ramsden. Surveying the meritocracy: the problems of intelligence and mobility in the studies of the Population Investigation Committee. Studies in History and Philosophy of Biological and Biomedical Sciences, 47 (part A)(??):130–141, September 2014. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848614000533.

Ramesh:2016:SCI

[Ram16a] Aditya Ramesh. Scientific commodities, imperial dreams. Studies in History and Philosophy of Biological and Biomedical Sciences, 60(??):88–91, December 2016. CODEN ???? ISSN 1369-8486 (print), 1879-2499 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S136984861630022X.

Ramsey:2016:CAU

[Ram16b] Grant Ramsey. Can altruism be unified? Studies in History and Philosophy of Biological and Biomedical Sciences, 56(??):32–38, April 2016. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848615001570.

Rapini:2004:CIP

[Rap04] Alessandro Rapini. Classes or individuals? The paradox of systematics revisited. Studies in History and Philosophy of Biological and Biomedical Sciences, 35(4):675–695, December 2004. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848604000706.

Ratcliffe:2000:FF

[Rat00] Matthew Ratcliffe. The function of function. Studies in History and Philosophy of Biological and Biomedical Sciences, 31 (1):113–133, March 2000. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848699000394.

Ratcliffe:2002:EBM

[Rat02]  Matthew Ratcliffe.  Evolution and belief: the missing

question. Studies in History and Philosophy of Biological and Biomedical Sciences, 33(1):133–150, March 2002. CODEN ???? ISSN 1369-8486 (print), 1879-2499 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S1369848601000383.

Rausing:2004:ND

[Rau04] Lisbet Rausing. Nemesis divina. Studies in History and Philosophy of Biological and Biomedical Sciences, 35(1):185–190, March 2004. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848603001183.

Razeto-Barry:2013:CAC

[RB13] Pablo Razeto-Barry. Complexity, adaptive complexity and the Creative View of natural selection. Studies in History and Philosophy of Biological and Biomedical Sciences, 44(3):312– 315, September 2013. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S136984861300037X.

Razeto-Barry:2011:PCE

[RBF11] Pablo Razeto-Barry and Ramiro Frick. Probabilistic causation and the explanatory role of natural selection. Studies in History and Philosophy of Biological and Biomedical Sciences, 42(3):344– 355, September 2011. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848611000203.

Rubin:2020:CCI

[RBOH20] Hannah Rubin, Justin P. Bruner, Cailin O’Connor, and Simon Huttegger. Communication without common interest: a signaling experiment. Studies in History and Philosophy of Biological and Biomedical Sciences, 83(??):Article 101295, October 2020. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848619301438.

Rees:2006:PAQ

[Ree06] Amanda Rees. A place that answers questions: primatological field sites and the making of authentic observations. Studies in History and Philosophy of Biological and Biomedical Sciences, 37 (2):311–333, June 2006. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848606000215.

Reisz:2004:PTN

[Rei04] E. G. Reisz. Picturing tropical nature. Studies in History and Philosophy of Biological and Biomedical Sciences, 35(4):777– 792, December 2004. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848603001225.

Reiss:2012:CSI

[Rei12] Julian Reiss. Causation in the sciences: an inferentialist account. Studies in History and Philosophy of Biological and Biomedical Sciences, 43(4):769–777, December 2012. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848612000611.

Renwick:2015:ERD

[Ren15] Chris Renwick. Essay review: For Darwin read Malthus. Political Descent: Malthus, Mutualism, and the Politics of Evolution in Victorian England, Piers Hale. University of Chicago Press, Chicago (2014). 464 pp. Price $45.00, cloth, ISBN 978-0-226-10849-0. Studies in History and Philosophy of Biological and Biomedical Sciences, 51(??):64–66, June 2015. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848615000588.

Resnik:2003:PPU

[Res03] David B. Resnik. Is the precautionary principle unscientific? Studies in History and Philosophy of Biological and Biomedical Sciences, 34(2):329–344, June 2003. CO-DEN ???? ISSN 1369-8486 (print), 1879-2499 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S1369848602000742.

Reutlinger:2012:GRI

[Reu12] Alexander Reutlinger. Getting rid of interventions. Studies in History and Philosophy of Biological and Biomedical Sciences, 43(4):787–795, December 2012. CODEN ???? ISSN 1369-8486 (print), 1879-2499 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S1369848612000623.

Reydon:2005:NSP

[Rey05] Thomas A. C. Reydon. On the nature of the species problem and the four meanings of ‘species’. Studies in History and Philosophy of Biological and Biomedical Sciences, 36(1):135–158, March 2005. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848604000949.

Reydon:2006:GKN

[Rey06] Thomas A. C. Reydon. Generalizations and kinds in natural science: the case of species. Studies in History and Philosophy of Biological and Biomedical Sciences, 37(2):230–255, June 2006. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S136984860600015X.

Reynolds:2010:RC

[Rey10] Andrew Reynolds. The redoubtable cell. Studies in History and Philosophy of Biological and Biomedical Sciences, 41(3):194– 201, September 2010. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848610000452.

Reydon:2012:HPE

[Rey12] Thomas A. C. Reydon. How-possibly explanations as genuine explanations and helpful heuristics: a comment on Forber. Studies in History and Philosophy of Biological and Biomedical Sciences, 43(1):302–310, March 2012. CO-DEN ???? ISSN 1369-8486 (print), 1879-2499 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S1369848611000914.

Reynolds:2014:DCH

[Rey14] Andrew S. Reynolds. The deaths of a cell: How language and metaphor influence the science of cell death. Studies in History and Philosophy of Biological and Biomedical Sciences, 48 (part B)(??):175–184, December 2014. CO-DEN ???? ISSN 1369-8486 (print), 1879-2499 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S1369848614000776.

Roll-Hansen:1998:SNS

[RH98] Nils Roll-Hansen. Studying natural science without nature? Reflections on the realism of so-called laboratory studies. Studies in History and Philosophy of Biological and Biomedical Sciences, 29 (1):165–187, March 1998. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848698000053.

Roll-Hansen:2017:LOT

[RH17] Nils Roll-Hansen. The life organic: the Theoretical Biology Club and the roots of epigenetics. Studies in History and Philosophy of Biological and Biomedical Sciences, 65(??):30–34, October 2017. CODEN ???? ISSN 1369-8486 (print), 1879-2499 (electronic).

URL /science/article/pii/S1369848617301085.

Rheinberger:2009:RSE

[Rhe09] Hans-Jörg Rheinberger. Recent science and its exploration: the case of molecular biology. Studies in History and Philosophy of Biological and Biomedical Sciences, 40(1):6–12, March 2009. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848608000903.

Rheinberger:2016:AIM

[Rhe16] Hans-Jörg Rheinberger. Afterword: Instruments as media, media as instruments. Studies in History and Philosophy of Biological and Biomedical Sciences, 57(??):161–162, June 2016. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848616300139.

Richards:2000:KBB

[Ric00] Robert J. Richards. Kant and Blumenbach on the Bildungstrieb: a historical misunderstanding. Studies in History and Philosophy of Biological and Biomedical Sciences, 31(1):11–32, March 2000. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848699000424.

Richards:2001:HDG

[Ric01] Martin Richards. How distinctive is genetic information? Studies in History and Philosophy of Biological and Biomedical Sciences, 32(4):663–687, December 2001. CODEN ???? ISSN 1369-8486 (print), 1879-2499 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S1369848601000279.

Richards:2008:AIE

[Ric08] Martin Richards. Artificial insemination and eugenics: celibate motherhood, eutelegenesis and germinal choice. Studies in History and Philosophy of Biological and Biomedical Sciences, 39 (2):211–221, June 2008. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848608000186.

Richards:2012:DPD

[Ric12] Robert J. Richards. Darwin’s principles of divergence and natural selection: Why Fodor was almost right. Studies in History and Philosophy of Biological and Biomedical Sciences, 43(1):256–268, March 2012. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848611000902.

Riegner:2013:ANA

[Rie13] Mark F. Riegner. Ancestor of the new archetypal biology: Goethe’s dynamic typology as a model for contemporary evolutionary developmental biology. Studies in History and Philosophy of Biological and Biomedical Sciences, 44(4):735–744, December 2013. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848613000873.

Rodriguez-Ocana:2000:FEP

[RO00] Esteban Rodríguez-Ocaña. Foreign expertise, political pragmatism and professional elite: The Rockefeller Foundation in Spain, 1919–39. Studies in History and Philosophy of Biological and Biomedical Sciences, 31(3):447–461, September 2000. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848600000145.

Roberts:2013:RFC

[Rob13] Paul Roberts. Renegotiating forensic cultures: Between law, science and criminal justice. Studies in History and Philosophy of Biological and Biomedical Sciences, 44(1):47–59, March 2013. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848612001045.

Rocha:2011:SSV

[Roc11] Leon Antonio Rocha. Scientia sexualis versus ars erotica: Foucault, van Gulik, Needham. Studies in History and Philosophy of Biological and Biomedical Sciences, 42(3):328–343, September 2011. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848611000185.

Rocha:2012:WSJ

[Roc12] Leon Antonio Rocha. The way of sex: Joseph Needham and Jolan Chang. Studies in History and Philosophy of Biological and Biomedical Sciences, 43(3):611–626, September 2012. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848612000568.

Roos:2013:EAT

[Roo13] Anna Marie Roos. The experimental approach towards a historiography of alchemy (reviewing L. M. Principe, The Secrets of Alchemy). Studies in History and Philosophy of Biological and Biomedical Sciences, 44(4):787–789, December 2013. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848613001143.

Rosario:2014:FOS

[Ros14] Vernon A. Rosario. Fustigating the “One-Sex-Body” thesis. Studies in History and Philosophy of Biological and Biomedical Sciences, 48 (part A)(??):112–114, December 2014. CODEN ???? ISSN 1369-8486 (print), 1879-2499 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S1369848614000880.

Rostvik:2017:CPS

[Røs17] Camilla Mørk Røstvik. The changing power of scientific institutions: the modern histories of Nature and The Royal Society. Studies in History and Philosophy of Biological and Biomedical Sciences, 61(??):46–49, February 2017. CODEN ???? ISSN 1369-8486 (print), 1879-2499 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S1369848616300565.

Rouyan:2019:RUM

[Rou19] Anahita Rouyan. Reforming uncultivated minds: the species transmutation debate and American science of life in the antebellum agricultural press, 1820–1859. Studies in History and Philosophy of Biological and Biomedical Sciences, 76(??):Article 101170, August 2019. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848617301371.

Rowbottom:2011:AIE

[Row11] Darrell P. Rowbottom. Approximations, idealizations and ‘experiments’ at the physics–biology interface. Studies in History and Philosophy of Biological and Biomedical Sciences, 42(2):145– 154, June 2011. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848610001147.

Ramsey:2016:ENT

[RP16] Grant Ramsey and Charles H. Pence. evoText: a new tool for analyzing the biological sciences. Studies in History and Philosophy of Biological and Biomedical Sciences, 57(??):83–87, June 2016. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S136984861630019X.

Robles-Pineros:2020:ISE

[RPLBMA20] Jairo Robles-Piñeros, David Ludwig, Geilsa Costa Santos Baptista, and Adela Molina-Andrade. Intercultural science education as a trading zone between traditional and academic knowledge. Studies in History and Philosophy of Biological and Biomedical Sciences, 84(??):Article 101337, December 2020. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848620301485.

Rees:2006:I

[RR06a] Amanda Rees and Gregory Radick. Introduction. Studies in History and Philosophy of Biological and Biomedical Sciences, 37 (2):269–272, June 2006. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848606000185.

Rosoff:2006:HDR

[RR06b] Philip M. Rosoff and Alex Rosenberg. How Darwinian reductionism refutes genetic determinism. Studies in History and Philosophy of Biological and Biomedical Sciences, 37(1):122–135, March 2006. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848605001007.

Rudolph:2011:SEH

[Rud11] John L. Rudolph. Science education: History at the edge. Studies in History and Philosophy of Biological and Biomedical Sciences, 42(2):270–273, June 2011. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848610001263.

Ruse:2000:TYT

[Rus00] Michael Ruse. Teleology: yesterday, today, and tomorrow? Studies in History and Philosophy of Biological and Biomedical Sciences, 31(1):213–232, March 2000. CO-DEN ???? ISSN 1369-8486 (print), 1879-2499 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S1369848699000461.

Ruse:2002:TEE

[Rus02] Michael Ruse. The temptations of evolutionary ethics. Studies in History and Philosophy of Biological and Biomedical Sciences, 33(4):753–760, December 2002. CODEN ???? ISSN 1369-8486 (print), 1879-2499 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S1369848602000195.

Ruse:2005:DMM

[Rus05] Michael Ruse. Darwinism and mechanism: metaphor in science. Studies in History and Philosophy of Biological and Biomedical Sciences, 36(2):285–302, June 2005. CO-DEN ???? ISSN 1369-8486 (print), 1879-2499 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S1369848605000208.

Ruse:2011:DPS

[Rus11] Michael Ruse. Is Darwinism past its “sell-by” date? The Origin of Species at 150. Studies in History and Philosophy of Biological and Biomedical Sciences, 42(1):5–11, March 2011. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848610001056. See [Dar59, Dar98, Dar08].

Ruse:2014:DVL

[Rus14] Michael Ruse. Darwin versus the Liberals: the third assault of the intelligent designers. Studies in History and Philosophy of Biological and Biomedical Sciences, 46(??):89–92, June 2014. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848614000272.

Ruse:2016:EBQ

[Rus16] Michael Ruse. Evolutionary biology and the question of teleology. Studies in History and Philosophy of Biological and Biomedical Sciences, 58(??):100–106, August 2016. CODEN ???? ISSN 1369-8486 (print), 1879-2499 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S1369848615001740.

Russell:2017:EMO

[Rus17] Edmund Russell. Evolution made to order: Plant breeding and technological innovation in Twentieth-Century America. Studies in History and Philosophy of Biological and Biomedical Sciences, 65(??):26–29, October 2017. CODEN ???? ISSN 1369-8486 (print), 1879-2499 (electronic). URL /science/article/pii/

S1369848617301061.

Rutkin:2010:MAG

[Rut10] H. Darrel Rutkin. Mysteries of attraction: Giovanni Pico della Mirandola, astrology and desire. Studies in History and Philosophy of Biological and Biomedical Sciences, 41(2):117–124, June 2010. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848610000166.

Ross:2016:KPI

[RW16] Lauren N. Ross and James F. Woodward. Koch’s postulates: an interventionist perspective. Studies in History and Philosophy of Biological and Biomedical Sciences, 59(??):35–46, October 2016. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848616300437.

Sachse:2014:NPB

[Sac14] Christian Sachse. The new puzzle of biological groups and individuals. Studies in History and Philosophy of Biological and Biomedical Sciences, 46(??):117–120, June 2014. CODEN ???? ISSN 1369-8486 (print), 1879-2499 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S1369848614000090.

Sagoff:2017:TEN

[Sag17] Mark Sagoff. Theoretical ecology has never been etiological: a reply to Donhauser. Studies in History and Philosophy of Biological and Biomedical Sciences, 63(??):64–69, June 2017. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848616301273. See [Don16].

Samuel:2013:ITV

[Sam13] Nina Samuel. Images as tools. On visual epistemic practices in the biological sciences. Studies in History and Philosophy of Biological and Biomedical Sciences, 44(2):225–236, June 2013. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848613000320.

Santesmases:2000:ITE

[San00] María Jesús Santesmases. From intestine transport to enzymatic regulation: The works of the Spanish biochemist Alberto Sols (1917–1989). Studies in History and Philosophy of Biological and Biomedical Sciences, 31(2):287–313, June 2000. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848699000291.

Santesmases:2002:NPI

[San02] María Jesús Santesmases. National politics and international trends: EMBO and the making of molecular biology in Spain (1960–1975). Studies in History and Philosophy of Biological and Biomedical Sciences, 33(3):473–487, September 2002. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848602000158.

Sankaran:2010:BRI

[San10] Neeraja Sankaran. The bacteriophage, its role in immunology: how Macfarlane Burnet’s phage research shaped his scientific style. Studies in History and Philosophy of Biological and Biomedical Sciences, 41(4):367–375, December 2010. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848610000749.

Sankaran:2012:HDR

[San12] Neeraja Sankaran. How the discovery of ribozymes cast RNA in the roles of both chicken and egg in origin-of-life theories. Studies in History and Philosophy of Biological and Biomedical Sciences, 43(4):741–750, December 2012. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848612000696.

Sankaran:2013:BSC

[San13] Neeraja Sankaran. Breaking with the self: Can continuity in immunology succeed? Studies in History and Philosophy of Biological and Biomedical Sciences, 44(2):242–246, June 2013. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848613000058.

Sankaran:2014:WVW

[San14a] Neeraja Sankaran. When viruses were not in style: Parallels in the histories of chicken sarcoma viruses and bacteriophages. Studies in History and Philosophy of Biological and Biomedical Sciences, 48 (part B)(??):189–199, December 2014. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848614000934.

Santesmases:2014:HAK

[San14b] María Jesús Santesmases. The human autonomous karyotype and the origins of prenatal testing: Children, pregnant women and early Down’s syndrome cytogenetics, Madrid 1962–1975. Studies in History and Philosophy of Biological and Biomedical Sciences, 47 (part A)(??):142–153, September 2014. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848614000545.

Sapp:1994:EAH

[Sap94] Jan Sapp. Evolution by association: a history of symbiosis. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 1994. ISBN 0-19-508821-2 (paperback), 0-19-508820-4 (hardcover). xvii + 255 pp. LCCN QH548

.S26 1994. URL http://www.loc.gov/catdir/enhancements/
fy0603/94006321-d.html;   http://www.loc.gov/catdir/
enhancements/fy0603/94006321-t.html;   http://www.loc.
gov/catdir/enhancements/fy0722/94006321-b.html.

Sapp:2007:SME

[Sap07] J. Sapp. The structure of microbial evolutionary theory. Studies in History and Philosophy of Biological and Biomedical Sciences, 38(4):780–795, December 2007. CO-DEN ???? ISSN 1369-8486 (print), 1879-2499 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S1369848607000568.

Sarkar:1998:BRE

[Sar98] Sahotra Sarkar. Book review: Evolution by association: A history of symbiosis: Jan Sapp, (New York and Oxford: Oxford University Press, 1994), xvii + 255 pp. ISBN 0-19-508820-4 cloth; 0-19-508821-2 paperback £19.95. Studies in History and Philosophy of Biological and Biomedical Sciences, 29(1):211–218, March 1998. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848698000107. See [Sap94].

Sartenaer:2013:NMD

[Sar13] Olivier Sartenaer. Neither metaphysical dichotomy nor pure identity:: Clarifying the emergentist creed. Studies in History and Philosophy of Biological and Biomedical Sciences, 44(3):365– 373, September 2013. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848613000423.

Sarkar:2014:EPT

[Sar14a] Sahotra Sarkar. Environmental philosophy: From theory to practice. Studies in History and Philosophy of Biological and Biomedical Sciences, 45(??):89–91, March 2014. CODEN ???? ISSN 1369-8486 (print), 1879-2499 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S1369848613001350.

Sarkar:2014:EPR

[Sar14b] Sahotra Sarkar. Environmental philosophy: Response to critics. Studies in History and Philosophy of Biological and Biomedical Sciences, 45(??):105–109, March 2014. CO-DEN ???? ISSN 1369-8486 (print), 1879-2499 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S1369848613001362.

Skyrms:2019:PCS

[SB19] Brian Skyrms and Jeffrey A. Barrett. Propositional content in signals. Studies in History and Philosophy of Biological and Biomedical Sciences, 74(??):34–39, April 2019. CODEN ???? ISSN 1369-8486 (print), 1879-2499 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S1369848618300918.

Svensson:2012:ALE

[SC12] Erik I. Svensson and Ryan Calsbeek, editors. The adaptive landscape in evolutionary biology. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2012. ISBN 0-19-959537-2 (hardcover), 0-19-959538-0 (paperback). xv + 319 + 2 pp. LCCN QH366.2 .A32 2012.

Sullivan-Clarke:2013:CEN

[SC13] Andrea Sullivan-Clarke. On the causal efficacy of natural selection: a response to Richards’ critique of the standard interpretation. Studies in History and Philosophy of Biological and Biomedical Sciences, 44(4):745–755, December 2013. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848613001258.

Stark:2014:SHS

[SC14] Laura Stark and Nancy D. Campbell. Stowaways in the history of science: the case of simian virus 40 and clinical research on federal prisoners at the US National Institutes of Health, 1960. Studies in History and Philosophy of Biological and Biomedical Sciences, 48 (part B)(??):218–230, December 2014. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848614000922.

Schickore:2003:PGA

[Sch03a] Jutta Schickore. The ‘philosophical grasp of the appearances’ and experimental microscopy: Johannes Müller’s microscopical research, 1824–1832. Studies in History and Philosophy of Biological and Biomedical Sciences, 34(4):569–592, December 2003. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848603000736.

Schlosser:2003:FNE

[Sch03b] Gerhard Schlosser. Functions-new essays in the philosophy of psychology and biology. Studies in History and Philosophy of Biological and Biomedical Sciences, 34(4):685–697, December 2003. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848603000797.

Schmidgen:2003:TNS

[Sch03c] Henning Schmidgen. Time and noise: the stable surroundings of reaction experiments, 1860–1890. Studies in History and Philosophy of Biological and Biomedical Sciences, 34(2):237–275, June 2003. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848603000244.

Schaffer:2010:ARM

[Sch10a] Simon Schaffer. The astrological roots of mesmerism. Studies in History and Philosophy of Biological and Biomedical Sciences, 41 (2):158–168, June 2010. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848610000208.

Schulz:2010:ITT

[Sch10b] Armin W. Schulz. It takes two: sexual strategies and game theory. Studies in History and Philosophy of Biological and Biomedical Sciences, 41(1):41–49, March 2010.

CODEN ???? ISSN 1369-8486 (print), 1879-2499 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S1369848609000739.

Schulz:2013:BRF

[Sch13a] Armin W. Schulz. The benefits of rule following: a new account of the evolution of desires. Studies in History and Philosophy of Biological and Biomedical Sciences, 44(4):595–603, December 2013. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848613001210.

Schyfter:2013:HDM

[Sch13b] Pablo Schyfter. How a ‘drive to make’ shapes synthetic biology. Studies in History and Philosophy of Biological and Biomedical Sciences, 44(4):632–640, December 2013. CO-DEN ???? ISSN 1369-8486 (print), 1879-2499 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S1369848613000782.

Scheffler:2014:FCV

[Sch14a] Robin Wolfe Scheffler. Following cancer viruses through the laboratory, clinic, and society. Studies in History and Philosophy of Biological and Biomedical Sciences, 48 (part B)(??):185–188, December 2014. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848614001393.

Scheffler:2014:MFS

[Sch14b] Robin Wolfe Scheffler. Managing the future: the Special Virus Leukemia Program and the acceleration of biomedical research. Studies in History and Philosophy of Biological and Biomedical Sciences, 48 (part B)(??):231–249, December 2014. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S136984861400140X.

Schulz:2015:ITA

[Sch15] Armin W. Schulz. Interdisciplinary thinking about mechanisms and causes. Studies in History and Philosophy of Biological and Biomedical Sciences, 50(??):94–97, April 2015. CODEN ???? ISSN 1369-8486 (print), 1879-2499 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S1369848615000394.

Scholl:2016:SDC

[Sch16a] Raphael Scholl. Spot the difference: Causal contrasts in scientific diagrams. Studies in History and Philosophy of Biological and Biomedical Sciences, 60(??):77–87, December 2016. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848616300450.

Schulz:2016:AEN

[Sch16b] Armin W. Schulz. Altruism, egoism, or neither: a cognitiveefficiency-based evolutionary biological perspective on helping behavior. Studies in History and Philosophy of Biological and Biomedical Sciences, 56(??):15–23, April 2016. CODEN ???? ISSN 1369-8486 (print), 1879-2499 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S1369848615001569.

Suarez-Diaz:2009:MEC

[SD09] Edna Suárez-Díaz. Molecular evolution: concepts and the origin of disciplines. Studies in History and Philosophy of Biological and Biomedical Sciences, 40(1):43–53, March 2009. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848608000940.

Suarez-Diaz:2014:IPM

[SD14] Edna Suárez-Díaz. Indigenous populations in Mexico: Medical anthropology in the work of Ruben Lisker in the 1960s. Studies in History and Philosophy of Biological and Biomedical Sciences, 47 (part A)(??):108–117, September 2014. CO-DEN ???? ISSN 1369-8486 (print), 1879-2499 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S136984861400051X.

Suarez-Diaz:2008:HOC

[SDAM08] Edna Suárez-Díaz and Victor H. Anaya-Muñoz. History, objectivity, and the construction of molecular phylogenies. Studies in History and Philosophy of Biological and Biomedical Sciences, 39(4):451–468, December 2008. CODEN ???? ISSN 1369-8486 (print), 1879-2499 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S1369848608000745.

Sechel:2012:MKI

[Sec12] Teodora Daniela Sechel. Medical knowledge and the improvement of vernacular languages in the Habsburg Monarchy: a case study from Transylvania (1770–1830). Studies in History and Philosophy of Biological and Biomedical Sciences, 43(3):720– 729, September 2012. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848612000283.

Sepkoski:2014:PST

[Sep14] David Sepkoski. Paleontology at the “high table”? Popularization and disciplinary status in recent paleontology. Studies in History and Philosophy of Biological and Biomedical Sciences, 45 (??):133–138, March 2014. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848613001684.

Sepkoski:2016:RLT

[Sep16] David Sepkoski. “Replaying Life’s Tape”: Simulations, metaphors, and historicity in Stephen Jay Gould’s view of life. Studies in History and Philosophy of Biological and Biomedical Sciences, 58(??):73–81, August 2016. CO-DEN ???? ISSN 1369-8486 (print), 1879-2499 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S136984861500182X.

Serpente:2011:CIS

[Ser11] Norberto Serpente. Cells from icons to symbols: Molecularizing cell biology in the 1980s. Studies in History and Philosophy of Biological and Biomedical Sciences, 42(4):403–411, December 2011. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848611000690.

Serpente:2016:JMI

[Ser16] Norberto Serpente. Justifying molecular images in cell biology textbooks: From constructions to primary data. Studies in History and Philosophy of Biological and Biomedical Sciences, 55(??):105–116, February 2016. CODEN ???? ISSN 1369-8486 (print), 1879-2499 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S1369848615001156.

Sesardic:2013:CAR

[Ses13] Neven Sesardic. Confusions about race: a new installment. Studies in History and Philosophy of Biological and Biomedical Sciences, 44(3):287–293, September 2013. CO-DEN ???? ISSN 1369-8486 (print), 1879-2499 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S1369848613000198.

Seth:2020:BRD

[Set20] Suman Seth. Book review: Difference and Disease by Suman Seth: Reply by the author. Studies in History and Philosophy of Biological and Biomedical Sciences, 80(??):Article 101218, April 2020. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848619301293. See [Sie20, Bra20, Góm20].

Stotz:2015:DDB

[SG15] Karola Stotz and Paul Griffiths. Dissecting developmental biology. Studies in History and Philosophy of Biological and Biomedical Sciences, 53(??):134–138, ???? 2015. CODEN ???? ISSN 1369-8486 (print), 1879-2499 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S1369848615001107.

Stotz:2004:HBC

[SGK04] Karola Stotz, Paul E. Griffiths, and Rob Knight. How biologists conceptualize genes: an empirical study. Studies in History and Philosophy of Biological and Biomedical Sciences, 35(4):647– 673, December 2004. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S136984860400069X.

Shanahan:2001:MCF

[Sha01] Timothy Shanahan. Methodological and contextual factors in the Dawkins/Gould dispute over evolutionary progress. Studies in History and Philosophy of Biological and Biomedical Sciences, 32(1):127–151, January 2001. CODEN ???? ISSN 1369-8486 (print), 1879-2499 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S136984860000025X.

Shavit:2004:SVP

[Sha04] Ayelet Shavit. Shifting values partly explain the debate over group selection. Studies in History and Philosophy of Biological and Biomedical Sciences, 35(4):697–720, December 2004. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848604000718.

Shapiro:2007:BSS

[Sha07] J. A. Shapiro. Bacteria are small but not stupid: cognition, natural genetic engineering and socio-bacteriology. Studies in History and Philosophy of Biological and Biomedical Sciences, 38(4): 807–819, December 2007. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848607000544.

Shanahan:2008:WDZ

[Sha08] Timothy Shanahan. Why don’t zebras have machine guns? Adaptation, selection, and constraints in evolutionary theory. Studies in History and Philosophy of Biological and Biomedical Sciences, 39(1):135–146, March 2008. CODEN ???? ISSN 1369-8486 (print), 1879-2499 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S1369848607000957.

Shanahan:2011:PID

[Sha11] Timothy Shanahan. Phylogenetic inertia and Darwin’s higher law. Studies in History and Philosophy of Biological and Biomedical Sciences, 42(1):60–68, March 2011. CODEN ???? ISSN 1369-8486 (print), 1879-2499 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S1369848610001068.

Shapiro:2014:DFE

[Sha14] Adam R. Shapiro. Darwin’s foil: the evolving uses of William Paley’s Natural Theology 1802–2005. Studies in History and Philosophy of Biological and Biomedical Sciences, 45(??):114–123, March 2014. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848613001726.

Shaw:2016:DGA

[Sha16] Jennifer Shaw. Documenting genomics: Applying archival theory to preserving the records of the Human Genome Project. Studies in History and Philosophy of Biological and Biomedical Sciences, 55(??):61–69, February 2016. CODEN ???? ISSN 1369-8486 (print), 1879-2499 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S1369848615001132.

Shelley:1999:MAE

[She99] Cameron Shelley. Multiple analogies in evolutionary biology. Studies in History and Philosophy of Biological and Biomedical Sciences, 30(2):143–180, June 1999. CO-DEN ???? ISSN 1369-8486 (print), 1879-2499 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S1369848698000302.

Shelley:2010:WTA

[She10] Cameron Shelley. Why test animals to treat humans? On the validity of animal models. Studies in History and Philosophy of Biological and Biomedical Sciences, 41(3):292–299, September 2010. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848610000361.

Sheldrake:2012:AHM

[She12a] Merlin Sheldrake. Albert Howard and the mycorrhizal association. Studies in History and Philosophy of Biological and Biomedical Sciences, 43(1):225–231, March 2012. CODEN ???? ISSN 1369-8486 (print), 1879-2499 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S1369848611001269.

Shelley:2012:D

[She12b] Cameron Shelley. Discussion. Studies in History and Philosophy of Biological and Biomedical Sciences, 43(1):311, March 2012. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848611000896.

Sheldon:2014:CDS

[She14] Myrna Perez Sheldon. Claiming Darwin: Stephen Jay Gould in contests over evolutionary orthodoxy and public perception, 1977–2002. Studies in History and Philosophy of Biological and Biomedical Sciences, 45(??):139–147, March 2014. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848613001271.

Sheredos:2017:CSG

[She17] Benjamin Sheredos. Communicating with scientific graphics: a descriptive inquiry into non-ideal normativity. Studies in History and Philosophy of Biological and Biomedical Sciences, 63(??):32– 44, June 2017. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848617300547.

Shickle:2006:CPW

[Shi06] Darren Shickle. The consent problem within DNA biobanks. Studies in History and Philosophy of Biological and Biomedical Sciences, 37(3):503–519, September 2006. CODEN ???? ISSN 1369-8486 (print), 1879-2499 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S1369848606000471.

Shmidt:2020:RSC

[Shm20] Victoria Shmidt. Race science in Czechoslovakia: Serving segregation in the name of the nation. Studies in History and Philosophy of Biological and Biomedical Sciences, 83(??):Article 101241, October 2020. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848618301365.

Siena:2020:BRD

[Sie20] Kevin Siena. Book review: Difference and Disease by Suman Seth: a View from London. Studies in History and Philosophy of Biological and Biomedical Sciences, 80(??):Article 101219, April 2020. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848619301256. See reply [Set20].

Silver:2001:CML

[Sil01] Lee M. Silver. Confused meanings of life, genes and parents. Studies in History and Philosophy of Biological and Biomedical Sciences, 32(4):647–661, December 2001. CO-DEN ???? ISSN 1369-8486 (print), 1879-2499 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S1369848601000322.

Simon:1999:NTH

[Sim99] Jonathan Simon. Naming and toxicity: a history of strychnine. Studies in History and Philosophy of Biological and Biomedical Sciences, 30(4):505–525, December 1999. CO-DEN ???? ISSN 1369-8486 (print), 1879-2499 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S1369848698000314.

Sinding:2004:SMF

[Sin04] Christiane Sinding. The specificity of medical facts: the case of diabetology. Studies in History and Philosophy of Biological and Biomedical Sciences, 35(3):545–559, September 2004. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848604000512.

Sinclair:2008:SIL

[Sin08] Alison Sinclair. Social imaginaries: the literature of eugenics. Studies in History and Philosophy of Biological and Biomedical Sciences, 39(2):240–246, June 2008. CO-DEN ???? ISSN 1369-8486 (print), 1879-2499 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S1369848608000216.

Skopek:2011:PEU

[Sko11] Jeffrey M. Skopek. Principles, exemplars, and uses of history in early 20th century genetics. Studies in History and Philosophy of Biological and Biomedical Sciences, 42(2):210–225, June 2011. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848610001093.

Sterner:2014:NSM

[SL14] Beckett Sterner and Scott Lidgard. The normative structure of mathematization in systematic biology. Studies in History and Philosophy of Biological and Biomedical Sciences, 46(??):44–54, June 2014. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848614000302.

Sleigh:1998:LDG

[Sle98] Charlotte Sleigh. Life, death and galvanism. Studies in History and Philosophy of Biological and Biomedical Sciences, 29(2):219– 248, September 1998. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848698000168.

Sleigh:2009:PBP

[Sle09] Charlotte Sleigh. Plastic body, permanent body: Czech representations of corporeality in the early twentieth century. Studies in History and Philosophy of Biological and Biomedical Sciences, 40(4):241–255, December 2009. CO-DEN ???? ISSN 1369-8486 (print), 1879-2499 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S1369848609000508.

Sloan:2006:KHN

[Slo06] Phillip R. Sloan. Kant on the history of nature: The ambiguous heritage of the critical philosophy for natural history. Studies in History and Philosophy of Biological and Biomedical Sciences, 37(4):627–648, December 2006. CO-DEN ???? ISSN 1369-8486 (print), 1879-2499 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S1369848606000690.

Sloan:2012:HWT

[Slo12] Phillip R. Sloan. How was teleology eliminated in early molecular biology? Studies in History and Philosophy of Biological and Biomedical Sciences, 43(1):140–151, March 2012. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848611000495.

Sloan:2013:SPH

[Slo13] Phillip R. Sloan. The species problem and history. Studies in History and Philosophy of Biological and Biomedical Sciences, 44 (2):237–241, June 2013. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848613000034.

Sloan:2014:ERK

[Slo14a] Phillip R. Sloan. The essence of race: Kant and Late Enlightenment Reflections. Studies in History and Philosophy of Biological and Biomedical Sciences, 47 (part A)(??):191–195, September 2014. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848614000752.

Sloane:2014:SNT

[Slo14b] Jesse D. Sloane. The state, the nation, and their limits: Recent publications on the history of Chinese medicine. Studies in History and Philosophy of Biological and Biomedical Sciences, 47 (part A)(??):218–223, September 2014. CO-DEN ???? ISSN 1369-8486 (print), 1879-2499 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S1369848614000065.

Sytsma:2012:TTT

[SLR12] Justin Sytsma, Jonathan Livengood, and David Rose. Two types of typicality: Rethinking the role of statistical typicality in ordinary causal attributions. Studies in History and Philosophy of Biological and Biomedical Sciences, 43(4):814–820, December 2012. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848612000659.

Sprevak:2004:MSP

[SM04] Mark Sprevak and Christina McLeish. Magic, semantics, and Putnam’s vat brains. Studies in History and Philosophy of Biological and Biomedical Sciences, 35(2):227–236, June 2004. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848604000184.

Skipper:2005:TAE

[SM05] Robert A. Skipper, Jr. and Roberta L. Millstein. Thinking about evolutionary mechanisms: natural selection. Studies in History and Philosophy of Biological and Biomedical Sciences, 36(2):327– 347, June 2005. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848605000221.

Smart:2006:PPA

[SM06] Andrew Smart and Paul Martin. The promise of pharmacogenetics: assessing the prospects for disease and patient stratification. Studies in History and Philosophy of Biological and Biomedical Sciences, 37(3):583–601, September 2006. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848606000525.

Sturm:2012:CDP

[SM12] Thomas Sturm and Annette Mülberger. Crisis discussions in psychology — new historical and philosophical perspectives. Studies in History and Philosophy of Biological and Biomedical Sciences, 43(2):425–433, June 2012. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848611001270.

Smart:2004:BVQ

[Sma04] J. J. C. Smart. The brain in the vat and the question of metaphysical realism. Studies in History and Philosophy of Biological and Biomedical Sciences, 35(2):237–247, June 2004. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848604000196.

Smead:2013:EAI

[Sme13] Rory Smead. Evolution and apparent irrationality (reviewing S. Okasha and K. Binmore, eds, Evolution and rationality: Decisions, co-operation and strategic behavior). Studies in History and Philosophy of Biological and Biomedical Sciences, 44(3):451– 454, September 2013. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848613000769.

Smith:2005:HPC

[Smi05] Roger Smith. The history of psychological categories. Studies in History and Philosophy of Biological and Biomedical Sciences, 36 (1):55–94, March 2005. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848604000962.

Smith:2009:EVB

[Smi09] Robyn Smith. The emergence of vitamins as bio-political objects during World War I. Studies in History and Philosophy of Biological and Biomedical Sciences, 40(3):179–189, September 2009. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848609000326.

Smit:2010:WWH

[Smi10] Harry Smit. Weismann, Wittgenstein and the homunculus fallacy. Studies in History and Philosophy of Biological and Biomedical Sciences, 41(3):263–271, September 2010. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848610000385.

Smith:2012:DEV

[Smi12] Justin E. H. Smith. Diet, embodiment, and virtue in the mechanical philosophy. Studies in History and Philosophy of Biological and Biomedical Sciences, 43(2):338–348, June 2012. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S136984861100118X.

Smith:2013:PRS

[Smi13] Elise Juzda Smith. Putting racial science in its place (reviewing A. Fabian, The skull collectors. Race, science, and America’s unburied dead and B. R. Brown, Until Darwin, science, human variety and the origins of race). Studies in History and Philosophy of Biological and Biomedical Sciences, 44(3):443–446, September 2013. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S136984861300040X.

Smith:2020:IMT

[Smi20] Roger Smith. Inhibition and metaphor of top-down organization. Studies in History and Philosophy of Biological and Biomedical Sciences, 83(??):Article 101253, October 2020. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848620300029.

Smoller:2010:TAS

[Smo10] Laura Ackerman Smoller. Teste Albumasare cum Sibylla: astrology and the Sibyls in medieval Europe. Studies in History and Philosophy of Biological and Biomedical Sciences, 41(2):76–89, June 2010. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848610000129.

Smocovitis:2014:DPE

[Smo14] Vassiliki Betty Smocovitis. Disciplining and popularizing: Evolution and its publics from the modern synthesis to the present. Studies in History and Philosophy of Biological and Biomedical Sciences, 45(??):111–113, March 2014. CODEN ???? ISSN 1369-8486 (print), 1879-2499 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S136984861300126X.

Smocovitis:2015:CVE

[Smo15] Vassiliki Betty Smocovitis. Commentary: The variation and evolution of plants: Historical perspectives. Studies in History and Philosophy of Biological and Biomedical Sciences, 50(??):77–79, April 2015. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848615000084.

Snobelen:2001:SMA

[Sno01]  Stephen David Snobelen.  Of stones, men and angels:

The competing myth of Isabelle Duncan’s pre-Adamite man (1860). Studies in History and Philosophy of Biological and Biomedical Sciences, 32(1):59–104, January 2001. CO-DEN ???? ISSN 1369-8486 (print), 1879-2499 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S1369848600000212.

Sober:2013:TFP

[Sob13] Elliott Sober. Trait fitness is not a propensity, but fitness variation is. Studies in History and Philosophy of Biological and Biomedical Sciences, 44(3):336–341, September 2013. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848613000162.

Socha:2020:SEB

[Soc20] Yuirubán Hernández Socha. Scientific encounters between Colombia and the United States analyzed through publishing practices in Caldasia journal: the birds of the Republic of Colombia as a publishing event. Studies in History and Philosophy of Biological and Biomedical Sciences, 82(??):Article 101289, August 2020. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848619301323.

Sokoloff:2002:RNB

[Sok02] Leon Sokoloff. Refugees from Nazism and the biomedical publishing industry. Studies in History and Philosophy of Biological and Biomedical Sciences, 33(2):315–324, July 2002. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S003936810200016X.

Solomon:2000:TGD

[Sol00] Susan Gross Solomon. ‘through a glass darkly’: The Rockefeller Foundation’s International Health Board and Soviet Public Health. Studies in History and Philosophy of Biological and Biomedical Sciences, 31(3):409–418, September 2000. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848600000108.

Solomon:2020:MNJ

[Sol20] Miriam Solomon. Medical Nihilism by Jacob Stegenga: After medical nihilism. Studies in History and Philosophy of Biological and Biomedical Sciences, 81(??):Article 101271, June 2020. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848620300352. See [Ful20, Hea20, Hea20,

Gab20, Ste20].

Sommer:2014:PRH

[Som14] Andreas Sommer. Psychical research in the history and philosophy of science. an introduction and review. Studies in History and Philosophy of Biological and Biomedical Sciences, 48 (part A)(??):38–45, December 2014. CO-DEN ???? ISSN 1369-8486 (print), 1879-2499 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S1369848614001204.

Soom:2012:MDM

[Soo12] Patrice Soom. Mechanisms, determination and the metaphysics of neuroscience. Studies in History and Philosophy of Biological and Biomedical Sciences, 43(3):655–664, September 2012. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848612000684.

Scott-Phillips:2016:CCE

[SP16] Thomas C. Scott-Phillips. Can cultural evolution bridge scientific continents? Studies in History and Philosophy of Biological and Biomedical Sciences, 57(??):170–173, June 2016. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848616000194.

Spary:2012:ICP

[Spa12] E. C. Spary. Introduction: Centre and periphery in the eighteenth-century Habsburg ‘medical empire’. Studies in History and Philosophy of Biological and Biomedical Sciences, 43(3): 684–690, September 2012. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848612000325.

Speake:2011:IMI

[Spe11] Stephen W. Speake. Infectious milk: issues of pathogenic certainty within ideational regimes and their biopolitical implications. Studies in History and Philosophy of Biological and Biomedical Sciences, 42(4):530–541, December 2011. CODEN ???? ISSN 1369-8486 (print), 1879-2499 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S1369848611000549.

Spencer:2013:IST

[Spe13] Quayshawn Spencer. Introduction to “Is There Space for Race in Evolutionary Biology?”. Studies in History and Philosophy of Biological and Biomedical Sciences, 44(3):247–249, September 2013. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848613000435.

Spencer:2014:URN

[Spe14] Quayshawn Spencer. The unnatural racial naturalism. Studies in History and Philosophy of Biological and Biomedical Sciences, 46(??):38–43, June 2014. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848614000284.

Spencer:2015:PRM

[Spe15] Quayshawn Spencer. Philosophy of race meets population genetics. Studies in History and Philosophy of Biological and Biomedical Sciences, 52(??):46–55, August 2015. CODEN ???? ISSN 1369-8486 (print), 1879-2499 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S1369848615000655.

Sera-Shriar:2011:EMR

[SS11] Efram Sera-Shriar. Ethnology in the metropole: Robert Knox, Robert Gordon Latham and local sites of observational training. Studies in History and Philosophy of Biological and Biomedical Sciences, 42(4):486–496, December 2011. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848611000677.

Sera-Shriar:2015:AOR

[SS15a] Efram Sera-Shriar. Arctic observers: Richard King, monogenism and the historicisation of Inuit through travel narratives. Studies in History and Philosophy of Biological and Biomedical Sciences, 51(??):23–31, June 2015. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848615000308.

Sera-Shriar:2015:HHD

[SS15b] Efram Sera-Shriar. Human history and deep time in nineteenthcentury British sciences: an introduction. Studies in History and Philosophy of Biological and Biomedical Sciences, 51(??):19–22, June 2015. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848615000357.

Surman:2014:NCH

[SSH14] Jan Surman, Katalin Stráner, and Peter Haslinger. Nomadic concepts in the history of biology. Studies in History and Philosophy of Biological and Biomedical Sciences, 48 (part B)(??): 127–129, December 2014. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848614000958.

Sepkoski:2017:ITG

[ST17] David Sepkoski and Marco Tamborini. Introduction: Towards a global history of paleontology: the paleontological reception of Darwin’s thought. Studies in History and Philosophy of Biological and Biomedical Sciences, 66(??):1–2, December 2017. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848617301632.

Stahnisch:2012:ENN

[Sta12] Frank W. Stahnisch. The emergence of Nervennahrung: Nerves, mind and metabolism in the long eighteenth century. Studies in History and Philosophy of Biological and Biomedical Sciences, 43 (2):405–417, June 2012. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848611001245.

Stark:2013:GAR

[Sta13] James F. Stark. Not by germs alone (reviewing C. Gradman and E. Forster, (trans.), Laboratory disease: Robert Koch’s medical bacteriology). Studies in History and Philosophy of Biological and Biomedical Sciences, 44(3):435–438, September 2013. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848613000393.

Steigerwald:2002:IJI

[Ste02a] Joan Steigerwald. Instruments of judgment: Inscribing organic processes in Late Eighteenth-Century Germany. Studies in History and Philosophy of Biological and Biomedical Sciences, 33 (1):79–131, March 2002. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S136984860100036X.

Sterrett:2002:DAB

[Ste02b] Susan G. Sterrett. Darwin’s analogy between artificial and natural selection: how does it go? Studies in History and Philosophy of Biological and Biomedical Sciences, 33(1):151–168, March 2002. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848601000395.

Sterelny:2004:PBM

[Ste04] Kim Sterelny. Po-bo man? Studies in History and Philosophy of Biological and Biomedical Sciences, 35(4):729–741, December 2004. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848604000731.

Stephens:2005:WCE

[Ste05a] Christopher Stephens. What can evolutionary theory teach us about human nature? Studies in History and Philosophy of Biological and Biomedical Sciences, 36(1):221–232, March 2005. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848604000998.

Sterelny:2005:AVL

[Ste05b] Kim Sterelny. Another view of life. Studies in History and Philosophy of Biological and Biomedical Sciences, 36(3):585– 593, September 2005. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848605000531.

Steigerwald:2006:IKT

[Ste06a] Joan Steigerwald. Introduction: Kantian teleology and the biological sciences. Studies in History and Philosophy of Biological and Biomedical Sciences, 37(4):621–626, December 2006. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848606000689.

Steigerwald:2006:KCN

[Ste06b] Joan Steigerwald. Kant’s concept of natural purpose and the reflecting power of judgement. Studies in History and Philosophy of Biological and Biomedical Sciences, 37(4):712–734, December 2006. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848606000732.

Stern:2008:YNB

[Ste08] Megan Stern. ‘yes:-no:-I have been sleeping-and now-now-I am dead’: undeath, the body and medicine. Studies in History and Philosophy of Biological and Biomedical Sciences, 39(3):347– 354, September 2008. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848608000435.

Stegmann:2010:RBN

[Ste10a] Ulrich Stegmann. Reply to Bence Nanay’s ‘Natural selection and the limited nature of environmental resources’. Studies in History and Philosophy of Biological and Biomedical Sciences, 41(4):420– 421, December 2010. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848610000701. See [Nan10].

Stegmann:2010:WCN

[Ste10b] Ulrich E. Stegmann. What can natural selection explain? Studies in History and Philosophy of Biological and Biomedical Sciences, 41(1):61–66, March 2010. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848609000715.

Steel:2011:EUF

[Ste11a] Daniel Steel. Extrapolation, uncertainty factors, and the precautionary principle. Studies in History and Philosophy of Biological and Biomedical Sciences, 42(3):356–364, September 2011. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848611000173.

Stegenga:2011:MAP

[Ste11b] Jacob Stegenga. Is meta-analysis the platinum standard of evidence? Studies in History and Philosophy of Biological and Biomedical Sciences, 42(4):497–507, December 2011. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848611000665.

Stegenga:2015:EMI

[Ste15a] Jacob Stegenga. Effectiveness of medical interventions. Studies in History and Philosophy of Biological and Biomedical Sciences, 54(??):34–44, December 2015. CODEN ???? ISSN 1369-8486 (print), 1879-2499 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S1369848615000850.

Stegenga:2015:ME

[Ste15b] Jacob Stegenga. Measuring effectiveness. Studies in History and Philosophy of Biological and Biomedical Sciences, 54(??): 62–71, December 2015. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S1369848615000837. See comment

[Juk17].

Steigerwald:2016:EIM

[Ste16] Joan Steigerwald. Entanglements of instruments and media in investigating organic life. Studies in History and Philosophy of Biological and Biomedical Sciences, 57(??):107–111, June 2016. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848616300127.

Sterelny:2017:CEC

[Ste17] Kim Sterelny. Cultural evolution in California and Paris. Studies in History and Philosophy of Biological and Biomedical Sciences, 62(??):42–50, April 2017. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848616301078.

Stegenga:2020:MNJ

[Ste20] Jacob Stegenga. Medical Nihilism by Jacob Stegenga: Reply by the author. Studies in History and Philosophy of Biological and Biomedical Sciences, 81(??):Article 101274, June 2020. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848620300285. See [Ful20, Sol20, Hea20,

Gab20].

Stich:2016:WTM

[Sti16] Stephen Stich. Why there might not be an evolutionary explanation for psychological altruism. Studies in History and Philosophy of Biological and Biomedical Sciences, 56(??):3–6, April 2016. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848615001557.

Stott:2004:THM

[Sto04] Rebecca Stott. Thomas Huxley: Making the ‘man of science’. Studies in History and Philosophy of Biological and Biomedical Sciences, 35(1):199–207, March 2004. CO-DEN ???? ISSN 1369-8486 (print), 1879-2499 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S1369848603001201.

Stolberg:2012:APF

[Sto12] Michael Stolberg. ‘abhorreas pinguedinem’: Fat and obesity in early modern medicine (c. 1500–1750). Studies in History and Philosophy of Biological and Biomedical Sciences, 43(2):370–378, June 2012. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S136984861100121X.

Strasser:2002:IMB

[Str02] Bruno J. Strasser. Institutionalizing molecular biology in postwar Europe: a comparative study. Studies in History and Philosophy of Biological and Biomedical Sciences, 33(3):515– 546, September 2002. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S136984860200016X.

Strasser:2012:DDS

[Str12] Bruno J. Strasser. Data-driven sciences: From wonder cabinets to electronic databases. Studies in History and Philosophy of Biological and Biomedical Sciences, 43(1):85–87, March 2012. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848611000859.

Strick:2014:CLC

[Str14] James Strick. The cycle of life concept, soil microbiology and soil science restored to the history of ecology. Studies in History and Philosophy of Biological and Biomedical Sciences, 48 (part A)(??):119–121, December 2014. CO-DEN ???? ISSN 1369-8486 (print), 1879-2499 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S1369848614001277.

Sturm:2012:BPK

[Stu12] Thomas Sturm. Bühler and Popper: Kantian therapies for the crisis in psychology. Studies in History and Philosophy of Biological and Biomedical Sciences, 43(2):462–472, June 2012. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848611001324.

Suarez:2001:SDC

[Suá01] Edna Suárez. Satellite-DNA: a case-study for the evolution of experimental techniques. Studies in History and Philosophy of Biological and Biomedical Sciences, 32(1):31–57, January 2001. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848600000261.

Sullivan:2020:TMB

[Sul20] Jacqueline Sullivan. Are there model behaviors for model organism research? Commentary on Nicole Nelson’s Model Behavior (Chicago 2018). Studies in History and Philosophy of Biological and Biomedical Sciences, 82(??):Article 101266, August 2020. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848620300054.

Sunderland:2011:MDS

[Sun11] Mary Evelyn Sunderland. Morphogenesis, Dictyostelium, and the search for shared developmental processes. Studies in History and Philosophy of Biological and Biomedical Sciences, 42(4):508– 517, December 2011. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848611000653.

Swazo:2005:RIR

[Swa05] Norman K. Swazo. Research integrity and rights of indigenous peoples: appropriating Foucault’s critique of knowledge/power. Studies in History and Philosophy of Biological and Biomedical Sciences, 36(3):568–584, September 2005. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848605000518.

Swenson:2015:MBW

[Swe15] Sarah A. Swenson. ‘from man to bacteria’: W. D. Hamilton, the theory of inclusive fitness, and the post-war social order. Studies in History and Philosophy of Biological and Biomedical Sciences, 49(??):45–54, February 2015. CODEN ???? ISSN 1369-8486 (print), 1879-2499 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S1369848614001769.

Swiatczak:2017:TEV

[Swi17] Bartlomiej Swiatczak. Towards an ecological view of immunity. Studies in History and Philosophy of Biological and Biomedical Sciences, 63(??):85–88, June 2017. CO-DEN ???? ISSN 1369-8486 (print), 1879-2499 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S1369848617300316.

Tabb:2016:DOB

[Tab16] Kathryn Tabb. Darwin at Orchis Bank: Selection after the Origin. Studies in History and Philosophy of Biological and Biomedical Sciences, 55(??):11–20, February 2016. CODEN ???? ISSN 1369-8486 (print), 1879-2499 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S1369848615001697.

Tamborini:2017:RDL

[Tam17] Marco Tamborini. The reception of Darwin in late nineteenthcentury German paleontology as a case of pyrrhic victory. Studies in History and Philosophy of Biological and Biomedical Sciences, 66(??):37–45, December 2017. CO-DEN ???? ISSN 1369-8486 (print), 1879-2499 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S1369848617300523.

Tambolo:2020:UMC

[Tam20] Luca Tambolo. An unappreciated merit of counterfactual histories of science. Studies in History and Philosophy of Biological and Biomedical Sciences, 81(??):Article 101183, June 2020. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848618300736.

Tansey:2003:HDH

[Tan03] E. M. Tansey. Henry Dale, histamine and anaphylaxis: reflections on the role of chance in the history of allergy. Studies in History and Philosophy of Biological and Biomedical Sciences, 34(3):455– 472, September 2003. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848603000517.

Taub:2006:PNE

[Tau06] Liba Taub. Preserving nature? Ecology, tourism and other themes in the national parks. Studies in History and Philosophy of Biological and Biomedical Sciences, 37(3):602–611, September 2006. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848606000537.

Taylor:2009:NRA

[Tay09] Peter J. Taylor. Nothing reliable about genes or environment: new perspectives on analysis of similarity among relatives in light of the possibility of underlying heterogeneity. Studies in History and Philosophy of Biological and Biomedical Sciences, 40(3):210– 220, September 2009. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848609000302.

Taylor:2015:ERD

[Tay15] Peter J. Taylor. Essay review: Distinctions that make a difference? Beyond versus: The struggle to understand the interaction of nature and nurture, James Tabery; MIT Press, Cambridge, MA, 2014, pp. xiii + 279, Price US$40.00, £29.95 hardback, ISBN: 978-0-262-02737-3. Studies in History and Philosophy of Biological and Biomedical Sciences, 51(??):70–76, June 2015. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848615000412.

Thoren:2016:SSS

[TB16] Henrik Thorén and Line Breian. Stepping stone or stumbling block? Mode 2 knowledge production in sustainability science. Studies in History and Philosophy of Biological and Biomedical Sciences, 56(??):71–81, April 2016. CODEN ???? ISSN 1369-8486 (print), 1879-2499 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S1369848615001636.

Tee:2018:MDA

[Tee18]  Sim-Hui Tee.  Mechanism diagrams and abstraction-by-

aggregation. Studies in History and Philosophy of Biological and Biomedical Sciences, 71(??):17–25, October 2018. CODEN ???? ISSN 1369-8486 (print), 1879-2499 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S1369848617301668.

Tee:2019:FEM

[Tee19] Sim-Hui Tee. Fictional experimental modeling in biology: In vivo representation. Studies in History and Philosophy of Biological and Biomedical Sciences, 74(??):1–6, April 2019. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848617300845.

Teira:2013:IEB

[Tei13] David Teira. On the impartiality of early British clinical trials. Studies in History and Philosophy of Biological and Biomedical Sciences, 44(3):412–418, September 2013. CODEN ???? ISSN 1369-8486 (print), 1879-2499 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S136984861300071X.

Teira:2019:PTM

[Tei19] David Teira. Placebo trials without mechanisms: How far can they go? Studies in History and Philosophy of Biological and Biomedical Sciences, 77(??):Article 101177, October 2019. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848618300116.

Templeton:2013:BRH

[Tem13] Alan R. Templeton. Biological races in humans. Studies in History and Philosophy of Biological and Biomedical Sciences, 44(3): 262–271, September 2013. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848613000460.

Teufel:2011:WCT

[Teu11] Thomas Teufel. Wholes that cause their parts: Organic selfreproduction and the reality of biological teleology. Studies in History and Philosophy of Biological and Biomedical Sciences, 42 (2):252–260, June 2011. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848610001226.

Thagard:1998:UBD

[Tha98a] Paul Thagard. Ulcers and bacteria I: discovery and acceptance. Studies in History and Philosophy of Biological and Biomedical Sciences, 29(1):107–136, March 1998. CODEN ???? ISSN 1369-8486 (print), 1879-2499 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S1369848698000065.

Thagard:1998:UBI

[Tha98b] Paul Thagard. Ulcers and bacteria II: Instruments, experiments, and social interactions. Studies in History and Philosophy of Biological and Biomedical Sciences, 29(2):317–342, September 1998. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848698000247.

Theunissen:2000:TRS

[The00] Bert Theunissen. Turning refracting into a science: F. C. Donders’ ‘scientific reform’ of lens prescription. Studies in History and Philosophy of Biological and Biomedical Sciences, 31(4):557– 578, December 2000. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848600000121.

Theunissen:2014:PAB

[The14] Bert Theunissen. Practical animal breeding as the key to an integrated view of genetics, eugenics and evolutionary theory: Arend L. Hagedoorn (1885–1953). Studies in History and Philosophy of Biological and Biomedical Sciences, 46(??):55–64, June 2014. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848614000338.

Thomson:2001:PHE

[Tho01] Elaine Thomson. Physiology, hygiene and the entry of women to the medical profession in Edinburgh c. 1869– c. 1900. Studies in History and Philosophy of Biological and Biomedical Sciences, 32(1):105–126, January 2001. CODEN ???? ISSN 1369-8486 (print), 1879-2499 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S1369848600000273.

Thomas:2006:YHE

[Tho06] Marion Thomas. Yerkes, Hamilton and the experimental study of the ape mind: from evolutionary psychiatry to eugenic politics. Studies in History and Philosophy of Biological and Biomedical Sciences, 37(2):273–294, June 2006. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848606000197.

Thompson:2015:LGC

[Tho15] Paul Thompson. Lessons from the Gaia controversy. Studies in History and Philosophy of Biological and Biomedical Sciences, 49(??):75–78, February 2015. CODEN ???? ISSN 1369-8486 (print), 1879-2499 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S1369848614001551.

Thomas:2016:BBP

[Tho16] Marion Thomas. Between biomedical and psychological experiments: the unexpected connections between the Pasteur Institutes and the study of animal mind in the second quarter of twentieth-century France. Studies in History and Philosophy of Biological and Biomedical Sciences, 55(??):29–40, February 2016. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848615001600.

Thompson:2017:DTR

[Tho17] R. Paul Thompson. Darwin and teleology: Redefinition or historicizing? Studies in History and Philosophy of Biological and Biomedical Sciences, 63(??):94–97, June 2017. CODEN ???? ISSN 1369-8486 (print), 1879-2499 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S1369848617300201.

Tjiattas:2001:IMC

[Tji01] Mary Tjiattas. Interdisciplinary methodology: The case of Kitcher’s Freud. Studies in History and Philosophy of Biological and Biomedical Sciences, 32(3):535–555, September 2001. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848601000152.

Toepfer:2012:TCR

[Toe12] Georg Toepfer. Teleology and its constitutive role for biology as the science of organized systems in nature. Studies in History and Philosophy of Biological and Biomedical Sciences, 43(1):113–119, March 2012. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S136984861100046X.

Toombs:2003:HEC

[Too03] Ethan Toombs. Harmony, explanatory coherence and the debate between the reticular theory and neuron theory of nerve cell structure: ECHO’s resolution of a quiet revolution. Studies in History and Philosophy of Biological and Biomedical Sciences, 34(4):615–632, December 2003. CO-DEN ???? ISSN 1369-8486 (print), 1879-2499 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S1369848603000748.

Topp:2007:PIT

[Top07] Leslie Topp. Psychiatric institutions, their architecture, and the politics of regional autonomy in the Austro–Hungarian monarchy. Studies in History and Philosophy of Biological and Biomedical Sciences, 38(4):733–755, December 2007. CODEN ???? ISSN 1369-8486 (print), 1879-2499 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S1369848607000441.

Torrens:2013:ERV

[Tor13] Erica Torrens. Essay review: Visualizing the order of nature: Trees of life: A visual history of evolution. Theodore W. Pietsch. The Johns Hopkins University Press, Baltimore (2012). pp. 358, Price $58.99, Hardcover, ISBN-13: 978-1-4214-0479-0. Studies in History and Philosophy of Biological and Biomedical Sciences, 44 (1):110–113, March 2013. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848612001288.

Tosh:2002:PEP

[Tos02] Nick Tosh. Possession, exorcism and psychoanalysis. Studies in History and Philosophy of Biological and Biomedical Sciences, 33(4):583–596, December 2002. CODEN ???? ISSN 1369-8486 (print), 1879-2499 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S1369848602000225.

Totelin:2007:SVH

[Tot07] Laurence M. V. Totelin. Sex and vegetables in the Hippocratic gynaecological treatises. Studies in History and Philosophy of Biological and Biomedical Sciences, 38(3):531–540, September 2007. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848607000325.

Tresch:2004:SPR

[Tre04] John Tresch. In a solitary place: Raymond Roussel’s brain and the French cult of unreason. Studies in History and Philosophy of Biological and Biomedical Sciences, 35(2):307–332, June 2004. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848604000238.

Tresker:2020:TCD

[Tre20a] Steven Tresker. Theoretical and clinical disease and the biostatistical theory. Studies in History and Philosophy of Biological and Biomedical Sciences, 82(??):Article 101249, August 2020. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S136984861930038X.

Tresker:2020:TCC

[Tre20b] Steven Tresker. A typology of clinical conditions. Studies in History and Philosophy of Biological and Biomedical Sciences, 83(??):Article 101291, October 2020. CO-DEN ???? ISSN 1369-8486 (print), 1879-2499 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S1369848619301645.

Trivino:2020:HEC

[TS20] Vanessa Triviño and Javier Suárez. Holobionts: Ecological communities, hybrids, or biological individuals? A metaphysical perspective on multispecies systems. Studies in History and Philosophy of Biological and Biomedical Sciences, 84(??):Article 101323, December 2020. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848620301345.

Tsou:2012:ICR

[Tso12] Jonathan Y. Tsou. Intervention, causal reasoning, and the neurobiology of mental disorders: Pharmacological drugs as experimental instruments. Studies in History and Philosophy of Biological and Biomedical Sciences, 43(2):542–551, June 2012. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848612000040.

Tulodziecki:2007:BTE

[Tul07] Dana Tulodziecki. Breaking the ties: epistemic significance, bacilli, and underdetermination. Studies in History and Philosophy of Biological and Biomedical Sciences, 38(3):627–641, September 2007. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848607000337.

Tulodziecki:2011:CSE

[Tul11] Dana Tulodziecki. A case study in explanatory power: John Snow’s conclusions about the pathology and transmission of cholera. Studies in History and Philosophy of Biological and Biomedical Sciences, 42(3):306–316, September 2011. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848611000197.

Turner:2000:FFI

[Tur00] Derek Turner. The functions of fossils: inference and explanation in functional morphology. Studies in History and Philosophy of Biological and Biomedical Sciences, 31(1):193–212, March 2000. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848699000436.

Turney:2018:BRF

[Tur18] Jon Turney. Book review: Frankenstein, Annotated for Scientists, Engineers and Creators of All Kinds, Mary Shelley, David H. Guston, Ed Finn, Jason Scott Robert (Eds.). MIT Press, Campbridge, MA (2017), 320, Price $19.95, £14.95 paperback, ISBN: 978-0-262-53328-7. Studies in History and Philosophy of Biological and Biomedical Sciences, 67(??):41–43, February 2018. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848618300062.

Turda:2020:SAE

[Tur20] Marius Turda. Subversive affinities: Embracing Soviet science in late 1940s Romania. Studies in History and Philosophy of Biological and Biomedical Sciences, 83(??):Article 101131, October 2020. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848617301516.

Uller:2014:TED

[UH14] Tobias Uller and Heikki Helanterä. Towards an evolutionary developmental biology of cooperation? Studies in History and Philosophy of Biological and Biomedical Sciences, 48 (part B)(??): 267–271, December 2014. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848614001538.

Ulett:2014:MCO

[Ule14] Mark A. Ulett. Making the case for orthogenesis: the popularization of definitely directed evolution (1890–1926). Studies in History and Philosophy of Biological and Biomedical Sciences, 45 (??):124–132, March 2014. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848613001714.

Vaesen:2014:CRH

[Vae14] Krist Vaesen. Chimpocentrism and reconstructions of human evolution (a timely reminder). Studies in History and Philosophy of Biological and Biomedical Sciences, 45(??):12–21, March 2014. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848613001787.

Vail:2015:TAN

[Vai15] David D. Vail. Toxicity abounds: New histories on pesticides, environmentalism, and Silent Spring. Studies in History and Philosophy of Biological and Biomedical Sciences, 53(??):118–121, ???? 2015. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848615001077.

Valles:2012:LPC

[Val12] Sean A. Valles. Lionel Penrose and the concept of normal variation in human intelligence. Studies in History and Philosophy of Biological and Biomedical Sciences, 43(1):281–289, March 2012. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848611000860.

VanDijck:2000:DCV

[Van00] José Van Dijck. Digital cadavers: the Visible Human Project as anatomical theater. Studies in History and Philosophy of Biological and Biomedical Sciences, 31(2):271–285, June 2000. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848699000205.

Varga:2018:RNK

[Var18] Somogy Varga. “Relaxed” natural kinds and psychiatric classification. Studies in History and Philosophy of Biological and Biomedical Sciences, 72(??):49–54, December 2018. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848617301772.

Vaz:2016:STR

[Vaz16] Nelson M. Vaz. Self-tolerance revisited. Studies in History and Philosophy of Biological and Biomedical Sciences, 55(??):128– 132, February 2016. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848615001673.

vandenBelt:2002:LFC

[vdB02] Henk van den Belt. Ludwik Fleck and the causative agent of syphilis: sociology or pathology of science? A rejoinder to Jean Lindenmann. Studies in History and Philosophy of Biological and Biomedical Sciences, 33(4):733–750, December 2002. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848602000201. See [Lin01, Lin02].

vandenBerg:2013:WRK

[vdB13] Hein van den Berg. The Wolffian roots of Kant’s teleology. Studies in History and Philosophy of Biological and Biomedical Sciences, 44(4):724–734, December 2013. CO-DEN ???? ISSN 1369-8486 (print), 1879-2499 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S1369848613000988.

vandenBerg:2018:BBC

[vdB18] Hein van den Berg. A blooming and buzzing confusion: Buffon, Reimarus, and Kant on animal cognition. Studies in History and Philosophy of Biological and Biomedical Sciences, 72(??): 1–9, December 2018. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S136984861730225X.

vanderHoogte:2014:SSC

[vdHP14] Arjo Roersch van der Hoogte and Toine Pieters. Science in the service of colonial agro-industrialism: the case of cinchona cultivation in the Dutch and British East Indies, 1852–1900. Studies in History and Philosophy of Biological and Biomedical Sciences, 47 (part A)(??):12–22, September 2014. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848614000703.

vanderMeer:2008:ESF

[vdM08] Theo van der Meer. Eugenic and sexual folklores and the castration of sex offenders in the Netherlands (1938–1968). Studies in History and Philosophy of Biological and Biomedical Sciences, 39 (2):195–204, June 2008. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848608000162.

Vecchi:2020:DOD

[Vec20] Davide Vecchi. DNA is not an ontologically distinctive developmental cause. Studies in History and Philosophy of Biological and Biomedical Sciences, 81(??):Article 101245, June 2020. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848618301031.

Velasco:2008:SCS

[Vel08] Joel D. Velasco. Species concepts should not conflict with evolutionary history, but often do. Studies in History and Philosophy of Biological and Biomedical Sciences, 39(4):407–414, December 2008. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848608000794.

Vernon:2001:TTR

[Ver01] Keith Vernon. A truly taxonomic revolution? Numerical taxonomy 1957–1970. Studies in History and Philosophy of Biological and Biomedical Sciences, 32(2):315–341, June 2001. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848601000073.

Vermeir:2004:PPP

[Ver04] Koen Vermeir. The ‘physical prophet’ and the powers of the imagination. Part I: a case-study on prophecy, vapours and the imagination (1685–1710). Studies in History and Philosophy of Biological and Biomedical Sciences, 35(4):561–591, December 2004. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848604000652.

Vermeir:2005:PPP

[Ver05] Koen Vermeir. The ‘physical prophet’ and the powers of the imagination. Part II: a case-study on dowsing and the naturalisation of the moral, 1685–1710. Studies in History and Philosophy of Biological and Biomedical Sciences, 36(1):1–24, March 2005. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848604000986.

Verhoeff:2014:SAF

[Ver14] Berend Verhoeff. Stabilizing autism: a Fleckian account of the rise of a neurodevelopmental spectrum disorder. Studies in History and Philosophy of Biological and Biomedical Sciences, 46 (??):65–78, June 2014. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848614000405.

vanHelvoort:2000:DSC

[vH00] Ton van Helvoort. A dispute over scientific credibility: The struggle for an independent institute for cancer research in pre-World War II Berlin. Studies in History and Philosophy of Biological and Biomedical Sciences, 31(2):315–354, June 2000. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848699000308.

vanHelvoort:2014:VCS

[vH14] Ton van Helvoort. ‘Virus & cancer studies’ — still fascinating after all these years (2014). Studies in History and Philosophy of Biological and Biomedical Sciences, 48 (part B)(??):258–259, December 2014. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848614001368.

Vincent:2013:DBS

[Vin13] Bernadette Bensaude Vincent. Discipline-building in synthetic biology. Studies in History and Philosophy of Biological and Biomedical Sciences, 44(2):122–129, June 2013. CODEN ???? ISSN 1369-8486 (print), 1879-2499 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S1369848613000216.

vonLinne:1737:GPE

[vL37] Carl von Linné. Genera plantarum eorumque characteres naturales secundum numerum, figuram, situm, and proportionem omnium fructificationis partium.: Genera plantarum. (Latin) [The genera of plants and their characters natural according to the number, the figure, the site, and proportion of all the parts of fructification: The genera of plants]. C. Wishoff, Lugduni Batavorum [New York], 1737. 8 + 384 + 20 pp. LCCN QK91 .G4.

Vlerick:2014:BCD

[Vle14] Michael Vlerick. Biological constraints do not entail cognitive closure. Studies in History and Philosophy of Biological and Biomedical Sciences, 48 (part A)(??):21–27, December 2014. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848614000892.

vanRegenmortel:2016:MVL

[vR16] M. H. V. van Regenmortel. The metaphor that viruses are living is alive and well, but it is no more than a metaphor. Studies in History and Philosophy of Biological and Biomedical Sciences, 59(??):117–124, October 2016. CODEN ???? ISSN 1369-8486 (print), 1879-2499 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S1369848616300115.

Villagomez-Resendiz:2020:MSE

[VR20] Radamés Villagómez-Reséndiz. Mapping styles of ethnobiological thinking in North and Latin America: Different kinds of integration between biology, anthropology, and TEK. Studies in History and Philosophy of Biological and Biomedical Sciences, 84(??):Article 101308, December 2020.

CODEN ???? ISSN 1369-8486 (print), 1879-2499 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S1369848619301864.

vanWyhe:2013:SAR

[vW13] John van Wyhe. “My appointment received the sanction of the Admiralty”: Why Charles Darwin really was the naturalist on HMS Beagle. Studies in History and Philosophy of Biological and Biomedical Sciences, 44(3):316–326, September 2013. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848613000368.

vanWyhe:2016:IRW

[vW16] John van Wyhe. The impact of A. R. Wallace’s Sarawak Law paper reassessed. Studies in History and Philosophy of Biological and Biomedical Sciences, 60(??):56–66, December 2016. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848616300681.

vanWyhe:2015:GWO

[vWK15] John van Wyhe and Peter C. Kjærgaard. Going the whole orang: Darwin, Wallace and the natural history of orangutans. Studies in History and Philosophy of Biological and Biomedical Sciences, 51(??):53–63, June 2015. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848615000370.

Wells:1976:OHB

[W^{+}76] H. G. (Herbert George) Wells et al. The outline of history: being a plain history of life and mankind. Scholarly Press, St. Clair Shores, MI, USA, fourth edition, 1976. ISBN 0-403-07204-2. xxv + 1395 + 46 pp. LCCN D21 .W4 1976. Four volumes. Written originally with the advice and editorial help of Ernest Barker and others; illustrated with line drawings by J. F. Horrabin, and with reproductions of many famous paintings.

Walsh:2000:CSN

[Wal00] D. M. Walsh. Chasing shadows: natural selection and adaptation. Studies in History and Philosophy of Biological and Biomedical Sciences, 31(1):135–153, March 2000.

CODEN ???? ISSN 1369-8486 (print), 1879-2499 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S1369848699000412.

Waller:2001:IHR

[Wal01] John C. Waller. Ideas of heredity, reproduction and eugenics in Britain, 1800–1875. Studies in History and Philosophy of Biological and Biomedical Sciences, 32(3):457–489, September 2001. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848600000285.

Walsh:2006:ONP

[Wal06] D. M. Walsh. Organisms as natural purposes: The contemporary evolutionary perspective. Studies in History and Philosophy of Biological and Biomedical Sciences, 37(4):771–791, December 2006. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848606000768.

Walker:2008:CSS

[Wal08] Tom Walker. Could sexual selection have made us psychological altruists? Studies in History and Philosophy of Biological and Biomedical Sciences, 39(1):153–162, March 2008. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848607000970.

Walsh:2012:MPC

[Wal12] Denis Walsh. Mechanism and purpose: a case for natural teleology. Studies in History and Philosophy of Biological and Biomedical Sciences, 43(1):173–181, March 2012. CODEN ???? ISSN 1369-8486 (print), 1879-2499 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S1369848611000525.

Walsh:2013:DMS

[Wal13] Denis M. Walsh. Descriptions and models: Some responses to Abrams. Studies in History and Philosophy of Biological and Biomedical Sciences, 44(3):302–308, September 2013. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848613000927.

Wright:2017:HBR

[WCB17] Cory Wright, Matteo Colombo, and Alexander Beard. HIT and brain reward function: a case of mistaken identity (theory). Studies in History and Philosophy of Biological and Biomedical Sciences, 64(??):28–40, August 2017. CODEN ???? ISSN 1369-8486 (print), 1879-2499 (electronic). URL /science/article/pii/

S1369848617300614.

Weber:1998:RGC

[Web98] Marcel Weber. Representing genes: classical mapping techniques and the growth of genetical knowledge. Studies in History and Philosophy of Biological and Biomedical Sciences, 29(2):295– 315, September 1998. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S136984869800017X.

Weber:2002:TTE

[Web02] Marcel Weber. Theory testing in experimental biology: the chemiosmotic mechanism of ATP synthesis. Studies in History and Philosophy of Biological and Biomedical Sciences, 33 (1):29–52, March 2002. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848601000164.

Weber:2011:EED

[Web11] Bruce H. Weber. Extending and expanding the Darwinian synthesis: the role of complex systems dynamics. Studies in History and Philosophy of Biological and Biomedical Sciences, 42 (1):75–81, March 2011. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S136984861000107X.

Wainer:2020:FVE

[WEHT20] Juan Manuel Garrido Wainer, Juan Felipe Espinosa, Natalia Hirmas, and Nicolás Trujillo. Free-viewing as experimental system to test the Temporal Correlation Hypothesis: a case of theorygenerative experimental practice. Studies in History and Philosophy of Biological and Biomedical Sciences, 83(??):Article 101307, October 2020. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S136984862030056X.

Weindling:2000:OAR

[Wei00] Paul Weindling. An overloaded ark? The Rockefeller Foundation and refugee medical scientists, 1933–45. Studies in History and Philosophy of Biological and Biomedical Sciences, 31(3):477– 489, September 2000. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848600000157.

Weiskopf:2020:RCE

[Wei20] Daniel A. Weiskopf. Representing and coordinating ethnobiological knowledge. Studies in History and Philosophy of Biological and Biomedical Sciences, 84(??):Article 101328, December 2020. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848620301394.

Wells:1921:OHB

[Wel21] H. G. (Herbert George) Wells. The outline of history, being a plain history of life and mankind. Macmillan Publishing Company, New York, NY, USA, 1921. ???? pp. LCCN D21 .W4 1920a.

Wells:1926:OHB

[Wel26] H. G. (Herbert George) Wells. The outline of history; being a plain history of life and mankind. Macmillan Publishing Company, New York, NY, USA, 1926. ???? pp. LCCN D21 .W4 1926a.

Wells:1931:OHB

[Wel31] H. G. (Herbert George) Wells. The outline of history, being a plain history of life and mankind. A Star book. Garden City Publishing Co., Inc., Garden City, NY, USA, 1931. xxi + 1 + 1174 pp. LCCN D21 .W4 1931a.

Wells:1956:OHB

[Wel56] H. G. (Herbert George) Wells. The outline of history, being a plain history of life and mankind. Garden City Books, Garden City, NY, USA, 1956. xvii + 966 pp. LCCN D21 .W4 1956.

Weldon:2014:MB

[Wel14] Stephen P. Weldon. Monkey business. Studies in History and Philosophy of Biological and Biomedical Sciences, 48 (part A)(??):115–118, December 2014. CODEN ???? ISSN 1369-8486 (print), 1879-2499 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S136984861400079X.

Wells:2020:KLE

[Wel20] Aaron Wells. Kant, Linnaeus, and the economy of nature. Studies in History and Philosophy of Biological and Biomedical Sciences, 83(??):Article 101294, October 2020. CODEN ???? ISSN 1369-8486 (print), 1879-2499 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S1369848618301729.

Wexler:2008:OCP

[Wex08] Rebecca Wexler. Onward, Christian penguins: wildlife film and the image of scientific authority. Studies in History and Philosophy of Biological and Biomedical Sciences, 39(3):273– 279, September 2008. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848608000368.

Winther:2015:MLF

[WGEN15] Rasmus Grønfeldt Winther, Ryan Giordano, Michael D. Edge, and Rasmus Nielsen. The mind, the lab, and the field: Three kinds of populations in scientific practice. Studies in History and Philosophy of Biological and Biomedical Sciences, 52(??):12–21, August 2015. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848615000114.

Walsby:2017:SCS

[WH17] A. E. Walsby and M. J. S. Hodge. Schrödinger’s code-script: not a genetic cipher but a code of development. Studies in History and Philosophy of Biological and Biomedical Sciences, 63(??):45– 54, June 2017. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848616301042.

White:2004:DHH

[Whi04] Paul White. Desmond/Huxley: the hot-blooded historian. although his world view ultimately sank into orthodoxy, he never lost his love of battle.: Huxley: From devil’s disciple to evolution’s high priest, Adrian Desmond; Penguin, London, 1998, pp. xxii + 820, Price £10.99 paperback, ISBN 0-14-017309-9. Studies in History and Philosophy of Biological and Biomedical Sciences, 35 (1):191–198, March 2004. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848603001195. See [Des98].

Whitney:2014:DNS

[Whi14] Kristoffer Whitney. Domesticating nature?: Surveillance and conservation of migratory shorebirds in the “Atlantic Flyway”. Studies in History and Philosophy of Biological and Biomedical Sciences, 45(??):78–87, March 2014. CO-DEN ???? ISSN 1369-8486 (print), 1879-2499 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S1369848613001337.

Widmer:2014:MBM

[Wid14] Alexandra Widmer. Making blood ‘Melanesian’: Fieldwork and isolating techniques in genetic epidemiology (1963– 1976). Studies in History and Philosophy of Biological and Biomedical Sciences, 47 (part A)(??):118–129, September 2014. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848614000521.

Wilson:1998:BRS

[Wil98] Bradley E. Wilson. Book review: Sociobiology, sex, and science: Holcomb, H. R., (Albany, NY: State University of New York Press, 1993), x + 447 pp., ISBN 0-7914-1260-1 paperback. Studies in History and Philosophy of Biological and Biomedical Sciences, 29(1):201–210, March 1998. CODEN ???? ISSN 1369-8486 (print), 1879-2499 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S1369848698000090. See [Hol93].

Wilson:1999:SCA

[Wil99] Fred Wilson. Some controversies about method in nineteenthcentury psychology. Studies in History and Philosophy of Biological and Biomedical Sciences, 30(1):91–127, March 1999. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848698000259.

Wilkinson:2000:BVG

[Wil00] Lise Wilkinson. Burgeoning visions of global public health: The Rockefeller Foundation, The London School of Hygiene and Tropical Medicine, and the ‘hookworm connection’. Studies in History and Philosophy of Biological and Biomedical Sciences, 31(3): 397–407, September 2000. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848600000182.

Wilkie:2001:GPW

[Wil01] Andrew O. M. Wilkie. Genetic prediction: What are the limits? Studies in History and Philosophy of Biological and Biomedical Sciences, 32(4):619–633, December 2001. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848601000280.

Wilson:2002:APE

[Wil02a] David A. H. Wilson. Animal psychology and ethology in Britain and the emergence of professional concern for the concept of ethical cost. Studies in History and Philosophy of Biological and Biomedical Sciences, 33(2):235–262, July 2002. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368102000134. See comments [Wil03].

Wilson:2002:DCE

[Wil02b] David Sloan Wilson. Darwin’s cathedral: evolution, religion, and the nature of society. University of Chicago Press, Chicago, IL, USA and London, UK, 2002. ISBN 0-226-90134-3. v + 268 pp. LCCN BL60 .W544 2002. URL http://www.

loc.gov/catdir/bios/uchi051/2002017375.html;   http:
//www.loc.gov/catdir/description/uchi052/2002017375.
html;   http://www.loc.gov/catdir/toc/uchi051/2002017375.
html.

Wilson:2003:APE

[Wil03] D. A. H. Wilson. Animal psychology and ethology in Britain and the emergence of professional concern for the concept of ethical cost [Studies in History and Philosophy of Biological and Biomedical Sciences, 33C/2 (2002), 235–261]. Studies in History and Philosophy of Biological and Biomedical Sciences, 34(1):201, March 2003. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848602000766. See [Wil02a].

Wilner:2006:DAS

[Wil06] Eduardo Wilner. Darwin’s artificial selection as an experiment. Studies in History and Philosophy of Biological and Biomedical Sciences, 37(1):26–40, March 2006. CO-DEN ???? ISSN 1369-8486 (print), 1879-2499 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S1369848605000968.

Wilmot:2007:BFC

[Wil07a] Sarah Wilmot. Between the farm and the clinic: agriculture and reproductive technology in the twentieth century. Studies in History and Philosophy of Biological and Biomedical Sciences, 38 (2):303–315, June 2007. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848607000118.

Wilmot:2007:PSA

[Wil07b] Sarah Wilmot. From ‘public service’ to artificial insemination: animal breeding science and reproductive research in early twentieth-century Britain. Studies in History and Philosophy of Biological and Biomedical Sciences, 38(2):411–441, June 2007. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848607000179.

Williams:2012:SAE

[Wil12a] Elizabeth A. Williams. Sciences of appetite in the Enlightenment, 1750–1800. Studies in History and Philosophy of Biological and Biomedical Sciences, 43(2):392–404, June 2012. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848611001233.

Wilson:2012:SCE

[Wil12b] Edward O. Wilson. The social conquest of Earth. Liveright Pub. Corporation, New York, NY, USA, 2012. ISBN 0-87140-413-3 (hardcover). viii + 330 pp. LCCN GN360 .W53 2012.

Wimsatt:2013:ABA

[Wim13] William C. Wimsatt. Articulating Babel: an approach to cultural evolution. Studies in History and Philosophy of Biological and Biomedical Sciences, 44(4):563–571, December 2013. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848613001167.

Winsor:2001:CLS

[Win01] Mary P. Winsor. Cain on Linnaeus: the scientist–historian as unanalysed entity. Studies in History and Philosophy of Biological and Biomedical Sciences, 32(2):239–254, June 2001. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848601000103.

Winter:2013:RFF

[Win13] Alison Winter. The rise and fall of forensic hypnosis. Studies in History and Philosophy of Biological and Biomedical Sciences, 44 (1):26–35, March 2013. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848612001057.

Witteveen:2015:TOM

[Wit15] Joeri Witteveen. “A temporary oversimplification”: Mayr, Simpson, Dobzhansky, and the origins of the typology/population dichotomy (part 1 of 2). Studies in History and Philosophy of Biological and Biomedical Sciences, 54(??):20–33, December 2015. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848615001351.

Witteveen:2016:TOM

[Wit16] Joeri Witteveen. “A temporary oversimplification”: Mayr, Simpson, Dobzhansky, and the origins of the typology/population dichotomy (part 2 of 2). Studies in History and Philosophy of Biological and Biomedical Sciences, 57(??):96–105, June 2016. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S136984861500134X.

Wilson:2008:MWM

[WL08] Duncan Wilson and Gaël Lancelot. Making way for molecular biology: institutionalizing and managing reform of biological science in a UK university during the 1980s and 1990s. Studies in History and Philosophy of Biological and Biomedical Sciences, 39 (1):93–108, March 2008. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848607000921.

Winther:2015:IGP

[WMN15] Rasmus Grønfeldt Winther, Roberta L. Millstein, and Rasmus Nielsen. Introduction: Genomics and philosophy of race. Studies in History and Philosophy of Biological and Biomedical Sciences, 52(??):1–4, August 2015. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848615000564.

Wolfe:2014:OOG

[Wol14] Charles T. Wolfe. The organism as ontological go-between: Hybridity, boundaries and degrees of reality in its conceptual history. Studies in History and Philosophy of Biological and Biomedical Sciences, 48 (part B)(??):151–161, December 2014. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848614000806.

Wong:2020:ECN

[Won20] T. Y. William Wong. Evolutionary contingency as nontrivial objective probability: Biological evitability and evolutionary trajectories. Studies in History and Philosophy of Biological and Biomedical Sciences, 81(??):Article 101246, June 2020. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848619300640.

Woods:2007:FCV

[Woo07] Abigail Woods. The farm as clinic: veterinary expertise and the transformation of dairy farming, 1930–1950. Studies in History and Philosophy of Biological and Biomedical Sciences, 38(2):462– 487, June 2007. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848607000192.

Wood:2015:DEH

[Woo15] Roger J. Wood. Darbishire expands his vision of heredity from Mendelian genetics to inherited memory. Studies in History and Philosophy of Biological and Biomedical Sciences, 53(??):16–39, ???? 2015. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848615000813.

Worboys:2007:WTB

[Wor07] Michael Worboys. Was there a bacteriological revolution in late nineteenth-century medicine? Studies in History and Philosophy of Biological and Biomedical Sciences, 38(1):20–42, March 2007. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848606000902.

Wouters:2003:FNB

[Wou03] Arno G. Wouters. Four notions of biological function. Studies in History and Philosophy of Biological and Biomedical Sciences, 34(4):633–668, December 2003. CODEN ???? ISSN 1369-8486 (print), 1879-2499 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S136984860300075X.

Wells:1971:OHB

[WPW71] H. G. (Herbert George) Wells, Raymond Postgate, and G. P. (George Philip) Wells. The outline of history, being a plain history of life and mankind. Doubleday, Garden City, NY, USA, new revised edition, 1971. xii + 1103 pp. LCCN D21 .W4 1971. Two volumes. Revised and brought up to date by Raymond Postgate and G. P. Wells. With maps and plans by J. F. Horrabin.

Wringe:2011:CIC

[Wri11] Bill Wringe. Cognitive individualism and the child as scientist program. Studies in History and Philosophy of Biological and Biomedical Sciences, 42(4):518–529, December 2011. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848611000550.

Worboys:2020:IMD

[WSP20] Michael Worboys, Julie-Marie Strange, and Neil Pemberton. The Invention of the Modern Dog by Michael Worboys, Julie-Marie Strange, and Neil Pemberton: Reply by the authors. Studies in History and Philosophy of Biological and Biomedical Sciences, 84(??):Article 101331, December 2020. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848620301424. See [Der20, Ank20].

Wylie:2010:SSS

[Wyl10] Caitlin Donahue Wylie. Setting a standard for a “silent” disease: defining osteoporosis in the 1980s and 1990s. Studies in History and Philosophy of Biological and Biomedical Sciences, 41(4):376– 385, December 2010. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848610000774.

Yale:2011:MCC

[Yal11] Elizabeth Yale. Marginalia, commonplaces, and correspondence: Scribal exchange in early modern science. Studies in History and Philosophy of Biological and Biomedical Sciences, 42(2):193–202, June 2011. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848610001238.

Yan:2017:BNS

[YH17] Karen Yan and Jonathon Hricko. Brain networks, structural realism, and local approaches to the scientific realism debate. Studies in History and Philosophy of Biological and Biomedical Sciences, 64(??):1–10, August 2017. CODEN ???? ISSN 1369-8486 (print), 1879-2499 (electronic). URL /science/article/

pii/S1369848617300808.

Yi:2018:TIP

[Yi18] Doogab Yi. Taming intellectual property in biotechnology. Studies in History and Philosophy of Biological and Biomedical Sciences, 68–69(??):73–77, April/June 2018. CO-DEN ???? ISSN 1369-8486 (print), 1879-2499 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S1369848618300530.

Yoeli-Tlalim:2010:TWW

[YT10] Ronit Yoeli-Tlalim. Tibetan ‘wind’ and ‘wind’ illnesses: towards a multicultural approach to health and illness. Studies in History and Philosophy of Biological and Biomedical Sciences, 41(4):318– 324, December 2010. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848610000671.

Yu:2017:CPR

[Yu17] Xiaobo Yu. Chinese paleontology and the reception of Darwinism in early twentieth century. Studies in History and Philosophy of Biological and Biomedical Sciences, 66(??):46–54, December 2017. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S136984861730050X.

Zammito:2006:TTN

[Zam06] John Zammito. Teleology then and now: The question of Kant’s relevance for contemporary controversies over function in biology. Studies in History and Philosophy of Biological and Biomedical Sciences, 37(4):748–770, December 2006. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848606000756.

Zammito:2012:LTR

[Zam12] John H. Zammito. The Lenoir thesis revisited: Blumenbach and Kant. Studies in History and Philosophy of Biological and Biomedical Sciences, 43(1):120–132, March 2012. CODEN ???? ISSN 1369-8486 (print), 1879-2499 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S1369848611000471.

Zawidzki:2015:CDD

[Zaw15] Tadeusz Wieslaw Zawidzki. As close to the definitive Dennett as we’re going to get. Studies in History and Philosophy of Biological and Biomedical Sciences, 50(??):98–102, April 2015. CODEN ???? ISSN 1369-8486 (print), 1879-2499

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848615000266.

Zimmer:2011:CPB

[ZDSF11] Karl Günter Zimmer, Max Delbrück, Phillip R. Sloan, and D. Brandon Fogel, editors. Creating a physical biology: the Three-Man Paper and early molecular biology. University of Chicago Press, Chicago, IL, USA and London, UK, 2011. ISBN 0-226-76782-5 (hardcover), 0-226-76783-3 (paperback). 319 pp. LCCN QH506 .C73 2011.

Zimmerman:2001:LBH

[Zim01] Andrew Zimmerman. Looking beyond history: the optics of German anthropology and the critique of humanism. Studies in History and Philosophy of Biological and Biomedical Sciences, 32(3): 385–411, September 2001. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848601000176.